A Bibliography of IEEE Transactions on Computers (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2025 Version 1.23

Title word cross-reference

(2m ± 1) [HGS83b]. (d, k) [MR82]. (N, K) [KK86]. 1 [wu87b]. 1/2 [ZW87]. 2 [FLN89, Jes80a, KR82, RY82]. 2(log_{2} N) − 1 [Lee85b]. 2/3 [ZW87]. 22n + 1 [STR87]. 2m [PNH88]. 3 [WH88]. 4 [KW81, TPSS85]. 5.8n log_{2} n [CW80]. 8 [Kal83]. [lg N + 1] [CT84a]. ^{2} [DD81, Eti80, FM89]. ^{2.5} [DM84]. 2p [SLS82]. ^{n} [HSE84]. AB [Bla83, Slo85]. D [Kak85, Kak83, LB88, SW82, WV87, WR81]. d > t [LB88]. ℓ [Hoc83]. GF (2^{m}) [Fen89, WTS^{+}85, YRT84, Zho88]. GF (P) [TKL86, HQR89]. GF [p^{n}] [Eng81]. K [GM87, NHAT89, Agg86, BL84, CI88, CE87, Dav89, Lee82, LC87b]. L/U [JK82]. λT [CGMP87]. M [Bla83, Mor80, Slo85, WH80a, Wus81, GH83b, HGS83b, PNH88]. m2 [MM83a]. N [CS87, Er84, Esf89, Mor80, Wus81, AP89, AG81b, Bha83a, GH83b, Kha82, LM87a]. n − 1/n [Nil84, Wen85, Can83]. O(2^{0.304n}) [Jia86]. O(log n) [BP85]. O(n) [Sip82]. O(n + k) [LD88]. O(N^{2}/ log^{3/2} N) [SR81a]. O(t^{3} + |E|) [Sul88]. P [Tha82, Tha84b, TKL86, YF88, Fro83, GNK86, KM83b, Wus82]. r [GH83a, HGS83a]. s^{2} [SSS89]. t [BV89, CA89, LB88, NGP86].

-Adic [Fro83, GNK86, KM83b]. -an [SGI89]. -ary [Er84]. -Bit [Kal83, Wus81, Mor80]. -Cube [Bha83a, AP89, CS87, Esf89]. -D [FLN89, KR82, wu87b]. -EC [LB88]. -Error [NGP86, BV89, LB88]. -Function [Tha84b]. -Functions [Tha82]. -Graph [WR81]. -Level [Wus82]. -Nearest [Lee82]. -Networks [MM83a]. -optimal [FM89]. -Ordered [WV87]. -Out-of- [GH83b, HGS83b, PNH88]. -out-of-n [BL84]. -Ranges [SW82]. -Sequence [WH80a]. -Sequences [Kak83]. -TDA-diagnosable [CA89]. -UED [LB88]. -unidirectional [LB88]. -Valued [KW81, YF88]. -way [SSS89, NHAT89].

1-out-of- [Kha82]. 1-out-of-3 [Gol84]. 10th [Ano80a]. 11 [LGH80]. 11/780 [AP85]. 1981 [Ano82-28, Ano82-29, Ano82-59]. 1982 [Ano82-30, Ano82-31, Ano82-32]. 1983 [Ano82x, Ano82u, Ano82w, Ano82b, Ano82c]. 1987 [Ano87z].

2 [KP88]. 2-Dimensional [LL84]. 249 [Ano80c]. 254 [Ano80d, Ano80e, Ano80f]. 256 [Ano80g]. 256K [Tan84]. 29 [Ano80b].

30 [Ano81a]. 31 [Ano82a]. 32 [Ano83a]. 32-bit [MBH89]. 33 [Ano84a]. 35 [Ano86b]. 36 [Ano87a]. 370 [Kob83, PMSB88]. 3rd [Ano82j].

4th [Ano80c, Ano83j].

64K [IC80].

780 [AP85]. 7th [Ano82d].

80 [Ano80h, Ano80i, Ano80j, Ano80k]. 81 [Ano81b, Ano81d]. 82 [Ano81c, Ano82e, Ano82g, Ano82h]. 83 [Ano82f, Ano83l].

A. [EA89a]. A/D [Yue80]. Abstract [MP83a, PL88, SBSM89, Wei88]. Abstractions [OK83, Oik87]. Acceleration [FS83]. accelerator [Mel89]. Acceptable [Cli80a]. Access [Bra83b, Bux83, CP82a, CP87b, CHL83, FT84, GFM83, GF87, HJ87, IS84, KP80a, KL82, LV82, MTMA85, PS85a, RG85, SN81, SH84, SR80, SR81b, Ten83, TV82, Cal88, CJ89, FHG88, GC89, KP88, MP89a, MG88, MW88, NK88a, PBL89].

Accessibility [AL85]. Accessing [ABK83]. Accumulator [SR86]. Accurate [LPI88, VM87b, GY89]. Achieved [Mar81]. Acknowledgment [Ano83c]. Acquisition [Cli80b]. ACRITH [Lan87]. Active [LM85, Uhr82]. Activity [ST86a, VZMBH89]. ADA [KUV85, VM87a]. Adaptive [HN84, JL85, Kum89, TRTN89]. add [BD89, Sto84]. Addendum [RTB81, Rob83]. Adder [MH80, Dor88]. Adders [BK82, CP87a, Cur80c, GHM87, LM82a, SM83, Bec88]. Adding [LD87]. Addition [FG82, Par88, TYY85]. Additional [Lau81]. Address [Nwa85]. Addressable [KK84b, SS82]. Addressing [De 82b, YTL87]. Adic [Fro83, GNK86, KM83b]. Adjacencies [HSE84]. Adjacent [RS85]. Adjusting [GC87]. Advance [Ano80h, Ano80i, Ano80j, Ano80k, Ano81b, Ano81d, Ano81c, Ano82e, Ano82g, Ano82h, Ano82f]. Advantage [NN87]. Advertisement [Ano84b, Ano84c]. Aerospace [PS83a]. Against [BS82]. Aggregation [BT86]. Ahead [KY89, Dor88]. Aided [RW84, Sin89]. Algebra [Sur81]. Algebraic [Che82b, Hon85, SY83, AL88, BC88, NH88]. Algebras [DB87, LL80]. Algorithm [ACGT84, ASP87, APD83, AT81, BA86, BSV83, BW80, BB87b, BS87a, Bla83, Car84, CE87, Che82a, CH85a, CP87b, DM84, Dah86, FTY87, Gaj81, GC80, Gel81, Goe81, GM87, HR87, HKR84, Hou87b, Hou87a, HRT^{+}84, Hua83, HA84, Jen81, Jia86, Jou86, Kam80, Kar84a, KK85, Kri84, KH83, Lak84, Lam83, LS82a, Lee87, LC87b, MTMA85, Mey84, MP83c, MI85, ML82, NN86, Niz84b, Niz84a, NS87, OE82, Pal86, Par81, RW81b, RB83b, RTB81, Rot86, Sas85, SBGS86, Slo85, Sto80b, WG80, WOH84, YM86b, AA88, AL88, ED88, JAM88, KPR88, LD88, LS88a, LD89, Man89, NH88, Pin89, RI89, RJ89, RS88b, San88a, San88b, Sin89, STR88, Sri89, SRC85, Sul88, TYY85, TR88, Wan89, WZ89, lYM88]. Algorithm-Based [BA86, HA84, AL88]. Algorithmic [Hoc87, Ram86a, Ram86b, ST86a, VZMBH89]. Algorithms [Ano85d, AN84, AS87b, Bry86, But81, Car83, CS86a, CL84, CF80a, CW83, Coy80, DL86, DK86, DS83, Dem82, DV87, DL87, Eng81, FS83, GV84, GK85, Gue86, Hon81, HC82, IKP86, JA85, KR82, KN84, Lam83, LM82b, LW86a, Li87, Lo88, LR80, Maj85, Man85, MP87, Mol82a, MF86, Nak86, NS81, NS82b, OP84, OOB85, PR81b, RJ82, SW84, Tha84a, Tha84b, Tsa81, Tsa83, VR86, Wei82, YGZ^{+}87, Zak85, AOES88, CI88, Che88b, CS89, Cra85, Fis88, Gel89, KT89b, LP89, MS88b, MS89, Muk89, PK88, Ram89, SS89c, WE83]. Alignment [FYAV87, LS88b]. All-to-All [Top89]. Allocation [CP82a, CS87, CL87, Cve87, LL83, MLT82, PB87, RW83, Sig82, Weg80, GC89, Gel89, KS89, San88a, San88b]. Allocations [KK84b]. ALOHA [Lam80, Tas83]. Alternative [Pat80, SD86a]. alternatives [ALL89]. ALU [PF82, Wu87a]. Ambiguity [Muk86]. American [Men84]. Among [DLC87, KS80a, KL82, YN83, dM88]. Analog [Cli80b, FTT^{+}80, PC81, Sto80b, THH80]. Analog-to-Digital [FTT^{+}80]. Analysis [AB80, AB82, ASKL81, AJ82, Agr83, AKS88, AO80, Ano81r, ASP87, AL81, AP85, BYH87, Bhu85, BT86, BFHW82, Bux83, CP82b, CI87, CFH81, Dav86, De 82b, DJ81, DK82, FH84, GW81, Har86, HI80, HS84, HT86, IO84, IDH86, JM87, KK87, KA89, KKS86, Lam80, Lee80, LV85, LC87a, MBC82, MTS89, Mol82a, Mol82b, Muz80, NS87, Pat82, Pra86b, PC81, SC87, Smi85, Smi80a, SS82, Sor85, TN81, TJB86, Tsa81, Tsa83, Wan82, WC83, WA85a, YA87, YBL89, BD89, BGY89, BL89, BI88, Cal88, CCF89, tCpS88, CMB88, CdL89, Coo89, Dub88, DT89, Fuk88, GH88, Gui89, HTT89, JS88, KR89b, Max88, NT88, Noe89, RKH88, STR88, ST88b, SL89, WE83, YKL88]. Analytic [HT83, RN88, SS80, TB86, Sch89]. Analytical [CM87b]. Analyzer [Hla86]. analyzers [Max88]. Ancestors [Tsi86]. AND-OR [LS89c]. Announcement [Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81b, Ano81d, Ano81c, Ano82i, Ano82e, Ano82g, Ano82h, Ano82f]. Announcing [Ano82j]. Annual [Ano82-60]. Anomalies [LW86a]. Anomalous [LMP82]. API [SK85, dJvdG88]. Application [Ano85y, Ano87x, CH82a, CL80, DTF80, KH83, LR80, SCP^{+}81, SL86, Sta85, TS84a, VHD82, Zhi84, Esf89, Kap89, NK88b, RN88, SL89, Thi89]. Applications [Ano80x, Ano81u, Ano81v, Ano82u, Ano82v, Ano82w, BCR83, CSR86, Cha80, CLH84, HZ81, Hua83, IC80, Jul80, KF80, Kan84, KK84c, Kon86, KAGER82, Lun87, MLB87, Pra80, PTC86, SLS82, Smi81a, Sto84, KW89, Kum88, LM88b]. Approach [Abr82, AM85, AKT86, Agr81, AGH^{+}82, ABT82, BSMS81, Bes86, Bha83a, Bha83b, Bra83b, CH85b, DM81, DSK87, Fla82, GM82b, IPM82, JM87, Kar81, KS80b, Kor86, LP83, MG86, NKY^{+}80, Pap83, RT85, RH87, Ros83, ST85, Sig82, SH81b, Sto80a, Til80, Tsa83, WC84, Weg80, WWS84, AK88, BW89a, JW89a, KK89, KW89, KS89, LZM89]. Approaches [SSF82, ST86a, LLJ89, VZMBH89]. Approximate [BGY89, NT88, Tow86]. Approximating [LR80]. Approximation [DLM86, Har86, KK85, Niz83]. Approximations [SGT86, Aky89]. Arbiter [BJ83, CAV86, SL89]. Arbitrary [HML84, WG80]. Arbitration [Cha87b, Gui89]. Architectural [Ano87c, HS88, RM83, VSHM82]. Architecture [AJ82, AAG^{+}87, Ano81r, Ano82-60, AG82, BS86, BP84, BFHW82, CP82b, FH83, GBG89, GS86, HA86, Haw85, Hen84, Hon85, HG87, IPM82, JTP85, KK82, KA87, KAGER82, Liu82, Liu84a, MLB87, PS83a, PR82, RM83, SBK85, SSB87, Shi82b, SWP86, SM82, SWK84, SG82, TL80, TRYS83, TLR83, VZ81, VBH^{+}81, WM84, CNO^{+}88, Con89, EA89b, Fen89, GW89, HTDR88, KHS88, KWFT88, MG88, NOYK88, PMSB88, Pol88, SGI89, TSM88, Tys88, MPPZ88]. Architectures [BPM^{+}86, DGS80, GS87, GK85, KK80a, LF80, Lop84, LF87, MBC82, MTG^{+}85, NJM83, NF84, NS87, Pra85a, Pra85b, Pra86a, RAE84, Red87a, Sei84, Str82, Veg84, WTS^{+}85, HA88, RTY^{+}88, SS89c, YJ89, YS89]. Area [Ame82, BPV83, BP85, Bux83, CGMP87, FT84, Kam87, KM87, KB84, LHPW85, LSW87, Pre83, RG85, WHT84, CBAP89, HA88, KP88, Ram88, RS88b, Sin88b, SR81a]. Area-Time [Pre83, WHT84]. Arithmetic [Agr80, AR83, Ano84f, AP85, BM86a, BR83a, Che82b, CVY83, CHH83, DDG80, GH83a, Gai85a, GNK86, HGS83a, HT80, HL86, HC82, Jen83, KUV85, KM84b, LZLH85, MK85, NJM83, NH85, Owe83, OI87b, SOA85, Sip84, SG80, Swa80, SCNS83, Tay82, TPSS85, Ulm83, ZB87, ZN80, Cos88, KM83a, Loz83, Man88, NPP88, Sco89]. ARQ [CL88]. Array [Agg86, AG81b, BVH83, Bha83a, Bok84, CIP87, Chu85, DV87, Ers85, Fla82, FHH^{+}83, FYK87, Jes80a, Jes80b, KH86, KAGER82, LV82, Len85, LY83, LA85, MG86, Nak86, Nwa85, Pap83, PBL89, RV84, Ree80, SF84, TOM81, Ten83, VRF84, WBA83, Zak84, Cos88, FLN89, Hos89, KJC89, PK88, ST88a, SL88a, WCS89]. Arrays [AC84, Aga80, AS85, BM86b, BNM86, BK84, DM81, ES80, FK85, FR85, FK81, GKS84, Gor87, JMKN86, KB84, Kor86, KS83a, LL85, LW85, MT87, MF86, OS87, O’L87, OP84, Ozg86, PR81a, PF83, RT85, RFS86, RB83a, RB83b, Ros83, Ros85b, SKF83, Sas81, SM87, UV81, Uhr82, VR86, VS86a, YWW86, ZH85, Ata88, FS88, JK89, KR89a, KT89b, LLJ89, LJL89, LP88, Man89, OT89, SR88a, Sin88b, VR89b].

Arrival [CS85]. Arrivals [Sha81]. Art [Kar82]. ary [Er84]. ASCA [NOYK88]. Aspects [Fen85, Ram86a, Ram86b]. Assertion [PS83b]. Assigning [MTMA85]. Assignment [FT84, FYK87, Lau81, LL84, Lo88, NKY^{+}80, PTT81, Rub81, Sas84, ST85, BS89a, CD88, CS89, DIN88, KK89, ST88a]. Assignments [Sin88a]. Associative [DL86, DGS80, Fet80, HNS84, SD87, Str82, CCWZ88, LD88]. associativity [HS89]. Astronomical [EHS80]. Asymmetric [HS81, Shi82a]. Asymptotically [KP80a]. Asynchronous [BM86a, CAV86, CG87, Hay81, HT82, Hol82a, Tha84a, Tha84b, WF83, WF80a, BC88, Coa88, San88a]. atomic [Blo88]. ATP [KP88]. ATP-2 [KP88]. attenuator [CG88]. Augmented [AS82b, MS82, LL88]. August [Ano87z]. Author [BH80, Des81, GLL81, Mil85, Nil86, RTW80, Ros86, SM80, Smi86, SCA81, SCN86, Weg82]. Authors [Ano80-45, Ano82-54, Ano82-55, Ano83-32, Ano83-33, Ano83-34, Ano83-35, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano84z, Ano84-27, Ano84-28, Ano84-29, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano85w, Ano86u, Ano86v, Ano86w, Ano86x, Ano86y, Ano86z, Ano86-27, Ano86-28, Ano86-29, Ano86-30, Ano86-31, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v]. Autocorrelation [Mor80, Wus81, Wus82]. Automata [PTC86, Wan81, WR81, HCMP89, HMC89, OM88, SKW88, ZMC89]. Automated [HP82, KW85, WW83]. Automatic [BCDM86, KS82a, SG83, ST86a, TJB86, WW83, VZMBH89]. Autonomous [MBN81]. Autoscale [TH82]. availability [GT88, MDG89]. Available [NF84]. AVL [Got81, Ell80]. Avoiding [Fos89]. AXE [OJ80].

Back [Car83]. Back-to-Back [Car83].

backend [BBW88]. backpointers [LCF89]. backtracking [JAM88]. Balanced [MI85, Vai84, Vai86, Wan82, Vai89]. Balancing [ABK83, BW89b, MI85]. Bandwidth [AM87, LVA82, LY83, LJ87, BN88]. Bank [Bai87]. Banyan [Fra81, Lea86, CM89, GH88, RK89]. Banyans [CM85]. Bar [SB83]. Base [LF80, LC87b, SK89]. Base- [LC87b]. Based [AB80, AB82, BA86, Bry86, CCSW85, CI87, CK87, CHM87, Gel81, GF87, HR87, HA84, IKI83, IR86, LC87a, Lop84, Mar86, MST85, MLB87, MG86, NMB83, NWLP82, Pap83, RFLS86, SS80, Sny81, SGMP85, WF83, WSR84, AL88, Dub88, GM88a, HIT88, LCS89, MKLC89, NOYK88, PS88, SBSM89, SD89b, Wei88]. bases [HTDR88]. Basis [DM83a, Mul85, OA83, Wan89]. Batcher [KH83]. Bayesian [Sta85]. BDDL [Sav86]. be [Fis88, Mar81]. BED [Dun85]. Behavior [Fen85, KC87a, LMP82, MP83a, Sha81, Wei82, BBG88]. behavioral [GY89]. Benchmark [WM88b, FM88]. Benchmarks [Men84]. Benes [Lee87, NS81, NS82b]. Berger [Don84, Pie87]. Berlekamp [HRT^{+}84]. Between [CW83, EDH80, HSE84, Kan85, DIY88, Muk89, WLC89, YF88]. Beyond [Moh85]. BGM [Mey84]. Bidirectional [Cle84, LP81, Sav86]. Bilateral [VS86a]. Bin [Cha83b]. Bin-Packing [Cha83b]. Binary [AR86, Abo88, Agr80, BDL83, BA82, Cha87a, Chu85, Cur80c, DS83, GH83a, Gna85, Gor87, HGS83a, HA86, HZ81, KK84a, LA85, MI85, NN87, RAE84, Ree80, SM83, Sas85, TY87, ZW87, AP89, BB89a, CJ89, HA88, Par88, SB88, SS89c, TYY85, Vai89, YS89]. Binary-Coded [GH83a, HGS83a]. Bisectional [GBG89]. Bit [BCR83, Bat82, Bis85, CL80, Gna83, HRT^{+}84, Kal83, Mor80, RB80, RS86b, Sip82, Sip84, SH81a, SH81b, SR82, TGJR88, VBH^{+}81, Van86, VSH89, Wu87a, Wus81, BS88, MBH89, Smi89a, Zho88]. Bit-Sequential [RS86b, Sip82, Sip84, SR82, Smi89a]. Bit-Serial [Bat82, Gna83, HRT^{+}84, RB80, BS88, Zho88]. Bit-Sliced [SH81a, SH81b, VBH^{+}81, Wu87a]. Bitonic [BP84, NHAT89]. Block [BAR87, De 82b, II81, Smi87, WK80, PBL89, USM89]. Block-Parallel [WK80]. Block-Structured [De 82b]. Blocking [Aky87, LJ87, Aky89]. Blocks [LS84, CC89, KW89]. Boards [NKY^{+}80]. Boolean [ATT81, AG81b, Bha83a, Bry86, FGPT89, Hay86a, Kuo87, LL80, LR80, Mar86, MTG83, RS84, Sas81, TT80, TTB85]. Born [BCA80]. Bose [OI87a]. bottleneck [Gel89]. Bottom [Cha83b]. Bottom-Left [Cha83b]. Bound [BD80, DC82, LW86a, GM88b, Lee89b, Sto89]. Bounded [IK82a, PP87, WB88]. Bounding [AM80, Mar87, MDG89, Pro88]. Bounds [Agg86, BA86, DO85, DO86, HL82, Lei85a, NM82, PB87, RS83b, Sri87, Loz83, PBL89]. BPC [NS82a]. Branch [ED87, LW86a]. Branch-and-Bound [LW86a]. break [Gef73]. Breakdowns [MK83b]. Bridging [AM85, Kar83, KS80a, KP80b, XS85, YN83, YN84]. Broadcast [DK86, FT84, GW81, Lam80, cLW88, Pro81, Sin88a, Tas83, Top89, WA85a, RS88a]. Broadcasting [CK87, SW84, Sto83, JH89]. Broadcasts [SC89]. Bruijn [SP89]. BSP [KS82b]. Bubble [BL80, BW81a, CLW80b, CLW80a, DGS80, KYSY80, LCW81, LV85, Str82]. BucketSort [LV85]. Buddy [PH86]. Buffer [LL81, Sha81]. Buffered [DJ81]. Build [GHM87]. Building [II81]. Building-Block [II81]. Built [AB86, AC83, AC84, BM86b, FMM84, FM87, KS86a, MS88a, TAF87, CKS88, RSK88]. Built-In [AB86, AC83, AC84, BM86b, FMM84, FM87, KS86a, TAF87, MS88a, CKS88, RSK88]. Burroughs [KS82b]. Burst [Adi84, Bos86, ZW87, Bla88]. Bus [ABK83, Bok84, Bux83, FT84, HV87, IO84, LVA82, LJ87, Mar84, MBC82, MG82, MBCG83, Pra85b, Pra86a, Tow86, BL89, CMB88, KK89, MAS85, RI89, YZ88]. Buses [Agg86, Dub88, WS88]. Byte [Che83, Che86a, Dao81, DV83, Dun85, IC80, KF82]. Byte-Organized [Dao81, Dun85]. Byte-Oriented [Che86a].

C [MA89, Ano80b, Ano81a, Ano82a, Ano83a, Ano84a, Ano86b, Ano87a, AG82, Gai88b]. C-29 [Ano80b]. C-30 [Ano81a]. C-31 [Ano82a]. C-32 [Ano83a]. C-33 [Ano84a]. C-35 [Ano86b]. C-36 [Ano87a]. C/D [Gai88b]. Cache [DB82, LGH80, Pat82, SG85a, Smi87, SR88c, Soh89, YPD83, Dub88, KM89b, LPI88, Thi89, WM88b, YBL89]. cache-based [Dub88]. Caches [BD83, HS89]. CADAC [CHH83]. Calculate [SL83a]. Calculating [Bla83, DG86, Slo85]. Calculation [Mos87, TKL86]. Calculations [TP87]. Calculus [TTB85]. Calibration [CMS82, VSSG88]. Call [Ano80l, Ano80x, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano81t, Ano81r, Ano81s, Ano81u, Ano81v, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano82x, Ano82u, Ano82v, Ano82w, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano83j, Ano83k, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano84d, Ano84e, Ano85a, Ano87c, Ano87b, Cas86, SP82]. Calls [Ano85b, Ano86c, Ano86d, Ano87d, RC89]. CAM [NOYK88]. CAM-based [NOYK88]. Cambridge [HN88]. Can [EA89a, KC86, Mar81]. Canonical [FTY87, Pag80, SR82]. Cantoni [EA89a]. Capabilities [CS86c, GM88a]. Capability [BC85, Che83, Lop84, Mil82, RT86, Sny81, SH87]. Capability-Based [Sny81]. Capacity [MM83a, DSH89]. cardinality [KM89a]. Cares [Bra83a]. Carried [RJ80]. Carry [CP87a, GHM87, Kal83, LM82a, Par88, Rhy84, Dor88]. Carry-free [Par88]. Carry-Save [LM82a]. Carry-Skip [GHM87]. Cascading [SRO84]. Case [BW80, Ell85, VSV81, GM88b, HIT88, STR88, SW88]. Cases [Kri87]. Cause [AB80, AB82]. CC [CM89]. CC-banyan [CM89]. CCD [BZV86, IC80]. CD [ASP87]. CD-based [SD89b]. CDM [GF87]. Cell [Hay80]. Cells [VS86a, TS88b]. Cellular [Fet80, OP84, OO87, PTC86, WGT81, WR81, HCMP89, HMC89, OT89, ZMC89]. Centennial [Ano85c]. Center [Ame82]. Central [OJ80, Kum89]. Certain [Coy80, RTJH86, WE83]. Chain [Tsa81, CdL89, SL89]. Chained [HNS84]. Chaining [BBP88, TYZ88]. Chains [BT86, DLM86]. Change [LF85, SC89, GM88b, SR88c]. Channel [AH86, CP87b, CHL83, GFM83, KL82, Lau81, LL84, PTT81, PL84, Rub81, DJ88, KP88]. Channel-Access [GFM83]. Channel-Assignment [Lau81, PTT81, Rub81]. Channels [IM87]. CHAOS [SGB87]. CHAOS-Kernel [SGB87]. Character [Wei80]. Characteristic [HSE84]. Characteristics [Kob84, MC86, CCF89]. Characterization [BS87b, CS85, ED87, YA87, SBSM89]. Characterizations [AM80]. Characterizing [Hoc83]. Charge [KT81]. Charge-Coupled [KT81]. Chaudhuri [OI87a]. Checker [BL84, FMM84, FM87, Gol84, Kha82, WH88]. Checkers [Eti80, GH83b, Gai85a, HGS83b, Jen83, KM84a, Pie87, Gai88c, JC88, ML88, NC88b, NPP88, PNH88]. Checking [BC85, CL80, FM87, Gai85a, Gai85b, Gol84, HGS83b, HML84, Jen83, Kha82, LM83, Lu82, Mil82, NK87b, RB83d, SD86b, SL83b, Wan81, DD88b, DD88c, Gai88a, Gai88c, MS88a, ML88, NPP88]. Checklist [Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano82-32]. Checkpoint [AMM86, HP87]. Checkpointing [SLL87, LM88a]. Chinese [Vu85]. Chip [CGMP87, Cha84, HT80, KS86a, LCW81, Maz87, PS80, SOH^{+}81, NC88a]. Choice [Smi87]. Chordal [AL81, CI88]. Chrestenson [Mor86]. Ciphers [Sie85b, Gef73]. Ciphertext [Sie85b]. Circuit [ABG85, BG86a, Kal83, KP80a, LZLH85, Mar81, SV88, SB85, Sch89]. Circuitry [Cur80b]. Circuits [AR86, AB86, AB80, AB82, Abr82, AF81, Ano80z, Ano85d, Arm80, BPV83, BCDM86, CD86, CL80, Dav86, Eti80, Fuj81, FT82, GCV80, Goe81, Hay81, Hay86a, JA85, KH86, LMP82, LSM81, Maj85, MC86, Mar87, Mar86, MP83c, Muz80, NK87a, Ozg86, RB83d, RR86, SSVV81, Sav80a, Sav80b, Sav80c, Sav81, SY83, SGMP85, Sur81, VR83, VSV81, Vui83, WW83, YM86a, BC88, BS89b, CHRR89, CBAP89, FW88, Gai88b, Jha88, Wal88]. Circular [Ers85]. Class [AC83, Ary85, BG86a, CM87b, FW81, Fla82, JLS80, Kan84, KKS86, MT87, PL83a, Pie87, Pra80, SL87, Sie85b, Sri87, SR80, WF80b, Che88b, Ski88, THL88, TS88a, VSSG88]. Classes [KM81, Van80, VS86b, SD89b, YM88a]. classical [Bos88]. Classification [Oru84, SMN82, Til80]. Clock [KO87, SR87, LZM89, SR88b]. Clocked [KSW88, WF83, RMCF88]. Clocking [Kes84, UT86, PP88]. Clocks [KSB85, VM88]. Closed [IO84, Mey82]. Closed-Form [IO84, Mey82]. Closedness [Fuj81]. Closure [KLL87]. Cluster [WL81]. Clustering [DLM86]. CMOS [Jha88, MC86, RR86, UV81, YM86a]. Code [Adi84, BS89c, BL84, Dao81, DV83, Gol84, Kha82, Muk87, SCP^{+}81, NC88b]. Coded [GH83a, HGS83a]. Codes [BK80, BR80, BP82, BR82, BR84, BL85, Bos86, Che82b, Che83, Che86a, CS87, DC87, Don84, Dun85, Er84, GH83a, Gai85a, HGS83a, HGS83b, KF82, Kan84, Kha82, Kri83, NGP86, OI87a, Pie87, Pra80, Red87b, RTJH86, SOA85, Shi82a, Smi84, TC84, VS80, ZW87, Bla88, BV89, Che88a, FHG88, Gai88b, LB88, NPP88, PNH88, THL88, WM88a]. Codewords [BK80]. coding [Kak85, SKW88]. Coefficients [Sus83]. Coherence [YYF85, YBL89]. Coherency [DB82]. Coherent [DLC87]. Column [FH86, Kan84]. Combinational [AB80, Abr82, Ano80z, CV83, CV84, Cri80, EZS82, FT82, Goe81, GH80a, KW85, LM86, Mar87, MM83b, MM84, Obe80, RR86, Sav80b, Sav80c, Sav81, Sto80a, VS86a, VR83, BS89b, CHRR89, DK89]. Combinatorial [CH82a, OS82, Par81, RT85, Vui83, WM84, FW88, RK89]. combined [DG89]. Combining [BBG88, PN85, Lee89b]. Comments [BC85, BW81b, Cas86, CV84, CC89, DK82, EA89a, Fet80, GY86, GNK86, Got81, HL86, Hur81, Hwa87, Kam80, Kor86, Kub82, LMP82, Lau81, MK83a, MG86, MA89, Nil84, Obe80, PTT81, PR81b, Rub81, Sig82, Sip82, Slo85, Smi89a, Str82, Ten83, VZMBH89, Wen85, Wor81, Wus81, YN83, YWW86]. Common [Tsi86]. Communicating [BR83b, DLC87, Len88]. Communication [CLH84, EH85, GV84, GBG89, GS86, Hof85, HG87, IK82b, IC82, KK80b, LL81, LL83, LS89a, PR82, RE80, Smi80b, Smi81b, Von83, Wit81, YZ88, FLN89, GC89, GZ88, JH89, WB88, YM88b]. Communication-Efficient [HG87]. Communications [Ano82x, DK86, Fra81, KK84a, LD87, RB82, Top89, Wan82, Gai89]. Commutativity [Wei88]. Commutativity-based [Wei88]. Compact [Cle84, HS84, KS86a].

Compaction [DLSM81, Fis81, GB83, IKI83]. Comparative [Cra85, MBC82, Max88, MWM80, Wah84]. Comparator [Bur84, Lee85a, Wei80, Bur82]. Comparators [HML84, TS84a]. Compare [LCW81]. Compare/Steer [LCW81]. Comparing [Uhr82]. Comparison [CH82a, DSK87, Fra81, OA83, SD87, HTDR88, YBL89]. Compatibility [HT86]. compatible [San88b]. COMPCON [Ano80d, Ano80e, Ano80f, Ano80y]. Competing [KL82]. Compilation [PKL80]. Compiler [MP87, Pol88, Coa88, CNO^{+}88]. Complement [BVH83, Cha87a, Sas85, Dad89]. Complete [ATT81, Chu85, MI85, TT80, VS86b, AA88, NMN89]. Completely [BR80, CM87c]. Completion [Lee85a, MH80]. Complex [DDG80, Moh85, STR87, SGT86, ST86b, TPSS85, BBG88, Fam88]. Complexity [AM87, Arm80, But81, CGMP87, CLW80b, Fuj81, FT82, GV84, KA84, Lei85a, Mar86, MS86, MT87, MAV84, Par87, PC81, Ric84, Sas81, SB85, Tho83b, Tsa81, VS86b, BRG89, CHRR89, FLN89, FS88, SAA89]. Component [JK80]. Components [Cli80a]. Composite [Dad80, Muz80]. compound [HX88]. Compression [Hla86, RS87, SR86, RSK88, TR88]. Compressors [Gaj80]. Compsac [Ano83l]. Compsac83 [Ano83m, Ano83n]. Computation [Abe84, BK84, Ebe87, Fro83, IC82, KUV85, Kon86, Kru83, LM85, Mor86, Mul85, SSF80, WH80b, ACGK88, Ban88, CCWZ88, Kum88, Loz83, SS89b, Smi89b, SKW88]. computation-intensive [Kum88]. Computational [Arm80, Cha84, LP84, LP85, Mel87, NJM83, MS89]. Computations [CGMP87, GNK86, KT87, HTK89, NS88]. Compute [SBGS86]. Computer [AR83, Ame82, AAG^{+}87, Ano80-32, Ano80-33, Ano80-34, Ano80-35, Ano80-36, Ano80-37, Ano80-38, Ano80-39, Ano80-40, Ano80-30, Ano80-41, Ano80-31, Ano80-42, Ano80-43, Ano80-44, Ano81r, Ano81w, Ano81x, Ano81-29, Ano81-30, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano81-37, Ano81-38, Ano81-44, Ano81-45, Ano82d, Ano82-33, Ano82-28, Ano82-34, Ano82-29, Ano82-35, Ano82-30, Ano82-36, Ano82-31, Ano82-37, Ano82-32, Ano82-38, Ano82-39, Ano82-40, Ano82-41, Ano82-42, Ano82-43, Ano82-44, Ano82-60, Ano82-61, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83r, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Ano83-29, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano86h, Ano86i]. Computer [Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano87-27, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, AG82, BIO82, Bai87, BW89b, BA84a, Bla83, BR86, Boo81, CS80a, CP82b, CA86, CFH81, CLP81, DG86, FTY87, GP86, Gav87, GT83, Ger82, GJ80, GT80, GGK^{+}83, GT87, GKS87, HTOS80, HL84, HL80, HRJ86, HN80, Hwa87, IM87, KK80b, KC87b, LMO84, LL83, LM82b, Liu84b, MTG^{+}85, MFW80, NS82a, Niz83, Niz84b, Niz84a, OA83, PGR86, Pra80, RGA85, RM86, RB83c, Ros85a, RW84, SS80, SW84, SL80, SL84, Slo85, SS81, SM82, Sta84, SG85b, SBMM87, TV82, Wan82, Wei80, WV80, YK82, YH84, YPD83, BGM88, Con89, DJ88, KHS88, KS89, MI89, MS89, San88a, San88b, SGI89]. computer [Thi89, WJ85]. Computers [CH84, EDH80, Hoc83, HI80, HG87, PS80, Sah84, Sez87, SWP86, Sto83, Uhr82, WA80, BP89, Car88, HCMP89, Par89, Wai88, Ano80b, Ano81a, Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82a, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83a, Ano83-31, Ano84a, Ano86b, Ano87a, Ano87o]. Computing [AP86, Ano81s, Ano82j, Ano83k, Ano85a, BK87, Cur80b, Fam87, GM82a, GMK85, GLS82, Hay84, Hon82, Hon85, HKR84, JMKN86, Kar81, LM88a, MLT82, McG80, Mey80, PS83a, RTB81, Ren84, ST85, Sta87, Sti80, TLR83, TCH^{+}86, Vai82, Vui83, WTS^{+}85, Bok88, RK89, SL88b, Tys88]. Concentrators [NM82]. Concept [KK86, Tha84b, NK88b]. Concepts [Ren84, RM83]. Concerning [Jes80b, OS82]. Concurrency [HT83, Li87, SLJ88, Wei88]. Concurrent [Ell80, Got81, IK85, JTP85, Lun87, MM88, NK88a, PF82, PF83, PS87, QKS85, Sei84, SSF80, LCF89, NH88, USM89]. Condensed [WM86]. Conditional [MH80]. Conditional-Sum [MH80]. Conditions [Kod81, VS86a]. CONET [WK81]. Conference [Ano80c, Ano80-47, Ano80-48, Ano80-49, Ano82j, Ano82d, Ano82v, Ano82b, Ano82c, Ano83j, Ano82u, Ano82w]. Configurable [BBB^{+}82, LW89]. Configuration [LP83]. Configurations [CHL83]. Configuring [Gav87]. Conflict [LSD^{+}89]. Conflicts [AS87a, KL82]. Congestion [LL81, Niz84b, Niz84a]. Connected [CCSW85, GFM83, LSSS85, MA86, MT87, NS82a, Oru84, Pre83, Sto83, BP89, Ban88, BF89, Car88, FL88, GM88b, UR88, Wai88]. Connecting [CS86c]. Connection [MAS84, WK81]. Connections [LVA82, LVF83, SM83]. Connectivity [Ama83, AS85, AS87b, Feu82, HP88, ISO85, Sav84]. Connector [CW80, CG88]. connector-switch-attenuator [CG88]. Conquer [GW84, HZ83]. Conservative [SGT86]. Conservativeness [MST85]. Considerations [KB84, Lun87, PS80, PMSB88]. Consistent [MTG^{+}85]. Consistently [GC80].

Constant [CGMP87, Li85, TW83]. Constrained [SG85b]. Constraint [CS80a, LY83]. Constraints [ZRS87, GZ88, Pro89]. Constructing [Blo88, CE87, HG87, Kri87]. Construction [CW80, JK80, OI87a, Miy89, WB88]. Constructions [DV83]. constructive [BW89a]. Contact [RT85]. Contain [Cli80a]. containers [Mey88]. Containing [WR84]. Containment [GM82b]. Content [KK84b, SS82]. Content-Addressable [KK84b, SS82]. Contention [Bai87, MBCG83, PN85]. Contents [Ano80-50]. Contiguous [BS83]. Contiguous-Element [BS83]. CONTIN [SG85b]. Continuous [GZ88]. Contract [Smi81b, Smi80b]. contractions [SO89]. Control [ASP87, ACD82, Bis85, CA86, CS80b, CFH81, DHS85, DC81a, DV83, Dun85, EHS80, EHM80, FH86, GC80, GF87, Hol82a, IM87, IK85, JTP85, JH80, KK84c, LL81, Lee87, Li87, LC87a, MTMA85, MSS82, Mic83, Oru84, OJ80, PS87, PS85b, RB83c, SS87, Smi80b, Smi81b, Sta85, ST87, TS86b, WF83, WS82b, CJ89, CKS88, GC89, HT88, Kum89, LCS89, RB88, San88a, San88b, SLJ88, SW89, Wei88]. Control-Unit [JH80]. Controllability [Sav83a]. Controlled [Bux83, Cli80b, CHH83, RJ80, YL81, AJ89]. Controlled-Precision [CHH83]. Controller [KYSY80]. Controllers [TL80]. conventional [EL87]. convergence [LP89]. Conversation [KY89]. Conversion [CSR86, Hua83, Kri83, Muk87, Ulm83, Yue80, EL87]. Convert [Fro83]. Converter [FTT^{+}80]. Convex [LS82a, MS88b]. Convolution [BPV83, HL80, Jen81, Lam83, TLR83, FLN89]. Convolutional [RTJH86]. Convolutionally [Met82]. Convolutions [Ers85]. Convolvers [Dan84, NJM83]. Cooley [NS87]. Cooperating [Klu83]. Cooperation [DLC87]. Cooperative [LF80]. Coordination [CS84]. Coping [LW86a, Nic89]. Copyright [Ano82-45, Ano82-46, Ano83-30, Ano84q, Ano84r, Ano84s, Ano84t, Ano85o, Ano86s, Ano86t, Ano87n]. CORDIC [HT80]. correct [MS88a]. Correcting [BP82, BR82, Che83, Che86a, ES80, HGS83a, KF82, NGP86, Red87b, Shi82a, TF82, ZW87, BV89, LB88, THL88]. Correcting/All [NGP86, BV89]. Correcting/Detecting [BR82, Pra80]. Correction [Adi84, ATT81, Ano80z, Ano81z, Ano81y, Ano81-27, Ano84f, Ano85c, Ano85d, Bur84, CH83, CF80b, DO86, Hou87b, Kar81, KM84b, LP85, MBR82, Niz84b, Par81, Pra86a, Ram83, Smi81b, TF82, TB82, TS87, VRP83, Dav89, Kak85, LCF89]. Corrections [Ram86b]. correctness [PL88]. Corrector [BAR87]. Correctors [Cli80a]. Correlation [RS85, SSF82]. Correlator [Cur80a]. Cosine [KR82, Kit80, WG80]. Cost [CS80a, Gai85a, KYSY80, RS83b, Ste83, LPI88, NPP88, WH88]. Cost-Effective [Ste83]. Cost-Performance [RS83b]. Counters [Cur80c, Dad80, DC81b, DC82]. Counting [Coh85, SW82]. Coupled [ASHK86, KT81, LF80, RW82, BL89]. coupling [CFP89, DIY88]. Cover [Bes86, PA81]. Coverage [AM80, DT89, Kha84, MST85, RT85, RT86, TAF87, WB81, GNH88]. Covering [YM85]. Covers [MS86]. CPAC [JTP85]. CPC [SGI89]. CPU [HS89, IR86, Smi87]. Crash [Hag86, TS86a, TS87]. CRAY [CS86b]. CRC [Wei80]. Creation [Lin82]. Criteria [Miy89]. Criterion [Kre87, ST85]. Cross [RB83b, SB83]. Cross-Bar [SB83]. Crossbar [Fra81, LVA82]. Crossover [SYK89]. CrossoverNet [SYJ89]. Crossovers [McC81]. Crosspoint [RB83a, RB88]. crosspoint-irredundant [RB88]. Crosspoints [NM82]. Cryptographic [MTMA85]. CSMA [ASP87, SD89b, YH84]. CSMA/CD [ASP87, SD89b]. CSMA/CD-based [SD89b]. Cube [AS82a, AG81b, Bha83a, CCSW85, CS87, NS82a, Oru84, OI87b, SL87, AP89, BF89, Esf89]. Cube-Class [SL87]. Cube-Connected [CCSW85, Oru84, BF89]. Cubic [Sto80b]. cubical [Ban88]. Cumulative [DG86]. Current [Kar82, TOM81, YM86a]. Cut [IM87, Kri84]. Cutset [HR87]. Cutting [Jou86]. Cycles [CCSW85, HSE84, Ban88, LHC89]. Cyclic [Red87b, Shi82a, TLR83, Che88a, RN88, SGI89, WM88a].

d [FYD84, FLN89, Gai88b, KR82, Rot86, Yue80, wu87b]. D-MESFET [HM89]. DAG [Gel81, Sri83]. DAG-Based [Gel81]. daisy [SL89]. DASD [Bra83b]. Data [AS82b, AO80, AK88, Ano80-46, Ano82x, BYH87, Bes83, BE87, CH84, Cli80b, Cur80a, Dem85, DGT84, Ell85, ED87, FYAV87, Fla82, Gau86, GKS87, HKSS86, Hla86, IK82b, IKI83, JK82, Jes80b, KBB86, LP81, LM87b, Lei84, LF80, Len85, LAS87, McG80, MS82, Mel87, Met83, PGR86, RM83, RG85, SM82, TB82, TKK86, WFL82, YYF85, CICR88, GW89, Gef73, HIT88, LL88, LCF89, PL88, RI89, RSK88, WG88, Wei88, Ano80g]. Data-Flow [Gau86, GW89, WG88]. Data-Rate [Cur80a]. Database [Ano81r, BYH87, BF89, BFHW82, CP82a, CL84, CP82b, CH82b, CH83, GP86, Hag86, Hon85, Pra86b, QI85, RFLS86, Sto84, WWS84, YK82]. Databases [Che82c, DGS80, GC80, Lee80, Li87, QI85, SBK85, Str82, DG89]. Dataflow [GT80, BS89c]. DB [LYI87]. DB/DC [LYI87]. DBEC [DC87]. dc [VSV81, LYI87]. DCFL [HM89]. deadline [HTT89]. Deadlines [RSZ89]. Deadlock [DS87, Gel81, IK82a, ED88]. Deadlock-Free [DS87, IK82a]. Deadlocks [RB89b]. DeBruijn [EH85]. Debugging [LMC87, MH81]. December [Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano82-32]. Decentralized [Sta85, Sta89, VW84]. Decimal [CHH83]. Decipherable [BK80]. Decision [AR86, Cha87a, Sta85, Abo88, Miy89, Sta89]. decision-making [Sta89]. Decisions [BF83]. Decodable [ZW87]. Decoder [MC86, RTJH86, STD^{+}85, SR88a]. Decoders [Liu84a, OI87a, Red87b]. Decoding [DC87, Vu85]. Decomposing [GH88]. Decomposition [Cri80, Ebe87, JK82, Kub82, OO87, Sas81, FW88, ZMC89]. Decoupled [SWP86]. Decoupling [Pro89]. Decrypting [Sie85b]. DED [Kan84, Dao81, Dun85, Gai88b]. Dedicated [BS82, RJ80, Agr88]. Deductive [Ozg86, Wal88]. Defect [WB81, MP89c]. Defects [FH86]. Defined [AG81a]. Definition [IPM82, JC88, KBB86]. Degradable [FR85, FM84, GT87, Mey80, BB89b]. Degradation [RW81a]. Degree [DF84, JS84]. Delay [BJ83, GW81, Har86, LS80, Niz83, Wan82, RMCF88, SM88]. delay-insensitive [RMCF88]. Delayed [Sil82, Sil83]. Delayed-Staging [Sil83]. delays [Cal88, MTS89, SR88b]. deletion [LS89b]. Delivery [WA85a]. Delta [DJ81]. DELtran [FH83]. Demand [FT84, WG88, KK89]. Demand-driven [WG88]. demands [NS88]. denotational [Bou89]. Dense [Dot84, LS82b, EA89b]. Density [Cur80b, GZ88]. dependability [DT89]. Dependence [IR86]. Dependencies [Che86b, ED87, RW81a, Nic89]. Dependency [IKI83]. Dependent [KH80, BN88, KR89b]. Dependent-Failure-Tolerant [KH80]. Derivation [CMS82, CM87c, Lam83]. Derive [Cha87a, Sas85]. Described [AR86]. Description [LM82b, MP83b, CG88]. Descriptions [MP83a]. Descriptive [PW81]. Design [AG81a, Agr83, Ano80z, BYH87, Bat80, BS89b, BA83, BNM86, Bur84, Cas86, CBAP89, CA80, CC89, CL80, Cur80a, DM83a, DS80, DC81a, DD88b, DD88c, Fet80, FK81, Fuj84, GH83b, GV84, GT82, Gol84, GH80a, Gui89, HS86, HML84, Hur81, II81, II83, IC80, IK82b, JS88, Jen83, KW81, KH80, Kha84, KLL87, LS84, LW85, Lin82, LV85, Liu82, LY83, Liu84a, Lun87, MK83a, MA82, MB80, MD86, MBN81, Mor86, Pal86, PR81a, PGR86, PS80, Pie87, PS83b, Por82, Pra83, RB83a, RB88, RB83d, RH87, RTJH86, RB82, Sal80, SKF83, SD86a, SD86b, Sav84, Sav80b, STD^{+}85, SF84, SLS82, SL84, SP82, SSF80, SGA81, SL83b, ST86a, SH81a, Sta88a, TS84a, TAF87, TJB86, VSHM82, WC84, WWS84, WP82, WC83, WW83, WF80a, ZG81]. Design [ZG87, AKS88, BD89, BW89a, Bur82, DG89, Gai88b, Gai88a, HR88, JC88, JP88, KR89a, cLW88, LB88, LZM89, MBH89, MA89, MKLC89, NK88b, NPP88, Noe89, OT89, PMSB88, PNH88, Pol88, Sas89, SR88a, Sta88b, TRH^{+}88, VZMBH89, WM88a, Wan89, YKL88]. Designed [Ree80]. Designing [GGK^{+}83, IKP86, KK80a]. Designs [Dot84, JH80, Kes84, MA82, Tan84, GY89]. Desktop [SG85b]. destination [KPR88]. Detectable [TY87]. Detecting [BP82, BR82, BL85, Bos86, Bra86, DTF80, KF82, NGP86, PS85a, Pra80, XS85, Bla88, BV89, CH89, LB88, THL88]. Detecting/Correcting [BP82, THL88]. Detection [Aga80, Ata85, Bha83b, BI86, Che83, CD86, CP87a, Coy80, Don84, FT82, IK85, Kar83, KS80a, Kar81, LSM81, OK83, PF82, PF83, Pra83, Ram86a, Ram86b, Sav80a, Sav80c, SL84, Ulm83, VRP83, VR83, Vu85, YN83, YWW86, BS88, DSH89, DK89, ED88, Jha88, LCF89, MM88, RI89, Reg88]. Determinacy [OS87]. determine [US88]. Determining [AT81, NN86, Sil82, TS86b, Irs88, Muk89].

Deterministic [BJS80, Mar84, OM88, Sig82, Weg80, DFC89]. developing [Kap89]. Development [Abb83, De 82a, KYSY80, McG80]. Developments [Fu80, RA80]. Device [WSR84]. Devices [KT81]. DFSP [HKSS86]. DFT [TRH^{+}88]. Diagnosability [NN86, SSVV81, SS86, SSB86, VSV81, YM87b, YM87a, KK88, SAA89, YM88b]. Diagnosable [CH81, DM84, DMY85, Dah86, MD86, SSF80, YM86b, YML86, CA89, Sul88]. Diagnosing [WGT81]. Diagnosis [AB80, AB82, Abr82, AG81b, Bha83a, But81, CFP89, CV84, DM83a, DM83b, DSK87, EZS82, FW81, HN84, HS81, HS85, HKR84, HC86, MM80, Man80, MK84, Mey84, SSF80, SAA87, VHD82, VH84, Zhi84, DSH89, HKR88, Hos89, KH88, LZM89, SAA89, lYM88]. Diagnostic [KK85]. Diagnostics [OJ80, GS89]. Diagonalization [WOH84]. Diagrams [AR86, CE87, Hoc87, Lee82]. Diameter [DF84, II81, II83, SBGS86]. Diameters [BR86, ISO85]. Dictionary [AK85, ORS82, SS85a, SL87, SA85]. Difference [LA85, MD86]. Different [SS85a, Max88]. Differential [BR83a, GLS82, Sto80b]. Digit [IO83, Owe83, ZB87, Par88]. Digital [AR86, Agr81, ADG^{+}85, BPM^{+}86, Bry84, Cli80b, Cur80a, FTT^{+}80, GM82b, HP82, Jen81, Joh80, KW81, KA87, KC87a, KS80b, LM87b, Lei84, Maj85, MS81, MB80, RB83c, Red87a, RW84, Shi82b, TL80, TRYS83, TTB85, UT86, ZN80, ZG81, Bou89, GY89, MG88, PP88]. Digraph [FYD84]. Digraphs [Cul80, HP88]. dimension [Thi89]. Dimensional [AC84, KA87, LL84, NJM83, TLR83, DIN88, HCMP89, KR89a, SS89b, SKW88, TSM88]. DIMOND [JK80]. Diogenes [Kor86, Ros83]. Direct [Bes86, Fro83, Hol82a, MG86, Pap83, PA81, AP89, DD88a].

Directed [II83, ISO85]. disambiguation [Nic89]. Discrete [CF80b, Dav80b, DD81, FYAV87, KR82, MG86, Mul85, Pap83, RTB81, SD87, TCH^{+}86, Bon83a, WG80]. disjoint [JC88, NC88b]. Disk [Kim86, CCF89, McC85, RB89a]. Disks [Lei84]. Displays [CH85a]. Distance [CW83, Coh85, PC89]. Distinguishing [Hon81]. Distributed [Ano82j, Ano82-58, Ano84f, Ano86c, ACD82, BYH87, CP82a, CCWZ88, CA80, Che82a, CL84, Che82c, CP87b, CA83, CA86, CVY83, CH82b, CH83, CLH84, CH85b, CL87, CS89, CN82, CMM87, DHS85, DK86, EDH80, Ell85, FH80, GM82a, GMK85, GC80, GP86, Gav87, GLS82, GFM83, GF87, GJ80, HS85, HKR84, HKR87, IK82b, KT87, Kar81, KW89, KO87, KH88, KSB87, KC87b, LL83, LF85, LYI87, LE80, Li87, LM85, Lo88, MLT82, MK83a, OK83, Pat80, PS87, PR82, RSZ89, Ram88, Ram89, RB89b, RSS85, SS80, SBK85, ST85, SC87, SC89, Sie81, Smi80b, Smi81b, Sta84, Sta87, Sta88a, SD87, VM87a, Wah84, WM85, WV80, Wor81, YH86, YGZ^{+}87, Zak85, ZN80, AA88, Bok88, Con89, HKR88, Hos89, KHS88, KK89, KWFT88, Kum89, KS89, LW88, MP89b, Mil88]. distributed [RKH88, San88a, SS89a, SD89a, Sin89, SRC85, Sta88b, lYM88]. Distributing [YTL87]. Distribution [IK82b, KK82, KSW88, Lak84, Jou89, MP89c]. Distributions [DG86, Har86, JL85, Sch89]. Distributive [JL85]. Distributor [GS86]. diversity [AK88]. Divide [GW84, HZ83, PF83, ZG87]. Divide-and-Conquer [GW84, HZ83]. Divider [ZG81]. Division [Bus83, KM83b, Li85, MPPZ88, Par87, PP87, Tsa83, ZH85]. Division-Free [Tsa83]. Do [Sav83a]. Domain [SGB87]. Domains [RB83d]. Don’t [Bra83a]. Double [KF82, Sav86, Uug81, MI89, RS88b]. Double-Edge-Triggered [Uug81]. double-loop [MI89, RS88b]. Doubly [DHH85]. DPM [Mil88]. Driven [Bra83b, WG88]. Dual [Ano81z, BPM^{+}86, DHR80, ST86a, HTDR88, VZMBH89, Wan89]. Dual-Mode [Ano81z, DHR80]. Due [RW81a]. Dyadic [Bes83]. Dynamic [AJ82, AL85, CP87b, Fen85, HKR84, IS84, KK80a, KK82, KK87, Kob83, Kob84, LL80, MP89b, NS88, Niz84b, Niz84a, Ree84, SD89a, SH84, ST87, UR88, WBA83, JS88]. Dynamic-Full-Access [SH84]. Dynamical [SSVV81]. Dynamically [BS86, BBB^{+}82, HTOS80, Hur81, Pra85a].

E/D [HM89]. E/D-MESFET [HM89]. earliest [HTT89]. Early [MH80]. Easily [Bha83b, SKF83, SF84, SH81a, YL81, JP88]. East [Ano81b, Ano81c, Ano82e, Ano82f]. EC [LB88]. Edge [AT81, Sto80a, Uug81, LS89b]. Edge-Tracing [Sto80a]. Edges [CW80]. Editor [Ano80-27, Ano80-28, Ano81-28, Ano82y, Ano82z, Ano82-27, Ano83o, Ano83p, Ano84g, Ano85e, Ano85f, Ano86e, Ano86f, Ano86g, Ano87e, Ano87f, Fen84b, Fen84a, Kar82]. Editors [Ano83q, BW81b, SJ84]. Effect [AB80, AB82, De 83, GKS87, Iye84, KP87, MA82, Jou89]. Effect-Cause [AB80]. Effective [LJ87, Ste83]. Effectiveness [BD83, Smi80a]. Effects [Cve87, DB82, SBMM87, MTS89]. Efficiency [But81, EZL89]. Efficient [Ano81y, Ano81-27, APD83, BH81, Bec88, Bha83a, CT84a, Cha83b, Dah86, Fam88, Fra83, FJ88, Gla80, Gor87, GHM87, HGS83b, Hon85, HG87, KK84a, KN84, Kon86, KH83, Lei85b, MS88b, NMB83, NPP88, Par86, PNH88, QKS85, RB80, Sam80, SD89b, SA85, Vu85, Bos88, JK89, Miy89, Sin88b, THL88]. Efficiently [TKL86]. Eight [Chu85]. Eighth [Ano82x]. election [AA88]. Elections [GM82a]. Electronic [Dem85, Sta88a, Sta88b]. Element [BS83, NJM83, SE87]. Elementary [Mul85]. Eliminating [ZB87]. Elimination [Sor85, Sri83, AJ89, HQR89]. Embedded [AS85, KM84a, SMV87]. Embedding [AR82, GKS84, SKW88, CC88]. Embeddings [Gor87, SB88]. Empirical [Kob86, MK83b, Sav83b, Coo89]. Employing [Cur80c, ZN80]. Empress [BBB^{+}82]. Emulation [BI86]. Enclosures [Vai82]. Encoded [Met82]. Encoder [HRT^{+}84, MC86, ST88b]. Encoder-Decoder [MC86]. Encoders [Liu82]. Encoding [MG86, MAV84, Pap83, ST86b]. Encryption [Kak85]. End [IM87, SYJ89, SYK89]. End-to-End [IM87]. Energy [Hay86b]. Engineering [Ano81-44, Ano81-45, Ano82-61, Kum88, Ano80c]. Enhance [SRO84, Soh89]. Enhanced [Har86, GNH88]. Enhancement [ASKL81, HA88, KR89a]. Enhancing [AKT86, Pol88]. Ensuring [KSB85]. Enumeration [EDH80, Goe81, YTY82]. Environment [BDL83, Man84, SSS^{+}83, BM89, UR88]. environments [LL88]. Equal [Weg85]. Equalities [Sam80]. Equations [ATT81, APD83, BR83a, ED83, GLS82, TT80, WK80, WH80b, AOES88, DD88a]. Equijoin [Che82c]. Equilibrium [Fuk88]. equipartitioning [OM88]. Equivalence [BJ83, Mur81]. Equivalent [GH80b]. Error [Adi84, BP82, BR82, BR84, BL85, Bos86, CMS82, Che83, Che86a, CI87, Cli80a, DV83, Dun85, ES80, Fro83, Gai85b, HGS83a, Jen83, KF82, Kar81, MBR82, NK87b, NK88b, NGP86, PF82, PF83, Por82, Pra80, Ram83, Red87b, RTJH86, SL84, Shi82a, TY87, TF82, TB82, Tsa81, Tsa83, Van86, VRP83, WE83, ZW87, BGM88, BV89, BS88, Gai88b, Kak85, LCF89, LB88, Loz83, MM88, MW88, SL88b, ST88b, THL88]. Error-Correcting [Che83, Che86a, ES80, HGS83a, Pra80, Red87b, Shi82a, TF82, ZW87].

Error-Correcting/Detecting [Pra80]. Error-Correction [Adi84, Kak85]. Error-Detectable [TY87]. Error-Detecting [BL85, Bos86]. Error-Detection [Che83]. Error-Trellis [RTJH86]. Errors [BS82, Don84, IR86, WR84]. Essential [Kre87, Kuo87, MS88c]. Estimates [HV87, MST85]. Estimating [GM88b, KA84, LCS89, KT89a]. Estimation [CLH84, GK85]. ETH [BBB^{+}82]. ETH-Multiprocessor [BBB^{+}82]. Euclidean [FYSK84, Man89]. Evaluating [ED87, GMK85, HS89, JH80, Kee83, Mar84, Mey80]. Evaluation [Ano83q, BM86a, GJ80, GT87, GB83, HL84, HNS84, Iye84, Kam80, KS80b, LS84, Maz87, MFW80, Niz84b, Niz84a, PGR86, RB82, SS81, SRC85, WS82b, Che88c, CC89, GT88, KP88, LPI88, Noe89, RB89a]. Even [KH83]. Evolution [Aup83]. Exact [Aky87, GNK86, HV87, TP87]. examination [McC85]. Example [Boc82]. Examples [Ren84]. Exception [Cri82]. Exchange [AS87b, DK82, KDJ85, KJ86, Ste83, TN81, TKK86, WF80c, WF81, BN88, LHC89, SR81a, TS88a]. Exchange-Type [Par80]. exclusion [Sin89]. Exclusive [CH85b]. Exclusive-Writer [CH85b]. Execution [BG84, BE87, Cra85, FH83, HP87, KY89, PS87, Veg84, VM87a, KW89, NH88, PP88]. Exhaustive [BCR83, Che88a, MTG83, TW83]. Existence [Hon81, Kod81]. EXMAN [PGR86]. Expected [ACGT84, MM83a]. Experience [Dem82, GLS82]. Experiences [FHH^{+}83]. Experiment [FH83, LE80]. Experimental [LS80, CI89, HM89, US88]. Experiments [DLSM81, LSM81, LM83, SD86b, SG83]. Explicit [Ulm83]. Exploit [Her87]. Exponential [Kar84b]. Expression [BBP88, VZ81, WS82b, Fos89].

Expressions [BM86a, Kee83, SOA85]. EXtended [PGR86]. Extending [Ano80d, Ano80e, Ano80f, Her87, Moh85, CCF89]. extension [SK89]. Extensions [Mah80]. External [BS82, BLP84]. Extra [AS82a, Bha83b, Pra83, Van86, GM89, LH88]. extra-stage [GM89]. Extraction [WS82b, RJ89]. Extremum [WM84]. Extremum-Search [WM84].

Fabricated [YM86a]. Facility

[Lan87, VBH^{+}81]. FACR [Hoc83]. factor [TRH^{+}88]. factorization [USM89]. Faddeeva [NH88]. Failure [BN88, CA83, HKR84, IBM82, Iye84, KH80, KS86b, KR89b, PC81, RC89, SL83a, SLJ88]. Failure/Load [IBM82]. Failures [IKT81, Sav80c, SD80, VI84, WR84, MI89, VM88, lYM88]. Fair [LS89b]. Fall [Ano80d, Ano80e, Ano80f, Ano80y]. Families [JS84, SS81]. Family [VR86, JK89, USM89]. Fan [PP87, WB88]. Fan-In [PP87, WB88]. Farey [Kri83]. Fast [Adi84, BG84, CF80a, CT84b, DL86, Fet80, Gna85, HGS83b, Hou87b, Jes80a, KR82, Kam80, Kod81, KM83b, Li85, Pie87, PR81b, RTB81, Sco89, SK89, SS89c, WHT84, YGZ^{+}87, wu87b, EA89b, Fen89, Fis88, HN88, Kap89, Par89, Sri89, TSM88, TRH^{+}88, TR88, WZ89, Hou87a, RJ82, TLR83]. Faster [Kal83]. Fat [Lei85b]. Fat-trees [Lei85b]. Fault [AB80, AB82, Abr82, AM85, AA88, AS82a, AM80, Aga80, AF81, Agr81, Agr88, AP86, AK87, Ano81z, Ano81s, Ano83k, Ano85a, AG81b, ACD82, BA86, BK87, BC85, Bha83a, Bha83b, BI86, CD86, CP87a, CV84, CH81, CS86c, CN82, Cos88, Coy80, Cri82, DM83b, DM84, Dah86, DSK87, DK89, Dao81, DHR80, DHS85, EZS82, EH85, FW81, FJ88, FT82, GCV80, GM88a, GT83, GBG89, GH80a, GS86, HA86, Haw85, Hay84, Hon82, HKR87, HA84, HC87, IK85, Iye84, IDH86, JA88, Kes84, Kha84, KP80b, KB84, Kor86, KSB85, KK86, KJC89, LS84, LF87, LP88, MS81, MM80, MA82, MSS82, Mey84, Mil82, MM83b, MM84, NK87a, NN86, OK83, Ozg86, Pra80, PR82, Pra83, Pra85a, Pra85b, Pra86a, RAE84, RV86, RT86, RB83d]. Fault [Red87a, Red87b, Ren84, Ros83, Ros85b, Sav80c, SMV87, SY83, SL80, SCP^{+}81, SSB87, SL83a, SH84, SL86, Sil82, Smi80a, Sti80, Tan84, TAF87, VR89a, VHD82, VH84, VR83, WA80, WB81, Wu87a, YN84, YM86b, YML86, YM87b, YM87a, YH86, Zhi84, AK88, AL88, Ban88, BD89, BW89a, Bos88, CCWZ88, CI89, CM88b, CM88c, CC89, Coa88, DSH89, DT89, Esf89, Gai88a, Gai88c, GS89, GNH88, HKR88, Hos89, Jha88, KWFT88, KM89a, KR89a, KSW89, KPR88, KR89b, KT89b, LLJ89, LJL89, LW89, LZM89, Mey88, ML88, NK88b, Reg88, Sin88b, Ski88, SAA89, Sul88, Tys88, TYZ88, UR88, VR89b, VM88, Wal88, lYM88, YM88b]. Fault-Diagnosis [FW81, Hos89]. Fault-Location [Wu87a]. Fault-Masking [SY83]. Fault-Secure [NK87a, NK88b]. Fault-Set [RB83d]. Fault-Tolerance [HKR87, SH84]. Fault-Tolerant [AS82a, AP86, Ano81s, Ano83k, Ano85a, ACD82, BK87, BC85, CN82, Dao81, EH85, FJ88, GT83, GBG89, GS86, HA86, Haw85, Hay84, Hon82, HC87, Iye84, IDH86, Kes84, KB84, Kor86, LS84, MS81, MA82, MSS82, Mil82, Pra80, PR82, Pra85a, Pra85b, Pra86a, RV86, Red87a, Red87b, Ren84, Ros83, Ros85b, SCP^{+}81, SSB87, SL83a, Sti80, Tan84, WA80, YH86, AA88, JA88, KJC89, LP88, VR89a, BD89, BW89a, CCWZ88, CM88b, CM88c, CC89, DT89, Gai88a, GS89, Hos89, KR89a, KPR88, KR89b, KT89b, LLJ89, LJL89, LW89, LZM89, Mey88, Ski88, Tys88, TYZ88, VR89b, VM88]. Fault/Transient [DM83a]. Faults [AM85, AMM86, BG86a, GH80b, Hay80, HC86, Kar83, KS80a, KP80b, LSM81, LM86, MS81, MK84, MM84, Oik87, Ozg86, PS85a, RT85, RB83a, RB83b, RR86, SKF83, SK85, Sav80a, SSB86, SN81, SR87, SB85, SR80, SR81b, SBMM87, XS85, YN83, YN84, YWW86, CFP89, KW89, MP89a, SM88, dJvdG88]. Faulty [BG86a, GMK85, KM89a]. favorite [Fis88]. Feedback [Che86b, Dav80a, KS80a, WM86, XS85, YN83, WM88a]. Fermat [TRYS83, TCH^{+}86]. FET [RR86]. Fetch [LMO84, BD89, Sto84]. fetch-and-add [BD89, Sto84]. Fetching [HI80]. FFT [Ano81y, HL86, Bon83b, Car83, CM88b, Gla80, JA88, NS87, SCNS83, TPSS85, WD84]. fiber [GC89]. Fibonacci [NN87]. Field [Eng81, NJM83, Red87a, FM89, HTDR88, TRH^{+}88, Wan89]. Fields [YRT84, Man89, Pin89]. File [KYSY80, LL83, RW83, TS83, Twi83]. Files [BL80, CH85b, Met83, BBW88, SS89a]. fill [AJ89]. fill-in [AJ89]. Filter [Pra86b]. Filtered [Wus82]. Filtering [Red87a, TRYS83]. Filters [KA87, FS88]. Finding [Agg86, Bok84, GW84, Kan85, Tsi86]. Finite [CIP87, CHM87, Eng81, Fro83, HT86, KM83a, Mah80, MK85, NJM83, Oik87, Red87a, SE87, WGT81, Wan81, WR84, YRT84, HTDR88, Man89, Pin89, TRH^{+}88, Wan89, KM84b]. Finite-Segment [Fro83]. Finite-State [CIP87, CHM87, Oik87, WGT81]. Finite-Turn [Wan81]. FIR [FS88]. Fire [Adi84]. Firefly [TSS88]. FISHNET [KHS88]. Fixed [CS80a, DV87, JS84, MF86]. Fixed-Size [DV87]. Flaws [SMN82]. flexible [SRC85, WZ89]. Flight [MSS82]. Flip [Cha83a, LMP82, Mey88, Uug81]. Flip-Flop [Cha83a]. Flip-Flops [LMP82, Uug81]. Flip-trees [Mey88]. Floating [AP85, FG82, ST88b, Loz83, Sco89, WE83]. Floating-point [ST88b, Loz83, WE83]. floating-slash [Sco89]. Flooding [Top89]. Flop [Cha83a]. Flops [LMP82, Uug81]. Flow [ACGT84, AO80, BYH87, BE87, CH84, CS80b, CFH81, DGT84, Gau86, GKS87, HKSS86, IM87, KBB86, KS83a, LP81, LM87b, LC87a, McG80, PGR86, RM83, SS87, SM82, TJB86, GW89, LD89, San88a, San88b, WG88]. fly [EL87]. Forces [VW84]. Foreword [Ano80-29, Gar80, Hay84]. fork [KA89, NT88]. fork-join [KA89]. fork/join [NT88]. Form [Aky87, Ano82-45, Ano82-46, Ano83-30, Ano84q, Ano84r, Ano84s, Ano84t, Ano85o, Ano86s, Ano86t, Ano87n, IO84, Mey82, STR87, Sri87, Aky89, CMB88, CdL89, dM88]. Formal [KBB86, KSW89, MSS82, Sny81, Sta88a, TF82, Sta88b]. Formalism [Tha84a]. formation [Fos89]. Formatting [WSR84]. Forms [FTY87, Pag80]. Formulas [Sam80]. Fornmts [FJW85]. Fortran [HL82]. Fortran-Like [HL82]. Forward [Gel81]. Four [CL80]. Four-Bit [CL80]. Fourier [Bon83a, CF80b, PR81b, RTB81, GS87, Mah80, TSM88, Tho83a, TCH^{+}86]. FP [WG88]. Fractal [SWK84, Thi89, Wai88]. fractal-like [Wai88]. Frame [LP81]. Framework [GJ80]. Free [DS87, Fro83, IK82a, LSD^{+}89, Ozg86, SC87, Tay83, Tsa83, Par88]. Frequently [YL81]. Full [Cur80c, MD86, SH84]. Full-Difference [MD86]. Fully [GFM83, Hua83, IO83, Owe83, GM88b]. Fully-Connected [GFM83]. Function [Ano81z, Bry86, Car84, Cha87a, DHR80, GK83, GF87, KP80a, Kuo87, Mor80, RFLS86, RB80, Sas85, Tha84b, WB81, Wus81, Wus82, KM89b, Noe89, PS88, WM88b]. Function-Independent [Ano81z, DHR80]. Functional [AR86, AKT86, BA84b, BF80, CG88, PS85a, PS88, SH81b, SR81b, UV81, Veg84, BI88, SL88a, SS88b]. Functionally [BYH87]. Functions [BIO82, BZV86, Bes86, BG86a, CM87c, Dav80b, DD81, Kam80, KF80, Kar81, KS82a, Kre87, LM87a, MG86, MTG83, Muk86, Mul85, Pap83, PA81, Por80, Por82, RS84, RS86a, Sas81, Sta81, Tha82, WHT84, Abo88, BB89a, FGPT89, JW89a, MKLC89, YM88a, YF88]. functions-regular [YM88a]. Funds [Sta88a, Sta88b]. Funds-Transfer [Sta88a, Sta88b]. Further [KP88, Rub81, Wen85]. Future [Hur84, PS80]. Fuzzy [HRJ86, KF80, YM86a, Zhi84]. Fuzzy-Valued [KF80].

G [CC89]. GaAs [MLB87, HM89]. GaAs-Based [MLB87]. Galois [Eng81, FM89]. Gamma [LH88, PR84]. GAMMON [BW89b]. Gate [BS87b, Hur81, NKY^{+}80]. Gates [LM87a, SRO84]. Gaussian [AJ89, HQR89, Sor85, Sri83]. GCD [BK84]. General [CHM87, HFPS82, Kre87, Liu84b, SL80, SS86, SLS82, Smi85, VSH89, BW89a, CCF89, Wai88]. General-Purpose [HFPS82, Liu84b, SLS82]. Generalization [CCSW85, YM87b]. Generalized [AK85, BA82, BA83, BA84a, CW80, CF80b, DC81b, DC82, Esf89, FM87, IKI83, LC87b, Sas81, SAA87, Tha84b, BBG88, FLN89, Len88]. Generate [Goe81, HO81]. Generating [Er84, Kuo87, LC87b, SWK84, VS80, Wai88, CHRR89]. Generation [AR86, AM85, AMM86, Ano85d, BCR83, Bes83, BF80, DB87, FS83, GK83, JA85, KS82a, LM82b, LA85, Nwa85, SK85, SG83, TW83, TC84, TA80, Che88a, Che88b, HMC89, PS88, dJvdG88]. Generator [Ano81-27, BH81, LS80, GNH88, Reg88]. Generators [AC83, Mar85, Sav83b, WH80a]. Generic [AM80]. Geometric [Til80]. Geometry [Cha84, LP84, LP85, MS89]. Given [DF84]. Global [Agg86, Bok84, Fis81, GB83, IKI83, Lee80, TJB86, SD89a]. Golomb [Rob83]. Golub [Moh85]. Good [Sav83a, Tri82]. Gracefully [FR85, BB89b]. Gradual [Lee89a]. grained [AJ88]. Grammars [CPG83]. Granularity [Cve87]. Graph [Agr83, Bry86, DV87, IKI83, JM87, KBB86, MR82, RS84, SN81, ST85, SBGS86, WR81]. Graph-Based [Bry86]. Graph-Theoretic [JM87]. Graphs [AK87, Cul80, DF84, II83, ISO85, JS84, LS82b, RFS86, SE87, Van80, CI88, Mey88, SO89]. Gray [CS87, Er84, VS80]. Greedy [DM83a, DM83b]. Grids [AR82, CC88]. Group [AK87, FMM84, PTC86, AK89]. Group-Parity [FMM84]. group-theoretic [AK89]. Grouping [LA85]. Groups [OO87]. growth [KM89b]. Guarantee [Sav83a]. guaranteed [GT88]. Guest [Ano83q, BW81b, Kar82, SJ84]. Guided [PK87].

H. [CC89]. Haar [RJ82]. Hadamard [Irs88]. Hamiltonian [LHC89]. Hand [RJ80]. Hand-Carried [RJ80]. Handling [Cri82, Gau86, Mur81]. handshake [SL89]. Hard [CMM87, DC81a, FH86, Gef73, SRC85, Sta89]. Hardware [AB86, Agr80, Boc82, Bur84, DM81, HP82, JO84, LS84, Lei85b, LM82b, MS81, Man84, MP83a, Muk89, PW81, PR81b, Pra86b, SD80, TP87, ACGK88, Bur82, CC89, KW89, SR88b]. Hardware-Efficient [Lei85b]. Hardwired [LC87b, Mic83]. Hartley [Hou87b, Hou87a]. Hash [Cle84]. Hashing [Bur84, HNS84, Bur82]. Having [AB86, Tam80]. HDLC [LP81, Wan82]. Head [SYJ89, CCF89, SYK89]. Head-End [SYJ89, SYK89]. Height [Vai84, Vai86]. Height-Balanced [Vai84, Vai86]. helix [Dav89]. Hensel [Kri83, Muk87]. heterogeneous [BF88, SW89]. Heuristic [Bes86, BG86b, Cha83b, DL87, Lo88, SW84, LD89]. heuristically [Sin89]. heuristically-aided [Sin89]. Heuristics [Kri87]. Hexagonal [GKS84]. Hidden [GF87, Gel89]. Hierarchical [Abr82, CS80b, EHM80, Fen85, MP83b, SLS82, TRTN89, tCpS88, Con89]. Hierarchies [CS80a, GT82, Sil82].

Hierarchy [MTMA85, Sil83]. High [AO80, Cha87b, CHL83, CLP81, Cur80b, Cur80a, De 82a, HFPS82, HP87, Kha84, LMO84, Maj85, NOYK88, OI87a, PKL80, Smi80b, Smi81b, SB80, SG80, TYY85, TY87, TAF87, UT86, WH80a, ZN80, ZG81, ZG87, Gai89, JP88, KHS88, LM88b, PP88, SBSM89, Soh89]. High-Level [AO80, Smi80b, Smi81b, LM88b, SBSM89]. High-Performance [CLP81, HP87, LMO84, Gai89, JP88, KHS88]. High-Speed [Cha87b, HFPS82, Maj85, OI87a, PKL80, TY87, UT86, WH80a, ZN80, ZG81, ZG87, NOYK88, TYY85, PP88]. Higher [Bus83]. Highly [Lun87, PL83b, SBMM87, Che88b, RJ89]. Highway [CFM^{+}86]. hit [WM88b]. HM [SLS82]. Hocquenghem [OI87a]. Homogeneous [Dim85, RFS86, KK88]. hop [Pro89]. Horizontal [DLSM81, IKI83, MAV84, PB87, VBH^{+}81]. Hot [PN85, YTL87]. Hot-Spot [YTL87]. HP [MPPZ88]. HPC [LF85]. HRFC [CS80b]. Hu [MA89]. Huffman [Par81]. hull [MS88b]. Hundred [Von83]. Hybrid [DM83b, MM80, RTB81, RG85, SSB86, YM87b, YM87a, YM88b]. Hyperbus [BA84a]. HYPERchannel [CG87, FH84]. Hypercube [BA84a, CS86a, DJ88, GS89, RS88a]. Hypercubes [SS88a, AOES88, CC88, JH89, Kat88, SB88]. Hypergraph [Ros85b]. Hypergraphs [FYSK84]. Hypernet [HG87]. Hypertree [GS81].

I/O [KB85, LY83, PW81, RB89a, Sto89, WSR84, XS85]. IBM [Kob83, PMSB88]. Ideas [BE87]. Identification [CH81, DM84, GH80b, YM86b, YML86, Sul88]. Identifying [Dah86]. IEEE [Boo81, Ano80-32, Ano80-33, Ano80-34, Ano80-35, Ano80-36, Ano80-37, Ano80-38, Ano80-39, Ano80-40, Ano80-30, Ano80-41, Ano80-31, Ano80-42, Ano80-43, Ano80-44, Ano80b, Ano81-29, Ano81-30, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano81-37, Ano81-38, Ano81a, Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82-45, Ano82-46, Ano82-33, Ano82-28, Ano82-34, Ano82-29, Ano82-35, Ano82-30, Ano82-36, Ano82-31, Ano82-37, Ano82-32, Ano82-38, Ano82-39, Ano82-40, Ano82-41, Ano82-42, Ano82-43, Ano82-44, Ano82a, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83-30, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83r, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Ano83-29, Ano83a, Ano83-31, Ano84q, Ano84r, Ano84s, Ano84t, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m]. IEEE [Ano84n, Ano84o, Ano84p, Ano84a, Ano85o, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano86s, Ano86t, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86b, Ano87n, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87a, Ano87o, AP85]. IF [BG84]. II [SSVV81, Smi89a]. Illiac [RK86, SS84]. Illumination [LS82a]. Image [Ano81r, AGH^{+}82, BFHW82, CP82b, HFPS82, KY82, KWR82, NWLP82, Par86, RS85, SSK^{+}81, SSF82, SS82, WS82a]. Imagery [WFL82]. Images [Gue86]. Imaginary [KM81]. Imase [HP88]. IMC [CLH84]. Impact [GCV80, GV84, SL84, Pol88]. Impacts [KY89]. Implementation [AAG^{+}87, AP85, BZV86, BB87b, BL84, CT84a, CSR86, Cha83b, Dav80b, De 82b, Ers85, Hol82a, HL86, HRT^{+}84, JM87, Jes80a, Jul80, KY82, KH83, Lan87, MP83c, MG86, Pap83, PR81b, RB80, SBGS86, SS81, SG85b, SCNS83, Tha84a, Tha84b, WHT84, Wei80, YK82, Zwa85, Bos88, CICR88, SS89c]. Implementations [KT85, Vu85, WGT81, WD84, Max88, PL88].

Implemented [Dao81, SL80, TS84a]. Implementing [SP88, TP87, YS89]. Implicants [MS86]. Implicating [DMY85]. Implications [VSHM82, ALL89]. Implicit [Goe81, Ulm83]. Implicit-Explicit [Ulm83]. implied [KM89a]. Importance [HCMP89]. Impossible [BG86a]. Improve [KC86, EA89a]. Improved [BD80, CE87, HL82, Kri84, KT89a, Lak84, PS85a, Dor88]. Improvement [CL84, Kan81, MA82]. Improving [Ano81y, Gla80, PH86, Sri87]. incentive [San88b]. Incomplete [Kat88, KK88]. Incompletely [Yam80]. Incorporate [KK84b]. Incorporating [BE87, FTT^{+}80]. increases [Coa88]. Incremental [Smi89b]. Independent [Ano81z, Ano81-27, BH81, DHR80, Jia86, RTY^{+}88]. Index [Ano80b, Ano81a, Ano82a, Ano83a, Ano84a, Ano86b, Ano87a]. Indexed [Bes83]. Indication [NK87b, Gai88c]. Indicator [Gai85b]. Inductive [PS83b]. Industrial [ACD82]. Inequalities [Sam80]. Inertial [BJ83]. Influence [MW88, RT86, RAP87]. Influences [SS85b]. Information [Agr81, Ano80-45, Ano82-54, Ano82-55, Ano83j, Ano83-32, Ano83-33, Ano83-34, Ano83-35, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano84z, Ano84-27, Ano84-28, Ano84-29, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano85w, Ano86u, Ano86v, Ano86w, Ano86x, Ano86y, Ano86z, Ano86-27, Ano86-28, Ano86-29, Ano86-30, Ano86-31, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, CA80, GF87, Her87, KS83a, MK83a, VSHM82, VHD82]. Inherent [Owe83]. Innovation [Ano82-56, Ano82-57]. Input [ASHK86, Gna83, Kar83, KS80a, LL81, Sas84, Sip82, YN83, Dad89, Max88]. Input/Output [ASHK86, Kar83]. Inputs [Bha83b, Cha87a, HML84, Kuo87, Pra83, Sas85, RB88]. insensitive [RMCF88]. Insertion [DR85, Ell80, Got81]. Instant [LMC87]. Instantaneous [MST85]. Instruction [AKT86, Ary85, De 83, DO85, DO86, FJW85, HI80, Kee83, Kob83, LMO84, NF84, SS87, SG85a, Sto84, VM87b, WS84, Coo89, Jou89]. Instruction-Fetching [HI80]. instruction-level [Jou89]. Instruction-Scheduling [Ary85]. Instrument [EHS80]. Instrumentation [SSS^{+}83]. Integer [BR83a, MPPZ88, PP87, Sav83b, STR87, ST86b]. Integers [Pre83]. Integral [TTB85]. Integrated [Cur80b, LZLH85, SSS^{+}83, WBA83, YM86a, KK89]. Integrated-Circuit [LZLH85]. Integration [LL85, SL80]. Integrity [Lei84, Lu82]. Intelligent [SYJ89, Wei80, SYK89]. intensive [Kum88]. Interactions [Sta81]. Interactive [LF80]. interchangeable [LS88b]. Interconnect [MA82]. Interconnection [AS82a, Agr83, AL85, BA83, Bhu85, CCSW85, CM87b, DS87, DHS85, Dot84, FW81, FS80, FS81, FWT82, GS81, Haw85, HZ81, HC87, JS84, KS83b, KSW88, LS82b, LSD^{+}89, Maz87, NM87, OS82, OP84, OOB85, PL83a, PN85, PK80, RV86, RB82, Sez87, SH84, Sie81, SH87, Val83, WF83, WA85b, WF80b, WF80c, WL81, YA87, AKS88, AK89, BD89, BT89, tCpS88, Che88c, DJ88, GM89, JW89b, KR89b, LS88a, LY89, Ski88, TYZ88, VR89a]. Interconnections [IS84, MB80, Pat81]. Interface [FTT^{+}80, RI89, RE80]. interfaces [RI89]. Interference [DK82, HV87, MBCG83, MAS84, Smi85, TN81, YPD82, Fuk88]. Interferences [Kan85]. Interleaved [RW81b]. Interleaving [Kim86]. Interlock [Mel87]. Intermittent [DM83a, LSM81, MS81, Sav80a, Sav80c, VR83]. Intermittent-Fault [DM83a]. Intermittent-Fault/Transient-Upset [DM83a]. Intermodule [CLH84]. Internal [HT83, Ozg86, XS85, Gai88c]. Internally [MM83b, RMCF88]. International [Ano80c, Ano80a, Ano82j, Ano82b, Ano82c, Ano82-60, Ano83b]. Internode [KK84a]. interpolating [Noe89]. Interpretation [LE80, WG88]. Interpreter [BM86a]. Interpretive [Sta81]. Interprocessor [BS87a, IC82, RB82]. Interruptions [Sha81]. interrupts [SP88]. Intersection [CW83, Til80]. Intersections [BW80, GW84, SW82, WLC89]. Interstitial [Sin88b]. interval [Loz83]. Introduction [AR83, AP86, Ano81-41, Ano82-58, Ano83q, Chu80, Kar82, LS81, LVW85, MR86, SJ84, Shi87, Sta87, Sti80, Tor87]. Invalidation [HS81]. Invariant [Ste83]. Inverses [TKL86, WTS^{+}85]. inversion [EA89b, Fen89]. Inverter [Nak87, Ozg86]. Inverter-Free [Ozg86]. Inverter-Minimum [Nak87]. Irredundant [LM86, SB80, RB88]. Ising [HCMP89]. Isolation [YML86]. Isomorphism [RW83, Sri89]. ISPS [VBH^{+}81]. Issue [Ano81t, Ano81r, Ano81s, Ano83k, Ano84e, Ano85c, Ano85a, Ano86c, Ano86d, Ano87c, Ano87z, CP82b, Shi87, Tor87, WS84]. Issues [Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83-31, AP85, VM87a]. Issuing [AKT86]. Iterated [Fra83]. Iteration [LC87b, RS86a]. Iterations [FYD84]. Iterative [AC84, AOES88, KM83b, Nak86, Par87, PR81a, DD88b, Man89, Mel89]. Itoh [HP88]. IV [RK86]. IV-Type [RK86].

Jacobi [LP89, WOH84]. January [Ano82-59]. Japanese [Men84]. JCIT [Ano83j]. Jerusalem [Ano83j]. Job [EHM80, FYK87, Sta85]. join [KA89, NT88]. Josephson [SGI89]. June [Ano82-59].

K. [CC89, MA89]. Kappa [KPR88]. Kernel [SGB87, Gai89]. key [CJ89]. Keys [MTMA85, Weg85]. Kleeman [EA89a]. Knapsack [Kar84a]. Known [MK84].

L [DD81, Eti80]. L. [EA89a]. Labeling [MTG^{+}85]. Ladders [CF80a]. LAN [SYJ89, SYK89]. Language [AO80, De 82a, KAGER82, PW81, ACGK88, JC88, SBSM89]. Languages [AO80, Ano87c, Cra85, De 82b, LM82b, QKS85, TF82, Veg84, Wan81]. LAN’s [ASP87]. Large [AF81, APD83, Coh85, DM81, DLM86, DF84, FK85, Hwa87, Met83, Pre83, QI85, RGA85, SL80, SR87, Wit81, WL81, YTL87, AJ88, AOES88, BBW88, DIY88, HA88, PKP89, Sin88b, YS89]. large-grained [AJ88]. Large-Scale [YTL87]. Latch [BJ83, Sav86, Fos89]. Latched [Cur80c]. Latency [CI87, SL86, CI89]. Latent [SBMM87]. laxity [HTT89]. Layer [CD88, PL84, Ric84]. Layered [KT85]. Layout [BK82, KH86, SOH^{+}81, SS85b, UV81, SR81a, TS88b]. Layouts [Tri82]. LBANC [Lam83]. learning [OM88]. Least [Sto80b, GM88b]. Lee [CC89]. Left [Cha83b]. Length [AL85, BDW86, CM87a, Dan83, SB84, Wus82]. lengths [dM88]. Lens [FS81]. Level [AO80, BIO82, BF83, BF80, Bry84, De 82a, HTOS80, IC80, LC87a, Mey84, OK83, PS83b, PS85b, Ram86a, Ram86b, Smi80b, Smi81b, SAA87, VM87b, WB81, Wus82, Jou89, KM89a, LM88b, SBSM89, WCS89]. LFSR [Hla86, WM86]. liable [Hay84]. Like [HL82, Wai88]. Likely [Dah86]. Lilith [Coo89]. Limit [Vui83]. Limitations [FWT82, Owe83, Rhy84]. Limited [KT87, LY83]. Limits [LL81]. Linda [ACGK88]. Line [FYD84, IO83, OE82, Owe83, RB83b, Sam80, Smi87, Sto80b, TY87, ZB87, KM89b, WE83]. Linear [APD83, BR80, Che86b, CD86, Cle84, DV87, Ers85, FHG88, Gaj81, GNK86, PC81, PP87, RV84, RFS86, SCP^{+}81, TC84, VR86, WK80, WM86, WM88a, WV87, WH80b, AL88, AOES88, BB89b, DD88a, KT89b, Mel89, NH88, PK88]. Linearly [MA86, FL88]. Lines [KS80a, YN83]. Link [IKT81, LC87a, Pra85b, Pra86a, Pro89]. Link-Level [LC87a]. Linked [DHH85]. links [YM88b]. List [Ano82-59, Ano84-30, Ano84-31, Ano85x, Ano86-32, Ano86a, Ano87o, Ano87w, Ano87y]. lists [Dav89]. LMS [MP83c]. Load [ABK83, BW89b, CA83, IBM82, KC87b, Lea86, SC89, WM85, Kum89, MTS89, SW89]. Load-Sharing [Lea86, SW89]. Loading [RW84]. Local [Ame82, Ano82d, BW89b, Bux83, Car84, CHL83, Coh85, DLSM81, Dav89, FT84, FYAV87, Gue86, Hwa87, Kam87, KM87, LCF89, LHPW85, RGA85, RG85, TV82, WFL82, YH84, YGZ^{+}87, KHS88, KK89, KP88, Ram88, RN88, RS88b, WJ85]. Local-Area [Bux83]. Locality [Kob86]. Locatability [Kre87]. Locatable [GH80a]. Located [Met83]. Locating [IKT81]. Location [CD86, DHS85, GP86, KS80a, LSM81, RB83b, Wu87a, YN83]. lock [CJ89]. Locked [KSB85]. Locking [Lee80]. Logarithm [HL86, Kar84b, LA85, LC87b, SCNS83]. Logarithmic [LZLH85, TGJR88, ST88b]. Logic [Aga80, Ano81z, Ano83b, BG86a, Boc82, BNM86, Bra83a, BCDM86, CH82a, Cra85, Cri80, Cur80b, Cur80c, DM81, DDG80, Dao81, DHR80, Feu82, Fuj81, FK81, FT82, Goe81, GH80b, Hay86a, HO81, Hur81, Hur84, KW85, KM86, KT81, LM82a, LM87a, LS80, LA85, MC86, MH80, MS86, MG86, Muk86, MR86, MH81, Ozg86, Pal86, Pap83, PR81a, Por82, RT85, RB83a, RB83b, RR86, RW84, SKF83, Sas81, SY83, SM87, SOH^{+}81, SGA81, SB85, Smi81a, Sto80a, SB80, TY87, TOM81, TC84, TJB86, WS84, WW83, YK82, YM86a, YWW86, ZW87, Zhi84, Bos88, Jha88, MKLC89, YM88a, Ano80a]. logic/fault [Bos88]. Logical [GCV80, GMG84, HFPS82, Kar87, Kre87, DG89]. Logics [Hay86b]. Long [NF84]. Look [AM80, GMG84, KY89, Dor88]. Look-Ahead [KY89, Dor88]. Look-Up [AM80, GMG84]. Lookahead [Rhy84]. Lookup [Par87]. Loop [BW81a, Hwa87, RGA85, WOH84, WI84, MI89, PK88, PKP89, RS88b]. Loop-Structured [WI84]. Looping [SC87]. Loops [AG81a, BG84, CF80a, HL82, Kob84, LAS87]. Loosely [LF80]. Loss [FG82]. LOTOS [CFM^{+}86]. Low [Gai85a, HTOS80, Kha84, KYSY80, TAF87, LPI88, NPP88]. Low-Cost [Gai85a, KYSY80, LPI88, NPP88]. Low-Level [HTOS80]. Lower [BNM86, BD80, DO85, DO86, NM82]. Lowest [Tsi86]. LRU [WM88b]. LSI [CL80, GCV80, LS80, RJ80, SOH^{+}81, Wei80].

M. [MA89]. Machine [Ano80-46, Ano80g, Ano81-27, AK85, BH81, BYH87, Bha83b, GMLV84, GY86, GM87, Hon85, ORS82, QI85, RTY^{+}88, SBSM89, ST86a, SA85, WR84, HM89, Jou89, NOYK88, VZMBH89, ACGK88]. Machine-Independent [Ano81-27, BH81, RTY^{+}88]. Machines [BH84, CIP87, CHM87, Cul80, DLSM81, DS80, Gaj81, Has84, HP87, IKP86, Klu83, LM83, MAV84, Oik87, PK80, Sal80, SD86a, SD86b, SS85a, SL87, SCP^{+}81, Tam80, UT87, WGT81, WC83, WR81, Yam80, BRG89, BF88, BS89c, Bur88, DD88c, Kap89, SGI89]. Macro [SGMP85]. Macro-Based [SGMP85]. MAFT [KWFT88]. Magnetic [BL80, BW81a, CLW80b, CLW80a, DGS80, Str82]. Main [ST87]. Maintaining [BL80]. Maintenance [Liu84b]. Major [BW81a]. Major/Minor [BW81a]. Majority [ZW87, YM88a, YF88]. making [Sta89]. Malicious [SR87, VM88]. Management [Ano81r, BFHW82, CP82b, QKS85, WFL82, ALL89, Bur88, LW88, RTY^{+}88]. Managing [TS83]. MANchester [PGR86]. MANIP [WM84]. Manipulation [Bry86]. Manipulations [AN84]. Manipulator [AS82b, MS82, LL88]. Mapping [Bok81, LA87, MF86, RFS86, SE87, KT89b, LW89]. March [SR81b]. Marginal [Cha83a]. Markov [BT86, MG82, MAS85]. Masking [KC86, RT86, SY83, EA89a]. Massive [GMLV84, GY86]. Massively [Bat80, HG87]. Matching [ACGK88, ST85, WBA83]. matrices [USM89]. Matrix [APD83, Ebe87, GNK86, HA84, HC82, LR80, Mel87, MT87, O’L87, RV84, ST87, Tha84a, Tsa81, Twi83, VRF84, VR86, Wan81, EA89b, Fam88, JK89, LP88, VR89b]. Matrix-Multiplication [VRF84]. Maximal [Wus82]. Maximal-Length [Wus82]. Maximally [MI85]. Maximizing [Sav80c]. Maximum [Agg86, Bok84, Jia86, Kon86, LS88b, Kap89, PP88, Pro88]. Mean [CdL89, GK83, Smi85, Sto80b]. Meaningful [YM88a]. Means [Hla86, Tha82, DSH89]. Measure [MS81, YM87a, GT88]. Measured [Cha83a]. Measurement [Ame82, CI87, De 82a, FH84, FHH^{+}83, IR86, PL83b, SL86, BM89, Mil88, SL88b]. Measurement-Based [CI87, IR86]. Measurements [AP85, WS88]. Measures [CGMP87, Man85, PC81, Smi80a, Sri87, Uhr82, Esf89, STR88]. Measuring [Kum88, NF84]. Mechanical [MSS82]. Mechanism [BI86, Cas86, Len85, SP82, GY89, Len88, LW88]. Mechanisms [BGM87, VM87b]. Membership [Ano85y, Ano87x, Til80]. Memoriam [Boo81]. Memories [BW81a, BR84, BD83, CLW80a, DL86, Fen85, HJ87, Hay80, KK84b, OJ85, PS85a, Pat82, RW81b, Ric86, SMV87, SN81, Smi87, SS82, SM82, SR80, SR81b, WvL85, BGM88, CCWZ88, FHG88, MP89a, MW88]. Memory [AS87a, Bai87, Bhu85, Bis85, BL80, BS83, Cha80, CS80a, Che86a, CS86b, CLW80b, CLP81, CM87b, CHM87, Dao81, Dav80b, DGS80, DV83, Dun85, ES80, Fet80, GMLV84, GK83, GY86, GGK^{+}83, Hag86, HNS84, HV87, IC80, IS84, Kan81, KF82, Kan84, KK84b, KS82a, KS86a, KYSY80, KL82, LV82, LV85, LJ87, MBCG83, Met82, MBR82, MAS84, NS87, Par86, Pat81, RW81a, RB83c, Sig82, Smi85, SD87, Str82, ST87, Tan84, Ten83, VRP83, Weg80, YPD82, AJ89, ALL89, CFP89, CI89, CMB88, FM88, Fuk88, LPI88, LD88, Noe89, PBL89, RTY^{+}88, SV88, SD89a, SGI89, Soh89]. Memory-Based [CHM87]. Memory-Resident [Hag86]. Merge [KH83, McC85, SSS89]. Merged [Swa80]. Mergesort [KB85]. Merging [AS87a, Bil89, Kru83, Par81]. Merit [Aup83]. Mersenne [TKL86]. Merwin [Boo81]. MESFET [HM89]. Mesh [LSSS85, MS89, Pre83, Sto83, BP89, Car88, SB88, UR88, Wai88, WCS89]. Mesh-Connected [LSSS85, Pre83, Sto83, BP89, UR88, Wai88]. Meshes [SE87]. Message [DS87, FJ88, WA85b, Len88]. message-passing [Len88]. Messages [LM85]. Metastable [KC87a]. Method [Ano81y, Bes83, Che82c, Fro83, FM84, GH83b, Gla80, JL85, Moh85, OI87a, PS85a, PS83b, PA81, RY82, SL83a, Yam80, Irs88, MKLC89, PC89, SS88b]. method-design [MKLC89]. Methodology [HL84, IK82b, KH86, JP88]. Methods [BAR87, De 82b, HR87, Mur81, RS87, SD86a, UT87, WK80, LPI88]. Metric [JH80]. metrical [GM88b]. Microarchitecture [AP85]. Microcode [DLSM81, Fis81, GB83, MP83b, SG83]. Microcomputer [FTT^{+}80, IC80, KH80]. Microcomputers [Wit81]. microeconomic [KS89]. Micronet [WV80]. Microprocessor [Abe84, BR83a, BCA80, Cli80b, CL80, De 83, EHS80, LF80, MLB87, NWLP82, RJ80, THH80, MBH89]. Microprocessor-Based [NWLP82]. Microprocessor-Controlled [RJ80]. Microprocessors [Ano80d, Ano80e, Ano80f, AG81b, AGL^{+}80, Bha83a, BA84b, MB80, SH81b, TA80, SS88b]. Microprogram [Abb83, Ano81-27, BH81, MH81]. Microprogrammable [HTOS80, ZN80]. Microprogrammed [BIO82, Bis85, IK85, MAV84, RB83c]. Microprogramming [RA80]. Microprograms [IKI83, TTT81]. MICROS [WV80]. might [Fis88]. migration [SD89a]. Migrations [Sta81]. MIMD [BBB^{+}82, GGK^{+}83, MS82, SSK^{+}81, UR88]. Min [Kri84]. Min-Cut [Kri84]. Minicomputers [AGL^{+}80]. Minimal [CM87c, DTF80, IKT81, SB88, SB80, YM85, RB88, RI89]. Minimally [Pag80]. Minimax [ST85]. Minimizability [FGPT89]. Minimization [Bes86, Bis85, Car84, GMG84, IC82, RB83c, RY82, Rot86, Zhi84, NMN89]. Minimize [BDW86, II81]. Minimizing [ACGT84, FTY87, Yam80]. Minimum [BP85, Bus83, CP87a, CW83, FS88, GK83, II83, LM87a, LSW87, MS86, Nak87, PC89, Pro81, SM83, Sie85a, WWSW87, CHRR89, CBAP89, HTT89, Kap89]. Minimum-Area [LSW87]. Minor [BW81a]. miss [Thi89]. Mixed [CSR86, DS80, Hua83]. Mixed-Mode [DS80]. Mixed-Radix [CSR86, Hua83]. Mobile [GFM83, GM88b]. Mode [Ano81z, DHR80, DS80, TOM81, YM86a]. Model [Ary85, Bry84, CMS82, Che82b, CM87b, CHM87, De 83, HT86, IR86, LF85, LE80, MLT82, Mar84, MP83a, Mey84, MP83b, Ros85b, SS80, SY83, SS86, SN81, SL84, Smi85, WHT84, WA80, WS82b, WH80b, WK81, AK89, AL88, BC88, tCpS88, KM89a, MAS85, Sch89]. Modeling [Ano85d, BL89, DGT84, FHH^{+}83, JA85, KM87, KP87, MBCG83, MP89c, Niz83, PS87, RE80, SL88b, Wei82, DT89, HIT88, SL89, WS88]. Models [Bra83b, CG87, DLM86, DG89, GCV80, Gav87, HT82, HT83, IDH86, KBB86, Li87, MG82, MAS84, SS85a, SGT86, Sny81, SD87, TB86, Tow86, BBG88, GZ88, KM89b, RN88, SS89b]. Modern [Lin82]. Modification [Mar81, Sto80b]. Modified [Car88, Don84, NH88, TS88a]. Modified-mesh [Car88]. Modular [Abr82, ABT82, AGL^{+}80, BK87, BS88, CAV86, FW88, HA86, JK80, KH80, Obe80, RV84, Sal80, SLJ88, Van86, VH84, PNH88]. Module [CH82a, NKY^{+}80, SD80, Wal88]. Modules [FYK87, SGA81, ST87, WF80a, RMCF88]. Moduli [GMG84]. Modulo [Bla83, Jul80, RY82, Slo85, WH88]. Modulo- [RY82]. modulus [SK89]. moments [dM88]. Monitoring [SS87]. monotone [CHRR89]. MOS [Ano85d, CV84, Ram86b, Bry84, CL80, DB87, EZS82, GCV80, JA85, NC88a, Ram86a, SOH^{+}81, TS84a, TS88b]. Most [Dah86]. Movable [GCW83]. Move [Len85, TL80]. Movements [Fla82]. Moving [Mos87]. MP [AG82, CS86b]. MP/C [AG82]. MPG [Ano81-27, BH81]. MRU [SR88c]. MSF [AS87b]. MTEC [EHS80]. Muller [DK89, FTY87, Pag80]. multi [Pro89]. multi-hop [Pro89]. Multiaccess [BW89b, DJ88, WJ85]. Multibus [LVF83]. Multicache [YYF85]. multicast [Gai89]. Multicomputer [AL82, BS86, IBM82, KK87, KK85, ST87, WM84, BF89, GM88a, KK88]. Multicomputers [NM87, VW84, RS88a]. Multicopy [Lee80]. Multicriteria [IPM82]. Multidimensional [CF80b, GS87, KS83a, Vai84, Vai86, Vai89, Ata88, PKP89]. Multigrid [CS86a]. Multihop [CP87b, CK87, CS89]. Multilanguage [BF88]. Multilevel [GT82, Kar87, Sil82, Sta81]. multimegabit [JP88]. Multimicrocomputer [RS83b, Ree84, WL81]. Multimicroprocessor [SLS82, SMN82]. multimodule [SL88b]. multinode [MI89]. Multipipeline [HX88, Sah84]. multipipelines [GZR89]. Multiple [AB80, AKT86, Aga80, AF81, Ano80a, Ano83q, Ano83b, BA86, BS87a, Bra83b, Cha87a, Cha80, CD86, CP87a, CF80a, Coy80, Cur80b, Cve87, Dav86, FT84, Hay86b, Hur84, HN80, IO84, Jha88, KT81, Kuo87, LVA82, LJ87, LM86, Mar84, MG82, Mor86, MR86, RT85, RS85, RAP87, RG85, SKF83, San89, Sas81, Sas85, SGA81, SD80, Tow86, VZ81, Van86, VSH89, WOH84, WW83, YPD83, Aky89, BB89a, BGY89, CMB88, Dub88, FW88, KK89, KP88, Max88, MG88, MAS85, RB89a, Sas89, SD89b, Tys88, YZ88]. Multiple- [VZ81]. Multiple-Access [RG85, MG88]. Multiple-Bit [Van86, VSH89]. Multiple-Bus [IO84, LVA82, LJ87, CMB88, KK89, MAS85, YZ88]. multiple-disk [RB89a]. multiple-input [Max88]. Multiple-Output [Dav86]. Multiple-Processor [BS87a, Cve87]. Multiple-Read [Cha80]. Multiple-Stream [YPD83]. Multiple-Valued [Ano80a, Ano83b, Cha87a, Hay86b, Hur84, KT81, Kuo87, Mor86, MR86, Sas81, Sas85, SGA81, WW83, BB89a, Sas89]. multiple-values [FW88]. Multiple-way [San89]. Multiplexed [NJM83]. Multiplexers [Pal86]. Multiplexing [Cli80b]. Multiplication [BD80, Jul80, LR80, MPPZ88, Mel87, Moh85, Nak86, RV84, Tsa81, VRF84, VR86, BS88, Fam88, JK89, Pin89, SS89c, TYY85, VR89b]. Multiplications [VSH89, WTS^{+}85]. Multiplicative [TKL86]. Multiplier [BVH83, Gna83, Gna85, HRT^{+}84, Joh80, MBH89, Pre83, RS86b, SRO84, STR87, Sip82, SR82, TY87, TH82, Tay82, Tay83, wu87b, FM89, NC88a, Smi89a, Zho88]. Multiplier/shifter [MBH89]. Multipliers [MSV80, SF84, YRT84, Che88b, Dad89, HTDR88, Wan89]. Multiply [PF83, ZG87]. Multiport [SMV87]. multiprecision [Sco89]. Multiprocessed [LAS87]. Multiprocessing [BDL83, WBA83]. Multiprocessor [AG82, BIO82, BYH87, BDW86, BD83, BBB^{+}82, CS86a, CS87, CN82, DSK87, DS87, Dem82, Dim85, GS81, HV87, HZ81, IO84, Jen81, KK84b, KN84, KA87, LS84, LW86b, LC87a, MBC82, MG82, MBCG83, Pra85b, Pra86a, Sah84, SSB87, SR87, Sig82, SSF80, SB83, SWK84, Tow86, Weg80, YH86, YPD82, Agr88, AJ89, BBW88, BM89, CMB88, CC89, DSH89, Fuk88, GZ88, HTK89, LW88, LW89, RTY^{+}88, SP89, TSS88, WG88, YBL89]. Multiprocessor-Based [LC87a]. Multiprocessor/Computer [AG82]. Multiprocessors [ASHK86, BB87a, CH84, Du85, DB82, KS83b, LVA82, Lun87, PKL80, Pat81, Pat82, RV86, SSS^{+}83, Smi85, Wah84, YTL87, AP89, ALL89, BL89, Dub88, FM88, JW89b, PC89, SV88, SD89a, VSSG88, WS88]. Multiprogrammed [CS80a]. Multiserver [MK83b]. Multistage [Agr83, AL85, BT89, FW81, KS83b, KSW88, LSD^{+}89, PL83a, PN85, PK80, SH87, WF80b, AKS88, tCpS88, Che88c, GM88a, GM89, KR89b, Lee89b, VR89a]. Multistep [Lee89a]. Multiterm [Bur84, Bur82]. Multiterminal [GL86]. Multithreshold [BZV86]. Multitrunk [CG87]. Multivalued [BZV86, CD86, Eti80, Mar86, PA81, Por82, RS86a, Ric86, Smi81a, Zhi84]. Multiversion [Her87]. Multiway [KB85]. mutual [Sin89]. MVA [Lam83]. MVL [Bes86]. MVS [Kob86, VI84].

n [BL84, DM84]. N. [Smi89a]. NAND [LM87a]. Native [UT87]. NBS [Ame82]. NC [CI88]. Near [GH83a, UT87, LD89]. Near-Native [UT87]. near-optimal [LD89]. Near-Perfect [GH83a]. Nearest [Chu85, Lee82, SE87]. Nearest-Neighbor [SE87]. Negabinary [Agr80, Yue80]. Negabinary-Binary [Agr80]. Negative [KM81]. Neighbor [Chu85, Lee82, SE87]. nested [PK88]. Net [SYK89, Smi80b, Smi81b]. Nets [GL86, Mol82b, BBG88, LM88b, SL89]. Network [AS82a, AS82b, Ame82, AL81, AL82, AS87b, BA84a, BP85, Bon83b, CP87b, CS86c, DLM86, EHM80, GF87, HT82, HZ81, II81, KM87, LSSS85, LL83, Lea86, Lee87, MS82, NS81, NS82b, Niz83, Niz84b, Niz84a, PR84, Pra83, Pra85a, Pra85b, Rud85, SS80, SSB87, Sez87, Ste83, TB86, WK81, WV80, WI84, WA85b, WF80c, WL81, WF81, WFL82, YL81, AP89, BP89, Ban88, BD89, Bil89, BT89, GM88a, HP89, KPR88, KR89b, LL88, LH88, LY89, LM89, LHC89, PL88, Pro88, Ram89, RS88b, SP89, San89, SR81a, Tys88, WJ85]. network-based [GM88a]. networking [HN88, HX88]. Networks [Agr83, AL85, Aky87, Ama83, BW89b, BG86a, BA83, Bhu85, BS87a, BR86, Bra86, BS87b, CCSW85, CH82a, CV83, CV84, CH84, CG87, CK87, CS80b, CFH81, CHL83, CM87b, DS87, DR85, DHS85, DJ81, DK82, Dot84, DHH85, EZS82, EH85, FW81, FT84, FYAV87, FF82, Fra81, FWT82, FH84, FJ88, GC87, Gel81, GFM83, GW81, GH80a, GH80b, Har86, Haw85, HO81, HT86, HC87, Hwa87, IKT81, IM87, IK82b, IO83, JLS80, JK80, Kam87, Kar87, KK87, KK80b, Kri84, KS83b, KSW88, Kub82, KDJ85, KJ86, LM82a, LM87a, Lam80, LL81, Lee85b, Lei85b, LC87a, LS89d, LHPW85, LM86, MM83a, Maz87, Mel87, MM83b, MM84, Nak87, NMB83, NM87, OS82, Obe80, Oru84, OOB85, OO87, PL83a, Par80, PN85, PK80, RGA85, RV86, RK86, RS83b, Ree84]. Networks [Rhy84, RG85, SW84, SL87, SS84, SH84, Sie80, Sie81, Sie85a, Sin88a, SB85, Sri87, SH87, TN81, Top89, TV82, Uhr82, Val83, Von83, WC84, WF83, Wit81, WF80b, YA87, YH86, YH84, YGZ^{+}87, AA88, AKS88, AK89, Aky89, BBG88, BGY89, Bil89, BN88, CG88, CH89, tCpS88, CL88, Che88c, CM89, CS89, CdL89, DK89, DD88b, Esf89, GH88, GM89, GM88b, HT88, JS88, JA88, JW89b, KHS88, KK89, KP88, LS88a, Lee89b, LSD^{+}89, MI89, MKLC89, Pro89, Ram88, RN88, RK89, San88a, San88b, Ski88, Sri89, TS88a, TYZ88, VR89a, WB88, dM88, Ano82d]. Neumann [BE87]. Ninth [Ano82-60]. NMOS [FTT^{+}80, THH80]. no [RM86]. Nodal [Niz83]. Node [EHM80, OK83]. Nodes [CP87b, GFM83, GF87]. Noisy [IM87]. Non [BB87a]. Non-Uniform [BB87a]. Nonbinary [Dao81, SOA85]. nondeterminism [Par89]. nonequivalent [AKS88]. Nonfault [Kre87]. Nonlinear [SSVV81, VSV81]. Nonminimality [Coy80]. Nonnumeric [BIO82]. nonrectangular [Che88c]. Nonrepairable [FM84]. Nonrestoring [Bus83]. Nonuniform [LP81, CMB88, Jou89]. NOR [LM87a]. normal [HTDR88, Wan89]. Normalization [ZB87, cLW88]. Note [Jou86, Mar85, NK87a, CCF89, CH89, Par89]. Notes [Par80]. Notice [Ano80-27, Ano80-28, Ano81-28, Ano82y, Ano82z, Ano82-27, Ano83o, Ano83p, Ano84g, Ano85e, Ano85f, Ano86e, Ano86f, Ano86g, Ano87e, Ano87f, Fen84b, Fen84a]. Novel [Bha83b, Bos88]. NP [NMN89]. NP-complete [NMN89]. Number [AS82b, CSR86, DDG80, DLM86, GMG84, Hua83, HML84, IK82a, Jen83, Jul80, Kod81, KM81, LM87a, Mar85, MK85, MP83c, OA83, Ram83, SM83, STR87, TGJR88, TRYS83, TCH^{+}86, WG80, GM89, HMC89, RB88, WLC89]. Numbers [Fro83, Ger82, KM83b, STR87, Ulm83, Dad89, Par88]. Numerical [GV84, Kar81, SB85, Tsa83, GT88]. NYU [GGK^{+}83].

O [KB85, LY83, PW81, RB89a, Sto89, WSR84, XS85]. Object [Fra83, RB89b]. Object-Oriented [RB89b]. Objects [SGB87, BR88]. Observability [Sav83a]. Observations [Mur81]. Obstacles [WWSW87]. Obtaining [MI85, RI89]. OCR [RJ80]. Odd [DDG80, Kan84, KH83, SOA85]. Odd-Even [KH83]. Odd-Valued [DDG80]. Odd-Weight-Column [Kan84]. ODE [BAR87]. Office [NWLP82]. omega [Bil89, Ste83]. On-Chip [KS86a, LCW81, Maz87]. On-Line [IO83, OE82, Owe83, RB83b, Sam80, Sto80b, TY87, ZB87, WE83]. On-the-fly [EL87]. One [AC84, CIP87, HCMP89, SKW88]. One-Dimensional [AC84, HCMP89, SKW88]. One-Way [CIP87]. Only [Dav80b, GK83, Sie85b]. onto [FYK87, SE87]. Open [CPG83, RR86, Jha88]. Operands [BVH83, PF82]. Operating [Ano87c, VI84, WV80]. Operation [GKS87]. Operational [DG86]. Operations [CP82a, Gue86, HA84, RFLS86, ST87, BF89, Ble89, NH88, SR88c, ZMC89]. optic [GC89]. Optical [Lei84, MP83c]. Optimal [Agg86, AS87a, Ary85, AH86, AT81, BPV83, BW80, BBP88, BP84, BP82, CA80, CW83, CH82b, CH83, DV87, Fam87, GT82, GCW83, GK85, KP80a, KLL87, Lau81, LS82a, LW85, MTMA85, MK83a, Man85, MB80, MD86, MG86, NS82a, Oik87, Pal86, Pap83, PTT81, Pre83, PL84, RW81b, RM86, Rub81, ST85, SLL87, Sin88a, Sri83, Sto89, TS86b, UV81, VR86, VR89b, WHT84, YKL88, Zak85, BP89, Bec88, FM89, KS89, LM88a, LD89, Sas89, SSS89, SW89]. Optimality [Val83]. Optimally [SSB87]. Optimization [AL85, ACGT84, CS80a, FYAV87, Gav87, HN80, KT85, PS83a, Sas84, Sil83, Tha82, TTT81, TKK86, YH84, BS89c, HS88, RI89]. optimizations [Pol88]. Optimized [TS88b]. Optimizing [BS83, Hoc83]. Optimum [HO81, JH89, Rob83, TR88]. OR-parallel [Cra85]. Order [CE87, HP87, ZMC89]. Ordered [Fet80, WV87, LD88]. Ordering [Li87]. Orderly [KL82]. Ordinary [BR83a].

Organization [CP87b, LVF83, Soh89]. Organizations [MWM80, SG85a]. Organized [Dao81, DV83, Dun85]. Oriented [Abr82, Che86a, LF87, RB89b]. Orrery [ADG^{+}85]. Orthogonal [NMB83, HTK89]. Orthogonally [MT87]. Other [O’L87, VS86b, Von83]. our [Ano80d, Ano80e, Ano80f]. Out-of-Order [HP87]. Outcomes [SS86]. Outerplanar [FJ88]. Output [ASHK86, Cur80a, Dav86, Gna83, Kar83, KS80a, Oik87, SD86b, Sas84, Sip82, YN83]. Outputs [Bha83b]. Overflow [Tay83]. Overflow-Free [Tay83]. Overhead [BNM86, JH80, Kha84, TAF87, CBAP89]. overlap [PP88]. Overlapped [VSH89]. Overview [Sie81].

P [VS86b]. Package [VRP83]. Packet [BCA80, CH84, CS80b, Gel81, GW81, KDJ85, Lam80, LL81, LC87a, LS89d, LHPW85, RG85, Tas83, CS89, Pro88, Pro89, YBL89]. Packet-Switched [KDJ85, YBL89]. Packet-Switching [GW81]. Packing [Cha83b]. page [SD89a]. paged [RTY^{+}88]. Paging [ASKL81, BF83]. Pair [Har86]. Pair-Wise [Har86]. Pairwise [DD88a, Sor85]. Papers [Ano80l, Ano80x, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80u, Ano80v, Ano80w, Ano81t, Ano81r, Ano81s, Ano81u, Ano81v, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano82x, Ano82u, Ano82v, Ano82w, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano83j, Ano83k, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano84d, Ano84e, Ano85b, Ano85a, Ano86c, Ano86d, Ano87c, Ano87d, Ano87b, Ano80t]. paradigm [SW89]. Parallel [AS87a, Ano82b, Ano82c, Ano82-58, Ano86c, APD83, BM86a, Bat80, Bat82, BAR87, BK82, BD83, BE87, Can83, CIP87, CT84b, Cur80c, Dad80, Dan84, DS83, DC81b, DC82, Ebe87, ED83, Gaj80, Gaj81, GV84, Gar80, GLS82, Gna85, GGK^{+}83, GK85, HJ87, HT82, HL82, HNS84, Hla86, Hoc83, HP89, HZ83, HMC89, Hua83, HN80, HG87, IC82, JK82, Kar84a, KN84, Kru83, KH83, LMC87, LA87, Lei85a, LW86a, MP89a, MA86, MTG^{+}85, MI85, ML82, NS81, NS82b, NMB83, Nil84, OJ85, PL83b, PS85b, PK87, RB80, SM83, SS89b, SJ84, Sie81, SSF82, SBGS86, Sip82, Sip84, Sri83, Sta87, TB86, TRYS83, TLR83, Tsi86, TS84b, Van80, VS86b, WK80, Wan81, Wei82, Wen85, WOH84, WvL85, WH80b, WD84, YA87, YTY82, wu87b, ACGK88]. parallel [Ban88, BRG89, BF88, Ble89, Bok88, Bos88, Car88, Che88b, Cra85, DD88a, EZL89, FL88, Fis88, FM88, HTK89, JAM88, LP89, MS88b, NC88a, NT88, NS88, Par89, PKP89, RJ89, SP89, SS89c, YJ89]. Parallel-Pipeline [TLR83, WD84]. Parallel-Pipelined [BD83]. parallel/pipelined [BRG89]. Parallelism [FH80, HTOS80, NF84, PB87, WS82b, Wor81, Zak84, AJ88, Jou89, Kum88, Pol88, PKP89]. Parallelization [NS87]. Parity [CH85a, FMM84, HO81, KM84a, LM87a, Met83, BS89b]. Parsers [TF82]. Parsing [CIP87]. Part [DF84, SSVV81, VSV81]. Partial [Fam87, GLS82, HQR89]. partially [JW89a]. Partition [SBK85]. Partition-Tolerant [SBK85]. Partitionable [BS86, LW86b, SSK^{+}81, LW88]. Partitioned [HC82, JL85]. Partitioning [AJ88, BB87a, Bok88, CM89, Cve87, Fam87, FWT82, Kri84, Kri87, Mar86, MF86, Sie80, WC84, PC89, San89]. Partitionings [RAP87]. PASM [SSK^{+}81, TS86b]. passing [Len88, RN88]. Passive [OK83]. Path [Abr82, AL85, DB87, GC89, HR87, KLL87, PL83a, RV86, BP89, RI89]. Path-Oriented [Abr82]. Paths [Bra86, Bra83b, Che82a, Kan85, Lak84, SBGS86, WWSW87, CH89, WLC89].

Pattern [Ano81r, BFHW82, CP82b, DB87, Fu80, Hay80, SK85, SDB84, SB84, SN81, SSK^{+}81, SR80, TW83, TC84, Tri82, WBA83, Che88a, DFC89, GNH88, MP89a, SM88, dJvdG88]. Pattern-Sensitive [Hay80, SN81, SR80, MP89a]. Patterns [BCR83, CS85, HO81]. PDP [LGH80]. PDP-11 [LGH80]. Perfect [GH83a, HT88]. perfect-shuffle [HT88]. Performability [FM84, GT87, IDH86, Mey80, MFW80, Mey82, STR88]. Performable [AS82b, GM89]. Performance [AP89, ASKL81, AKT86, AAG^{+}87, Ano83q, BSV83, BA83, BP84, BM89, tCpS88, CL88, Che88c, CG87, CLP81, Cve87, De 82b, De 83, Du85, ES80, ED87, Fra81, FHH^{+}83, GS89, GT80, GKS87, HJ87, HL84, HV87, HIT88, HP87, IO84, JH80, Kam87, KY89, KC86, KP87, KS83b, KJ86, KB85, Lam80, LMO84, LF80, Li87, LC87a, LHPW85, Mar84, MBC82, Mol82b, Niz84b, Niz84a, NS87, PH86, PL83b, Pat81, PGR86, Pra86b, RS83b, Ree84, RAP87, SL84, SC87, SRO84, SB83, SS81, Sri87, SD87, Tas83, UT87, VSSG88, WS88, YH84, ALL89, BL89, CMB88, EA89a, Gai89, GH88, GZ88, HTT89, JP88, Jou89, KHS88, KP88, KR89b, LPI88, Lee89b, MAS85, Soh89, WH88, YZ88, Zwa85]. Performance-measurement [BM89]. Performing [RW84]. Periodic [WvL85]. Periodically [Can83, Nil84, Wen85]. Peripheral [FTT^{+}80]. permissible [MKLC89]. Permutation [HT86, Kub82, Lee85b, NS81, NS82b, OO87, Sie80, SH87, OT89, Sri89]. Permutations [AS82b, NS82a, RK86, YL81, GM89, USM89]. Permuting [CM85, CM89]. Personal [CLP81, LMO84]. personalized [JH89]. Perspective [Sta84]. Petri [BBG88, LM88b, Mol82b, SL89]. PFFT [Ano84f, CVY83]. Phase [Hoc87, KSB85, Sas84, Val83, ED88, SKW88]. Phase-Locked [KSB85]. Phenomenon [RT86]. Physical [GCV80, WWS84, DG89]. PICAP [AGH^{+}82]. PICCOLO [YK82]. Picture [YK82]. Piecewise [RM83]. Pin [FWT82]. Ping [SC87]. Ping-Pong-Type [SC87]. Pipeline [Ano81y, ED87, Gla80, GM87, Pat80, STD^{+}85, TLR83, WD84, Bok88, MG88, SR88a, SGI89, TRH^{+}88, Tys88]. Pipelined [AJ82, Ano84f, BD83, BJS80, CVY83, Cur80c, Gaj81, HI80, RW81b, VZ81, WS84, BRG89, SGI89, SP88]. pipelined-memory [SGI89]. Pipelines [NH85]. Pipelining [JMKN86, Jen81]. pipes [Zwa85]. Pivoting [Sor85, HQR89]. Pixels [RS85]. PLA [BL84, Fuj84, Kha84, RT86, RH87, Sas84, TS84a, TAF87, YM85]. Placement [NKY^{+}80]. Planar [CW83, McC81]. Plane [Lee82]. Planned [Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83-31]. Planning [LP83]. PLA’s [CHRR89, BB89a, ML88, RB88, Sas89, HR88]. Plural [UT87]. PM221 [LY89]. PM2I [SS84]. PMC [KM89a]. PN [Mor80, Wus81]. Point [AP85, FG82, RB83b, SW84, Vai82, Fuk88, Loz83, ST88b, WE83]. Point-to-Point [SW84]. Points [Von83, WG80]. Poisson [Hoc83, Sha81]. Policies [BK87, SG85a, Sto89]. Policy [CS80b, BP89]. Polygon [AT81]. Polygons [CW83, LS82a]. Polylogic [Por80]. Polymorphic [LM89, LW86b]. Polymorphic-torus [LM89]. Polynomial [BK84, BD80, HSE84, LS89d, Muk87, STR87, TLR83, ZH85]. Polyprocessor [Man84]. Pong [SC87]. Pool [KK84c]. Poolpo [KK84c]. Portion [SOH^{+}81]. ports [WB88]. Positive [KM81]. Power [Vui83, RK89]. Practical [AM85, KN84, ML82, PK87]. PRAM [AS87b, Gel89]. Precedence [CL87]. Precise [Abe84, LA85, SP88]. Precision [CHH83, KM84b, MPPZ88, MK85, KM83a].

Predicate [CS84]. Prediction [FMM84, FM87, GT83, Thi89, VSSG88]. Predictor [BAR87]. Preemptive [ZRS87, GY89]. Preface [Fen84c, Hon82]. Prefetch [De 83]. Preliminary [Ano82-60, GB83, Men84]. Preloading [TS84b]. Presence [MK84, SSB86, SR87, WWSW87, YN84, VM88]. Preset [LM83]. President [Boo81]. Presortedness [Man85]. Press [Ano82-29, Ano80-30, Ano80-31, Ano82-28, Ano82-30, Ano82-31, Ano82-32]. Prevention [Gel81]. Prime [Jul80, LV82, MS86, Ten83, TKL86, TRH^{+}88]. Primes [Kuo87]. Primitive [ZMC89, BD89, Ble89]. Primitives [BF80, Hon85]. Principles [TB82]. Printed [NKY^{+}80, Sch89]. Prior [Ano83c]. Priority [Sha81, NK88a, SD89b]. Private [BD83, Pat82]. Probabilistic [CPG83, GM87]. Probabilistically [Dah86]. Probabilities [Mar87, KT89a, LS89c]. Probability [Cri80, LJ87, Sav80c, WR84]. Probing [Cle84]. Problem [Bok81, CPG83, Cve87, DLC87, FYD84, FYAV87, Jia86, Kar84a, Lak84, Lau81, LL84, LP86, MTG^{+}85, MR82, Muk87, OS82, PTT81, RAP87, Rub81, Smi80b, Smi81b, YYF85, YM85, ZB87, Bos88, CD88, NMN89, OM88, ST88a, SW88]. Problems [AS85, AH86, BB87a, CW83, FH80, FT82, KK80a, KLL87, Sav84, Til80, VS86b, WM84, Wor81, BRG89, Bok88, LS89b, Ram89, SAA89]. Procedure [AO80, Cas86, HO81, KT85, RC89, RW84, SP82, VR83]. Procedures [KKS86, LW86b, SR80, Wan82]. Proceedings [Ano80d, Ano80e, Ano80f, Ano80g, Ano80-47, Ano80-48, Ano80-49]. Process [EHM80, KK84c, LD87, Ros85a, SL84, YM86a, YGZ^{+}87]. Processes [IK82a, KL82, Len88]. Processing [Ano80-46, Ano80g, Ano82-58, Ano86c, AGH^{+}82, Bat82, Bes83, BPM^{+}86, CL84, CH82b, CH83, CLH84, CH85b, Gar80, GS87, GMG84, HT82, HZ83, HN80, KN84, KY82, KWR82, LA87, LM87b, ML82, NMB83, NWLP82, Par86, SJ84, Sie81, SSK^{+}81, SSF82, TB86, THH80, TS84b, Zak84, DIY88, FL88, HX88, MG88, SP89, Ano82b, Ano82c]. Processor [Ano83q, Ano84f, BA86, Bat80, Bhu85, BS87a, Bok84, CS87, CVY83, CT84b, CM87b, Cve87, Dot84, FL88, FS80, FK85, Fla82, FR85, FHH^{+}83, HKSS86, HT80, Hen84, HFPS82, HL82, JTP85, JS84, KS82a, KB84, KSB87, KS82b, KAGER82, LS82b, Len85, Liu84b, LJ87, Man84, Mar84, MA86, MWM80, NJM83, Nwa85, OS82, Pat81, PB87, Pra85a, RW81b, RAP87, Ree80, Ros85b, SE87, SS87, Shi82b, TGJR88, VZ81, VBH^{+}81, WOH84, YA87, ZN80, CM88b, Cos88, Gel89, Kum89, PKP89, Sin88b, TSM88, Tys88]. Processor-Memory [Bhu85, CM87b, Pat81]. Processor-Memory-Switch [KS82a]. Processors [AKT86, Agg86, ACGT84, Ary85, Bis85, BJS80, BE87, CT84a, FYK87, GNK86, Jes80a, Jes80b, KK84c, Kor86, Lu82, MA86, RB80, Ros83, Sip84, WD84, Ata88, DIY88, FLN89, HP89, Hos89, KJC89, MM88, NK88b, SL88a, SP88, Soh89, WCS89]. Product [Aky87, Aky89, Bla83, CMB88, MT87, Slo85, Sri87, Tsa81, CdL89, dM88]. Product-form [CMB88]. Professional [Ano81w, Ano81x]. Profile [Kob83]. Program [ASKL81, Ano82-60, De 82a, SM82, SG85b, SO89]. Programmable [Aga80, BNM86, DM81, DC81a, FK81, Hur81, LA85, MG86, Ozg86, Pap83, RT85, RB83a, RB83b, SKF83, Sas81, SM87, Shi82b, TOM81, Wei80, YWW86]. Programming [Ano87c, CMM87, OO87, QKS85, BB89b, BF88]. Programs [BC85, CL84, ED87, HT83, KK82, LMC87, LM87b, Mil82, Tha82, VM87a, Mil88, Thi89, WG88]. Progressive [LYI87]. Prolog [NOYK88]. Proof [AM87, VSH89, LP89]. Proofs [Sam80]. propagating [NK88b]. Propagation [HC86, SMV87, SL88b]. Properties [CM85, CF80b, KM86, KF80, PTC86, RS86a, SS88a, Ste83, CM89, KM89a]. Proposed [Mar85, Veg84]. Prospects [Smi81a]. protect [Gef73]. protected [BGM88]. Protection [BS82, Met82, Sny81, Van86]. Protocol [CHL83, GFM83, GF87, Hol82b, KT85, Lam80, RG85, Smi81b, TRTN89, YH84, KP88, Smi80b]. Protocols [Her87, Hof85, LS89a, RE80, RW82, TV82, Coa88, Gui89, LCS89, SD89b, YBL89]. Prototype [MC86]. Provably [SGT86]. Providing [Van86]. Proving [KW85, ST86a, WW83, KSW89, PL88, VZMBH89]. PROWAY [CFM^{+}86]. Pseudo [Hay86a]. Pseudo-Boolean [Hay86a]. Pseudoexhaustive [GNH88, McC84, WM86, HR88]. Pseudoparallel [AJ82]. Pseudorandom [BM86b, CM87a, WCM87]. Publications [Ano80-32, Ano80-33, Ano80-34, Ano80-35, Ano80-36, Ano80-37, Ano80-38, Ano80-39, Ano80-40, Ano80-41, Ano80-42, Ano80-43, Ano80-44, Ano81-29, Ano81-30, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano81-37, Ano81-38, Ano82-33, Ano82-34, Ano82-35, Ano82-36, Ano82-37, Ano82-38, Ano82-39, Ano82-40, Ano82-41, Ano82-42, Ano82-43, Ano82-44, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Ano83-29, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m]. PUMPS [BFHW82]. Purpose [HFPS82, Liu84b, SL80, SLS82, Wai88]. pyramids [MS88c].

Q [RMCF88]. Q-modules [RMCF88]. Quadratic [STR87, TCH^{+}86].

Quadratic-Polynomial [STR87]. Quality [Sav83b]. Quantitative [GJ80, OA83, SD87]. Quantization [Niz83]. Quarter [Joh80]. Quasi [IM87, Tam80]. Quasi-Cut-Through [IM87]. Quasi-Stable [Tam80]. Quaternary [Cur80c, Dao81, MC86]. Queries [Che82c, YGZ^{+}87]. Query [CL84, CH82b, CH83]. queue [KA89, dM88]. Queueing [Aky87, CFH81, DLM86, Har86, HT82, HT83, Lee80, Niz83, SS80, Sri87, TB86, Aky89, BBG88, CdL89, SW89, dM88]. queues [NK88a, NT88]. Quick [KS86b]. Quicksort [Weg85]. Quotient [FF82].

R [Smi89a, Lam80, Tas83]. R-ALOHA [Lam80, Tas83]. Race [Ram86a, Ram86b]. Rademacher [HS84]. Radio [CP87b, CK87, CS89, Pro88, Pro89]. Radix [Arm80, Bus83, CSR86, GH83a, HGS83a, Hua83, Jes80a, KM81, SOA85, TPSS85]. Radix- [GH83a, HGS83a, TPSS85]. Rail [Kha82, TY87]. Railroad [SMN82]. RAM [FH86, Reg88, SK85, dJvdG88]. RAMs [DFC89, JP88]. Random [DTF80, DFC89, Feu82, Mar85, PS85a, Sav83b, SDB84, SB84, SM88, SN81, Sha81, SOH^{+}81, SR80, SR81b, TV82, VR83, FGPT89, FHG88, HMC89, MP89a, MW88]. Random-Access [SR80, MP89a]. randomized [JAM88]. Range [Jou86]. Ranges [SW82]. ranked [ZMC89]. Ranking [Zak85]. Rapid [IS84]. Raster [CH85a]. Rate [Cur80a, MSV80, TS86b, ZW87, San88a, San88b, TSM88]. Rate-Multipliers [MSV80]. Rates [Iye84]. ratio [Thi89]. Rational [Fro83, KM84b, MK85, KM83a]. Rationals [Kri83]. Ratios [Sil82]. Reach [Ano80d, Ano80e, Ano80f]. Read [Cha80, Dav80b, GK83, SS82]. Read-Only [Dav80b, GK83]. Read/Write [SS82]. Real [AJ82, Ano86d, Ano87z, ABT82, ACD82, CL87, CMM87, Ger82, KO87, KSB87, KC87b, MLB87, PS87, SGB87, Shi87, SLL87, SC89, SMN82, VM87b, WFL82, HTT89, HIT88, KW89, SRC85, Sta89, WS88]. Real-Time [AJ82, Ano86d, Ano87z, ABT82, ACD82, CL87, CMM87, KO87, KSB87, KC87b, MLB87, PS87, SGB87, Shi87, SLL87, SC89, SMN82, VM87b, WFL82, HTT89, KW89, SRC85, Sta89]. Realistic [DO85, DO86]. Realization [IC80, Por80, SCP^{+}81]. Realizations [CH82a, RR86]. realize [BB89a]. Realizing [TYZ88, UT87, JW89a]. reallocation [Sta89]. really [Gef73]. Rearrangeability [Lee85b]. REBUS [ACD82]. Reciprocity [Agr80]. recoded [Par88]. Recognition [Fu80, SSK^{+}81, Van80, Ram89, YJ89]. Recognizers [Tri82, Fos89]. recognizing [CI88]. Recomputing [PF82]. reconfigurability [RI89]. Reconfigurable [GZR89, KK84a, LF87, RM86, TKK86, TS84b, WV80, LS88a]. Reconfiguration [CN82, Kan81, LW86b, OOB85, RM86, SD80, WCS89, Hos89, HA88, RS88b, UR88]. Reconfigure [KK87]. reconfiguring [LLJ89]. Recovery [Hag86, KKS86, KS86b, LGH80, LS84, LYI87, OJ80, TS86a, TS87, VI84, YH86, CC89, Hos89, KW89, LM88a, MW88, SLJ88, US88]. Rectangle [GL86, GW84]. Rectangles [BW80]. Rectangular [AR82, CM85, Kan85, CC88, WLC89]. Rectilinear [WWSW87]. Recurrence [Gaj81]. recurrences [PC89]. Redistribution [CA83]. Reduce [Owe83]. Reduced [JO84]. Reducing [BR86, LAS87]. Reduction [Klu83, LVF83, NH85, SS89a, DD88a]. Redundancy [BK87, Bra83a, Fam87, Kak83, KH80, KC86, KP87, MS81, RB83d, SD80, Van86, Agr88, BW89a, EA89a, JS88, RI89, Sin88b, WCS89]. Redundant [Can83, Nil84, PL83a, RV86, SB80, TY87, Wen85, Wu87a, EL87, SK89, TYY85].

Redundant-Path [RV86]. Reed

[DK89, DC87, FTY87, HRT^{+}84, Liu82, Liu84a, OI87a, Pag80, STD^{+}85, SR88a]. Referee [Ano86a, Ano87o, Ano87y]. Referees [Ano82-59, Ano84-30, Ano84-31, Ano85x, Ano86-32, Ano87w]. reference [KM89b]. references [CMB88]. Reflected [Er84]. regeneration [LCS89]. regeneration-based [LCS89]. Region [BS83, LS82a]. Register [BR84, CF80a, Dan83, Dav80a, ES80, HSE84, Mor80, PS83b, PS85b, TS83, WM86, Wus81, WM88a]. Registers [Che86b, Blo88]. Registration [RS85]. Regular [AL82, BK82, CM85, Haw85, ISO85, Muk86, BP89, Fos89, YM88a]. Regularization [SG85b]. relabeling [GW89]. Related [PB87]. Relation [Mar86]. Relational [Che82c, DGS80, Hon85, QI85, Str82]. Relations [CL87]. Relationship [IBM82, YF88, Zhi84]. Relationships [Agr80, dM88]. Relay [GF87]. Reliability [Ano81y, BGM87, Can83, CMS82, ES80, GT83, Gla80, HR87, Iye84, Kan81, KS82a, KS80b, MS81, MST85, MBR82, Nil84, RKH88, SGT86, SB80, WA80, Wen85, BT89, BGM88, LCS89]. Reliable [Ano80-47, Ano80-48, Ano80-49, Ano81s, Ano83k, Cas86, Cha87b, Hon82, Hwa87, RGA85, RS88a, SP82, SBMM87]. Reliably [GC80]. Remainder [Vu85]. remapping [NS88]. Remark [Coy80]. Remote [Cas86, RC89, SP82]. Remotely [Ano80-46, Ano80g, Met83]. Removal [Fam87]. Reorderings [O’L87]. Repair [HKR84, HP87, MM80, Man80, RM86, CFP89]. Repairable [DG86, MDG89]. Repaired [Can83, Nil84, Wen85]. Repetitive [Wan81]. Replacement [BF83, FH86, Obe80, SG85a]. Replay [LMC87]. replica [LCS89]. Replicated [CH85b, GC80, Met83]. Reply [BH80, Des81, GLL81, Mil85, Nil86, RTW80, Ros86, SM80, Smi86, SCA81, SCN86, Weg82, EA89a]. Reporting [BW80, SW82]. Representation [CV83, FH80, FYSK84, Ger82, HRJ86, PA81, PK80, TY87, Wor81]. Representational [Bou89]. representations [EL87]. Representing [De 83]. requesting [tCpS88]. Requests [RW81b]. required [BB89a]. Requirements [MA82, RSZ89]. Resampling [WS82a]. Reservation [CHL83, KP88]. Resident [Hag86]. Residue [CSR86, GMG84, Hua83, Jen83, MP83c, MG86, Pap83, Ram83, STR87, ST86b, TH82, Tay82, Tay83, TCH^{+}86, Ulm83, Vu85, Cos88, WH88]. Residue-Based [MG86, Pap83]. Resolution [KL82, RB89b]. Resource [HN80, JW89b, RSZ89, Wah84, WJ85, ZRS87, KS89, LW88, NS88]. Resources [KT87, Uhr82]. Respect [Oik87]. Response [GMK85, LP81]. Responses [Cha83a]. Restricted [Jou86]. Restructurable [Pra85a]. Restructuring [LJL89]. Result [Bus83]. Results [IBM82, Jes80b, MR82, Men84, KP88, WH88]. Retransmission [ASP87]. Retrieval [Fet80, CCWZ88, LD88]. Retrofitting [AP85]. Reverse [WF80c]. Reverse-Exchange [WF80c]. Reviewers [Ano87z]. Revisited [NN87]. Rewritable [TOM81]. Rhyne [Smi89a]. Richard [Boo81]. Ring [AL81, DHH85, JLS80, KM87, LHPW85, Ban88, BGY89, CL88, HP89, HN88, RS88b]. Ring-Structured [JLS80]. Rings [KK87, Mah80]. Ripple [CP87a]. RISC [HM89, TS83]. RNet [CMM87]. RNS [SK89, TPSS85]. Robust [Rud85, TS86a, TS87, VRF84]. Role [Liu84b]. rollback [LM88a, US88]. ROM [OI87a]. Root [OE82, ZG87, RJ89]. Rooting [Maj85]. Rotation [AH86, Fra83]. Rounding [Lee89a, Sco89]. Routines [Li85]. Routing [AH86, CS80b, DS87, DL87, EH85, FJ88, GL86, GW81, Jes80b, LSD^{+}89, MS82, NS81, OP84, PL84, RS83a, Ric84, SC87, TRTN89, Val83, BP89, BS89a, CD88, DIN88, HT88, LD89, SW89, VR89a]. Roving [BI86]. Row [DL87, FH86, RS83a, BS89a, DIN88]. Row/Column [FH86]. Rulers [Rob83]. Run [Nic89]. Run-time [Nic89]. Running [GK83].

Same [ST86a, VZMBH89]. Sampled [Sto80b]. sampling [HCMP89]. Sara [RE80]. Satellite [BCA80]. Satellite-Born [BCA80]. satisfying [WB88]. Save [LM82a]. SbED [Kan84]. Scale [LL85, SL80, YTL87, PKP89]. Scan [CH85a, HS86, SD86a, MA89]. Scanning [VSH89]. Scans [Ble89]. Scene [AJ82]. Schedule [BDW86]. Schedules [AB86]. Scheduling [Ary85, BSV83, BS87a, BDW86, BJS80, DS83, EHM80, Fis81, GF87, GKS87, GB83, KN84, KK85, KS86b, LM87b, Man84, MA86, PK87, RW81b, RSZ89, Ros85a, Sah84, Sig82, Sri83, Sta85, VW84, VM87b, Weg80, ZRS87, BRG89, CNO^{+}88, CKS88, GY89, HTT89, Pro89, SRC85, WJ85]. Scheduling-Function-Based [GF87]. Scheme [Adi84, ABK83, Bur84, CLW80a, Hag86, HJ87, KY89, KS86a, LSD^{+}89, Oru84, PK87, RB82, SD80, YTY82, Bur82, HA88, LW89, Sin88b]. Schemes [FT84, KH83, Lee80, MS82, TS84b, UT86, WvL85, WV87, CL88]. Schur [Ebe87]. Science [Ano81-44, Ano81-45, Ano82-61]. Scientific [Abe84, KUV85, KS82b, HTK89, Kum88]. scientific/engineering [Kum88]. Scope [Moh85]. Search [BDL83, BW81a, CT84a, DL86, Ell80, Got81, NN87, Pra86b, WM84, Sto89]. searches [Sto89]. Searching [DGS80, Kru83, Str82]. SEC [Dao81, Dun85, Gai88b, Kan84]. SEC/DED [Gai88b]. Secondary [LV85, TS86b]. Section [Mic83]. Secure [NK87a, ML88, NK88b]. secure/propagating [NK88b]. Security [KS83a]. seek [CCF89]. seek-time [CCF89]. Segment [Fro83, LP86]. Segmented [Rob85]. selecting [Miy89]. Selection [BS83, Car84, GM87, GMG84, WC84, SS89a]. selector [Kap89]. Self [BCR83, BC85, CL80, DM81, DMY85, FMM84, FM87, Gai85a, Gai85b, GC87, Gol84, HGS83b, HS85, HT88, HML84, Jen83, Kha82, KM84a, LS89a, Mic83, Mil82, NK87b, NS81, Pie87, PK87, RB83d, SSF80, SL83b, TS84a, TAF87, WGT81, BGW89, CKS88, DD88b, DD88c, Gai88a, Gai88c, HKR88, LZM89, MS88a, ML88, NPP88, RSK88, Wan89]. Self-Adjusting [GC87]. Self-Checking [BC85, CL80, FM87, Gai85a, Gai85b, Gol84, HGS83b, HML84, Jen83, Kha82, Mil82, NK87b, RB83d, SL83b, DD88b, DD88c, Gai88a, Gai88c, ML88, NPP88]. Self-Diagnosable [SSF80]. Self-Diagnosing [WGT81]. Self-Diagnosis [HS85, LZM89]. self-dual [Wan89]. self-fault [HKR88]. Self-Implicating [DMY85]. Self-Routing [NS81, HT88]. Self-Scheduling [PK87]. self-stabilize [BGW89]. Self-Stabilizing [BC85, Mil82]. Self-Synchronizing [LS89a]. Self-Test [Mic83, TAF87, CKS88, MS88a, RSK88]. Self-Testing [BCR83, DM81, FMM84, KM84a, Pie87, TS84a]. SELRAM [Kap89]. semantics [Bou89]. semi [MAS85]. semi-Markov [MAS85]. Semiconductor [Che86a, FH86, PS85a, SR80, SR81b, MP89a]. Semijoin [CL84]. Semisystolic [Ers85]. Sensed [Ano80-46, Ano80g]. Sensitive [Hay80, SN81, SR80, MP89a]. Separability [WWS84]. Separable [Smi84]. separate [Gai88c]. Separating [BR80]. Sequence [Mor80, WH80a, Wus81, Reg88]. Sequences [Coh85, DTF80, Hon81, Kak83, Kob83, Kon86, PS85b, Wus82, Kak85, Man88]. sequency [Irs88]. Sequential [AB82, Bha83b, BCDM86, CH82a, Cul80, DS80, Has84, Hay81, IKP86, Kar84b, LSM81, LM83, Mor80, NK87a, Pra83, RS86b, Sal80, SD86a, SD86b, Sav80a, SGA81, Sip82, Sip84, SR82, Tam80, Twi83, VHD82, VH84, Wan81, Wus81, Yam80, DD88c, Len88, Smi89a, Wal88, YKL88]. Sequential/Parallel [Wan81]. Serial [Bat82, Dan84, Gna83, Gna85, Hla86, HRT^{+}84, OJ85, RB80, Uhr82, wu87b, BS88, Dad89, Zho88]. serial-input [Dad89]. Serial-Parallel [Gna85, wu87b]. Serial/Parallel [Dan84]. Server [Sha81]. Servers [Har86, Aky89]. Service [CFM^{+}86, Sha81, TS86b, RN88, WB88]. Set [CP87a, Dah86, GM82b, Jia86, Kee83, KA84, NS82b, RB83d, Til80, Coo89, SAA89]. Sets [AF81, DO85, DO86, FJW85, HRJ86, IKT81, KP80b, RT86, Sto80a, CHRR89, DD88a, KM89a]. Several [WHT84]. Shared [GGK^{+}83, Mar84, NS87, YPD83, AJ89, ALL89, FM88, SV88, WS88]. Shared-Memory [NS87, ALL89, SV88]. Sharing [KC87b, Lea86, SC89, Wah84, WM85, JW89b, MTS89, SW89]. Shift [BR84, Che86b, CF80a, CH85a, Dan83, Dav80a, HSE84, Mor80, WM86, Wus81, WM88a]. Shift-Register [BR84, CF80a, Dan83]. Shifted [PF82]. shifter [MBH89]. Shin [CC89]. Short [BG86a, LHPW85, SB85]. Short-Circuit [SB85]. Short-Packet [LHPW85]. Shortest [Che82a, KLL87, Lak84, SBGS86, WWSW87, BP89]. shortest-path [BP89]. Shuffle [AS87b, DK82, KDJ85, KJ86, Par80, Ste83, TN81, WF81, BN88, HT88, LHC89, SR81a, TS88a]. Shuffle-Exchange [AS87b, KDJ85, KJ86, Ste83, WF81, BN88, LHC89, SR81a]. Shuffle/Exchange [DK82, Par80, TN81, TS88a]. Shuffle/Exchange-Type [Par80]. Shuffling [SS84]. SIFT [MSS82, MFW80]. Sigma [Sez87]. Sign [HL86, SCNS83, Ulm83, Vu85]. Sign/Logarithm [HL86, SCNS83]. Signal [BPM^{+}86, HKSS86, KY82, Lee85a, LM87b, LS89c, Mar87, Shi82b, THH80, ZN80, KT89a, MG88]. Signals [Sto80b].

Signature [Dav86, Has84, Hla86, LM86, Smi80a, Max88, MS88a]. Signatured [SS87]. Signatures [MM83b, MM84]. Signed [RS86b, Par88, Smi89a]. signed-digit [Par88]. Significance [FG82]. SIMD [LL88, NS82a, PK80, SS84, Sez87, SSK^{+}81, WS82a, WR81]. SIMD/MIMD [SSK^{+}81]. similarity [Muk89]. Simple [AG81a, HO81, Lam83, Oru84, RW83, Tsa83, VR83, dM88, Irs88]. Simplex [BB87b]. Simplification [MS86]. Simplified [SL83a, Ste83]. Simulated [GT80]. Simulating [MS88c]. Simulation [AM85, BS87b, CS86b, DJ81, GC87, JLS80, LL81, Ozg86, Ram86a, Ram86b, SWP86, VBH^{+}81, YJ89, BD89, Bos88, Con89, GY89, SV88, Wal88]. Simulator [Abb83, Bry84]. Simulators [MH81]. Single [AF81, Cha80, DL87, EHM80, Hay80, HC87, IKT81, KF82, MBC82, MBR82, MM84, NJM83, Oik87, PS80, PK80, RS83a, RT86, Ram83, Ric84, SD86b, Sav80a, Sav80c, Shi82a, Sri87, VR83, BS89a, BGY89, BGM88, BL89, CJ89, DIN88, KJC89, LD89, NC88a, SAA89, TS88a]. Single-and [PK80]. single-bus [BL89]. Single-Cell [Hay80]. Single-Chip [PS80]. Single-Error [MBR82, BGM88]. single-flow [LD89]. single-key-lock [CJ89]. Single-Layer [Ric84]. Single-Node [EHM80]. Single-Output [SD86b]. single-row [BS89a, DIN88]. Single-Stage [HC87, TS88a]. single-track [KJC89]. Single-Write [Cha80]. Situations [DM83b, MM80]. Size [Bon83b, Bus83, DV87, KA84, MF86, Smi87, BB89a, Pro88, US88]. Sizes [LP81]. Skew [Ers85]. Skewed [HJ87]. Skewing [Kan81, WV87]. Skip [GHM87]. Slash [MK85, Sco89]. Sliced [SH81a, SH81b, VBH^{+}81, Wu87a]. Slicing [LSW87]. Slide [PW81]. Sliding [LC87a]. Sliding-Window [LC87a]. Slotted [ASP87, KM87]. SM3 [BS86]. Small [ISO85, DIY88, Par89]. Smith [MA89].

Sneak [Bra86, CH89]. Society [Ano80-30, Ano80-31, Ano82-28, Ano82-30, Ano82-31, Ano82-32, Ano80-32, Ano80-33, Ano80-34, Ano80-35, Ano80-36, Ano80-37, Ano80-38, Ano80-39, Ano80-40, Ano80-41, Ano80-42, Ano80-43, Ano80-44, Ano81-29, Ano81-30, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano81-37, Ano81-38, Ano82-33, Ano82-34, Ano82-29, Ano82-35, Ano82-36, Ano82-37, Ano82-38, Ano82-39, Ano82-40, Ano82-41, Ano82-42, Ano82-43, Ano82-44, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83r, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Ano83-29, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano87-27, Ano87g]. Society [Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Boo81]. Soft [KC87b, WR84, lYM88]. Software [Ano80c, Ano81w, Ano81x, Cri82, McG80, MH81, VI84, AK88, KW89, Ano80-47, Ano80-48, Ano80-49]. solitons [SKW88]. Solomon [DC87, HRT^{+}84, Liu82, Liu84a, OI87a, STD^{+}85, SR88a]. Solution [Aky87, ATT81, APD83, BR83a, CPG83, ED83, FM84, GLS82, IO84, Lau81, LP86, ML82, Muk87, PTT81, Rub81, TT80, WH80b, AOES88, BBG88, CMB88, DD88a, Mel89, SW88]. Solutions [BG86b, Mey82, OM88]. Solver [Hoc83, Smi80b, Smi81b]. Solvers [DLC87]. Solving [Gaj81, Jia86, MTG^{+}85, WM84, WK80]. Some [Ama83, DLSM81, Fen85, Jes80b, MWM80, MR82, Mur81, Von83, WC83]. Sort [KB85, RFLS86, FM88, NHAT89, SSS89]. Sort-Based [RFLS86]. Sorted [BL80]. Sorter [LCW81, CM88c]. Sorting [AS87a, Ano84e, BBW88, BP84, BP85, BLP84, CF80a, CLW80b, Dem85, DO85, DO86, JO84, JL85, Kru83, KH83, LSSS85, Lei85a, LVW85, Man85, OJ85, RFLS86, RSS85, Rud85, Sie85a, Tho83b, WC83, YTY82, Zak85, Bil89, HP89, Kap89, McC85, Ram88, SP89, SS89b]. Space [Car83, FJ88, FYSK84, LP83, RB83d, TR88]. Space-Efficient [FJ88]. Spanning [Chu85, DJ88, WWSW87]. Spare [DSH89]. Sparse [AJ89, APD83, Mel87, AOES88, DD88a, FS88, Mel89]. Spatial [LP83, VSHM82, MP89c]. Speaking [Aup83]. Special [Ano81t, Ano81r, Ano81s, Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83k, Ano83-31, Ano84e, Ano85a, Ano86c, Ano86d, Ano87c, Ano87z, CP82b, RFLS86, Shi87, Tor87, SW88, YM88a]. Special-Function [RFLS86]. Specialized [VS80]. Specific [NM87]. Specification [Boc82, BR83b, CFM^{+}86, Hof85, KUV85, MSS82, Sta88a, PL88, Sta88b]. Specifications [Ano80-47, Ano80-48, Ano80-49]. Specified [CM87c, Yam80]. Spectra [Muz80]. Spectral [LM86, MM83b, MM84]. Spectrum [HS84, Mor86]. Speed [But81, Cha87b, CHL83, HFPS82, Maj85, OI87a, PKL80, SG80, TY87, UT86, WH80a, ZN80, ZG81, ZG87, NOYK88, PP88, TYY85]. Speed-Efficiency-Complexity [But81]. Speedup [ASHK86, EZL89, PB87]. split [KP88]. Spot [PN85, YTL87]. Square [AR82, Joh80, Maj85, OE82, ZG87]. Square-Rooting [Maj85]. Squares [Sto80b, Smi89b]. Squaring [Kar84b]. Stability [Tas83]. stabilize [BGW89]. Stabilizing [BC85, Mil82]. Stable [Tam80]. stack [KM89b]. Stage [AS82a, HC87, Lee85b, GM89, LH88, TS88a]. Staged [ABG85]. Staging [Sil82, Sil83]. Stamping [Her87]. Standard [YM86a, HTDR88]. Star [CHL83, Kam87, Kum89, WFL82]. Stars [KK87]. startup [Cal88]. State [CIP87, CHM87, Eng81, GK85, Hay81, Hla86, HT86, Kar82, Oik87, SC89, ST86a, WGT81, WR84, VZMBH89]. State-Change [SC89]. State-of-the-Art [Kar82]. Statements [BG84]. States [Tam80]. Static [LM87b, Maz87, RW84, Ski88]. statically [Nic89]. Statistical [BSV83, IBM82]. Statistically [Aup83]. Statistics [De 82a]. Status [Hur84]. Steer [LCW81]. Steering [Bis85]. Stencils [RAP87]. Stiff [BT86]. Stochastic [ACGT84, IDH86, LM88b, Mol82b, BBG88]. Storage [Bur88, CLW80a, GT82, HJ87, KP80a, LV85, QKS85, Sie85a, Sil82, TS86a, TS87, TS86b, WvL85, PBL89]. Store [Gel81]. Store-and-Forward [Gel81]. Stored [Hay81, Twi83]. Stores [BF83]. Strader [Smi89a]. Strategies [CN82, FS80, HI80, KDJ85, MTG^{+}85, RM86, SC87, TS83, LM88a]. Strategy [BW89b, BB87a, FS81, LA87, SGMP85, Val83, McC85, MP89b, YKL88]. Stream [Sie85b, YPD83]. Streams [Hla86, SS87]. String [AN84, Bur84, Bur82]. strings [Muk89]. Strongly [ML88, NK87a, NC88b, JC88, NK88b]. Structural [Kak83, Lu82]. Structure [Dim85, Gau86, JK82, LF85, Met83, TB82, WvL85, WA85b, WL81, Hos89]. Structured [BR83b, De 82b, JLS80, SGA81, WI84]. Structures [AG81a, AS85, BA84a, Bon83a, DMY85, Ell85, GKS84, HKR87, KS82a, LSW87, TS86a, TS87, WF83, WS82b, Wit81, CICR88, LCF89]. Stuck [CD86, Kar83, KP80b, LM86, MM84, RB83d, RR86, YN84, Jha88]. Stuck-At [Kar83, KP80b, MM84, RB83d, YN84, LM86]. Stuck-Fault [CD86]. Stuck-Open [RR86, Jha88]. Students [Ano81-44, Ano81-45, Ano82-61]. Study [BSV83, CS86b, Cra85, Ell85, IBM82, JMKN86, Kam87, KK80b, Kob86, LL81, MWM80, Mel87, MK83b, SWP86, VI84, VS86b, Wah84, CI89, HIT88, LLJ89, STR88, US88]. Subject [MK83b]. sublogarithmic [Gel89]. subsequences [Man88]. Subsystems [Bra83b]. Subtraction [FG82]. Subtree [LF87]. Successively [Sri87]. Sufficient [AMM86]. Suffix [BG86b]. Suitable [Muk86]. suited [SGI89]. Sum [Kon86, MH80, Mos87, SM87, FHG88]. summation [CM88a]. Sums [BSMS81, CM87c, Fam87, Smi89b]. super [CM88a]. Supercomputer [DGT84, GBG89, IK82b, Lin82]. Supercomputers [Men84, PK87, WS84, GZR89]. Supercomputing [Lei85b, Tor87]. Supersystem [IPM82]. Supersystems [AS82a, Ano81t, ABT82, KK80a, Kar82, SG82]. Support [Ano87c, SGB87]. Supporting [DGS80, Ree84, Str82, BD89]. Surfaces [SWK84, Wai88]. Survey [LP84, LP85, RA80, Ric86, Veg84, WP82, MM88]. SVD [LP89]. SW [CM85, RK89]. SW-banyan [RK89]. SW-Banyans [CM85]. Switch [Bry84, KS82a, LC87a, Ram86a, Ram86b, CG88]. Switch-Level [Bry84, Ram86a, Ram86b]. Switchable [ST87]. Switched [CS80b, KDJ85, YBL89]. switches [KJC89]. Switching [ABG85, BK87, BCA80, Car84, CH84, GW81, IM87, JK80, KF80, KK80b, Kob86, KDJ85, LS89d, Muz80, Obe80, OJ80, Par80, Por80, RS86a, Sur81, WI84]. Symbolic [Abb83, HR87, KS82a]. Symmetric [Kit80, MTG83, OO87, WA85b, YM85, AK89, JW89a]. Symmetry [Ata85]. Symposium [Ano80a, Ano80g, Ano82x, Ano82-60, Ano83b]. Synchronization [Cha87b, KO87, Lee80, LAS87, MP87, SR87, VM88, LZM89, NT88, SR88b, SW88]. Synchronized [Kim86]. Synchronizer [BJ83, LMP82]. Synchronizers [KC86, EA89a]. Synchronizing [BR88, FK85, Hon81, LS89a].

Synchronizing/Distinguishing [Hon81]. Synchronous [AB82, Du85, LD87, LM87b, RB82, Sal80, YPD82, PL88]. Syndrome [Ano80z, BSMS81, Mar81, RTJH86, Sav80b, Sav81, TR88]. Syndrome-Testability [Mar81]. Syndrome-Testable [Ano80z, Sav80b]. Syndrome-Testing [Sav81]. Syndrome-Untestable [Sav81]. Syndromes [SM87]. Synthesis [BZV86, Bra83a, CH82a, Cri80, DD81, DC81b, DC82, Eng81, HP82, KW85, KK87, Mol82a, Tha82, Tha84b, VR86, LS88a, WM88b]. Synthesizing [PK88, Che88b]. SYREL [HR87]. Systaltic [OS87]. System [AJ82, Ano81-27, AGH^{+}82, ACD82, BH81, BS86, BS82, BA82, BBB^{+}82, Bux83, CS86b, CFH81, CLP81, Cli80b, CMM87, DM83a, Dah86, De 82a, EHS80, GM82a, GMK85, GLS82, GMG84, Hag86, HN84, HC86, IC80, KH80, Kan81, Kan84, KK82, Kes84, LV82, Liu84b, Lun87, Man80, Mar84, MBCG83, MD86, MS82, MSS82, Mey84, MP83c, ML82, Mor86, NN86, NWLP82, OJ80, Par86, PL83b, PS87, RJ80, Ros85a, SYJ89, SBK85, SS86, SSB86, SR87, SOH^{+}81, SMN82, STR87, SSK^{+}81, SL83b, SAA87, Sta81, Sta88a, ST87, Tas83, TGJR88, TS86b, VSHM82, VI84, WOH84, WV80, WA85a, WFL82, YYF85, BBG88, BF89, BB89b, CJ89, FL88, HTT89, HN88, KM89a, KH88, Kum89, LW88, LW89, Mil88, SYK89, SL88b, Sta89, Sta88b, YJ89, Kob83, Lan87, PMSB88, Zwa85]. System/370 [Kob83, PMSB88, Lan87]. Systematic [Bla88, BL85, NGP86, OT89]. Systematically [Ree80]. Systems [ASKL81, Ano83q, Ano86d, Ano87c, Ano87z, AL82, BA86, Bat82, BW89b, BK87, BR83b, BR80, BS83, BPM^{+}86, BD83, Bry84, Can83, CS80a, CA80, CL84, Che86a, CA83, CA86, CH82b, CH83, CLH84, CH85b, CL87, CLW80b, CH81, CN82, Cve87, DM84, DMY85, DG86, DV83, Dun85, ED83, FH80, FM84, Gaj81, Gau86, GP86, Gav87, GT83, GBG89, GM82b, GJ80, GT87, HS81, HS85, HZ81, HKR84, HC82, IK82a, IK82b, IO84, Iye84, IDH86, Jen81, KW81, KF82, Kar81, KK84b, KK80b, KC87a, KO87, KS80b, KM81, KP87, KK85, KSB87, KC87b, LF85, LYI87, LJ87, Lo88, MLT82, MK83a, MS81, MG82, MB80, MK85, Mey80, MK83b, ML82, NK87b, Nil84, OK83, OA83, PH86, PS83a, PR82, Ram83, RM86, RAP87, RB89b, RW84, SSVV81]. Systems [SL83a, ST85, Shi87, SC89, SSF80, SB83, Sny81, SH81a, Sta84, SBMM87, Ten83, TB86, Tow86, TCH^{+}86, TTB85, TS84b, UT86, VH84, VSV81, WM85, Wen85, Wor81, WL81, YM86b, YML86, YPD83, YPD82, Agr88, AOES88, Bou89, BGW89, CCF89, CMB88, CA89, DSH89, DT89, EZL89, Fuk88, Gai88a, GC89, GM88a, GS89, HMC89, HKR88, KK88, KS89, LM88a, LPI88, Mel89, MP89b, MAS85, MDG89, PP88, PBL89, PKP89, RKH88, RB89a, RN88, SW89, Sin89, SRC85, Sul88, US88, WJ85, YZ88, lYM88, Ano82j]. Systolic [AN84, AS85, BK84, BPM^{+}86, CM88a, CICR88, Cha84, Gue86, HQR89, IKP86, KY82, KLL87, LSSS85, LL85, LW85, MT87, MF86, Mor86, O’L87, Sav84, SSS89, YRT84, ZH85, wu87b, CM88c, Cos88, EA89b, FS88, KR89a, KT89b, LLJ89, LJL89, cLW88, LP88, Mel89, PL88, ST88a, SR88a, SO89, SS89c, Zho88].

T. [Smi89a]. t1 [YML86]. t1-Diagnosable [YML86]. t1/t1 [YML86]. t1/t1-Diagnosable [YML86]. Table [AM80, Ano80-50, Bus83, GMG84, Niz84b, Niz84a, Par87]. Tables [Cle84]. Tag [KB85, KPR88]. Tagged [Lop84, BS89c, GW89]. tagged-token [BS89c]. tags [HT88]. Taken [Mor80, Wus81]. Tandem [CFH81]. Target [AMM86]. Task [BSV83, CL87, Kob86, LF80, Lo88, MLT82, Man84, PS87, ST85, TB86, TS84b, VW84, VM87b, SRC85, Sta89, US88]. Task/Processor [Man84]. Tasks [BDW86, HT82, KS86b, RSZ89, SLL87, CCWZ88, Sto89]. TBED [DC87]. TDA [CA89]. Technique [BT86, DLM86, Fis81, McC84, RW82, SGMP85, Van86, VS80, WM86, Bos88, HT88, RSK88]. Techniques [DHS85, KW85, NH85, Owe83, PKL80, TCH^{+}86, WW83, AJ88, CMB88, KSW89, KT89a, SS89a]. Technology [Ano83j, DD81, DC81a, Lin82, Smi81a, SG82, NC88a]. Telecommunications [Ama83]. Telescope [EHS80]. Temporal [Boc82, BCDM86, SL89]. Temporary [QKS85]. Terminals [GCW83, LS88b]. Ternary [BS87b, HS86, MA89, MSV80, Muk86, YM88a]. Test [AR86, AB86, AC83, AM85, AMM86, AF81, Ano85d, BF80, CP87a, CM87a, CKS88, DB87, Fuj81, FS83, HO81, IKT81, JA85, KT85, KP80b, KK85, KA84, Kri87, Lan87, LS80, LM82b, LSM81, Mar85, MBN81, McC84, Mic83, PC81, RT86, RS87, SK85, Sav83b, SB84, SS86, Sto80a, SR80, SR81b, TW83, TC84, TA80, TAF87, VR83, WM86, dJvdG88, Abo88, CHRR89, Che88a, GNH88, MS88a, PS88, RSK88]. Test-Experiments [LSM81]. Testability [BNM86, Fuj84, GCV80, HS86, Mar81, PC81, RB83a, Sav83a, SDB84, SS85b, SB80, VS86a, WP82, MA89, SM88]. Testable [AG81a, Ano80z, Bha83b, Kha84, Kor86, Pag80, PR81a, RR86, RH87, Ros83, Sal80, SKF83, SD86b, Sav80b, SF84, SH81a, BS89b, CBAP89, HR88, JP88]. Testing [AC83, AC84, AF81, Agr81, AL85, Ano81z, BSMS81, BCR83, BA84b, CM87a, Cli80a, CS84, DM81, DHR80, DTF80, Dav80a, FMM84, FM87, Has84, Hay80, HS84, KM84a, KS86a, LS89d, LM86, McC84, MH81, Pie87, Rob85, Sav80c, Sav81, SR86, SM87, SGMP85, SH81b, Sus83, TS84a, WCM87, WM86, Bec88, DFC89, MP89a, MP89b, SL88a, SS88b, Sri89, YKL88]. Tests [BM86b, FK81, Goe81, Kar83, YN84, KK88]. th [CE87]. th-Order [CE87]. Their [Agr80, AK87, CH82a, FJW85, GCV80, RAP87, Zhi84, Mar86, Pol88, RN88, SS89c, SL89]. Theorem [KW85, ST86a, WW83, KSW89, VZMBH89, Vu85]. Theorem-Proving [KW85, KSW89]. Theoretic [Agr81, JM87, Jul80, Kod81, AK89]. Theoretical [Agr83]. Theory [BR82, DS80, Hol82b, HT86, LB88, PP88, SS80, Sie80, SL83b, SAA87, Sta85, VHD82, PS88]. Theory-Based [SS80]. think [Fis88]. thinning [WZ89]. thread [ALL89]. Three [Hla86, PL84, SS86, SD87, TY87, WF80a, CD88]. Three-Layer [PL84, CD88]. Three-Rail [TY87]. Three-State [Hla86]. Three-Valued [SS86, WF80a]. Threshold [MTG83, YF88, ZMC89]. Throughput [Bra83b, Dub88, LP81, Wan82]. Throughput-Driven [Bra83b]. ti [YM86b]. ti-Diagnosable [YM86b]. Tight [Lei85a]. Tightly [ASHK86, RW82, BL89]. tightly-coupled [BL89]. Time [AM87, AJ82, Ano86d, Ano87z, ABT82, ACD82, BPV83, BP85, CS85, Car83, CL87, CMM87, DG86, GMK85, GK83, Har86, Her87, HL82, Kam80, KO87, KSB87, KC87b, LP81, LD87, MLB87, MAS84, PS87, Pre83, PP87, RB83d, SGB87, Shi87, SLL87, SC89, SMN82, VM87b, WHT84, WBA83, WFL82, Wu87a, ZRS87, Bec88, CCF89, HTT89, KW89, Nic89, SRC85, Sta89, WS88]. Time-Bandwidth [AM87]. time-shared [WS88]. Time-Space [Car83]. Time-Stamping [Her87]. Times [ACGT84, KSW88, SL83a]. Timestamp [Li87]. Timestamp-Ordering [Li87]. Timing [BI88, JM87, VM87b, VM87a]. Titles [Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano82-32]. TMR [Gai88a]. Today [Gar80]. Token [BGW89, GW89, SM82, BS89c, CL88, RN88, RS88b]. token-passing [RN88]. Tolerance [AK87, BA86, CS86c, Cri82, HKR87, HA84, KSB85, KK86, LF87, MI89, RAE84, SL80, SH84, Agr88, AK88, AL88, Coa88, Cos88, Esf89, GM88a, KWFT88, KSW89, Sin88b].

Tolerant [AS82a, AP86, Ano81s, Ano83k, Ano85a, ACD82, BK87, BC85, CN82, DM83a, Dao81, EH85, FJ88, GT83, GBG89, GS86, HA86, Haw85, Hay84, Hon82, HC87, Iye84, IDH86, KH80, Kes84, KB84, Kor86, LS84, MS81, MA82, MSS82, Mil82, Por82, Pra80, PR82, Pra85a, Pra85b, Pra86a, RV86, Red87a, Red87b, Ren84, Ros83, Ros85b, SBK85, SCP^{+}81, SSB87, SL83a, Sti80, Tan84, WA80, YH86, AA88, Ban88, BD89, BW89a, CCWZ88, CM88b, CM88c, CC89, DT89, Gai88a, GS89, Hos89, JA88, KR89a, KPR88, KR89b, KT89b, KJC89, LLJ89, LJL89, LW89, LZM89, LP88, Mey88, Ski88, Tys88, TYZ88, VR89a, VR89b, VM88]. Tolerating [SD80]. Tomography [SG80]. Tools [HT86, BM89]. Topological [SS88a]. Topologies [Hwa87, RGA85, BP89]. Topology [GS81, Bil89, Kum89]. Torus [Von83, LM89]. Totally [Gai85a, Gai85b, Gai88c, Gol84, GH80a, HGS83b, HML84, Kha82, LM83, NK87b, RB83d, SL83b, Gai88a, ML88, NPP88]. Tours [Cul80]. Trace [Fis81, GB83, CNO^{+}88]. Tracing [Sto80a]. Track [SMN82, KJC89]. Tradeoff [KK80b]. Tradeoffs [But81, Car83, DIY88, KSB87, Lin82, MBH89]. Traffic [BS87a, NM87, WA85b]. Traffic-Specific [NM87]. TRAM [JP88]. Transaction [LYI87, DIY88]. Transactions [Ano80b, Ano81a, Ano81-39, Ano81-40, Ano81-42, Ano81-43, Ano82a, Ano82-47, Ano82-48, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano83a, Ano83-31, Ano84a, Ano86b, Ano87a, Ano87o, BR88]. transduction [MKLC89]. Transfer [LHPW85, PS83b, PS85b, Sta88a, Sta88b]. Transform [CF80b, GS87, HFPS82, Hou87b, KR82, Kit80, PR81b, RTB81, TCH^{+}86, TSM88, Bon83a, Hou87a, RJ82, TLR83, WG80]. Transformations [ASKL81]. Transforms [Jes80a, Jul80, Kod81, Mah80, Tho83a, TRYS83, DK89]. Transient [BT86, CMS82]. Transient-Upset [DM83a]. Transistor [Bra86, CH89]. Transition [DTF80, WR84, Reg88]. Transitive [KLL87]. Translation [AO80]. Transmission [SR88b, SRO84]. Transparency [RC89]. Transpose [O’L87]. Transposition [Twi83]. Treating [Muk86]. treatment [KW89]. Tree [AS85, BW81a, BH84, CV83, CK87, CLW80a, GKS84, GM87, HZ81, HKR87, LF87, RAE84, SGA81, WA85b, Abo88, Bur88, CICR88, HA88, TYY85, YS89]. Tree-Based [CK87]. Trees [BBP88, CT84a, Cha87a, Chu85, DS83, Ell80, Gor87, Got81, HA86, KK84a, KK87, MI85, NMB83, Pro81, Vai84, Vai86, WWSW87, CI88, Lei85b, LS89c, Mey88, Miy89, Vai89]. Trellis [RTJH86]. Trends [Ano80x, Ano81u, Ano81v, Ano82u, Ano82v, Ano82w]. Triangles [MD86]. Triangular [ED83, ML82, RB82]. triangularizations [LP88]. tridiagonal [USM89]. Triggered [Uug81]. Triggering [Cha83a]. Triple [KH80, Van86]. Triple-Modular [KH80]. Trivalent [LS82b]. True [BVH83]. Truth [GMG84]. Truth-Table [GMG84]. TSC [BL84, Eti80, GH83b, Gai88b, PNH88]. Tukey [NS87]. Turn [Wan81]. Tutorial [RA80, Ano80h, Ano80i, Ano80j, Ano80k, Ano81b, Ano81d, Ano81c, Ano82e, Ano82g, Ano82h, Ano82f]. Two [BIO82, BF83, BVH83, Bon83a, BK80, ED88, IC80, Kes84, Kha82, KA87, Kon86, MWM80, NJM83, ST86a, TLR83, Val83, Blo88, CCF89, Dad89, DIN88, KR89a, LLJ89, LP89, Muk89, SS89b, TSM88, WLC89, WCS89, YJ89]. Two-Dimensional [KA87, NJM83, TLR83, DIN88, KR89a, SS89b, TSM88]. two-head [CCF89]. Two-Level [BIO82, BF83, IC80, WCS89]. Two-Phase [Val83, ED88]. Two-Processor [MWM80]. Two-Rail [Kha82]. two-writer [Blo88]. Type [Her87, Par80, RK86, SC87]. Types [Sta81, PL88, Wei88].

UED [LB88]. Ultra [SG80]. Ultra-High-Speed [SG80]. Ultracomputer [GGK^{+}83]. Ultrahigh [GT83]. Unate [MM83b, PS88]. Unavoidability [KC87a]. Unbuffered [KJ86]. Uncertainty [Hay86b, HRJ86]. Underlying [Sie80]. Undetectability [KP80b]. Undetectable [Oik87, RB83a, YN84]. Unicomputers [Hoc83]. Unidirectional [BP82, BR82, BR84, BL85, Bos86, Don84, NGP86, Bla88, BV89, LB88, THL88]. Unification [Cra85, VS86b]. Unified [Fla82, RS87, Til80, WA80]. Uniform [ACGT84, BB87a, CF80a, PK80, KW89]. Unilateral [AC84]. uniprocessor [RTY^{+}88]. Uniquely [BK80]. Unit [AKT86, CHH83, DC81a, FTT^{+}80, JH80, KM84b, LMO84, RFLS86, ZG87, KM83a, Noe89]. Units [GH83a, Hol82a, IK85, LZLH85, ZG81]. Universal [But81, CH82a, FK81, Fuj84, Kar83, Lei85b, SGA81]. Universal-Logic-Module [CH82a]. Universality [WF81, TS88a]. Unknown [JL85]. Unordered [Smi84]. unpredictable [Nic89]. unreliable [YM88b]. Unrestricted [RB83d]. Unsorted [SA85]. unsymmetric [DD88a]. Untestable [Sav81]. Unusual [Wei82]. Updating [CH85b, GC80]. Upper [DC82, GM88b]. Upset [DM83a]. Upsets [GM82b]. Use [FMM84, FM87, HS84, KYSY80, MH81, Loz83]. Used [PK80, YL81]. user [KH88]. Uses [GF87]. Using [AR86, Ano80d, Ano80e, Ano80f, AGL^{+}80, BYH87, BVH83, BR83a, BCDM86, Cha87a, CS87, Cle84, Cri80, DB87, DD81, DC81a, GK83, GNK86, HJ87, HNS84, HRT^{+}84, HRJ86, IKP86, KW85, KH80, Kha82, KS86a, LS84, LF87, MD86, MP83c, Mol82b, NJM83, OI87a, Pal86, PA81, Pra83, RB83d, SYJ89, SD86b, SS87, ST85, SOA85, STR87, Sie85b, TY87, TC84, TPSS85, Tha84b, TRYS83, TCH^{+}86, Uhr82, VBH^{+}81, Van86, WBA83, BW89a, BI88, CCWZ88, Che88a, CC89, HCMP89, HMC89, HTDR88, HT88, Kak85, KSW89, KJC89, LCF89, MM88, RB88, SYK89, SR88a, SK89, WM88a, Wan89, WM88b]. Utilization [BCA80, RW81a, FL88]. Utilizing [PKP89].

V [Zwa85]. V-System [Zwa85]. Validation [Hol82b, RW84, RW82]. Validity [KP80b]. Value [Smi85, CdL89]. Valued [Ano80a, Ano83b, Cha87a, Cur80b, DDG80, Hay86b, Hur84, KW81, KF80, KT81, Kuo87, Mor86, MR86, Sas81, Sas85, SS86, SGA81, WW83, WF80a, BB89a, Sas89, YF88]. Values [BVH83, QKS85, FW88]. Variability [Iye84]. Variable [Bon83b, Dan83, MAS84, Sas84, Miy89]. Variable-Length [Dan83]. Variables [LM87a, YM85]. Variants [Dor88]. Variation [CS85]. varying [NS88]. VAX [AP85]. VAX-11 [AP85]. VAX-11/780 [AP85]. Vector [Ary85, Bai87, HJ87, HN80, NH85, Cal88, GZR89, HX88, PMSB88]. Vector-Reduction [NH85]. Vectorization [Mos87]. Vectors [TW83, WV87, Irs88]. Verification [BCDM86, McC84, MSS82, PS83b, PS85b, RE80, BC88, KSW89]. Verifying [Sus83]. Versatile [Len85, SP89]. Version [Ste83]. Versus [Arm80, GCV80, DFC89, EZL89]. Vertical [LP86, PB87, Sta81]. Very [CT84b, NF84, PR81b, QI85, SL80]. Via [BS89a, LM85, MG86, Pap83, DIN88, NMN89, TYZ88]. video [TSM88]. Videotex [WA85a]. View [RS87, Smi81a]. Virtual [BS83, BH84, UT87, Bur88, LCF89, RTY^{+}88]. Visibility [AT81, LP86]. Vision [HRJ86]. VLIW [CNO^{+}88]. VLSI [MA89, Ram86b, AB86, AM87, AK85, BSMS81, BCR83, BPV83, BB87b, BB89b, BP84, BP85, Bon83a, Bon83b, BLP84, BK84, BPM^{+}86, CGMP87, CBAP89, Che88b, Cli80a, CKS88, DC81a, Fen89, FK85, Fra81, FWT82, FM89, GC87, GS87, GKS84, Gor87, Hen84, HZ81, HMC89, HRT^{+}84, HTDR88, HS86, HC82, JO84, KB84, KP87, Kri84, KH86, Liu82, LY83, Liu84a, LP86, MA82, Mol82a, NMB83, ORS82, Pre83, PTC86, PL88, Ram86a, RTJH86, Ros85b, SP89, SS89b, SS85a, Sei84, STD^{+}85, SR88a, SF84, Sin88b, SBGS86, Soh89, SA85, TYY85, Tay82, Tho83a, Tho83b, VR89b, Vui83, WHT84, WTS^{+}85, WCS89, WSR84, WD84, YTY82, YS89, YKL88]. VLSI-Based [WSR84]. VLSI/WSI [WCS89]. Voice [RG85]. Voice-Packet [RG85]. Vol [Ano80b, Ano81a, Ano82a, Ano83a, Ano84a, Ano86b, Ano87a]. Voronoi [CE87, Lee82]. Voting [BGM87]. VSLI [WF83].

Wafer [LL85]. Wafer-Scale [LL85]. Waiting [KSW88]. Walsh [HS84, Sus83]. Warp [AAG^{+}87, WBA83]. Watch [CH85a]. Watchdog [Lu82, MM88]. Wave [VW84]. Wavefront [KAGER82]. Waveguide [MP83c]. Way [CIP87, GHM87, NHAT89, San89, SSS89]. Weather [DGT84]. Week [Ano80h, Ano80i, Ano80j, Ano80k, Ano81b, Ano81d, Ano81c, Ano82e, Ano82g, Ano82h, Ano82f]. Weight [Kan84, TW83]. Weighted [BSMS81]. West [Ano81d, Ano82g, Ano82h]. Where [TKL86]. Which [Cli80a]. wide [Mey88]. Window [LC87a, TV82]. Wired [KM86]. Wiring [GCW83, LSW87, NKY^{+}80, Sch89]. Wise [Har86]. Without [AS87a, MM80, Van86, Agr88]. Word [MAV84, NF84, YJ89]. Words [Mor80, Wus81]. Wordwidth [RB83c]. Work [Ano83c]. Workload [CS85, IR86, MW88]. Workloads [Ree84]. workstation [TSS88]. Worst [BW80, GM88b]. Write [Cha80, SS82]. Writer [CH85b, Blo88]. WSI [HS88, WCS89, YS89]. x [HSE84, CS86b, CH85a]. X-MP [CS86b]. X.21 [RE80].

Y. [CC89]. Y.-H. [CC89]. Yield [KB84, KP87, MA82, HS88, HA88, KR89a, Soh89].

ZMOB [KWR82].

References

Abu-Amara:1988:FTD

[AA88]  H. H. Abu-Amara.  Fault-

tolerant distributed algorithm for election in complete networks. IEEE Transactions on Computers, 37(4):449–453, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2189.

Annaratone:1987:WCA

[AAG^{+}87] Marco Annaratone, Emmanuel Arnould, Thomas Gross, H. T. Kung, Monica Lam, Onat Menzilcioglu, and Jon A. Webb. The Warp computer: Architecture, implementation, and performance. IEEE Transactions on Computers, C-36 (12):1523–1538, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009502.

Abramovici:1980:MFD

[AB80] M. Abramovici and M. A.

Breuer. Multiple fault diagnosis in combinational circuits based on an effect-cause analysis. IEEE Transactions on Computers, C-29(6):451–460, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675604.

Abramovici:1982:FDS

[AB82] M. Abramovici and M. A. Breuer. Fault diagnosis in synchronous sequential circuits based on an effect–cause analysis. IEEE Transactions on Computers, C-31(12):1165–1172, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675940.

Abadir:1986:TSV

[AB86] M. S. Abadir and M. A. Breuer. Test schedules for VLSI circuits having built-in test hardware. IEEE Transactions on Computers, C-35(4):361–367, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676771.

Abbott:1983:SSM

[Abb83] C. Abbott. A symbolic simulator for microprogram development. IEEE Transactions on Computers, C-32(8):770–774, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676320.

Aberth:1984:PSC

[Abe84] Oliver Aberth. Precise scientific computation with a microprocessor. IEEE Transactions on Computers, C-33(8):685–690, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009354.

Arango:1985:SCS

[ABG85] M. Arango, H. Badr, and D. Gelernter. Staged circuit switching. IEEE Transactions on Computers, C-34(2):174–180, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676555.

Afshari:1983:LBB

[ABK83] P. V. Afshari, S. C. Brueli, and R. Y. Kain. On the load balancing bus accessing scheme. IEEE Transactions on Computers, C-32(8):766–770, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676319.

Aborhey:1988:BDT

[Abo88] S. Aborhey. Binary decision tree test functions. IEEE Transactions on Computers, 37(11):1461–1465, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8719.

Abramovici:1982:HPO

[Abr82]  M. Abramovici.  A hierar-

chical, path-oriented approach to fault diagnosis in modular combinational circuits. IEEE Transactions on Computers, C-31(7):672–677, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676064.

Arnold:1982:MAR

[ABT82] R. G. Arnold, R. O. Berg, and J. W. Thomas. A modular approach to real-time supersystems. IEEE Transactions on Computers, C-31(5):385–398, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676016.

Aboulhamid:1983:CTG

[AC83] M. E. Aboulhamid and E. Cerny. A class of test generators for built-in testing. IEEE Transactions on Computers, C-32 (10):957–959, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676141.

Aboulhamid:1984:BTO

[AC84] E. M. Aboulhamid and E. Cerny. Built-in testing of one-dimensional unilateral iterative arrays. IEEE Transactions on Computers, C-33(6):560–564, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676481.

Ayache:1982:RFT

[ACD82] J.-M. Ayache, J.-P. Courtiat, and M. Diaz. REBUS, A fault-tolerant distributed system for industrial real-time control. IEEE Transactions on Computers, C-31(7):637–647, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676061.

Ahuja:1988:MLH

[ACGK88] S. Ahuja, N. J. Carriero, D. H. Gelernter, and V. Krishnaswamy. Matching language and hardware for parallel computation in the Linda Machine. IEEE Transactions on Computers, 37(8):921–929, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2244.

Agrawala:1984:SOA

[ACGT84] A. K. Agrawala, E. G. Coffman, Jr., M. R. Garey, and S. K. Tripathi. A stochastic optimization algorithm minimizing expected flow times on uniform processors. IEEE Transactions on Computers, C-33(4):351–356, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676440.

Applegate:1985:DO

[ADG^{+}85] J. H. Applegate, M. R. Douglas, Y. Gursel, P. Hunter, C. L. Seitz, and G. J. Sussman. A digital orrery. IEEE Transactions on Computers, C-34(9):822–831, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676638.

Adi:1984:FBE

[Adi84] Wael Adi. Fast burst errorcorrection scheme with fire code. IEEE Transactions on Computers, C-33(7):613–618, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009334.

Agarwal:1981:MFT

[AF81] V. K. Agarwal and A. S. F. Fung. Multiple fault testing of large circuits by single fault test sets. IEEE Transactions on Computers, C-30(11):855–865, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675716.

Abraham:1981:DTS

[AG81a] J. A. Abraham and D. D.

Gajski. Design of testable structures defined by simple loops. IEEE Transactions on Computers, C-30(11):875–884, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675718.

Armstrong:1981:FDB

[AG81b] J. R. Armstrong and F. G.

Gray. Fault diagnosis in a Boolean n cube array of microprocessors. IEEE Transactions on Computers, C-30(8):587–590, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675844.

Arden:1982:MCM

[AG82] B. W. Arden and R. Ginosar. MP/C: a multiprocessor/computer architecture. IEEE Transactions on Computers, C-31(5):455–473, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676022.

Agarwal:1980:MFD

[Aga80] V. K. Agarwal. Multiple fault detection in programmable logic arrays. IEEE Transactions on Computers, C-29(6):518–522, June 1980. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675612.

Aggarwal:1986:OBF

[Agg86] A. Aggarwal. Optimal bounds for finding maximum on array of processors with k global buses. IEEE Transactions on Computers, C-35(1):62–64, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676658.

Antonsson:1982:PSA

[AGH^{+}82] D. Antonsson, B. Gudmundsson, T. Hedblom, B. Kruse, A. Linge, P. Lord, and T. Ohlsson. PICAP — A system approach to image processing. IEEE Transactions on Computers, C-31(10):997–1000, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675908.

Arulpragasam:1980:MMU

[AGL^{+}80] J. A. Arulpragasam, R. A. Giggi, R. F. Lary, D. T. Sullivan, and Chin-Chang Wu. Modular minicomputers using microprocessors. IEEE Transactions on Computers, C-29 (2):149–160, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675539.

Agrawal:1980:NBA

[Agr80] D. P. Agrawal. On negabinarybinary arithmetic relationships and their hardware reciprocity. IEEE Transactions on Computers, C-29(11):1032–1035, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675502.

Agrawal:1981:ITA

[Agr81] V. D. Agrawal. An information theoretic approach to digital fault testing. IEEE Transactions on Computers, C-30(8):582–587, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675843.

Agrawal:1983:GTA

[Agr83] D. P. Agrawal. Graph theoretical analysis and design of multistage interconnection networks. IEEE Transactions on Computers, C-32(7):637–648, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676295.

Agrawal:1988:FTM

[Agr88]  P. Agrawal.  Fault toler-

ance in multiprocessor systems without dedicated redundancy.

IEEE Transactions on Computers, 37(3):358–362, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2174.

Atallah:1986:ORP

[AH86] M. J. Atallah and S. E. Hambrusch. Optimal rotation problems in channel routing. IEEE Transactions on Computers, C-35(9):843–847, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676846.

Agrawal:1982:PPS

[AJ82] D. P. Agrawal and R. Jain. A pipelined pseudoparallel system architecture for real-time dynamic scene analysis. IEEE Transactions on Computers, C-31(10):952–962, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675904.

Agrawal:1988:PTL

[AJ88] R. Agrawal and H. V. Jagadish. Partitioning techniques for large-grained parallelism. IEEE Transactions on Computers, 37(12):1627–1634, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9739.

Alaghband:1989:SGE

[AJ89] G. Alaghband and H. F. Jordan. Sparse Gaussian elimination with controlled fill-in on a shared memory multiprocessor. IEEE Transactions on Computers, 38(11):1539–1557, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42123.

Atallah:1985:GDM

[AK85] M. J. Atallah and S. Rao Kosaraju. A generalized dictionary machine for VLSI. IEEE Transactions on Computers, C-34(2):151–155, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676551.

Akers:1987:GGT

[AK87] S. B. Akers and B. Krishnamurthy. On group graphs and their fault tolerance. IEEE Transactions on Computers, C-36(7):885–888, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676983.

Ammann:1988:DDA

[AK88] P. E. Ammann and J. C.

Knight. Data diversity: an approach to software fault tolerance. IEEE Transactions on Computers, 37(4):418–425, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2185.

Akers:1989:GTM

[AK89] S. B. Akers and B. Krishnamurthy. A group-theoretic model for symmetric interconnection networks. IEEE Transactions on Computers, 38(4):555–566, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21148.

Agrawal:1988:ADN

[AKS88] D. P. Agrawal, S.-C. Kim, and N. K. Swain. Analysis and design of nonequivalent multistage interconnection networks. IEEE Transactions on Computers, 37(2):232–237, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2154.

Acosta:1986:IIA

[AKT86] R. D. Acosta, J. Kjelstrup, and H. C. Torng. An instruction issuing approach to enhancing performance in multiple functional unit processors. IEEE Transactions on Computers, C-35(9):815–828, September 1986. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676841.

Akyildiz:1987:EPF

[Aky87] I. F. Akyildiz. Exact product form solution for queueing networks with blocking. IEEE Transactions on Computers, C-36(1):122–125, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009458.

Akyildiz:1989:PFA

[Aky89] I. F. Akyildiz. Product form approximations for queueing networks with multiple servers and blocking. IEEE Transactions on Computers, 38(1):99–114, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8732.

Arden:1981:ACR

[AL81] B. W. Arden and Hikyu Lee.

Analysis of chordal ring network. IEEE Transactions on Computers, C-30(4):291–295, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675777.

Arden:1982:RNM

[AL82] B. W. Arden and Hikyu Lee.

A regular network for multicomputer systems. IEEE Transactions on Computers, C-31(1):60–69, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675886.

Agrawal:1985:DAT

[AL85] D. P. Agrawal and Ja-Song Leu. Dynamic accessibility testing and path length optimization of multistage interconnection networks. IEEE Transactions on Computers, C-34(3):255–266, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676568.

Anfinson:1988:LAM

[AL88] C. J. Anfinson and F. T. Luk. A linear algebraic model of algorithm-based fault tolerance. IEEE Transactions on Computers, 37(12):1599–1604, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9736.

Anderson:1989:PIT

[ALL89] T. E. Anderson, E. D. Lazowska, and H. M. Levy. The performance implications of thread management alternatives for shared-memory multiprocessors. IEEE Transactions on Computers, 38(12):1631–1644, December 1989. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40843.

Agarwal:1980:GFC

[AM80] V. K. Agarwal and G. M. Masson. Generic fault characterizations for table look-up coverage bounding. IEEE Transactions on Computers, C-29(4):288–299, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675567.

Abramovici:1985:PAF

[AM85] M. Abramovici and P. R.

Menon. A practical approach to fault simulation and test generation for bridging faults. IEEE Transactions on Computers, C-34(7):658–663, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676604.

Abu-Mostafa:1987:TBP

[AM87] Y. S. Abu-Mostafa. On the time-bandwidth proof in VLSI complexity. IEEE Transactions on Computers, C-36 (2):239–240, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676888.

Amar:1983:CST

[Ama83] D. Amar. On the connectivity of some telecommunications networks. IEEE Transactions on Computers, C-32(5):512–519, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676267.

Amer:1982:MCN

[Ame82]  P. D. Amer.  A measure-

ment center for the NBS local area computer network. IEEE Transactions on Computers, C-31(8):723–729, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676078.

Abramovici:1986:CFS

[AMM86] M. Abramovici, P. R. Menon, and D. T. Miller. Checkpoint faults are not sufficient target faults for test generation. IEEE Transactions on Computers, C-35(8):769–771, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676831.

Apostolico:1984:SAS

[AN84] A. Apostolico and A. Negro. Systolic algorithms for string manipulations. IEEE Transactions on Computers, C-33(4):361–364, April 1984. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676442.

Anonymous:1980:ISM

[Ano80a] Anonymous. The 10th International Symposium on Multiple-Valued Logic. IEEE Transactions on Computers, C-29(4): 333, April 1980. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675580.

Anonymous:1980:IIT

[Ano80b] Anonymous. 1980 index IEEE Transactions on Computers vol. C-29. IEEE Transactions on Computers, C-29(12):1163, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675520.

Anonymous:1980:ICSg

[Ano80c] Anonymous. 249 4th International Conference on Software Engineering. IEEE Transactions on Computers, C-29(6):549–c, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675621.

Anonymous:1980:PCFa

[Ano80d] Anonymous. 254: Proceedings:

COMPCON 79 Fall “Using Microprocessors — Extending our Reach”. IEEE Transactions on Computers, C-29(4):333–b, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675577.

Anonymous:1980:PCFb

[Ano80e] Anonymous. 254: Proceedings: COMPCON 79 Fall “Using Microprocessors — Extending our Reach”. IEEE Transactions on Computers, C-29(6):549–a, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675619.

Anonymous:1980:PCFc

[Ano80f] Anonymous. 254: Proceedings: COMPCON 79 Fall “Using Microprocessors — Extending our Reach”. IEEE Transactions on Computers, C-29(7):678–a, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675643.

Anonymous:1980:PSM

[Ano80g] Anonymous. 256 Proceedings: Symposium on Machine Processing of Remotely Sensed Data. IEEE Transactions on Computers, C-29(4):333–c, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675578.

Anonymous:1980:AATa

[Ano80h] Anonymous. Advance announcement: Tutorial Week 80. IEEE Transactions on Computers, C-29(8):759, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675671.

Anonymous:1980:AATb

[Ano80i]  Anonymous.  Advance an-

nouncement: Tutorial Week 80. IEEE Transactions on Computers, C-29(9):844, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675687.

Anonymous:1980:AATc

[Ano80j]  Anonymous.  Advance an-

nouncement: Tutorial Week 80. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675487.

Anonymous:1980:AATd

[Ano80k] Anonymous. Advance announcement: Tutorial Week 80. IEEE Transactions on Computers, C-29(11):1035, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675505.

Anonymous:1980:ACP

[Ano80l] Anonymous. Announcement & call for papers. IEEE Transactions on Computers, C-29(12): 1163, December 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPa

[Ano80m] Anonymous. Call for papers. IEEE Transactions on Computers, C-29(1):61, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPb

[Ano80n] Anonymous. Call for papers. IEEE Transactions on Computers, C-29(1):61, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPc

[Ano80o] Anonymous. Call for papers. IEEE Transactions on Computers, C-29(1):61, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPd

[Ano80p] Anonymous. Call for papers. IEEE Transactions on Computers, C-29(2):211, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPe

[Ano80q] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(2):211, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPf

[Ano80r] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(3):267, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPg

[Ano80s] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(3):267, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPh

[Ano80t] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(5):416, May 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPi

[Ano80u] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(6):549, June 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPj

[Ano80v] Anonymous. Call for papers.

IEEE Transactions on Computers, C-29(8):759, August 1980.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPk

[Ano80w] Anonymous. Call for papers. IEEE Transactions on Computers, C-29(12):1163, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CPT

[Ano80x] Anonymous. Call for papers trends and applications. IEEE Transactions on Computers, C-29(12):1163, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1980:CF

[Ano80y] Anonymous. Compcon 80 Fall. IEEE Transactions on Computers, C-29(7):678, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675646.

Anonymous:1980:CST

[Ano80z] Anonymous. Correction to “Syndrome-Testable Design of Combinational Circuits”. IEEE Transactions on Computers, C-29(11):1012–1013, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675496.

See [Sav80b].

Anonymous:1980:ENa

[Ano80-27] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-29(1):1, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675450.

Anonymous:1980:ENb

[Ano80-28] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-29(9):761, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675674.

Anonymous:1980:F

[Ano80-29] Anonymous. Foreword. IEEE Transactions on Computers, C-29(2):66–67, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675528.

Anonymous:1980:ICSk

[Ano80-30] Anonymous. IEEE Computer Society Press. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675485.

Anonymous:1980:ICSm

[Ano80-31] Anonymous. IEEE Computer Society Press. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675485.

Anonymous:1980:ICSa

[Ano80-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(1):61, January 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675463.

Anonymous:1980:ICSb

[Ano80-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(2):211, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675549.

Anonymous:1980:ICSc

[Ano80-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(3):267, March 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675561.

Anonymous:1980:ICSd

[Ano80-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(4):333, April 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675579.

Anonymous:1980:ICSe

[Ano80-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(5):416, May 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675596.

Anonymous:1980:ICSf

[Ano80-37] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(6):549, June 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675622.

Anonymous:1980:ICSh

[Ano80-38] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(7):678, July 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675645.

Anonymous:1980:ICSi

[Ano80-39] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(8):759, August 1980. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1675670.

Anonymous:1980:ICSj

[Ano80-40] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(9):844, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675686.

Anonymous:1980:ICSl

[Ano80-41] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675486.

Anonymous:1980:ICSn

[Ano80-42] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675486.

Anonymous:1980:ICSo

[Ano80-43] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(11):1035, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675504.

Anonymous:1980:ICSp

[Ano80-44] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-29(12):1163, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675523.

Anonymous:1980:IA

[Ano80-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-29(1):61, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675464.

Anonymous:1980:MPR

[Ano80-46] Anonymous. Machine processing of remotely sensed data. IEEE Transactions on Computers, C-29(11):1035–a, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675503.

Anonymous:1980:PCSa

[Ano80-47] Anonymous. Proceedings of the Conference on Specifications of Reliable Software. IEEE Transactions on Computers, C-29(4):333–a, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675576.

Anonymous:1980:PCSb

[Ano80-48] Anonymous. Proceedings of the Conference on Specifications of Reliable Software. IEEE Transactions on Computers, C-29(6):549–b, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675620.

Anonymous:1980:PCSc

[Ano80-49] Anonymous. Proceedings of the Conference on Specifications of Reliable Software. IEEE Transactions on Computers, C-29(7):678–b, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675644.

Anonymous:1980:TC

[Ano80-50] Anonymous. Table of contents. IEEE Transactions on Computers, C-29(2):65, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675526.

Anonymous:1981:IIT

[Ano81a] Anonymous. 1981 index IEEE Transactions on Computers vol. C-30. IEEE Transactions on Computers, C-30(12):1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675745.

Anonymous:1981:AATa

[Ano81b] Anonymous. Advance announcement: Tutorial Week East 81. IEEE Transactions on Computers, C-30(3):241, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675764.

Anonymous:1981:AATc

[Ano81c] Anonymous. Advance announcement: Tutorial Week East 82. IEEE Transactions on Computers, C-30(12):1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675749.

Anonymous:1981:AATb

[Ano81d] Anonymous. Advance announcement: Tutorial Week West 81. IEEE Transactions on Computers, C-30(9):705, September 1981. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675875.

Anonymous:1981:CPa

[Ano81e] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(3):241, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPb

[Ano81f] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(3):241, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPc

[Ano81g] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(4):301, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPd

[Ano81h] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(4):301, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPe

[Ano81i] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(5):372, May 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPf

[Ano81j] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(6):455, June 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPg

[Ano81k] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(6):455, June 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPh

[Ano81l] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(7):523, July 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPi

[Ano81m] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(7):523, July 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPj

[Ano81n] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(8):613, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPk

[Ano81o] Anonymous. Call for papers.

IEEE Transactions on Computers, C-30(8):613, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPl

[Ano81p] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(9):705, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPm

[Ano81q] Anonymous. Call for papers. IEEE Transactions on Computers, C-30(12):1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPSb

[Ano81r] Anonymous. Call for papers special issue on computer architecture for pattern analysis and image database management. IEEE Transactions on Computers, C-30(8):613, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPSc

[Ano81s] Anonymous. Call for papers special issue on reliable and fault-tolerant computing. IEEE Transactions on Computers, C-30(8):613, August 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPSa

[Ano81t] Anonymous. Call for papers special issue on supersystems. IEEE Transactions on Computers, C-30(4):301, April 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPTa

[Ano81u] Anonymous. Call for papers trends and applications. IEEE Transactions on Computers, C-30(10):819, October 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CPTb

[Ano81v] Anonymous. Call for papers trends and applications. IEEE Transactions on Computers, C-30(11):904, November 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1981:CSPa

[Ano81w] Anonymous. Computer software professional. IEEE Transactions on Computers, C-30(3):241–c, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675766.

Anonymous:1981:CSPb

[Ano81x] Anonymous. Computer software professional. IEEE Transactions on Computers, C-30(4):301–c, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675782.

Anonymous:1981:CEM

[Ano81y] Anonymous. Correction to “An Efficient Method for Improving Reliability of a Pipeline FFT”. IEEE Transactions on Computers, C-30(10):819, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675705.

See [Gla80].

Anonymous:1981:CDM

[Ano81z] Anonymous. Correction to “Dual-Mode Logic for Function-Independent Fault Testing”. IEEE Transactions on Computers, C-30(10):819, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675704.

See [DHR80].

Anonymous:1981:CMS

[Ano81-27] Anonymous. Correction to “The MPG System: A Machine-Independent Efficient Microprogram Generator”. IEEE Transactions on Computers, C-30(10):818, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675703.

See [BH81].

Anonymous:1981:EN

[Ano81-28] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-30(3):169–171, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675752.

Anonymous:1981:ICSa

[Ano81-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(3):241, March 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675767.

Anonymous:1981:ICSb

[Ano81-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(4):301, April 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675783.

Anonymous:1981:ICSc

[Ano81-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(5):372, May 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675800.

Anonymous:1981:ICSd

[Ano81-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(6):455, June 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675821.

Anonymous:1981:ICSe

[Ano81-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(7):523, July 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675833.

Anonymous:1981:ICSf

[Ano81-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(8):613, August 1981. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675855.

Anonymous:1981:ICSg

[Ano81-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(9):705, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675874.

Anonymous:1981:ICSh

[Ano81-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(10):819, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675707.

Anonymous:1981:ICSi

[Ano81-37] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(11):904, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675724.

Anonymous:1981:ICSj

[Ano81-38] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-30(12):1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675748.

Anonymous:1981:ITCa

[Ano81-39] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-30(11):904, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675722.

Anonymous:1981:ITCb

[Ano81-40] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-30(12):1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675746.

Anonymous:1981:I

[Ano81-41] Anonymous. Introduction. IEEE Transactions on Computers, C-30(11):821–822, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675711.

Anonymous:1981:PSIa

[Ano81-42] Anonymous. Planned special issues IEEE Transactions on Computers. IEEE Transactions on Computers, C-30(9):704, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675871.

Anonymous:1981:PSIb

[Ano81-43] Anonymous. Planned special issues IEEE Transactions on Computers. IEEE Transactions on Computers, C-30(10):819, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675706.

Anonymous:1981:SCSa

[Ano81-44] Anonymous. Students in computer science and engineering! IEEE Transactions on Computers, C-30(9):705, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675873.

Anonymous:1981:SCSb

[Ano81-45] Anonymous. Students in computer science and engineering! IEEE Transactions on Computers, C-30(11):904, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675723.

Anonymous:1982:IIT

[Ano82a] Anonymous. 1982 index IEEE Transactions on Computers vol. C-31. IEEE Transactions on Computers, C-31(12):1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675948.

Anonymous:1982:ICPa

[Ano82b] Anonymous. 1983 International Conference on Parallel Processing. IEEE Transactions on Computers, C-31(11):1130, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675931.

Anonymous:1982:ICPb

[Ano82c] Anonymous. 1983 International Conference on Parallel Processing. IEEE Transactions on Computers, C-31(12):1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675952.

Anonymous:1982:CLC

[Ano82d] Anonymous. 7th Conference on Local Computer Networks. IEEE Transactions on Computers, C-31(9):917, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676111.

Anonymous:1982:AATa

[Ano82e] Anonymous. Advance announcement: Tutorial Week East 82. IEEE Transactions on Computers, C-31(2):177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675970.

Anonymous:1982:AATd

[Ano82f] Anonymous. Advance announcement: Tutorial Week East 83. IEEE Transactions on Computers, C-31(11):1130, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675934.

Anonymous:1982:AATb

[Ano82g] Anonymous. Advance announcement: Tutorial Week West 82. IEEE Transactions on Computers, C-31(8):805, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676091.

Anonymous:1982:AATc

[Ano82h] Anonymous. Advance announcement: Tutorial Week West 82. IEEE Transactions on Computers, C-31(9):917, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676114.

Anonymous:1982:A

[Ano82i] Anonymous. Announcement.

IEEE Transactions on Computers, C-31(6):477, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676029.

Anonymous:1982:AIC

[Ano82j] Anonymous. Announcing The 3rd International Conference on Distributed Computing Systems. IEEE Transactions on Computers, C-31(6):568, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676045.

Anonymous:1982:CPa

[Ano82k] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(1):87, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPb

[Ano82l] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(2):177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPc

[Ano82m] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(4):343, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPd

[Ano82n] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(5):474, May 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPe

[Ano82o] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(6):568, June 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPf

[Ano82p] Anonymous. Call for papers. IEEE Transactions on Computers, C-31(6):568, June 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPg

[Ano82q] Anonymous. Call for papers.

IEEE Transactions on Computers, C-31(8):805, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPh

[Ano82r] Anonymous. Call for papers.

IEEE Transactions on Computers, C-31(8):805, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPi

[Ano82s] Anonymous. Call for papers.

IEEE Transactions on Computers, C-31(10):1031, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPj

[Ano82t] Anonymous. Call for papers.

IEEE Transactions on Computers, C-31(10):1031, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPTa

[Ano82u] Anonymous. Call for papers 1983 Trends and Applications Conference. IEEE Transactions on Computers, C-31(9): 917, September 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPTb

[Ano82v] Anonymous. Call for papers 1983 trends and applications conference. IEEE Transactions on Computers, C-31(11):1130, November 1982. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPTc

[Ano82w] Anonymous. Call for papers 1983 Trends and Applications Conference. IEEE Transactions on Computers, C-31(12):1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:CPEa

[Ano82x] Anonymous. Call for papers Eighth Data Communications Symposium — 1983. IEEE Transactions on Computers, C-31(12):1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:ENa

[Ano82y] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-31(1):1, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675878.

Anonymous:1982:ENb

[Ano82z] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-31(4):269, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675991.

Anonymous:1982:ENc

[Ano82-27] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-31(7):573–574, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676052.

Anonymous:1982:ICSb

[Ano82-28] Anonymous. IEEE Computer Society Press checklist of new titles December 1981. IEEE Transactions on Computers, C-31(2):177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675967.

Anonymous:1982:ICSd

[Ano82-29] Anonymous. IEEE Computer Society press checklist of new titles December 1981. IEEE Transactions on Computers, C-31(3):266, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675986.

Anonymous:1982:ICSf

[Ano82-30] Anonymous. IEEE Computer Society Press checklist of new titles December 1982. IEEE Transactions on Computers, C-31(4):343, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676007.

Anonymous:1982:ICSh

[Ano82-31] Anonymous. IEEE Computer Society Press checklist of new titles December 1982. IEEE Transactions on Computers, C-31(5):474, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676025.

Anonymous:1982:ICSj

[Ano82-32] Anonymous. IEEE Computer Society Press checklist of new titles December 1982. IEEE Transactions on Computers, C-31(6):568, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676046.

Anonymous:1982:ICSa

[Ano82-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(1):87, January 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675896.

Anonymous:1982:ICSc

[Ano82-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(2):177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675969.

Anonymous:1982:ICSe

[Ano82-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(3):266, March 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675987.

Anonymous:1982:ICSg

[Ano82-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(4):343, April 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676008.

Anonymous:1982:ICSi

[Ano82-37] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(5):474, May 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676026.

Anonymous:1982:ICSk

[Ano82-38] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(6):568, June 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676048.

Anonymous:1982:ICSl

[Ano82-39] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(7):706, July 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676072.

Anonymous:1982:ICSm

[Ano82-40] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(8):805, August 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676092.

Anonymous:1982:ICSn

[Ano82-41] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(9):917, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676113.

Anonymous:1982:ICSo

[Ano82-42] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(10):1031, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675915.

Anonymous:1982:ICSp

[Ano82-43] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(11):1130, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675933.

Anonymous:1982:ICSq

[Ano82-44] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-31(12):1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675951.

Anonymous:1982:ICFa

[Ano82-45] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-31(1):87, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675894.

Anonymous:1982:ICFb

[Ano82-46] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-31(7):706, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676071.

Anonymous:1982:ITCa

[Ano82-47] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(1):86, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675892.

Anonymous:1982:ITCb

[Ano82-48] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(2):177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675966.

Anonymous:1982:ITCc

[Ano82-49] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(3):266, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675985.

Anonymous:1982:ITCd

[Ano82-50] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(4):342, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676005.

Anonymous:1982:ITCe

[Ano82-51] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(5):474, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676023.

Anonymous:1982:ITCf

[Ano82-52] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(6):568, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676044.

Anonymous:1982:ITCg

[Ano82-53] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-31(8):805, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676090.

Anonymous:1982:IAa

[Ano82-54] Anonymous. Information for authors. IEEE Transactions on Computers, C-31(1):87, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675897.

Anonymous:1982:IAb

[Ano82-55] Anonymous. Information for authors. IEEE Transactions on Computers, C-31(7):706, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676073.

Anonymous:1982:Ia

[Ano82-56] Anonymous. Innovation. IEEE Transactions on Computers, C-31(3):266, March 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1675988.

Anonymous:1982:Ib

[Ano82-57] Anonymous. Innovation. IEEE Transactions on Computers, C-31(5):474, May 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676027.

Anonymous:1982:IPD

[Ano82-58] Anonymous. Introduction: Parallel and distributed processing. IEEE Transactions on Computers, C-31(11):1033–1035, November 1982. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675919.

Anonymous:1982:LRJ

[Ano82-59] Anonymous. List of referees January 1, 1980–June 1981. IEEE Transactions on Computers, C-31(1):82–85, January 1982. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1982:PPN

[Ano82-60] Anonymous. Preliminary program ninth annual international symposium on computer architecture. IEEE Transactions on Computers, C-31(4):343, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676009.

Anonymous:1982:SCS

[Ano82-61] Anonymous. Students in computer science and engineering! IEEE Transactions on Computers, C-31(1):87, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675895.

Anonymous:1983:IIT

[Ano83a] Anonymous. 1983 index IEEE Transactions on Computers vol.

C-32. IEEE Transactions on Computers, C-32(12):1212, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676190.

Anonymous:1983:ISM

[Ano83b] Anonymous. The 1983 international symposium on multiplevalued logic. IEEE Transactions on Computers, C-32(2): 207, February 1983. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1676210.

Anonymous:1983:APW

[Ano83c] Anonymous. Acknowledgment of prior work. IEEE Transactions on Computers, C-32(9):880, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676342.

Anonymous:1983:CPa

[Ano83d] Anonymous. Call for papers. IEEE Transactions on Computers, C-32(2):207, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPb

[Ano83e] Anonymous. Call for papers. IEEE Transactions on Computers, C-32(3):328, March 1983.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPc

[Ano83f] Anonymous. Call for papers.

IEEE Transactions on Computers, C-32(4):422, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPe

[Ano83g] Anonymous. Call for papers.

IEEE Transactions on Computers, C-32(5):520, May 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPf

[Ano83h] Anonymous. Call for papers.

IEEE Transactions on Computers, C-32(6):598, June 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPg

[Ano83i] Anonymous. Call for papers.

IEEE Transactions on Computers, C-32(6):598, June 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPJ

[Ano83j] Anonymous. Call for papers 4th Jerusalem Conference on Information Technology (JCIT). IEEE Transactions on Computers, C-32(4):422, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:CPS

[Ano83k] Anonymous. Call for papers special issue on reliable and fault-tolerant computing. IEEE Transactions on Computers, C-32(6):598, June 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1983:Cc

[Ano83l]  Anonymous.  Compsac 83.

IEEE Transactions on Computers, C-32(10):976, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676146.

Anonymous:1983:Ca

[Ano83m] Anonymous. Compsac83. IEEE Transactions on Computers, C-32(8):784, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676325.

Anonymous:1983:Cb

[Ano83n] Anonymous. Compsac83. IEEE Transactions on Computers, C-32(9):880, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676344.

Anonymous:1983:ENa

[Ano83o] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-32(1):1, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676117.

Anonymous:1983:ENb

[Ano83p] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-32(7):601–602, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676291.

Anonymous:1983:GEI

[Ano83q] Anonymous. Guest Editors’ introduction performance evaluation of multiple processor systems. IEEE Transactions on Computers, C-32(1):2–3, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676118.

Anonymous:1983:ICSf

[Ano83r] Anonymous. IEEE Computer Society. IEEE Transactions on Computers, C-32(6):598, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676287.

Anonymous:1983:ICSa

[Ano83s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(1):98, January 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676129.

Anonymous:1983:ICSb

[Ano83t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(2):207, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676209.

Anonymous:1983:ICSc

[Ano83u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(3):328, March 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676228.

Anonymous:1983:ICSd

[Ano83v] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(4):422, April 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676249.

Anonymous:1983:ICSe

[Ano83w] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(5):520, May 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676269.

Anonymous:1983:ICSg

[Ano83x] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(7):696, July 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676303.

Anonymous:1983:ICSh

[Ano83y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(8):784, August 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676324.

Anonymous:1983:ICSi

[Ano83z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(9):880, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676343.

Anonymous:1983:ICSj

[Ano83-27] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(10):976, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676145.

Anonymous:1983:ICSk

[Ano83-28] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(11):1072, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676163.

Anonymous:1983:ICSl

[Ano83-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-32(12):1212, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676191.

Anonymous:1983:ICF

[Ano83-30] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-32(1):98, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676128.

Anonymous:1983:ITC

[Ano83-31] Anonymous. IEEE Transactions on Computers planned special issues. IEEE Transactions on Computers, C-32(5):520, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676268.

Anonymous:1983:IAa

[Ano83-32] Anonymous. Information for authors. IEEE Transactions on Computers, C-32(1):98, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676130.

Anonymous:1983:IAb

[Ano83-33] Anonymous. Information for authors. IEEE Transactions on Computers, C-32(7):696, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676304.

Anonymous:1983:IAc

[Ano83-34] Anonymous. Information for authors. IEEE Transactions on Computers, C-32 (11):1072, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676164.

Anonymous:1983:IAd

[Ano83-35] Anonymous. Information for authors. IEEE Transactions on Computers, C-32(12):1212, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676192.

Anonymous:1984:IIT

[Ano84a] Anonymous. 1984 index IEEE Transactions on Computers vol. C-33. IEEE Transactions on Computers, C-33(12):1265, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676397.

Anonymous:1984:Aa

[Ano84b] Anonymous. Advertisement. IEEE Transactions on Computers, C-33(1):108, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009320.

Anonymous:1984:Ab

[Ano84c] Anonymous. Advertisement. IEEE Transactions on Computers, C-33(1):110, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009322.

Anonymous:1984:CP

[Ano84d] Anonymous. Call for papers. IEEE Transactions on Computers, C-33(10):948, October 1984.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1984:CPS

[Ano84e] Anonymous. Call for papers special issue on sorting. IEEE Transactions on Computers, C-33(4):376, April 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1984:CPD

[Ano84f]  Anonymous.  Correction to

“A Pipelined Distributed Arithmetic PFFT Processor”. IEEE Transactions on Computers, C-33(3):288, March 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=

1676429. See [CVY83].

Anonymous:1984:EN

[Ano84g] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-33(9):773, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676493.

Anonymous:1984:ICSa

[Ano84h] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(2):200, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676413.

Anonymous:1984:ICSb

[Ano84i] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(3):288, March 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676430.

Anonymous:1984:ICSc

[Ano84j] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(4):376, April 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676449.

Anonymous:1984:ICSd

[Ano84k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(5):464, May 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676465.

Anonymous:1984:ICSe

[Ano84l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(6):588, June 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676488.

Anonymous:1984:ICSf

[Ano84m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(9):860, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676505.

Anonymous:1984:ICSg

[Ano84n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(10):948, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676363.

Anonymous:1984:ICSh

[Ano84o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(11):1048, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676382.

Anonymous:1984:ICSi

[Ano84p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-33(12):1265, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676399.

Anonymous:1984:ICFa

[Ano84q] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-33(1):109, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009321.

Anonymous:1984:ICFb

[Ano84r]  Anonymous.  IEEE copy-

right form. IEEE Transactions on Computers, C-33 (1):111–112, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009323.

Anonymous:1984:ICFc

[Ano84s] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-33(7):683–684, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009347.

Anonymous:1984:ICFd

[Ano84t] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-33(12):1265, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676398.

Anonymous:1984:IAa

[Ano84u] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(2):200, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676414.

Anonymous:1984:IAb

[Ano84v] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(3):288, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676431.

Anonymous:1984:IAc

[Ano84w] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(4):376, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676450.

Anonymous:1984:IAd

[Ano84x] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(5):464, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676466.

Anonymous:1984:IAe

[Ano84y] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(6):588, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676489.

Anonymous:1984:IAf

[Ano84z] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(9):860, September 1984. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1676506.

Anonymous:1984:IAg

[Ano84-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(10):948, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676364.

Anonymous:1984:IAh

[Ano84-28] Anonymous. Information for authors. IEEE Transactions on Computers, C-33 (11):1048, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676383.

Anonymous:1984:IAi

[Ano84-29] Anonymous. Information for authors. IEEE Transactions on Computers, C-33(12):1265, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676400.

Anonymous:1984:LRa

[Ano84-30] Anonymous. List of referees.

IEEE Transactions on Computers, C-33(7):590–591, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009331.

Anonymous:1984:LRb

[Ano84-31] Anonymous. List of referees. IEEE Transactions on Computers, C-33(11):951, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:CPS

[Ano85a] Anonymous. Call for papers special issue on fault-tolerant computing. IEEE Transactions on Computers, C-34(5): 488, May 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:CP

[Ano85b] Anonymous. Calls for papers. IEEE Transactions on Computers, C-34(4):392, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009394.

Anonymous:1985:CCI

[Ano85c] Anonymous. Correction to centennial issue. IEEE Transactions on Computers, C-34(1):98, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676524.

Anonymous:1985:CMT

[Ano85d] Anonymous. Correction to “Modeling and Test Generation Algorithms for MOS Circuits”. IEEE Transactions on Computers, C-34(7):680, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676609.

See [JA85].

Anonymous:1985:ENa

[Ano85e] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-34(1):1, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676509.

Anonymous:1985:ENb

[Ano85f] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-34(7):585–587, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676596.

Anonymous:1985:ICSb

[Ano85g] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(1):99, January 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676526.

Anonymous:1985:ICSd

[Ano85h] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(2):196, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676559.

Anonymous:1985:ICSf

[Ano85i] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(3):292, March 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676575.

Anonymous:1985:ICSj

[Ano85j] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(5):488, May 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676592.

Anonymous:1985:ICSm

[Ano85k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(7):680, July 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676610.

Anonymous:1985:ICSo

[Ano85l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(8):776, August 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676629.

Anonymous:1985:ICSq

[Ano85m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(9):872, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676648.

Anonymous:1985:ICSs

[Ano85n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-34(11):1068, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676542.

Anonymous:1985:ICF

[Ano85o] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-34(1):99, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676525.

Anonymous:1985:IAa

[Ano85p] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(1):99, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676527.

Anonymous:1985:IAb

[Ano85q] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(2):196, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAc

[Ano85r] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(3):292, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAd

[Ano85s] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(5):488, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAe

[Ano85t] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(7):680, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAf

[Ano85u] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(8):776, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAg

[Ano85v] Anonymous. Information for authors. IEEE Transactions on Computers, C-34(9):872, September 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1985:IAh

[Ano85w] Anonymous. Information for authors. IEEE Transactions on Computers, C-34 (11):1068, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676543.

Anonymous:1985:LR

[Ano85x] Anonymous. List of referees. IEEE Transactions on Computers, C-34(4):295, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009379.

Anonymous:1985:MA

[Ano85y] Anonymous. Membership application. IEEE Transactions on Computers, C-34(1):2, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676510.

Anonymous:1986:RL

[Ano86a] Anonymous. 1985 referee list.

IEEE Transactions on Computers, C-35(6):486–488, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009425.

Anonymous:1986:IIT

[Ano86b] Anonymous. 1986 index IEEE Transactions on Computers vol. C-35. IEEE Transactions on Computers, C-35(12): 1093–1102, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676720.

Anonymous:1986:CPSa

[Ano86c] Anonymous. Calls for papers special issue on parallel and distributed processing. IEEE Transactions on Computers, C-35(8):772, August 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:CPSb

[Ano86d] Anonymous. Calls for papers special issue on real-time systems. IEEE Transactions on Computers, C-35(9):852, September 1986. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:ENa

[Ano86e] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-35(4):285, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676758.

Anonymous:1986:ENb

[Ano86f]  Anonymous.  Editor’s no-

tice. IEEE Transactions on Computers, C-35(6):485, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009424.

Anonymous:1986:ENc

[Ano86g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-35(7):581–582, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676798.

Anonymous:1986:ICSb

[Ano86h] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(1):95, January 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676672.

Anonymous:1986:ICSd

[Ano86i] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(2):188, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676740.

Anonymous:1986:ICSf

[Ano86j] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(3):284, March 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676754.

Anonymous:1986:ICSh

[Ano86k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(4):388, April 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676777.

Anonymous:1986:ICSj

[Ano86l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(5):484, May 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676794.

Anonymous:1986:ICSn

[Ano86m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(7):675, July 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676815.

Anonymous:1986:ICSp

[Ano86n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(8):772, August 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676833.

Anonymous:1986:ICSr

[Ano86o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(9):852, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676850.

Anonymous:1986:ICSt

[Ano86p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(10):932, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676691.

Anonymous:1986:ICSv

[Ano86q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(11):1012, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676705.

Anonymous:1986:ICSx

[Ano86r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-35(12):1103, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676722.

Anonymous:1986:ICFa

[Ano86s] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-35(1):95, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676671.

Anonymous:1986:ICFb

[Ano86t] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-35(7):675, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676814.

Anonymous:1986:IAa

[Ano86u] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(1):95, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAb

[Ano86v] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(2):188, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAc

[Ano86w] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(3):284, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAd

[Ano86x] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(4):388, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAe

[Ano86y] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(5):484, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAf

[Ano86z] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(7):675, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAg

[Ano86-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(8):772, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAh

[Ano86-28] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(9):852, September 1986. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAi

[Ano86-29] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(10):932, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAj

[Ano86-30] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(11):1012, November 1986. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:IAk

[Ano86-31] Anonymous. Information for authors. IEEE Transactions on Computers, C-35(12):1103, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1986:LR

[Ano86-32] Anonymous. List of referees. IEEE Transactions on Computers, C-35(4):287, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676760.

Anonymous:1987:IIT

[Ano87a] Anonymous. 1987 index IEEE Transactions on Computers vol. C-36. IEEE Transactions on Computers, C-36(12): 1539–1550, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009503.

Anonymous:1987:CPb

[Ano87b] Anonymous. Call for papers. IEEE Transactions on Computers, C-36(8):1006, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009525.

Anonymous:1987:CPS

[Ano87c] Anonymous. Call for papers special issue on architectural support for programming languages and operating systems. IEEE Transactions on Computers, C-36(7):898, July 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:CPa

[Ano87d] Anonymous. Calls for papers. IEEE Transactions on Computers, C-36(1):126, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009459.

Anonymous:1987:ENa

[Ano87e] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-36(1):1, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009443.

Anonymous:1987:ENb

[Ano87f] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-36(6):645–646, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676957.

Anonymous:1987:ICSd

[Ano87g] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(2):256, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676894.

Anonymous:1987:ICSf

[Ano87h] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(3):384, March 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676915.

Anonymous:1987:ICSh

[Ano87i] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(4):516, April 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1676932.

Anonymous:1987:ICSj

[Ano87j] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(5):644, May 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676953.

Anonymous:1987:ICSl

[Ano87k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(6):772, June 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676971.

Anonymous:1987:ICSn

[Ano87l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(7):899, July 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676989.

Anonymous:1987:ICSt

[Ano87m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-36(10):1264, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676871.

Anonymous:1987:ICF

[Ano87n] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-36(7):899, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676988.

Anonymous:1987:ITCf

[Ano87o] Anonymous. IEEE Transactions on Computers referee list for 1986–1987. IEEE Transactions on Computers, C-36(6):646–649, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676958.

Anonymous:1987:IAb

[Ano87p] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(2):256, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAc

[Ano87q] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(3):384, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAd

[Ano87r] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(4):516, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAe

[Ano87s] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(5):644, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAf

[Ano87t] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(6):772, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAg

[Ano87u] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(7):899, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:IAh

[Ano87v] Anonymous. Information for authors. IEEE Transactions on Computers, C-36(10):1264, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1987:LR

[Ano87w] Anonymous. List of referees. IEEE Transactions on Computers, C-36(4):387, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676920.

Anonymous:1987:MA

[Ano87x] Anonymous. Membership application. IEEE Transactions on Computers, C-36(12):1551, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009505.

Anonymous:1987:RL

[Ano87y] Anonymous. Referee list. IEEE Transactions on Computers, C-36(12):1395, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009492.

Anonymous:1987:RSI

[Ano87z] Anonymous. Reviewers special issue on real-time systems August 1987. IEEE Transactions on Computers, C-36(8):903, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009513.

Anonymous:1987:CS

[Ano87-27] Anonymous. THE Computer Society. IEEE Transactions on Computers, C-36(10):1264, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676870.

Allan:1980:FAP

[AO80] S. J. Allan and A. E. Oldehoeft.

A flow analysis procedure for the translation of high-level languages to a data flow language. IEEE Transactions on Computers, C-29(9):826–831, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675682.

Aykanat:1988:IAS

[AOES88] C. Aykanat, F. Ozguner, F. Ercal, and P. Sadayappan. Iterative algorithms for solution of large sparse systems of linear equations on hypercubes. IEEE Transactions on Computers, 37(12):1554–1568, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9733.

Aspinwall:1985:RVM

[AP85] D. B. Aspinwall and Y. N. Patt.

Retrofitting the VAX-11/780 microarchitecture for IEEE floating point arithmetic — implementation issues, measurements, and analysis. IEEE Transactions on Computers, C-34(8):692–708, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676615.

Akers:1986:FTC

[AP86] S. B. Akers and D. K. Pradhan. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-35(4):285–287, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676759.

Abraham:1989:PDB

[AP89] S. Abraham and K. Padmanabhan. Performance of the direct binary n-cube network for multiprocessors. IEEE Transactions on Computers, 38(7):1000–1011, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30851.

Arnold:1983:EPA

[APD83] C. P. Arnold, M. I. Parr, and M. B. Dewe. An efficient parallel algorithm for the solution of large sparse linear matrix equations. IEEE Transactions on Computers, C-32(3):265–273, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676218.

Aleliunas:1982:ERG

[AR82] R. Aleliunas and A. L. Rosenberg. On embedding rectangular grids in square grids. IEEE Transactions on Computers, C-31(9):907–913, September 1982. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676109.

Agrawal:1983:ICA

[AR83] D. P. Agrawal and T. R. N.

Rao. Introduction: Computer arithmetic. IEEE Transactions on Computers, C-32(4):329–330, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676232.

Abadir:1986:FTG

[AR86] M. S. Abadir and H. K. Reghbati. Functional test generation for digital circuits described using binary decision diagrams. IEEE Transactions on Computers, C-35(4):375–379, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676774.

Armstrong:1980:CCC

[Arm80] J. R. Armstrong. The complexity of computational circuits versus radix. IEEE Transactions on Computers, C-29 (10):937–941, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675480.

Arya:1985:OIS

[Ary85] S. Arya. An optimal instructionscheduling model for a class of vector processors. IEEE Transactions on Computers, C-34(11):981–995, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676531.

Adams:1982:ESC

[AS82a] G. B. Adams III and H. J. Siegel. The extra stage cube: a fault-tolerant interconnection network for supersystems. IEEE Transactions on Computers, C-31(5):443–454, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676021.

Adams:1982:NPP

[AS82b] G. B. Adams III and H. J. Siegel. On the number of permutations performable by the augmented data manipulator network. IEEE Transactions on Computers, C-31(4):270–277, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675992.

Ashtaputre:1985:SAE

[AS85] S. Ashtaputre and C. D. Savage. Systolic arrays with embedded tree structures for connectivity problems. IEEE Transactions on Computers, C-34(5):483–484, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676589.

Akl:1987:OPM

[AS87a] Selim G. Akl and Nicola Santoro. Optimal parallel merging and sorting without memory conflicts. IEEE Transactions on Computers, C-36 (11):1367–1369, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009478.

Awerbuch:1987:NCM

[AS87b] B. Awerbuch and Y. Shiloach.

New connectivity and MSF algorithms for shuffle-exchange network and PRAM. IEEE Transactions on Computers, C-36(10):1258–1263, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676869.

Abu-Sufah:1986:IOS

[ASHK86] Walid Abu-Sufah, Harlan E.

Husmann, and David J. Kuck. On input/output speedup in tightly coupled multiprocessors. IEEE Transactions on Computers, C-35(6):520–530, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009429.

Abu-Sufah:1981:PEP

[ASKL81] W. Abu-Sufah, D. J. Kuck, and D. H. Lawrie. On the performance enhancement of paging systems through program analysis and transformations. IEEE Transactions on Computers, C-30(5):341–356, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675792.

Apostolopoulos:1987:ANR

[ASP87] T. K. Apostolopoulos, E. D. Sykas, and E. N. Protonotarios. Analysis of a new retransmission control algorithm for slotted CSMA/CD LAN’s. IEEE Transactions on Computers, C-36(6):692–701, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676962.

Avis:1981:OAD

[AT81] D. Avis and G. T. Toussaint. An optimal algorithm for determining the visibility of a polygon from an edge. IEEE Transactions on Computers, C-30(12):910–914, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675729.

Atallah:1985:SD

[Ata85] M. J. Atallah. On symmetry detection. IEEE Transactions on Computers, C-34(7):663–666, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676605.

Atallah:1988:MAP

[Ata88] M. J. Atallah. On multidimensional arrays of processors. IEEE Transactions on Computers, 37(10):1306–1309, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5995.

Almer:1981:CCS

[ATT81] D. Almer, M. A. Tapia, and J. H.

Tucker. Correction to “Complete Solution of Boolean Equations”. IEEE Transactions on Computers, C-30(10):812, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675701.

See [TT80].

Aupperle:1983:MES

[Aup83] E. M. Aupperle. Merit’s evolution — statistically speaking. IEEE Transactions on Computers, C-32(10):881–902, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676133.

Bhuyan:1982:GBS

[BA82] L. N. Bhuyan and D. P. Agarwal. On the generalized binary system. IEEE Transactions on Computers, C-31(4):335–338, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676003.

Bhuyan:1983:DPG

[BA83] L. N. Bhuyan and D. P. Agrawal. Design and performance of generalized interconnection networks. IEEE Transactions on Computers, C-32(12):1081–1090, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676168.

Bhuyan:1984:GHH

[BA84a] L. N. Bhuyan and D. P. Agrawal. Generalized hypercube and hyperbus structures for a computer network. IEEE Transactions on Computers, C-33(4):323–333, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676437.

Brahme:1984:FTM

[BA84b] D. Brahme and J. A. Abraham. Functional testing of microprocessors. IEEE Transactions on Computers, C-33(6):475–485, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676471.

Banerjee:1986:BAB

[BA86] P. Banerjee and J. A. Abraham.

Bounds on algorithm-based fault tolerance in multiple processor systems. IEEE Transactions on Computers, C-35(4):296–306, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676762.

Bailey:1987:VCM

[Bai87] D. H. Bailey. Vector computer memory bank contention. IEEE Transactions on Computers, C-36(3):293–298, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676901.

Banerjee:1988:CRC

[Ban88] P. Banerjee. The cubical ring connected cycles: a fault tolerant parallel computation network. IEEE Transactions on Computers, 37(5):632–636, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4617.

Birta:1987:PBP

[BAR87] L. G. Birta and O. Abou-Rabia. Parallel block predictor– corrector methods for ODE’s. IEEE Transactions on Computers, C-36(3):299–311, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676902.

Batcher:1980:DMP

[Bat80]  K. E. Batcher.  Design of

a massively parallel processor. IEEE Transactions on Computers, C-29(9):836–840, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675684.

Batcher:1982:BSP

[Bat82]  K. E. Batcher.  Bit-serial

parallel processing systems. IEEE Transactions on Computers, C-31(5):377–384, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676015.

Berger:1987:PSN

[BB87a] Marsha J. Berger and Shahid H. Bokhari. A partitioning strategy for non-uniform problems on multiprocessors. IEEE Transactions on Computers, C-36(5):570–580, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676942.

ICASE Report No. 85-55, Nov. 1985.

Bertossi:1987:VIS

[BB87b] A. A. Bertossi and M. A. Bonuccelli. A VLSI implementation of the simplex algorithm. IEEE Transactions on Computers, C-36(2):241–247, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676889.

Bender:1989:SPR

[BB89a] E. A. Bender and J. T.

Butler. On the size of PLAs required to realize binary and multiple-valued functions. IEEE Transactions on Computers, 38(1):82–98, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8731.

Bertossi:1989:GDV

[BB89b] A. A. Bertossi and M. A. Bonuccelli. A gracefully degradable VLSI system for linear programming. IEEE Transactions on Computers, 38(6):853–861, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24294.

Buehrer:1982:EME

[BBB^{+}82] R. E. Buehrer, H. Brundiers, H. Benz, B. Bron, H. Friess, W. Haelg, H. J. Halin, A. Isacson, and M. Tadian. The ETHmultiprocessor Empress: a dynamically configurable MIMD system. IEEE Transactions on Computers, C-31(11):1035–1044, November 1982. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675920.

Balbo:1988:CQN

[BBG88] G. Balbo, S. C. Bruell, and S. Ghanta. Combining queueing networks and generalized stochastic Petri nets for the solution of complex models of system behavior. IEEE Transactions on Computers, 37 (10):1251–1268, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5986.

Bernstein:1988:OCE

[BBP88] D. Bernstein, H. Boral, and R. Y. Pinter. Optimal chaining in expression trees. IEEE Transactions on Computers, 37(11):1366–1374, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8702.

Beck:1988:SLF

[BBW88] M. Beck, D. Bitton, and W. K. Wilkinson. Sorting large files on a backend multiprocessor. IEEE Transactions on Computers, 37(7):769–778, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2222.

Best:1985:CSS

[BC85] E. Best and F. Cristian. Comments on “Self-Stabilizing Programs: The Fault-Tolerant Capability of Self-Checking Programs”. IEEE Transactions on Computers, C-34(1):97–98, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676522.

See [Mil82].

Berthet:1988:AMA

[BC88] C. Berthet and E. Cerny.

An algebraic model for asynchronous circuits verification. IEEE Transactions on Computers, 37(7):835–847, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2229.

Burnell:1980:MUS

[BCA80] J. F. Burnell, S. C. Crist, and M. Arozullah. Microprocessor utilization in satelliteborn packet switching. IEEE Transactions on Computers, C-29(2):206–208, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675546.

Browne:1986:AVS

[BCDM86] M. C. Browne, E. M. Clarke, D. L. Dill, and B. Mishra. Automatic verification of sequential circuits using temporal logic. IEEE Transactions on Computers, C-35(12):1035–1044, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676711.

Barzilai:1983:EGB

[BCR83] Z. Barzilai, D. Coppersmith, and A. L. Rosenberg. Exhaustive generation of bit patterns with applications to VLSI self-testing. IEEE Transactions on Computers, C-32(2):190–194, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676202.

Brown:1980:ILB

[BD80] M. R. Brown and D. P. Dobkin. An improved lower bound on polynomial multiplication. IEEE Transactions on Computers, C-29(5):337–340, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675583.

Briggs:1983:EPC

[BD83] F. A. Briggs and M. Dubois.

Effectiveness of private caches in multiprocessor systems with parallel-pipelined memories. IEEE Transactions on Computers, C-32(1):48–59, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676123.

Banerjee:1989:DAS

[BD89] P. Banerjee and A. Dugar.

The design, analysis and simulation of a fault-tolerant interconnection network supporting the fetch-and-add primitive. IEEE Transactions on Computers, 38(1):30–46, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8728.

Baer:1983:BSM

[BDL83] J.-L. Baer, Hung-Chang Du, and R. E. Lander. Binary search in a multiprocessing environment. IEEE Transactions on Computers, C-32(7):667–677, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676298.

Blazewicz:1986:SMT

[BDW86] J. Blazewicz, M. Drabowski, and J. Weglarz. Scheduling multiprocessor tasks to minimize schedule length. IEEE Transactions on Computers, C-35(5):389–393, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676781.

Buehrer:1987:IDF

[BE87] Richard Buehrer and Kattamuri Ekanadham. Incorporating data flow ideas into von Neumann processors for parallel execution. IEEE Transactions on Computers, C-36(12):1515–1522, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009501.

Becker:1988:ETO

[Bec88]  B. Becker.  Efficient test-

ing of optimal time adders. IEEE Transactions on Computers, 37(9):1113–1121, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2262.

Besslich:1983:MGP

[Bes83] P. W. Besslich. A method for the generation and processing of dyadic indexed data. IEEE Transactions on Computers, C-32(5):487–494, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676260.

Besslich:1986:HMM

[Bes86]  P. W. Besslich.  Heuristic

minimization of MVL functions: a direct cover approach. IEEE Transactions on Computers, C-35(2):134–144, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676731.

Breuer:1980:FLP

[BF80] M. A. Breuer and A. D.

Friedman. Functional level primitives in test generation. IEEE Transactions on Computers, C-29(3):223–235, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675554.

Babaoglu:1983:TLR

[BF83] O. Babaoglu and D. Ferrari.

Two-level replacement decisions in paging stores. IEEE Transactions on Computers, C-32(12):1151–1159, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676176.

Bisiani:1988:MPP

[BF88] R. Bisiani and A. Forin. Multilanguage parallel programming of heterogeneous machines. IEEE Transactions on Computers, 37(8):930–945, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2245.

Baru:1989:DOC

[BF89] C. K. Baru and O. Frieder. Database operations in a cubeconnected multicomputer system. IEEE Transactions on Computers, 38(6):920–927, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24307.

Briggs:1982:PAP

[BFHW82] F. A. Briggs, King-Sun Fu, Kai Hwang, and B. W. Wah. PUMPS architecture for pattern analysis and image database management. IEEE Transactions on Computers, C-31 (10):969–983, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675906.

Banerjee:1984:FEL

[BG84] U. Banerjee and D. D. Gajski. Fast execution of loops with IF statements. IEEE Transactions on Computers, C-33 (11):1030–1033, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676377.

Bhattacharya:1986:ICF

[BG86a] B. B. Bhattacharya and B. Gupta.

On the impossible class of faulty functions in logic networks under short circuit faults. IEEE Transactions on Computers, C-35(1):85–90, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676666.

Bilgory:1986:HSS

[BG86b] A. Bilgory and D. D. Gajski.

A heuristic for suffix solutions. IEEE Transactions on Computers, C-35(1):34–42, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676655.

Barbara:1987:RVM

[BGM87] D. Barbara and H. Garcia-Molina. The reliability of voting mechanisms. IEEE Transactions on Computers, C-36 (10):1197–1208, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676860.

Blaum:1988:RSE

[BGM88] M. Blaum, R. Goodman, and R. McEliece. The reliability of single-error protected computer memories. IEEE Transactions on Computers, 37(1):114–119, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=75143.

Brown:1989:TSS

[BGW89] G. M. Brown, M. G. Gouda, and C.-L. Wu. Token systems that self-stabilize. IEEE Transactions on Computers, 38(6):845–852, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24293.

Bhuyan:1989:AAS

[BGY89] L. N. Bhuyan, D. Ghosal, and Q. Yang. Approximate analysis of single and multiple ring networks. IEEE Transactions on Computers, 38(7):1027–1040, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30853.

Bossen:1980:AR

[BH80] D. C. Bossen and S. J. Hong. Author’s reply. IEEE Transactions on Computers, C-29(8):759, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675667.

Baba:1981:MSM

[BH81] T. Baba and H. Hagiwara.

The MPG system: a machineindependent efficient microprogram generator. IEEE Transactions on Computers, C-30(6):373–395, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675804.

See correction [Ano81-27].

Burton:1984:VTM

[BH84] F. W. Burton and M. M. Huntbach. Virtual tree machines. IEEE Transactions on Computers, C-33(3):278–280, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676425.

Bhat:1983:EAF

[Bha83a] K. V. S. Bhat. An efficient approach for fault diagnosis in a Boolean n-cube array of microprocessors. IEEE Transactions on Computers, C-32(11):1070–1071, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676161.

Bhattacharyya:1983:NAF

[Bha83b] A. Bhattacharyya. On a novel approach of fault detection in an easily testable sequential machine with extra inputs and extra outputs. IEEE Transactions on Computers, C-32(3):323–325, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676226.

Bhuyan:1985:APM

[Bhu85]  L. N. Bhuyan.  An analy-

sis of processor-memory interconnection networks. IEEE Transactions on Computers, C-34(3):279–283, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676571.

Breuer:1986:REF

[BI86] M. A. Breuer and A. A. Ismaeel. Roving emulation as a fault detection mechanism. IEEE Transactions on Computers, C-35(11):933–939, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676695.

Brand:1988:TAU

[BI88] D. Brand and V. S. Iyengar. Timing analysis using functional analysis. IEEE Transactions on Computers, 37 (10):1309–1314, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5996.

Bilardi:1989:MSN

[Bil89] G. Bilardi. Merging and sorting networks with the topology of the omega network. IEEE Transactions on Computers, 38(10):1396–1403, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35835.

Baba:1982:TLM

[BIO82] T. Baba, K. Ishikawa, and K. Okuda. A two-level microprogrammed multiprocessor computer with nonnumeric functions. IEEE Transactions on Computers, C-31(12): 1142–1156, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675938.

Biswas:1985:BSM

[Bis85] N. N. Biswas. On bit steering in the minimization of the control memory of microprogrammed processors. IEEE Transactions on Computers, C-34(11):1057–1061, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676539.

Barros:1983:EAS

[BJ83] J. C. Barros and B. W. Johnson. Equivalence of the arbiter, the synchronizer, the latch, and the inertial delay. IEEE Transactions on Computers, C-32(7):603–614, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676292.

Bruno:1980:DSP

[BJS80] J. Bruno, J. W. Jones III, and Kimming So. Deterministic scheduling with pipelined processors. IEEE Transactions on Computers, C-29(4):308–316, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675569.

Book:1980:UDC

[BK80] R. V. Book and Sai Choi Kwan. On uniquely decipherable codes with two codewords. IEEE Transactions on Computers, C-29(4):324–325, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675571.

Brent:1982:RLP

[BK82] R. P. Brent and H. T. Kung. A regular layout for parallel adders. IEEE Transactions on Computers, C-31(3):260–264, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675982.

Brent:1984:SVA

[BK84] Richard P. Brent and H. T.

Kung. Systolic VLSI arrays for polynomial GCD computation. IEEE Transactions on Computers, C-33(8):731–736, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009358.

Berg:1987:SPM

[BK87] Menachem Berg and Israel Koren. On switching policies for modular redundancy faulttolerant computing systems. IEEE Transactions on Computers, C-36(9):1052–1062, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009536.

Bongiovanni:1980:MSF

[BL80] G. Bongiovanni and F. Luccio. Maintaining sorted files in a magnetic bubble memory. IEEE Transactions on Computers, C-29(10):855–863, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675468.

Bose:1984:PIC

[BL84] B. Bose and Der Jei Lin. PLA implementation of kout-of-n code TSC checker. IEEE Transactions on Computers, C-33(6):583–588, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676487.

Bose:1985:SUE

[BL85] B. Bose and Der Jei Lin. Systematic unidirectional errordetecting codes. IEEE Transactions on Computers, C-34 (11):1026–1032, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676535.

Bodnar:1989:MPA

[BL89] B. L. Bodnar and A. C. Liu. Modeling and performance analysis of single-bus tightly-coupled multiprocessors. IEEE Transactions on Computers, 38(3):464–470, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21134.

Blakely:1983:CAC

[Bla83]  G. R. Blakely.  A com-

puter algorithm for calculating the product AB modulo M. IEEE Transactions on Computers, C-32(5):497–500, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676262.

See comments [Slo85].

Blaum:1988:SUB

[Bla88] M. Blaum. Systematic unidirectional burst detecting codes. IEEE Transactions on Computers, 37(4):453–457, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2190.

Blelloch:1989:SPP

[Ble89]  G. E. Blelloch.  Scans as

primitive parallel operations. IEEE Transactions on Computers, 38(11):1526–1538, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42122.

Bloom:1988:CTW

[Blo88] B. Bloom. Constructing twowriter atomic registers. IEEE Transactions on Computers, 37 (12):1506–1514, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9729.

Bonuccelli:1984:ESV

[BLP84] M. A. Bonuccelli, E. Lodi, and L. Pagli. External sorting in VLSI. IEEE Transactions on Computers, C-33 (10):931–934, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676356.

Baccelli:1986:API

[BM86a] F. Baccelli and P. Mussi. An asynchronous parallel interpreter for arithmetic expressions and its evaluation. IEEE Transactions on Computers, C-35(3):245–256, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676748.

Bardell:1986:PAB

[BM86b] P. H. Bardell and W. H. McAnney. Pseudorandom arrays for built-in tests. IEEE Transactions on Computers, C-35(7):653–658, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676810.

Burkhart:1989:PMT

[BM89] H. Burkhart and R. Millen. Performance-measurement tools in a multiprocessor environment. IEEE Transactions on Computers, 38(5):725–737, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24274.

Bisbee:1988:FDB

[BN88] C. R. Bisbee and V. P. Nelson. Failure dependent bandwidth in shuffle-exchange networks. IEEE Transactions on Computers, 37(7):853–858, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2231.

Bozorgui-Nesbat:1986:LOD

[BNM86] S. Bozorgui-Nesbat and E. J.

McCluskey. Lower overhead design for testability of programmable logic arrays. IEEE Transactions on Computers, C-35(4):379–383, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676775.

Bochmann:1982:HST

[Boc82] G. V. Bochmann. Hardware specification with temporal logic: An example. IEEE Transactions on Computers, C-31(3):223–231, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675978.

Bokhari:1981:MP

[Bok81] S. H. Bokhari. On the mapping problem. IEEE Transactions on Computers, C-30(3):207–214, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675756.

Bokhari:1984:FMA

[Bok84] S. H. Bokhari. Finding maximum on an array processor with a global bus. IEEE Transactions on Computers, C-33(2):133–139, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676405.

Bokhari:1988:PPP

[Bok88] S. H. Bokhari. Partitioning problems in parallel, pipeline, and distributed computing. IEEE Transactions on Computers, 37(1):48–57, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=75137.

Bongiovanni:1983:TVS

[Bon83a] G. Bongiovanni. Two VLSI structures for the Discrete Fourier Transform. IEEE Transactions on Computers, C-32(8):750–754, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676314.

Bongiovanni:1983:VNV

[Bon83b] G. Bongiovanni. A VLSI network for variable size FFT’s. IEEE Transactions on Computers, C-32(8):756–760, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676316.

Booth:1981:MRM

[Boo81] T. L. Booth. In memoriam:

Richard E. Merwin, President of the IEEE Computer Society. IEEE Transactions on Computers, C-30(12):909, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675728.

Bose:1986:BUE

[Bos86]  B. Bose.  Burst unidirec-

tional error-detecting codes. IEEE Transactions on Computers, C-35(4):350–353, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676768.

Bose:1988:NTE

[Bos88]  P. Bose.  A novel tech-

nique for efficient parallel implementation of a classical logic/fault simulation problem.

IEEE Transactions on Computers, 37(12):1569–1577, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9734.

Boute:1989:RDS

[Bou89]  R. T. Boute.  Representa-

tional and denotational semantics of digital systems. IEEE Transactions on Computers, 38(7):986–999, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30850.

Bose:1982:OUE

[BP82] B. Bose and D. K. Pradhan. Optimal unidirectional error detecting/correcting codes. IEEE Transactions on Computers, C-31(6):564–568, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676043.

Bilardi:1984:ABS

[BP84] Gianfranco Bilardi and Franco P. Preparata. An architecture for bitonic sorting with optimal VLSI performance. IEEE Transactions on Computers, C-33(7):646–651, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009338.

Bilardi:1985:MAV

[BP85] Gianfranco Bilardi and Franco P.

Preparata. A minimum area VLSI network for O(log n) time sorting. IEEE Transactions on Computers, C-34(4):336–343, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009384.

Badr:1989:OSP

[BP89] H. G. Badr and S. Podar. An optimal shortestpath routing policy for network computers with regular mesh-connected topologies. IEEE Transactions on Computers, 38(10):1362–1371, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35831.

Bridges:1986:DSA

[BPM^{+}86] G. E. Bridges, W. Pries, R. D.

McLeod, M. Yunik, P. G. Gulak, and H. C. Card. Dual systolic architectures for VLSI digital signal processing systems. IEEE Transactions on Computers, C-35(10):916–923, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676684.

Baudet:1983:ATO

[BPV83] G. M. Baudet, F. P. Preparata, and J. E. Vuillemin. Area–time optimal VLSI circuits for convolution. IEEE Transactions on Computers, C-32(7):684–688, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676300.

Bose:1980:SCS

[BR80] B. Bose and T. R. N. Rao. Separating and completely separating systems and linear codes. IEEE Transactions on Computers, C-29(7):665–668, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675640.

Bose:1982:TUE

[BR82] B. Bose and T. R. N. Rao. Theory of unidirectional error correcting/detecting codes. IEEE Transactions on Computers, C-31(6):521–530, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676034.

Bazelow:1983:MSO

[BR83a] A. R. Bazelow and J. Raamot. On the microprocessor solution of ordinary differential equations using integer arithmetic. IEEE Transactions on Computers, C-32(2):204–207, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676207.

Bochmann:1983:SSC

[BR83b] G. V. Bochmann and M. Raynal. Structured specification of communicating systems. IEEE Transactions on Computers, C-32(2):120–133, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676197.

Bose:1984:UEC

[BR84] B. Bose and T. R. N. Rao.

Unidirectional error codes for shift-register memories. IEEE Transactions on Computers, C-33(6):575–578, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676485.

Bokhari:1986:RDC

[BR86] S. H. Bokhari and A. D.

Raza. Reducing the diameters of computer networks. IEEE Transactions on Computers, C-35(8):757–761, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676827.

Badrinath:1988:STO

[BR88] B. R. Badrinath and K. Ramamritham. Synchronizing transactions on objects. IEEE Transactions on Computers, 37(5):541– 547, May 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        4605.

Brand:1983:RDC

[Bra83a] D. Brand. Redundancy and don’t cares in logic synthesis. IEEE Transactions on Computers, C-32(10):947–952, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676139.

Brandwajn:1983:MDS

[Bra83b] A. Brandwajn. Models of DASD subsystems with multiple access paths: a throughput-driven approach. IEEE Transactions on Computers, C-32(5):451–463, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676256.

Brand:1986:DSP

[Bra86] D. Brand. Detecting sneak paths in transistor networks. IEEE Transactions on Computers, C-35(3):274–278, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676750.

Bernstein:1989:CSP

[BRG89] D. Bernstein, M. Rodeh, and I. Gertner. On the complexity of scheduling problems for parallel/pipelined machines. IEEE Transactions on Computers, 38(9):1308–1313, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29469.

Bryant:1984:SLM

[Bry84]  R. E. Bryant.  A switch-

level model and simulator for MOS digital systems. IEEE Transactions on Computers, C-33(2):160–177, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676408.

Bryant:1986:GBA

[Bry86] R. E. Bryant. Graph-based algorithms for Boolean function manipulation. IEEE Transactions on Computers, C-35(8):677–691, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676819.

Bellon:1982:PAE

[BS82] C. Bellon and G. Saucier. Protection against external errors in a dedicated system. IEEE Transactions on Computers, C-31(4):311–317, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675997.

Bowyer:1983:OCE

[BS83] K. W. Bowyer and C. F. Starmer. Optimizing contiguouselement region selection for virtual memory systems. IEEE Transactions on Computers, C-32(12):1201–1203, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676185.

Baru:1986:ASD

[BS86] C. K. Baru and S. Y. W. Su. The architecture of SM3: a dynamically partitionable multicomputer system. IEEE Transactions on Computers, C-35(9):790–802, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676839.

Bianchini:1987:ITS

[BS87a] R. P. Bianchini, Jr. and J. P. Shen. Interprocessor traffic scheduling algorithm for multiple-processor networks. IEEE Transactions on Computers, C-36(4):396–409, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676922.

Brzozowski:1987:CTS

[BS87b] Janusz A. Brzozowski and Carl-Johan Seger. A characterization of ternary simulation of gate networks. IEEE Transactions on Computers, C-36 (11):1318–1327, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009471.

Brosnan:1988:MED

[BS88] T. J. Brosnan and N. R.

Strader II. Modular error detection for bit-serial multiplication. IEEE Transactions on Computers, 37(9):1043–1052, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2255.

Bhasker:1989:ASR

[BS89a] J. Bhasker and S. Sahni. Via assignment in single-row routing. IEEE Transactions on Computers, 38(1):142–148, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8737.

Bhattacharya:1989:DPT

[BS89b] B. B. Bhattacharya and S. C. Seth. Design of parity testable combinational circuits. IEEE Transactions on Computers, 38(11):1580–1584, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=42129.

Bohm:1989:COT

[BS89c] A. P. W. Bohm and J. Sargeant. Code optimization for taggedtoken dataflow machines. IEEE Transactions on Computers, 38 (1):4–14, January 1989. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        8726.

Barzilai:1981:WSS

[BSMS81] Z. Barzilai, J. Savir, G. Markowsky, and M. G. Smith. The weighted syndrome sums approach to VLSI testing. IEEE Transactions on Computers, C-30(12):996–1000, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675744.

Bashir:1983:SSP

[BSV83] A. F. Bashir, V. Susarla, and K. Vairavan. A statistical study of the performance of a task scheduling algorithm.

IEEE Transactions on Computers, C-32(8):774–777, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676321.

Bobbio:1986:ATT

[BT86] A. Bobbio and K. S. Trivedi.

An aggregation technique for the transient analysis of stiff Markov chains. IEEE Transactions on Computers, C-35 (9):803–814, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676840.

Blake:1989:MIN

[BT89] J. T. Blake and K. S. Trivedi.

Multistage interconnection network reliability. IEEE Transactions on Computers, 38(11): 1600–1604, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42134.

Burkowski:1982:HHS

[Bur82] Forbes J. Burkowski. A hardware hashing scheme in the design of a multiterm string comparator. IEEE Transactions on Computers, C-31 (9):825–834, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676098.

See correction [Bur84].

Burkowski:1984:CHH

[Bur84]  F. J. Burkowski.  Correc-

tion to “A Hardware Hashing Scheme in the Design of a Multiterm String Comparator”. IEEE Transactions on Computers, C-33(4):375, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676447.

See [Bur82].

Burton:1988:SMV

[Bur88] F. W. Burton. Storage management in virtual tree machines. IEEE Transactions on Computers, 37(3):321–328, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2169.

Bushard:1983:MTS

[Bus83]  L. B. Bushard.  A min-

imum table size result for higher radix nonrestoring division. IEEE Transactions on Computers, C-32(6):521–526, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676273.

Butler:1981:SEC

[But81] J. T. Butler. Speed-efficiency-complexity tradeoffs in universal diagnosis algorithms. IEEE Transactions on Computers, C-30(8):590–596, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675845.

Bux:1983:ALA

[Bux83] W. Bux. Analysis of a localarea bus system with controlled access. IEEE Transactions on Computers, C-32(8):760–763, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676317.

Blaum:1989:ECA

[BV89] M. Blaum and H. Van Tilborg.

On t-error correcting/all unidirectional error detecting codes. IEEE Transactions on Computers, 38(11):1493–1501, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42121.

Bandeira:1983:TCA

[BVH83] N. Bandeira, K. Vaccaro, and J. A. Howard. A two’s complement array multiplier using true values of the operands. IEEE Transactions on Computers, C-32(8):745–747, August 1983. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676312.

Bentley:1980:OWC

[BW80] J. L. Bentley and D. Wood. An optimal worst case algorithm for reporting intersections of rectangles. IEEE Transactions on Computers, C-29(7):571–577, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675628.

Bongiovanni:1981:TSM

[BW81a] G. Bongiovanni and C. K. Wong. Tree search in major/minor loop magnetic bubble memories. IEEE Transactions on Computers, C-30(8):537–545, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675838.

Butler:1981:GEC

[BW81b] J. T. Butler and A. S. Wojcik. Guest Editors’ comments. IEEE Transactions on Computers, C-30(9):617–618, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675859.

Barbour:1989:GCA

[BW89a] A. E. Barbour and A. S. Wojcik. A general constructive approach to fault-tolerant design using redundancy. IEEE Transactions on Computers, 38(1):15–29, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8727.

Baumgartner:1989:GLB

[BW89b] K. M. Baumgartner and B. W.

Wah. GAMMON: a load balancing strategy for local computer systems with multiaccess networks. IEEE Transactions on Computers, 38(8):1098–1109, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30865.

Baba:1987:DFD

[BYH87] T. Baba, S. B. Yao, and A. R.

Hevner. Design of a functionally distributed, multiprocessor database machine using data flow analysis. IEEE Transactions on Computers, C-36(6):650–666, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676959.

Barr:1986:SMM

[BZV86] M. H. Abd-El Barr, S. G. Zaky, and Z. G. Vranesic. Synthesis of multivalued multithreshold functions for CCD implementation. IEEE Transactions on Computers, C-35(2):124–133, February 1986. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676730.

Chen:1980:ODD

[CA80] P. P.-S. Chen and J. Akoka. Optimal design of distributed information systems. IEEE Transactions on Computers, C-29(12):1068–1080, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675512.

Chou:1983:LRU

[CA83] T. C. K. Chou and J. A. Abraham. Load redistribution under failure in distributed systems. IEEE Transactions on Computers, C-32(9):799–808, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676329.

Chou:1986:DCC

[CA86] Timothy C. K. Chou and Jacob A. Abraham. Distributed control of computer systems. IEEE Transactions on Computers, C-35(6):564–567, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009433.

Chu:1989:TDS

[CA89] S.-C. Chu and J. R. Armstrong. t-TDA-diagnosable systems. IEEE Transactions on Computers, 38(6):914–920, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24306.

Calahan:1988:AVS

[Cal88] D. A. Calahan. An analysis of vector startup access delays. IEEE Transactions on Computers, 37(9):1134–1137, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2266.

Cantarella:1983:RPR

[Can83] R. G. Cantarella. The reliability of periodically repaired n − 1/n parallel redundant systems. IEEE Transactions on Computers, C-32(6):597–598, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676284.

See comments [Nil84, Wen85].

Carlson:1983:TST

[Car83] D. A. Carlson. Time-space tradeoffs on back-to-back FFT algorithms. IEEE Transactions on Computers, C-32(6):585–589, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676281.

Caruso:1984:LSA

[Car84] Giuseppe Caruso. A local selection algorithm for switching function minimization. IEEE Transactions on Computers, C-33(1):91–97, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009316.

Carlson:1988:MMC

[Car88] D. A. Carlson. Modified-mesh connected parallel computers. IEEE Transactions on Computers, 37(10):1315–1321, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5998.

Casey:1986:CDR

[Cas86] L. M. Casey. Comments on “The Design of a Reliable Remote Procedure Call Mechanism”. IEEE Transactions on Computers, C-35(3):283–284, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676753.

See [SP82].

Calvo:1986:AMA

[CAV86] J. Calvo, J. I. Acha, and M. Valencia. Asynchronous modular arbiter. IEEE Transactions on Computers, C-35(1):67–70, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676660.

Chalasani:1989:DTV

[CBAP89] P. R. Chalasani, S. Bhawmik, A. Acharya, and P. Palchaudhuri. Design of testable VLSI circuits with minimum area overhead. IEEE Transactions on Computers, 38 (10):1460–1462, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35841.

Chan:1988:ERG

[CC88] M. Y. Chan and F. Y. L.

Chin. On embedding rectangular grids in hypercubes. IEEE Transactions on Computers, 37(10):1285–1288, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5991.

Ciciani:1989:CDE

[CC89] B. Ciciani and G. Cantone. Comments on “Design and evaluation of a fault-tolerant multiprocessor using hardware recovery blocks” by Y.-H. Lee and K. G. Shin. IEEE Transactions on Computers, 38(9):1336–1337, September 1989. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29476.

See [LS84].

Calderbank:1989:NEA

[CCF89] A. R. Calderbank, E. G. Coffman, Jr., and L. Flatto. A note extending the analysis of two-head disk systems to more general seek-time characteristics. IEEE Transactions on Computers, 38(11):1584–1586, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=42130.

Carlsson:1985:INB

[CCSW85] G. E. Carlsson, J. E. Cruthirds, H. B. Sexton, and C. G. Wright. Interconnection networks based on a generalization of cube-connected cycles. IEEE Transactions on Computers, C-34(8):769–772, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676627.

Char:1988:DFT

[CCWZ88] J. M. Char, V. Cherkassy, H. Wechsler, and G. L. Zimmerman. Distributed and faulttolerant computation for retrieval tasks using distributed associative memories. IEEE Transactions on Computers, 37 (4):484–490, April 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
2196.

Chen:1986:MSF

[CD86] Chuen-Liang Chen and Min-Wen Du. Multiple stuckfault detection and location in multivalued linear circuits. IEEE Transactions on Computers, C-35(12):1068–1071, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676714.

Chang:1988:LAP

[CD88] K. C. Chang and H. C.

Du. Layer assignment problem for three-layer routing. IEEE Transactions on Computers, 37(5):625–632, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4616.

Conway:1989:MVA

[CdL89] A. E. Conway, E. de Souza e Silva, and S. S. Lavenberg. Mean value analysis by chain of product form queueing networks. IEEE Transactions on Computers, 38(3):432–442, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21129.

Chazelle:1987:IAC

[CE87] Bernard Chazelle and Herbert Edelsbrunner. An improved algorithm for constructing kthorder Voronoi diagrams. IEEE Transactions on Computers, C-36(11):1349–1354, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009474.

Chin:1980:FSA

[CF80a] F. Y. Chin and K. S. Fok. Fast sorting algorithms on uniform ladders (multiple shift-register loops). IEEE Transactions on Computers, C-29(7):618–631, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675633.

Corsini:1980:CPM

[CF80b] P. Corsini and G. Frosini. Correction to “Properties of the Multidimensional Generalized Discrete Fourier Transform”. IEEE Transactions on Computers, C-29(8):759, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675668.

Chu:1981:ATQ

[CFH81] W. W. Chu, G. Fayolle, and D. G. Hibbits. An analysis of a tandem queueing system for flow control in computer networks. IEEE Transactions on Computers, C-30(5):318–324, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675789.

Carchiolo:1986:LSP

[CFM^{+}86] V. Carchiolo, A. Faro, O. Mirabella, G. Pappalardo, and G. Scollo. A LOTOS specification of the PROWAY highway service. IEEE Transactions on Computers, C-35(11):949–968, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676697.

Chang:1989:DRM

[CFP89] M.-F. Chang, W. K. Fuchs, and J. H. Patel. Diagnosis and repair of memory with coupling faults. IEEE Transactions on Computers, 38(4):493–500, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21142.

Chlamtac:1987:PMA

[CG87] I. Chlamtac and O. Ganz.

Performance models of asynchronous multitrunk HYPERchannel networks. IEEE Transactions on Computers, C-36 (2):138–146, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676876.

Cerny:1988:FDC

[CG88] E. Cerny and J. Gecesi. Functional description of connectorswitch-attenuator networks. IEEE Transactions on Computers, 37(1):111–114, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=75142.

Card:1987:CMV

[CGMP87] Howard C. Card, P. Glenn Gulak, Robert D. McLeod, and Werner Pries. (λT) complexity measures for VLSI computations in constant chip area. IEEE Transactions on Computers, C-36(1):112–117, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009456.

Chwa:1981:FID

[CH81] Kyung-Yong Chwa and S. L. Hakimi. On fault identification in diagnosable systems. IEEE Transactions on Computers, C-30(6):414–422, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675807.

Chen:1982:CUL

[CH82a] X. Chen and S. L. Hurst. A comparison of universal-logic-module realizations and their application in the synthesis of combinatorial and sequential logic networks. IEEE Transactions on Computers, C-31(2):140–147, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675959.

Chu:1982:OQP

[CH82b] W. W. Chu and P. Hurley.

Optimal query processing for distributed database systems. IEEE Transactions on Computers, C-31(9):835–850, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676099.

Chu:1983:COQ

[CH83] W. W. Chu and P. Hurley. Correction to “Optimal Query Processing for Distributed Database Systems”. IEEE Transactions on Computers, C-32(9):878, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676340.

Chin:1984:PSN

[CH84] Chi-Yuan Chin and Kai Hwang.

Packet switching networks for multiprocessors and data flow computers. IEEE Transactions on Computers, C-33 (11):991–1003, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676373.

Chlamtac:1985:SXP

[CH85a] M. Chlamtac and I. Harary. The shift X parity watch algorithm for raster scan displays. IEEE Transactions on Computers, C-34(7):666–673, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676606.

Chu:1985:EWA

[CH85b] Wesley W. Chu and Joseph Hellerstein. The exclusive-writer approach to updating replicated files in distributed processing systems. IEEE Transactions on Computers, C-34(6):489–500, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009401.

Chakravarty:1989:NDS

[CH89] S. Chakravarty and H. B. Hunt III. A note on detecting sneak paths in transistor networks. IEEE Transactions on Computers, 38(6):861–864, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24295.

Chang:1980:MRS

[Cha80] S. S. L. Chang. Multiple-read single-write memory and its applications. IEEE Transactions on Computers, C-29(8):689–694, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675650.

Chaney:1983:MFF

[Cha83a] T. J. Chaney. Measured flip-flop responses to marginal triggering. IEEE Transactions on Computers, C-32(12):1207–1209, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676187.

Chazelle:1983:BLP

[Cha83b] B. Chazelle. The bottom-left bin-packing heuristic: An efficient implementation. IEEE Transactions on Computers, C-32(8):697–707, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676307.

Chazelle:1984:CGS

[Cha84] B. Chazelle. Computational geometry on a systolic chip. IEEE Transactions on Computers, C-33(9):774–785, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676494.

Chan:1987:UDT

[Cha87a] A. H. Chan. Using decision trees to derive the complement of a binary function with multiple-valued inputs. IEEE Transactions on Computers, C-36(2):212–214, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676883.

Chapiro:1987:RHS

[Cha87b] D. M. Chapiro. Reliable high-speed arbitration and synchronization. IEEE Transactions on Computers, C-36 (10):1251–1255, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676867.

Chen:1982:DAS

[Che82a] C. C. Chen. A distributed algorithm for shortest paths. IEEE Transactions on Computers, C-31(9):898–899, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676107.

Chen:1982:AMA

[Che82b] Chung Ho Chen. An algebraic model of arithmetic codes. IEEE Transactions on Computers, C-31(4):318–321, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675998.

Cheung:1982:MEQ

[Che82c] To-Yat Cheung. A method for equijoin queries in distributed relational databases. IEEE Transactions on Computers, C-31(8):746–751, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676081.

Chen:1983:ECC

[Che83]  Chin-Long Chen.  Error-

correcting codes with byte errordetection capability. IEEE Transactions on Computers, C-32(7):615–621, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676293.

Chen:1986:BOE

[Che86a]  C. L. Chen.  Byte-oriented

error-correcting codes for semiconductor memory systems. IEEE Transactions on Computers, C-35(7):646–648, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676807.

Chen:1986:LDL

[Che86b] C. L. Chen. Linear dependencies in linear feedback shift registers. IEEE Transactions on Computers, C-35(12):1086–1088, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676718.

Chen:1988:ETP

[Che88a]  C. L. Chen.  Exhaustive

test pattern generation using cyclic codes. IEEE Transactions on Computers, 37 (2):225–228, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2152.

Chen:1988:GCM

[Che88b] M. C. Chen. The generation of a class of multipliers: synthesizing highly parallel algorithms in VLSI. IEEE Transactions on Computers, 37(3):329–338, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2170.

Cherkassky:1988:PEN

[Che88c]  V. S. Cherkassky.  Perfor-

mance evaluation of nonrectangular multistage interconnection networks. IEEE Transactions on Computers, 37 (10):1269–1272, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5987.

Cohen:1983:CCP

[CHH83] Marty S. Cohen, T. E. Hull, and V. Carl Hamacher. CADAC: a controlled-precision decimal arithmetic unit. IEEE Transactions on Computers, C-32 (4):370–377, April 1983. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1676238.

Chu:1983:RCA

[CHL83] W. W. Chu, W. Haller, and K. K. Leung. Reservation channel access protocol for high speed local networks with star configurations. IEEE Transactions on Computers, C-32(8):763–766, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676318.

Coraor:1987:GMM

[CHM87] L. D. Coraor, P. T. Hulina, and O. A. Morean. A general model for memory-based finite-state machines. IEEE Transactions on Computers, C-36(2):175–184, February 1987.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676880.

Chakravarty:1989:CGM

[CHRR89] S. Chakravarty, H. B. Hunt III, S. S. Ravi, and D. J. Rosenkrantz. The complexity of generating minimum test sets for PLA’s and monotone combinational circuits. IEEE Transactions on Computers, 38(6):865–869, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24296.

Chu:1980:I

[Chu80]  W. W. Chu.  Introduction.

IEEE Transactions on Computers, C-29(12):1037–1038, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675508.

Chughtai:1985:CBS

[Chu85] M. Ashraf Chughtai. Complete binary spanning trees of the eight nearest neighbor array. IEEE Transactions on Computers, C-34(6):547–549, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009406.

Chillarege:1987:MBA

[CI87] R. Chillarege and R. K. Iyer.

Measurement-based analysis of error latency. IEEE Transactions on Computers, C-36(5):529–537, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676937.

Chandrasekharan:1988:NAR

[CI88] N. Chandrasekharan and S. S.

Iyengar. NC algorithms for recognizing chordal graphs and k trees. IEEE Transactions on Computers, 37(10):1178–1183, October 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5979.

Chillarege:1989:ESM

[CI89] R. Chillarege and R. K. Iyer. An experimental study of memory fault latency. IEEE Transactions on Computers, 38(6):869–874, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24297.

Chang:1988:STI

[CICR88] J. H. Chang, O. H. Ibarra, M. J.

Chung, and K. K. Rao. Systolic tree implementation of data structures. IEEE Transactions on Computers, 37(6):727–735, June 1988. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2211.

Chang:1987:PPO

[CIP87] Jik H. Chang, Oscar H. Ibarra, and Michael A. Palis. Parallel parsing on a one-way array of finite-state machines. IEEE Transactions on Computers, C-36(1):64–75, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009449.

Chang:1989:BSK

[CJ89] C. K. Chang and T. M. Jiang. A binary single-key-lock system for access control. IEEE Transactions on Computers, 38(10):1462–1466, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35842.

Chlamtac:1987:TBB

[CK87] I. Chlamtac and S. Kutten. Tree-based broadcasting in multihop radio networks. IEEE Transactions on Computers, C-36(10):1209–1223, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676861.

Craig:1988:TSC

[CKS88] G. L. Craig, C. R. Kine, and K. K. Saluja. Test scheduling and control for VLSI builtin self-test. IEEE Transactions on Computers, 37(9): 1099–1109, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2260.

Crouzet:1980:DSC

[CL80] Y. Crouzet and C. Landrault.

Design of self-checking MOS– LSI circuits: Application to a four-bit microprocessor. IEEE Transactions on Computers, C-29(6):532–537, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675615.

Chen:1984:IAS

[CL84] A. L. P. Chen and V. O. K.

Li. Improvement algorithms for semijoin query processing programs in distributed database systems. IEEE Transactions on Computers, C-33 (11):959–967, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676370.

Chu:1987:TAP

[CL87] W. W. Chu and L. M.- T. Lan. Task allocation and precedence relations for distributed real-time systems. IEEE Transactions on Computers, C-36(6):667–679, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676960.

Cheng:1988:PAS

[CL88] W.-Y. Cheng and J. W. S. Liu. Performance of ARQ schemes on token ring networks. IEEE Transactions on Computers, 37(7):826–834, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2228.

Cleary:1984:CHT

[Cle84]  John G. Cleary.  Compact

hash tables using bidirectional linear probing. IEEE Transactions on Computers, C-33 (9):828–834, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676499.

Chu:1984:EIC

[CLH84] Wesley W. Chu, Min-Tsung Lan, and Joseph Hellerstein. Estimation of intermodule communication (IMC) and its applications in distributed processing systems. IEEE Transactions on Computers, C-33(8):691–699, August 1984. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009355.

Cliff:1980:ATV

[Cli80a] R. A. Cliff. Acceptable testing of VLSI components which contain error correctors. IEEE Transactions on Computers, C-29(2):125–134, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675536.

Cliff:1980:DMA

[Cli80b] R. A. Cliff. Digital multiplexing of analog data in a microprocessor controlled data acquisition system. IEEE Transactions on Computers, C-29 (2):200–202, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675544.

Clark:1981:MSH

[CLP81] D. W. Clark, B. W. Lampson, and K. A. Pier. The memory system of a highperformance personal computer. IEEE Transactions on Computers, C-30(10):715–733, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675691.

Chung:1980:TSS

[CLW80a] K. M. Chung, F. Luccio, and C. K. Wong. A tree storage scheme for magnetic bubble memories. IEEE Transactions on Computers, C-29 (10):864–874, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675469.

Chung:1980:CSM

[CLW80b] Kin-Man Chung, F. Luccio, and C. K. Wong. On the complexity of sorting in magnetic bubble memory systems. IEEE Transactions on Computers, C-29(7):553–563, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675626.

Lin:1988:BNS

[cLW88] Ferng ching Lin and I-Chen Wu. Broadcast normalization in systolic design. IEEE Transactions on Computers, 37(11):1428–1434, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8711.

Cherkassky:1985:PPR

[CM85] Vladimir Cherkassky and Miroslaw Malek. On permuting properties of regular rectangular SWbanyans. IEEE Transactions on Computers, C-34(6):542–546, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009405.

Chin:1987:TLP

[CM87a] C. K. Chin and E. J. Mc-Cluskey. Test length for pseudorandom testing. IEEE Transactions on Computers, C-36 (2):252–256, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676892.

Conterno:1987:AMC

[CM87b] R. Conterno and R. Melen.

An analytical model for a class of processor-memory interconnection networks. IEEE Transactions on Computers, C-36(11):1374–1378, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009480.

Cutler:1987:DMS

[CM87c] R. B. Cutler and S. Muroga.

Derivation of minimal sums for completely specified functions. IEEE Transactions on Computers, C-36(3):277–292, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676900.

Capello:1988:SSS

[CM88a] P. R. Capello and W. L. Miranker. Systolic super summation. IEEE Transactions on Computers, 37(6):657–677, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2205.

Choi:1988:FTF

[CM88b] Y.-H. Choi and M. Malek. A fault-tolerant FFT processor. IEEE Transactions on Computers, 37(5):617–621, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4614.

Choi:1988:FTS

[CM88c] Y.-H. Choi and M. Malek. A fault-tolerant systolic sorter. IEEE Transactions on Computers, 37(5):621–624, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4615.

Cherkassky:1989:PPP

[CM89] V. Cherkassky and M. Malek. Partitioning and permuting properties of CC-banyan networks. IEEE Transactions on Computers, 38(2):274–278, February 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
16504.

Chiola:1988:PFS

[CMB88] G. Chiola, M. A. Marsan, and G. Balbo. Product-form solution techniques for the performance analysis of multiplebus multiprocessor systems with nonuniform memory references. IEEE Transactions on Computers, 37(5):532–540, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4604.

Coulas:1987:RHR

[CMM87] Michael F. Coulas, Glenn H.

MacEwen, and Genevieve Marquis. RNet: a hard real-time distributed programming system. IEEE Transactions on Computers, C-36(8):917–932, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009515.

Castillo:1982:DCT

[CMS82] X. Castillo, S. R. McConnel, and D. P. Siewiorek. Derivation and calibration of a transient error reliability model. IEEE Transactions on Computers, C-31(7):658–671, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676063.

Clarke:1982:DRS

[CN82] E. M. Clarke and C. N. Nikolaou. Distributed reconfiguration strategies for faulttolerant multiprocessor systems. IEEE Transactions on Computers, C-31(8):771–784, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676083.

Colwell:1988:VAT

[CNO^{+}88] R. P. Colwell, R. P. Nix, J. J. O’Donnell, D. B. Papworth, and P. K. Rodman. A VLIW architecture for a trace scheduling compiler. IEEE Transactions on Computers, 37(8):967–979, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2247.

Coan:1988:CIF

[Coa88]  B. A. Coan.  A compiler

that increases the fault tolerance of asynchronous protocols. IEEE Transactions on Computers, 37(12):1541–1553, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=9732.

Cohn:1985:CSL

[Coh85] Martin Cohn. Counting sequences with large local distance. IEEE Transactions on Computers, C-34(6):562, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009410.

Concepcion:1989:HCA

[Con89]  A. I. Concepcion.  A hi-

erarchical computer architecture for distributed simulation. IEEE Transactions on Computers, 38(2):311–319, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16512.

Cook:1989:EAL

[Coo89] R. P. Cook. An empirical analysis of the Lilith instruction set. IEEE Transactions on Computers, 38(1):156–158, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8740.

Cosentino:1988:FTS

[Cos88] R. J. Cosentino. Fault tolerance in a systolic residue arithmetic processor array. IEEE Transactions on Computers, 37(7):886– 890, July 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
2239.

Coy:1980:RNC

[Coy80] W. Coy. A remark on the nonminimality of certain multiple fault detection algorithms. IEEE Transactions on Computers, C-29(8):757–759, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675666.

Ceri:1982:AOD

[CP82a] S. Ceri and G. Pelagatti. Allocation of operations in distributed database access. IEEE Transactions on Computers, C-31(2):119–129, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675957.

Chien:1982:SIC

[CP82b] Y. T. Chien and E. A. Parrish. Special issue on computer architecture for pattern analysis and image database management. IEEE Transactions on Computers, C-31(10):921–922, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675900.

Cheng:1987:MTS

[CP87a] Wu-Tung Cheng and J. H. Patel. A minimum test set for multiple fault detection in ripple carry adders. IEEE Transactions on Computers, C-36(7):891–895, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676985.

Chlamtac:1987:DNO

[CP87b] I. Chlamtac and S. S. Pinter. Distributed nodes organization algorithm for channel access in a multihop dynamic radio network. IEEE Transactions on Computers, C-36(6):728–737, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676965.

Chaudhuri:1983:SOP

[CPG83] R. Chaudhuri, Son Pham, and O. N. Garcia. Solution of an open problem on probabilistic grammars. IEEE Transactions on Computers, C-32(8):748–750, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676313.

Crammond:1985:CSU

[Cra85] Jim Crammond. A comparative study of unification algorithms for OR-parallel execution of logic languages. IEEE Transactions on Computers, C-34(10): 911–917, October 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). Compares directory trees, hash windows, and variable importation.

Crist:1980:SCL

[Cri80] S. C. Crist. Synthesis of combinational logic using decomposition and probability. IEEE Transactions on Computers, C-29(11):1013–1016, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675497.

Cristian:1982:EHS

[Cri82] F. Cristian. Exception handling and software fault tolerance. IEEE Transactions on Computers, C-31(6):531–540, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676035.

Chanson:1980:OMH

[CS80a] S. T. Chanson and P. S. Sinha. Optimization of memory hierarchies in multiprogrammed computer systems with fixed cost constraint. IEEE Transactions on Computers, C-29(7):611–618, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675632.

Chu:1980:HRF

[CS80b] W. W. Chu and M. Y.-C. Shen. A hierarchical routing and flow control policy (HRFC) for packet switched networks. IEEE Transactions on Computers, C-29(11):971–977, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675491.

Cuny:1984:TCP

[CS84] J. E. Cuny and L. Snyder. Testing the coordination predicate. IEEE Transactions on Computers, C-33(3):201–208, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676417.

Calzarossa:1985:CVT

[CS85] M. Calzarossa and G. Serazzi. A characterization of the variation in time of workload arrival patterns. IEEE Transactions on Computers, C-34(2):156–162, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676552.

Chan:1986:MAH

[CS86a] T. F. Chan and Y. Saad. Multigrid algorithms on the hypercube multiprocessor. IEEE Transactions on Computers, C-35(11):969–977, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676698.

Cheung:1986:SSC

[CS86b] Tony Cheung and James E. Smith. A simulation study of the CRAY X-MP memory system. IEEE Transactions on Computers, C-35(7):613–622, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676802.

Ciminiera:1986:CNF

[CS86c] L. Ciminiera and A. Serra. A connecting network with fault tolerance capabilities. IEEE Transactions on Computers, C-35(6):578–580, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009436.

Chen:1987:PAC

[CS87] Ming-Syan Chen and Kang G. Shin. Processor allocation in an N-cube multiprocessor using Gray codes. IEEE Transactions on Computers, C-36 (12):1396–1407, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009493.

Cidon:1989:DAA

[CS89] I. Cidon and M. Sidi. Distributed assignment algorithms for multihop packet radio networks. IEEE Transactions on Computers, 38(10):1353–1361, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35830.

Chakraborti:1986:IMR

[CSR86] N. B. Chakraborti, J. S.

Soundararajan, and A. L. N. Reddy. An implementation of mixed-radix conversion for residue number applications. IEEE Transactions on Computers, C-35(8):762–764, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676829;
 http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35258.

Carey:1984:EIS

[CT84a] M. J. Carey and C. D. Thompson. An efficient implementation of search trees on [lg N + 1] processors. IEEE Transactions on Computers, C-33 (11):1038–1041, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676379.

Christ:1984:VFP

[CT84b] N. H. Christ and A. E. Terrano. A very fast parallel processor. IEEE Transactions on Computers, C-33(4):344–350, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676439.

Cull:1980:TGD

[Cul80]  P. Cull.  Tours of graphs,

digraphs, and sequential machines. IEEE Transactions on Computers, C-29(1):50–54, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675456.

Current:1980:HDR

[Cur80a] K. W. Current. A high datarate digital output correlator design. IEEE Transactions on Computers, C-29(5):403–405, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675592.

Current:1980:HDI

[Cur80b] K. W. Current. High density integrated computing circuitry with multiple valued logic. IEEE Transactions on Computers, C-29(2):191–195, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675542.

Current:1980:PBP

[Cur80c] K. W. Current. Pipelined binary parallel counters employing latched quaternary logic full adders. IEEE Transactions on Computers, C-29(5):400–403, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675591.

Chiang:1983:TRC

[CV83] Kuang-Wei Chiang and Z. G.

Vranesic. A tree representation of combinational networks. IEEE Transactions on Computers, C-32(3):315–319, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676224.

Chiang:1984:CFD

[CV84] Kuang-Wei Chiang and Z. G.

Vranesic. Comments on “Fault Diagnosis of MOS Combinational Networks”. IEEE Transactions on Computers, C-33(10):947, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676361.

See [EZS82].

Cvetanovic:1987:EPP

[Cve87] Z. Cvetanovic. The effects of problem partitioning, allocation, and granularity on the performance of multiple-processor systems. IEEE Transactions on Computers, C-36(4):421–432, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676924.

Chow:1983:PDA

[CVY83] P. Chow, Z. Vranesic, and Jui Lin Yen. A pipelined distributed arithmetic PFFT processor. IEEE Transactions on Computers, C-32(12): 1128–1136, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676173.

See correction [Ano84f].

Chung:1980:CGC

[CW80] K. M. Chung and C. K. Wong. Construction of a generalized connector with 5.8n log_{2} n edges. IEEE Transactions on Computers, C-29(11):1029–1032, November 1980. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675501.

Chin:1983:OAI

[CW83] F. Chin and Cao An Wang. Optimal algorithms for the intersection and the minimum distance problems between planar polygons. IEEE Transactions on Computers, C-32 (12):1203–1207, December 1983.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676186.

Dadda:1980:CPC

[Dad80]  L. Dadda.  Composite par-

allel counters. IEEE Transactions on Computers, C-29 (10):942–946, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675481.

Dadda:1989:SIM

[Dad89]  L. Dadda.  On serial-input

multipliers for two’s complement numbers. IEEE Transactions on Computers, 38(9): 1341–1345, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29478.

Dahbura:1986:EAI

[Dah86] A. T. Dahbura. An efficient algorithm for identifying the most likely fault set in a probabilistically diagnosable system. IEEE Transactions on Computers, C-35(4):354–356, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676769.

Danielsson:1983:VLS

[Dan83] P. E. Danielsson. A variablelength shift-register. IEEE Transactions on Computers, C-32(11):1067–1069, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676159.

Danielsson:1984:SPC

[Dan84]  Per E. Danielsson.  Serial/

parallel convolvers. IEEE Transactions on Computers, C-33(7):652–667, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009339.

Dao:1981:SDN

[Dao81] T. T. Dao. SEC–DED nonbinary code for fault-tolerant byte-organized memory implemented with quaternary logic. IEEE Transactions on Computers, C-30(9):662–666, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675864.

David:1980:TFS

[Dav80a] R. David. Testing by feedback shift register. IEEE Transactions on Computers, C-29(7):668–673, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675641.

Davio:1980:ROM

[Dav80b] M. Davio. Read-only memory implementation of discrete functions. IEEE Transactions on Computers, C-29 (10):931–934, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675478.

David:1986:SAM

[Dav86] R. David. Signature analysis for multiple-output circuits. IEEE Transactions on Computers, C-35(9):830–837, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676843.

Davis:1989:LCH

[Dav89] I. J. Davis. Local correction of helix(k) lists. IEEE Transactions on Computers, 38(5):718–724, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24273.

Dubois:1982:ECC

[DB82] M. Dubois and F. A. Briggs.

Effects of cache coherency in multiprocessors. IEEE Transactions on Computers, C-31 (11):1083–1099, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675925.

Damper:1987:MTP

[DB87] R. I. Damper and N. Burgess. MOS test pattern generation using path algebras. IEEE Transactions on Computers, C-36(9):1123–1128, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009546.

Dervisoglu:1981:HPC

[DC81a] B. I. Dervisoglu and D. J. Criscione. A hard programmable control unit design using VLSI technology. IEEE Transactions on Computers, C-30(10):800–810, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675698.

Dormido:1981:SGP

[DC81b] S. Dormido and M. A. Canto. Synthesis of generalized parallel counters. IEEE Transactions on Computers, C-30 (9):699–703, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675870.

Dormido:1982:UBS

[DC82] S. Dormido and M. A. Canto.

An upper bound for the synthesis of generalized parallel counters. IEEE Transactions on Computers, C-31(8):802–805, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676088.

Deng:1987:DDT

[DC87] Robert H. Deng and Daniel J.

Costello. Decoding of DBEC– TBED Reed–Solomon codes. IEEE Transactions on Computers, C-36(11):1359–1363, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009476.

Davio:1981:SDF

[DD81] M. Davio and J.-P. Deschamps.

Synthesis of discrete functions using I^{2}L technology. IEEE Transactions on Computers, C-30(9):653–661, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675863.

Davis:1988:PRD

[DD88a] T. A. Davis and E. S. Davidson. Pairwise reduction for the direct, parallel solution of sparse, unsymmetric sets of linear equations. IEEE Transactions on Computers, 37(12):

1648–1654, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9742.

Dhawan:1988:DSCa

[DD88b] S. Dhawan and R. C. De Vries. Design of self-checking iterative networks. IEEE Transactions on Computers, 37(9): 1121–1125, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2263.

Dhawan:1988:DSCb

[DD88c] S. Dhawan and R. C. De Vries. Design of self-checking sequential machines. IEEE Transactions on Computers, 37 (10):1280–1284, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5989.

Dao:1980:CNA

[DDG80] T. T. Dao, M. Davio, and C. Gossart. Complex number arithmetic with odd-valued logic. IEEE Transactions on Computers, C-29(7):604–611, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675631.

DePrycker:1982:DMS

[De 82a] M. De Prycker. On the development of a measurement system for high level language program statistics. IEEE Transactions on Computers, C-31 (9):883–891, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676103.

DePrycker:1982:PAI

[De 82b] M. De Prycker. A performance analysis of the implementation of addressing methods in block-structured languages. IEEE Transactions on Computers, C-31(2):155–163, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675961.

DePrycker:1983:REI

[De 83] M. De Prycker. Representing the effect of instruction prefetch in a microprocessor performance model. IEEE Transactions on Computers, C-32 (9):868–872, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676337.

Deminet:1982:EMA

[Dem82] J. Deminet. Experience with multiprocessor algorithms. IEEE Transactions on Computers, C-31(4):278–288, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675993.

Demuth:1985:EDS

[Dem85] Howard B. Demuth. Electronic data sorting. IEEE Transactions on Computers, C-34(4):296–310, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009380.

Despain:1981:AR

[Des81] A. M. Despain. Author’s reply. IEEE Transactions on Computers, C-30(6):452, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675813.

Delorme:1984:LGG

[DF84] C. Delorme and G. Farhi. Large graphs with given degree and diameter—Part I. IEEE Transactions on Computers, C-33(9):857–860, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676504.

David:1989:RPT

[DFC89] R. David, A. Fuentes, and B. Courtois. Random pattern testing versus deterministic testing of RAMs. IEEE Transactions on Computers, 38(5):637–650, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24267.

DeSouzaeSilva:1986:CCO

[DG86] E. De Souza e Silva and H. R.

Gail. Calculating cumulative operational time distributions of repairable computer systems. IEEE Transactions on Computers, C-35(4):322–332, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676765.

Dewan:1989:MCL

[DG89] R. M. Dewan and B. Gavish. Models for the combined logical and physical design of databases. IEEE Transactions on Computers, 38(7):955–967, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30848.

Doty:1980:MBM

[DGS80] K. L. Doty, J. D. Greenblatt, and Stanley Y. W. Su. Magnetic bubble memory architectures for supporting associative searching of relational databases. IEEE Transactions on Computers, C-29(11):957–970, November 1980. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675490.

Dennis:1984:MWD

[DGT84] Jack B. Dennis, Guang-Rong Gao, and Kenneth W. Todd. Modeling the weather with a data flow supercomputer. IEEE Transactions on Computers, C-33(7):592–603, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009332.

Du:1985:DLR

[DHH85] D. Z. Du, D. F. Hsu, and F. K. Hwang. Doubly linked ring networks. IEEE Transactions on Computers, C-34 (9):853–855, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676641.

Dasgupta:1980:DML

[DHR80] S. Dasgupta, C. R. P. Hartmann, and L. D. Rudolph. Dual-mode logic for functionindependent fault testing. IEEE Transactions on Computers, C-29(11):1025–1029, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675500.

See correction [Ano81z].

Davis:1985:FLT

[DHS85] Nathaniel J. Davis IV, William Tsun-Yuk Hsu, and Howard Jay Siegel. Fault location techniques for distributed control interconnection networks. IEEE Transactions on Computers, C-34(10):902–910, October 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.engr.colostate.edu/~hj/
journals/21.pdf.

Dimopoulos:1985:SHM

[Dim85] N. J. Dimopoulos. On the structure of the homogeneous multiprocessor. IEEE Transactions on Computers, C-34 (2):141–150, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676550.

Du:1988:TDA

[DIN88] H. C. Du, O. H. Ibarra, and J. F.

Naveda. On two-dimensional via assignment for single-row routing. IEEE Transactions on Computers, 37(6):721–727, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2210.

Dias:1988:TBC

[DIY88] D. M. Dias, B. R. Iyer, and P. S. Yu. Tradeoffs between coupling small and large processors for transaction processing. IEEE Transactions on Computers, 37(3):310–320, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2168.

Dias:1981:ASB

[DJ81] D. M. Dias and J. R. Jump. Analysis and simulation of buffered delta networks. IEEE Transactions on Computers, C-30(4):273–282, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675775.

Dowd:1988:SMC

[DJ88] P. W. Dowd and K. Jabbour. Spanning multiaccess channel hypercube computer interconnection. IEEE Transactions on Computers, 37(9): 1137–1142, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2267.

deJong:1988:TPG

[dJvdG88] P. de Jong and A. J. van de Goor. Test pattern generation for API faults in RAM. IEEE Transactions on Computers, 37(11):1426–1428, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8710.

Dias:1982:CIA

[DK82] D. M. Dias and M. Kumar. Comments on “Interference Analysis of Shuffle/Exchange Networks”. IEEE Transactions on Computers, C-31(6):546–547, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676037.

Dechter:1986:BCD

[DK86] R. Dechter and L. Kleinrock. Broadcast communications and distributed algorithms. IEEE Transactions on Computers, C-35(3):210–219, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676745.

Damarla:1989:FDC

[DK89] T. R. Damarla and M. Karpovsky. Fault detection in combinational networks by Reed– Muller transforms. IEEE Transactions on Computers, 38(6): 788–797, June 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
24287.

Davis:1986:FSA

[DL86] W. A. Davis and De-Lei Lee.

Fast search algorithms for associative memories. IEEE Transactions on Computers, C-35(5):456–461, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676788.

Du:1987:HAS

[DL87] D. H.-C. Du and Lee-Chin Hsu Liu. Heuristic algorithms for single row routing. IEEE Transactions on Computers, C-36(3):312–320, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676903.

Durfee:1987:CCA

[DLC87] Edmund H. Durfee, Victor R. Lesser, and Daniel D. Corkill. Coherent cooperation among communicating problem solvers. IEEE Transactions on Computers, C-36(11):1275–1291, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009468.

DeSouzaeSilva:1986:CAT

[DLM86] E. De Souza e Silva, S. S. Lavenberg, and R. R. Muntz. A clustering approximation technique for queueing network models with a large number of chains. IEEE Transactions on Computers, C-35(5):419–430, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676784.

Davidson:1981:SEL

[DLSM81] S. Davidson, D. Landskov, B. D.

Shriver, and P. W. Mallett. Some experiments in local microcode compaction for horizontal machines. IEEE Transactions on Computers, C-30(7):460–477, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675826.

Daehin:1981:HAS

[DM81] W. Daehin and J. Mucha.

A hardware approach to selftesting of large programmable logic arrays. IEEE Transactions on Computers, C-30 (11):829–833, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675713.

Dahbura:1983:GDB

[DM83a] A. T. Dahbura and G. M.

Masson. Greedy diagnosis as the basis of an intermittentfault/transient-upset tolerant system design. IEEE Transactions on Computers, C-32 (10):953–957, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676140.

Dahbura:1983:GDH

[DM83b] A. T. Dahbura and G. M. Masson. Greedy diagnosis of hybrid fault situations. IEEE Transactions on Computers, C-32(8):777–782, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676322.

Dahbura:1984:FIA

[DM84] A. T. Dahbura and G. M. Masson. An 0(n ^{2.5} ) fault identification algorithm for diagnosable systems. IEEE Transactions on Computers, C-33(6):486–492, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676472.

deSouzaeSilva:1988:SRA

[dM88] E. de Souza e Silva and R. R. Muntz. Simple relationships among moments of queue lengths in product form queueing networks. IEEE Transactions on Computers, 37(9):1125–1129, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2264.

Dahbura:1985:SIS

[DMY85] A. T. Dahbura, G. M. Masson, and Che-Liang Yang. Selfimplicating structures for diagnosable systems. IEEE Transactions on Computers, C-34(8):718–723, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676617.

Dittert:1985:LBS

[DO85] Eric Dittert and Michael J.

O’Donnell. Lower bounds for sorting with realistic instruction sets. IEEE Transactions on Computers, C-34(4):311–317, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009381.

See correction [DO86].

Dittert:1986:CLB

[DO86] E. Dittert and M. J. O’Donnell.

Correction to “Lower Bounds for Sorting with Realistic Instruction Sets”. IEEE Transactions on Computers, C-35(10): 932, October 1986. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1676690. See [DO85].

Dong:1984:MBC

[Don84] Hao Dong. Modified Berger codes for detection of unidirectional errors. IEEE Transactions on Computers, C-33(6):572–575, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676484.

Doran:1988:VIC

[Dor88] R. W. Doran. Variants of an improved carry look-ahead adder. IEEE Transactions on Computers, 37(9):1110–1113, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2261.

Doty:1984:NDD

[Dot84] K. W. Doty. New designs for dense processor interconnection networks. IEEE Transactions on Computers, C-33(5):447–450, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676461.

Davio:1985:IN

[DR85] M. Davio and C. Ronse. Insertion networks. IEEE Transactions on Computers, C-34(6):565–570, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009412.

Dervisoglu:1980:TDM

[DS80] B. I. Dervisoglu and H. A. Sholl. Theory and design of mixed-mode sequential machines. IEEE Transactions on Computers, C-29(7):639–648, July 1980. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675635.

Dekel:1983:BTP

[DS83] E. Dekel and S. Sahni. Binary trees and parallel scheduling algorithms. IEEE Transactions on Computers, C-32(3):307–315, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676223.

Dally:1987:DFM

[DS87] W. J. Dally and C. L. Seitz.

Deadlock-free message routing in multiprocessor interconnection networks. IEEE Transactions on Computers, C-36(5):547–553, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676939.

Dahbura:1989:SCM

[DSH89] A. T. Dahbura, K. K. Sabnani, and W. J. Hery. Spare capacity as a means of fault detection and diagnosis in multiprocessor systems. IEEE Transactions on Computers, 38(6):881–891, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24300.

Dahbura:1987:CAM

[DSK87] A. T. Dahbura, K. K. Sabnani, and L. L. King. The comparison approach to multiprocessor fault diagnosis. IEEE Transactions on Computers, C-36(3):373–378, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676912.

Dugan:1989:CMD

[DT89] J. B. Dugan and K. S. Trivedi. Coverage modeling for dependability analysis of fault-tolerant systems. IEEE Transactions on Computers, 38(6):775–787, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24286.

David:1980:MDT

[DTF80] R. David and P. Thevenod-Fosse. Minimal detecting transition sequences: Application to random testing. IEEE Transactions on Computers, C-29(6):514–518, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675611.

Du:1985:PSM

[Du85] Hung-Chang Du. On the performance of synchronous multiprocessors. IEEE Transactions on Computers, C-34(5):462–466, May 1985. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676585.

Dubois:1988:TAC

[Dub88] M. Dubois. Throughput analysis of cache-based multiprocessors with multiple buses. IEEE Transactions on Computers, 37(1):58–70, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75139.

Dunning:1985:SBD

[Dun85]  Larry A. Dunning.  SEC–

BED–DED codes for error control in byte-organized memory systems. IEEE Transactions on Computers, C-34(6):557–562, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009409.

Dunning:1983:CCE

[DV83] L. A. Dunning and M. R.

Varanasi. Code constructions for error control in byte organized memory systems. IEEE Transactions on Computers, C-32(6):535–542, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676275.

Doshi:1987:OGA

[DV87] K. A. Doshi and P. J. Varman. Optimal graph algorithms on a fixed-size linear array. IEEE Transactions on Computers, C-36(4):460–470, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676928.

El-Amawy:1989:CRC

[EA89a]  A. El-Amawy.  Comments,

with reply, on “Can redundancy and masking improve the performance of synchronizers?” by L. Kleeman and A. Cantoni. IEEE Transactions on Computers, 38(5):750–753, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24278.

See [KC86].

El-Amawy:1989:SAF

[EA89b] A. El-Amawy. A systolic architecture for fast dense matrix inversion. IEEE Transactions on Computers, 38(3):449–455, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21131.

Eberlein:1987:SDM

[Ebe87]  P. J. Eberlein.  On the

Schur decomposition of a matrix for parallel computation.

IEEE Transactions on Computers, C-36(2):167–174, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676879.

Evans:1983:PST

[ED83] D. J. Evans and R. C. Dunbar. The parallel solution of triangular systems of equations. IEEE Transactions on Computers, C-32(2):201–204, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676206.

Emma:1987:CBD

[ED87] P. G. Emma and E. S. Davidson. Characterization of branch and data dependencies in programs for evaluating pipeline performance. IEEE Transactions on Computers, C-36(7):859–875, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676981.

Elmagarmid:1988:TPD

[ED88] A. K. Elmagarmid and A. K.

Datta. Two-phase deadlock detection algorithm. IEEE Transactions on Computers, 37(11):1454–1458, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8717.

El-Dessouki:1980:DEB

[EDH80] O. I. El-Dessouki and W. H. Huen. Distributed enumeration on between computers. IEEE Transactions on Computers, C-29(9):818–825, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675681.

Esfahanian:1985:FTR

[EH85] A.-H. Esfahanian and S. L. Hakimi. Fault-tolerant routing in DeBruijn communication networks. IEEE Transactions on Computers, C-34 (9):777–788, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676633.

Engelberg:1980:JSS

[EHM80] G. P. Engelberg, J. A. Howard, and D. A. Mellichamp. Job scheduling in a single-node hierarchical network for process control. IEEE Transactions on Computers, C-29(8):710–719, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675653.

Ellis:1980:MMS

[EHS80] M. J. Ellis, G. R. Hovey, and T. E. Stapinski. MTEC: a microprocessor system for astronomical telescope and instrument control. IEEE Transactions on Computers, C-29 (2):208–211, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675547.

Ercegovac:1987:FCR

[EL87] M. D. Ercegovac and T. Lang.

On-the-fly conversion of redundant into conventional representations. IEEE Transactions on Computers, C-36(7):895–897, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676986.

Ellis:1980:CSI

[Ell80] C. S. Ellis. Concurrent search and insertion in AVL trees. IEEE Transactions on Computers, C-29(9):811–817, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675680.

See comments [Got81].

Ellis:1985:DDS

[Ell85]  Carla Schlatter Ellis.  Dis-

tributed data structures: a case study. IEEE Transactions on Computers, C-34(12):1178–1185, December 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

English:1981:SFS

[Eng81] W. R. English. Synthesis of finite state algorithms in a Galois field GF[p^{n}]. IEEE Transactions on Computers, C-30(3):225–229, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675759.

Er:1984:GAR

[Er84] M. C. Er. On generating the N-ary reflected Gray codes. IEEE Transactions on Computers, C-33(8):739–741, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009360.

Ersoy:1985:SAI

[Ers85] O. Ersoy. Semisystolic array implementation of circular, skew circular, and linear convolutions. IEEE Transactions on Computers, C-34(2):190–196, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676558.

Elkind:1980:RPE

[ES80] S. A. Elkind and D. P. Siewiorek. Reliability and performance of error-correcting memory and register arrays. IEEE Transactions on Computers, C-29 (10):920–927, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675475.

Esfahanian:1989:GMF

[Esf89] A.-H. Esfahanian. Generalized measures of fault tolerance with application to N-cube networks. IEEE Transactions on Computers, 38(11):1586–1591, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42131.

Etiemble:1980:MCT

[Eti80]  D. Etiemble.  Multivalued

I^{2}L circuits for TSC checkers. IEEE Transactions on Computers, C-29(6):537–540, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675616.

Eager:1989:SVE

[EZL89] D. L. Eager, J. Zahorjan, and E. D. Lazowska. Speedup versus efficiency in parallel systems. IEEE Transactions on Computers, 38(3):408–423, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21127.

El-Ziq:1982:FDM

[EZS82] Y. M. El-Ziq and S. Y. H. Su. Fault diagnosis of MOS combinational networks. IEEE Transactions on Computers, C-31(2):129–139, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675958.

See comments [CV84].

Fam:1987:OPR

[Fam87] A. T. Fam. Optimal partitioning and redundancy removal in computing partial sums. IEEE Transactions on Computers, C-36(10):1137–1143, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676854.

Fam:1988:ECM

[Fam88] A. T. Fam. Efficient complex matrix multiplication. IEEE Transactions on Computers, 37(7):877–879, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2236.

Feng:1984:ENb

[Fen84a]  T. Y. Feng.  Editor’s no-

tice. IEEE Transactions on Computers, C-33(7):589, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009330.

Feng:1984:ENa

[Fen84b] Tse-Yun Feng. Editor’s notice. IEEE Transactions on Computers, C-33(1):1, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009309.

Feng:1984:P

[Fen84c] Tse-Yun Feng. Preface. IEEE Transactions on Computers, C-33(12):1049, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676386.

Fenwick:1985:SAD

[Fen85] Peter M. Fenwick. Some aspects of the dynamic behavior of hierarchical memories. IEEE Transactions on Computers, C-34(6):570–573, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009413.

Feng:1989:VAF

[Fen89] G.-L. Feng. A VLSI architecture for fast inversion in GF(2^{m}). IEEE Transactions on Computers, 38(10):1383–1386, October 1989. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35833.

Fet:1980:CDF

[Fet80] Ya. I. Fet. Comments on “A Design of a Fast Cellular Associative Memory for Ordered Retrieval”. IEEE Transactions on Computers, C-29(8):756–757, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675664.

Feuer:1982:CRL

[Feu82]  M. Feuer.  Connectivity of

random logic. IEEE Transactions on Computers, C-31(1):29–33, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675882.

Fishburn:1982:QN

[FF82] J. P. Fishburn and R. A. Finkel. Quotient networks. IEEE Transactions on Computers, C-31(4):288–295, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675994.

Feldstein:1982:LSF

[FG82] A. Feldstein and R. Goodman. Loss of significance in floating point subtraction and addition. IEEE Transactions on Computers, C-31(4):328–335, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676002.

Fleisher:1989:MRB

[FGPT89] H. Fleisher, J. Giraldi, R. Phoenix, and M. Tavel. Minimizability of random Boolean functions. IEEE Transactions on Computers, 38(4):593–595, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21151.

Flynn:1980:PRP

[FH80] M. J. Flynn and J. L. Hennessy.

Parallelism and representation problems in distributed systems. IEEE Transactions on Computers, C-29(12):1080–1086, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675513.

See comments [Wor81].

Flynn:1983:EAD

[FH83] M. J. Flynn and L. W. Hoevel. Execution architecture: The DELtran experiment. IEEE Transactions on Computers, C-32(2):156–175, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676200.

Franta:1984:MAH

[FH84] W. R. Franta and J. R. Heath. Measurement and analysis of HYPERchannel networks. IEEE Transactions on Computers, C-33(3):249–260, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676422.

Fuja:1986:RCR

[FH86] T. Fuja and C. Heegard. Row/ column replacement for the control of hard defects in semiconductor RAM’s. IEEE Transactions on Computers, C-35 (11):996–1000, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676701.

Fuja:1988:LSC

[FHG88] T. Fuja, C. Heegard, and R. Goodman. Linear sum codes for random access memories. IEEE Transactions on Computers, 37(9):1030–1042, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2254.

Fromm:1983:EPM

[FHH^{+}83] H. Fromm, U. Hercksen, U. Herzog, K.-H. John, R. Klar, and W. Kleinoder. Experiences with performance measurement and modeling of a processor array. IEEE Transactions on Computers, C-32(1):15–31, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676120.

Fisher:1981:TST

[Fis81] J. A. Fisher. Trace scheduling:

a technique for global microcode compaction. IEEE Transactions on Computers, C-30(7):478–490, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675827.

Fisher:1988:YFP

[Fis88]  D. C. Fisher.  Your fa-

vorite parallel algorithms might not be as fast as you think. IEEE Transactions on Computers, 37(2):211–213, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2150.

Frederickson:1988:SEF

[FJ88] Greg. N. Frederickson and Ravi Janardan. Space-efficient and fault-tolerant message routing in outerplanar networks. IEEE Transactions on Computers, 37(12):1529–1541, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=9731.

Flynn:1985:IST

[FJW85] M. J. Flynn, J. D. Johnson, and S. P. Wakefield. On instruction sets and their fornmts. IEEE Transactions on Computers, C-34(3):242–254, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676567.

Fujiwara:1981:DPL

[FK81] H. Fujiwara and K. Kinoshita. A design of programmable logic arrays with universal tests. IEEE Transactions on Computers, C-30(11):823–828, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675712.

Fisher:1985:SLV

[FK85] A. L. Fisher and H. T. Kung. Synchronizing large VLSI processor arrays. IEEE Transactions on Computers, C-34(8):734–740, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676619.

Fellows:1988:PUL

[FL88] M. R. Fellows and M. A. Langston. Processor utilization in a linearly connected parallel processing system. IEEE Transactions on Computers, 37(5):594–603, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4610.

Flanders:1982:UAC

[Fla82] P. M. Flanders. A unified approach to a class of data movements on an array processor. IEEE Transactions on Computers, C-31(9):809–819, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676096.

Fang:1989:CCG

[FLN89] Zhixi Fang, Xiaobo Li, and L. M.

Ni. On the communication complexity of generalized 2-D convolution on array processors. IEEE Transactions on Computers, 38(2):184–194, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16495.

Furchtgott:1984:PSM

[FM84] D. G. Furchtgott and J. F.

Meyer. A performability solution method for degradable nonrepairable systems. IEEE Transactions on Computers, C-33(6):550–554, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676479.

Fujiwara:1987:SCG

[FM87] Eiji Fujiwara and Kohji Matsuoka. A self-checking generalized prediction checker and its use for built-in testing. IEEE Transactions on Computers, C-36(1):86–93, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009451.

Francis:1988:BPS

[FM88] R. S. Francis and I. D. Mathieson. A benchmark parallel sort for shared memory multiprocessors. IEEE Transactions on Computers, 37(12): 1619–1626, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=9738.

Furer:1989:OGF

[FM89] M. Furer and K. Mehlhorn. AT^{2}-optimal Galois field multiplier for VLSI. IEEE Transactions on Computers, 38(9): 1333–1336, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29475.

Fujiwara:1984:STG

[FMM84] E. Fujiwara, N. Mutoh, and K. Matsuoka. A self-testing group-parity prediction checker and its use for built-in testing. IEEE Transactions on Computers, C-33(6):578–583, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676486.

Foster:1989:ALF

[Fos89] M. J. Foster. Avoiding latch formation in regular expression recognizers. IEEE Transactions on Computers, 38(5):754–756, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24279.

Fortes:1985:GDP

[FR85] J. A. B. Fortes and C. S.

Raghavendra. Gracefully degradable processor arrays. IEEE Transactions on Computers, C-34(11):1033–1044, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676536.

Franklin:1981:VPC

[Fra81] M. A. Franklin. VLSI performance comparison of banyan and crossbar communications networks. IEEE Transactions on Computers, C-30(4):283–291, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675776.

Franklin:1983:EIR

[Fra83] W. R. Franklin. Efficient iterated rotation of an object. IEEE Transactions on Computers, C-32(11):1064–1067, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676158.

Froment:1983:EFC

[Fro83] A. Froment. Error free computation: a direct method to convert finite-segment p-adic numbers into rational numbers. IEEE Transactions on Computers, C-32(4):337–343, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676234.

Finkel:1980:PIS

[FS80] R. A. Finkel and M. H. Solomon. Processor interconnection strategies. IEEE Transactions on Computers, C-29(5):360–371, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675586.

Finkel:1981:LIS

[FS81] R. A. Finkel and M. H. Solomon.

The lens interconnection strategy. IEEE Transactions on Computers, C-30(12):960–965, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675735.

Fujiwara:1983:ATG

[FS83] H. Fujiwara and T. Shimono.

On the acceleration of test generation algorithms. IEEE Transactions on Computers, C-32(12):1137–1144, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676174.

Ferrari:1988:MCF

[FS88] L. A. Ferrari and P. V. Sankar.

Minimum complexity FIR filters and sparse systolic arrays. IEEE Transactions on Computers, 37(6):760–764, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2219.

Fujiwara:1982:CFD

[FT82] H. Fujiwara and S. Toida. The complexity of fault detection problems for combinational logic circuits. IEEE Transactions on Computers, C-31(6):555–560, June 1982. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676041.

Fine:1984:DAM

[FT84] M. Fine and F. A. Tobagi. Demand assignment multiple access schemes in broadcast bus local area networks. IEEE Transactions on Computers, C-33(12):1130–1159, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676391.

Funabashi:1980:NMP

[FTT^{+}80] T. Funabashi, K. Takagi, T. Tsukada, H. Nakamura, and M. Hara. An NMOS microcomputer peripheral interface unit incorporating an analogto-digital converter. IEEE Transactions on Computers, C-29(2):102–107, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675533.

Fleisher:1987:CAM

[FTY87] H. Fleisher, M. Tavel, and J. Yeager. A computer algorithm for minimizing Reed– Muller canonical forms. IEEE Transactions on Computers, C-36(2):247–250, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676890.

Fu:1980:RDPa

[Fu80] King-Sun Fu. Recent developments in pattern recognition. IEEE Transactions on Computers, C-29(10):845–854, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675467.

Fujiwara:1981:CTC

[Fuj81] H. Fujiwara. On closedness and test complexity of logic circuits. IEEE Transactions on Computers, C-30(8):556–562, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675840.

Fujiwara:1984:NPD

[Fuj84] Hideo Fujiwara. A new PLA design for universal testability. IEEE Transactions on Computers, C-33(8):745–750, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009363.

Fukuda:1988:EPA

[Fuk88] A. Fukuda. Equilibrium point analysis of memory interference in multiprocessor systems. IEEE Transactions on Computers, 37(5):585–593, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4609.

Feng:1981:FDC

[FW81] Tse-Yun Feng and Chuan-Lin Wu. Fault-diagnosis for a class of multistage interconnection networks. IEEE Transactions on Computers, C-30 (10):743–758, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675693.

Fang:1988:MDC

[FW88] K.-Y. Fang and A. S. Wojcik. Modular decomposition of combinatorial multiplevalues circuits. IEEE Transactions on Computers, 37 (10):1293–1301, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5993.

Franklin:1982:PLP

[FWT82] M. A. Franklin, D. F. Wann, and W. J. Thomas. Pin limitations and partitioning of VLSI interconnection networks. IEEE Transactions on Computers, C-31(11):1109–1116, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675927.

Fiol:1987:DOP

[FYAV87] M. A. Fiol, J. L. A. Yebra, I. Alegre, and M. Valero. Discrete optimization problem in local networks and data alignment. IEEE Transactions on Computers, C-36(6):702–713, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676963.

Fiol:1984:LDI

[FYD84] M. A. Fiol, J. L. A. Yebra, and I. A. De Miquel. Line digraph iterations and the (d, k) digraph problem. IEEE Transactions on Computers, C-33(5):400–403, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676455.

Fukunaga:1987:AJM

[FYK87] K. Fukunaga, S. Yamada, and T. Kasai. Assignment of job modules onto array processors. IEEE Transactions on Computers, C-36(7):888–891, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676984.

Fukunaga:1984:RHE

[FYSK84] K. Fukunaga, S. Yamada, H. S. Stone, and T. Kasai. A representation of hypergraphs in the Euclidean space.

IEEE Transactions on Computers, C-33(4):364–367, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676443.

Gaitanis:1985:TSCa

[Gai85a]  N. Gaitanis.  Totally self-

checking checkers for lowcost arithmetic codes. IEEE Transactions on Computers, C-34(7):596–601, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676598.

Gaitanis:1985:TSCb

[Gai85b] N. Gaitanis. A totally selfchecking error indicator. IEEE Transactions on Computers, C-34(8):758–761, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676624.

Gaitanis:1988:DTS

[Gai88a] N. Gaitanis. The design of totally self-checking TMR faulttolerant systems. IEEE Transactions on Computers, 37(11): 1450–1454, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8716.

Gaitanis:1988:DTE

[Gai88b] N. Gaitanis. The design of TSC error C/D circuits for SEC/ DED codes. IEEE Transactions on Computers, 37(3):258–265, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2162.

Gaitanis:1988:TSC

[Gai88c]  N. Gaitanis.  Totally self-

checking checkers with separate internal fault indication. IEEE Transactions on Computers, 37(10):1206–1213, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5982.

Gait:1989:KHP

[Gai89] J. Gait. A kernel for highperformance multicast communications. IEEE Transactions on Computers, 38 (2):218–226, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16498.

Gajski:1980:PC

[Gaj80] D. D. Gajski. Parallel compressors. IEEE Transactions on Computers, C-29(5):393–398, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675589.

Gajski:1981:ASL

[Gaj81] D. D. Gajski. An algorithm for solving linear recurrence systems on parallel and pipelined machines. IEEE Transactions on Computers, C-30(3):190–206, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675755.

Garcia:1980:FPP

[Gar80] O. N. Garcia. Foreword parallel processing today. IEEE Transactions on Computers, C-29(9):762–763, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675675.

Gaudlot:1986:SHD

[Gau86] Jean-Luc Gaudlot. Structure handling in data-flow systems. IEEE Transactions on Computers, C-35(6):489–502, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009426.

Gavish:1987:OMC

[Gav87]  B. Gavish.  Optimization

models for configuring distributed computer systems. IEEE Transactions on Computers, C-36(7):773–793, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676975.

Grishman:1983:PET

[GB83] R. Grishman and Su Bogong. A preliminary evaluation of trace scheduling for global microcode compaction. IEEE Transactions on Computers, C-32(12):1191–1194, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676181.

Ghafoor:1989:BFT

[GBG89] A. Ghafoor, T. R. Bashkow, and I. Ghafoor. Bisectional fault-tolerant communication architecture for supercomputer systems. IEEE Transactions on Computers, 38(10): 1425–1446, October 1, 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35837.

Gardarin:1980:DCA

[GC80] G. Gardarin and W. W. Chu.

A distributed control algorithm for reliably and consistently updating replicated databases. IEEE Transactions on Computers, C-29(12):1060–1068, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675511.

Gecsei:1987:SAN

[GC87] Jan Gecsei and Eduard Cerny. Self-adjusting networks for VLSI simulation. IEEE Transactions on Computers, C-36(9): 1114–1120, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009544.

Ganz:1989:PAA

[GC89] A. Ganz and I. Chlamtac. Path allocation access control in fiber optic communication systems. IEEE Transactions on Computers, 38(10):1372–1382, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=35832.

Galiay:1980:PVL

[GCV80] J. Galiay, Y. Crouzet, and M. Vergniault. Physical versus logical fault models MOS LSI circuits: Impact on their testability. IEEE Transactions on Computers, C-29(6):527–531, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675614.

Gopal:1983:OWM

[GCW83] I. S. Gopal, D. Coppersmith, and C. K. Wong. Optimal wiring of movable terminals. IEEE Transactions on Computers, C-32(9):845–858, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676333.

Geffe:1973:HPD

[Gef73] P. Geffe. How to protect data with ciphers that are really hard to break. Electronics, 46(1): 99–101, 1973. ISSN 0883-4989. This cipher was later broken by [Sie85b].

Gelernter:1981:DBA

[Gel81] D. Gelernter. A DAG-based algorithm for prevention of store-and-forward deadlock in packet networks. IEEE Transactions on Computers, C-30 (10):709–715, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675690.

Gelder:1989:PPA

[Gel89] A. V. Gelder. PRAM processor allocation: a hidden bottleneck in sublogarithmic algorithms. IEEE Transactions on Computers, 38(2):289–292, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16507.

Gerrity:1982:CRR

[Ger82] G. W. Gerrity. Computer representation of real numbers. IEEE Transactions on Computers, C-31(8):709–714, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676076.

Gold:1987:SFB

[GF87] Y. I. Gold and W. R. Franta. A scheduling-function-based distributed access protocol that uses CDM to relay control information in a network with hidden nodes. IEEE Transactions on Computers, C-36(5):562–569, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676941.

Gold:1983:DCA

[GFM83] Y. I. Gold, W. R. Franta, and S. Moran. A distributed channel-access protocol for fullyconnected networks with mobile nodes. IEEE Transactions on Computers, C-32 (2):133–147, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676198.

Gottlieb:1983:NUD

[GGK^{+}83] A. Gottlieb, R. Grishman, C. P. Kruskal, K. P. McAuliffe, L. Rudolph, and M. Snir. The NYU Ultracomputer — designing an MIMD shared memory parallel computer. IEEE Transactions on Computers, C-32(2):175–189, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676201.

Goundan:1980:DTF

[GH80a] A. Goundan and J. P. Hayes.

Design of totally fault locatable combinational networks. IEEE Transactions on Computers, C-29(1):33–44, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675454.

Goundan:1980:IEF

[GH80b] A. Goundan and J. P. Hayes.

Identification of equivalent faults in logic networks. IEEE Transactions on Computers, C-29 (11):978–985, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675492.

Gaitanis:1983:NPC

[GH83a] N. Gaitanis and C. Halatsis.

Near-perfect codes for binarycoded radix-r arithmetic units. IEEE Transactions on Computers, C-32(5):494–497, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676261.

Gaitanis:1983:NDM

[GH83b] N. Gaitanis and C. Halatsis. A new design method for m-out-of-n TSC checkers. IEEE Transactions on Computers, C-32(3):273–283, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676219.

Garg:1988:DBN

[GH88] U. Garg and Y.-P. Huang. Decomposing banyan networks for performance analysis. IEEE Transactions on Computers, 37(3):371–376, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2177.

Guyot:1987:WBE

[GHM87] A. Guyot, B. Hochet, and J.-M. Muller. A way to build efficient carry-skip adders. IEEE Transactions on Computers, C-36(10):1144–1152, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676855.

Gonzalez:1980:FQE

[GJ80] M. J. Gonzalez, Jr. and B. W.

Jordan, Jr. A framework for the quantitative evaluation of distributed computer systems. IEEE Transactions on Computers, C-29(12):1087–1094, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675514.

Garcia:1983:MMR

[GK83] G. H. Garcia and W. J. Kubitz. Minimum mean running time function generation using read-only memory. IEEE Transactions on Computers, C-32(2):147–156, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676199.

Graham:1985:PAA

[GK85] J. H. Graham and T. F.

Kadela. Parallel algorithms and architectures for optimal state estimation. IEEE Transactions on Computers, C-34 (11):1061–1068, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676540.

Gordon:1984:ETS

[GKS84] Dan Gordon, Israel Koren, and Gabriel M. Silberman.

Embedding tree structures in VLSI hexagonal arrays. IEEE Transactions on Computers, C-33(1):104–107, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009319.

Granski:1987:EOS

[GKS87] Michael Granski, Israel Koren, and Gabriel M. Silberman. The effect of operation scheduling on the performance of a data flow computer. IEEE Transactions on Computers, C-36(9):1019–1029, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009533.

Gonzalez:1986:RMN

[GL86] Teofilo F. Gonzalez and Sing-Ling Lee. Routing multiterminal nets around a rectangle. IEEE Transactions on Computers, C-35(6):543–549, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009431.

Glass:1980:EMI

[Gla80]  J. M. Glass.  An efficient

method for improving reliability of a pipeline FFT. IEEE Transactions on Computers, C-29(11):1017–1020, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675498.

See correction [Ano81y].

Gupta:1981:AR

[GLL81] U. I. Gupta, D. T. Lee, and J. Y.-T. Leung. Authors’ reply. IEEE Transactions on Computers, C-30(6):455, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675819.

Gelenbe:1982:EPS

[GLS82] E. Gelenbe, A. Lichnewsky, and A. Staphylopatis. Experience with the parallel solution of partial differential equations on a distributed computing system. IEEE Transactions on Computers, C-31(12):1157–1164, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675939.

Garcia-Molina:1982:EDC

[GM82a] H. Garcia-Molina. Elections in a distributed computing system. IEEE Transactions on Computers, C-31(1):48–59, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675885.

Glaser:1982:CSA

[GM82b] R. E. Glaser and G. M. Masson. The containment set approach to upsets in digital systems. IEEE Transactions on Computers, C-31(7):689–692, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676068.

Greenberg:1987:PPA

[GM87] A. G. Greenberg and U. Manber. A probabilistic pipeline algorithm for K selection on the tree machine. IEEE Transactions on Computers, C-36(3):359–362, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676908.

Gazit:1988:FTC

[GM88a] I. Gazit and M. Malek. Fault tolerance capabilities in multistage network-based multicomputer systems. IEEE Transactions on Computers, 37(7):788–798, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2224.

Gold:1988:EMC

[GM88b] Y. I. Gold and S. Moran. Estimating metrical change in fully connected mobile networks — a least upper bound on the worst case. IEEE Transactions on Computers, 37(9):

1156–1162, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2272.

Gazit:1989:NPP

[GM89] I. Gazit and M. Malek. On the number of permutations performable by extra-stage multistage interconnection networks. IEEE Transactions on Computers, 38(2):297–302, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16509.

Guest:1984:RNS

[GMG84] C. C. Guest, M. M. Mirsalehi, and T. K. Gaylord. Residue number system truth-table lookup processing — moduli selection and logical minimization. IEEE Transactions on Computers, C-33(10):927–931, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676355;
http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35232.

Garcia-Molina:1985:ERT

[GMK85] H. Garcia-Molina and J. Kent.

Evaluating response time in a faulty distributed computing system. IEEE Transactions on Computers, C-34 (2):101–109, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676546.

Garcia-Molina:1984:MMM

[GMLV84] H. Garcia-Molina, R. J. Lipton, and J. Valdes. A massive memory machine. IEEE Transactions on Computers, C-33(5):391–399, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676454.

See comments [GY86].

Gnanasekaran:1983:BSI

[Gna83] R. Gnanasekaran. On a bitserial input and bit-serial output multiplier. IEEE Transactions on Computers, C-32 (9):878–880, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676341.

Gnanasekaran:1985:FSP

[Gna85]  R. Gnanasekaran.  A fast

serial-parallel binary multiplier. IEEE Transactions on Computers, C-34(8):741–744, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676620.

Golan:1988:PTP

[GNH88] P. Golan, O. Novak, and J. Hlavicka. Pseudoexhaustive test pattern generator with enhanced fault coverage. IEEE Transactions on Computers, 37(4):496–500, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2198.

Gorgui-Naguib:1986:CMP

[GNK86] R. N. Gorgui-Naguib and R. A.

King. Comments on “Matrix Processors Using p-Adic Arithmetic for Exact Linear Computations”. IEEE Transactions on Computers, C-35 (10):928–930, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676686.

Goel:1981:IEA

[Goe81] P. Goel. An implicit enumeration algorithm to generate tests for combinational logic circuits. IEEE Transactions on Computers, C-30(3):215–222, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675757.

Golan:1984:DTS

[Gol84] P. Golan. Design of totally selfchecking checker for 1-out-of-3 code. IEEE Transactions on Computers, C-33(3):285, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676427.

Gordon:1987:EEB

[Gor87] Dan Gordon. Efficient embeddings of binary trees in VLSI arrays. IEEE Transactions on Computers, C-36 (9):1009–1018, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009532.

Gottlieb:1981:CCS

[Got81] A. Gottlieb. Comments on “Concurrent Search and Insertion in AVL Trees”. IEEE Transactions on Computers, C-30(10):812, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675700.

See [Ell80].

Gavish:1986:CDL

[GP86] B. Gavish and H. Pirkul. Computer and database location in distributed computer systems. IEEE Transactions on Computers, C-35(7):583–590, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676799.

Goodman:1981:HMI

[GS81] J. R. Goodman and C. H. Sequin. Hypertree: a multiprocessor interconnection topology. IEEE Transactions on Computers, C-30(12):923–933, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675731.

Guha:1986:FTD

[GS86] S. Guha and A. Sen. On fault-tolerant distributor communication architecture. IEEE Transactions on Computers, C-35(3):281–283, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676752.

Gertner:1987:VAM

[GS87] Izidor Gertner and Moshe Shamash. VLSI architectures for multidimensional Fourier transform processing. IEEE Transactions on Computers, C-36(11):1265–1274, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009467.

Ghafoor:1989:PFT

[GS89] A. Ghafoor and P. Sole. Performance of fault-tolerant diagnostics in the hypercube systems. IEEE Transactions on Computers, 38(8):1164–1172, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30870.

Gostelow:1980:PSD

[GT80] K. P. Gostelow and R. E. Thomas. Performance of a simulated dataflow computer. IEEE Transactions on Computers, C-29(10):905–919, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675474.

Geist:1982:ODM

[GT82] R. M. Geist and K. S. Trivedi. Optimal design of multilevel storage hierarchies. IEEE Transactions on Computers, C-31(3):249–260, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675981.

Geist:1983:URP

[GT83] R. M. Geist and K. S. Trivedi. Ultrahigh reliability prediction for fault-tolerant computer systems. IEEE Transactions on Computers, C-32 (12):1118–1127, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676172.

Goyal:1987:EPD

[GT87] A. Goyal and A. N. Tantawi.

Evaluation of performability for degradable computer systems. IEEE Transactions on Computers, C-36(6):738–744, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676966.

Goyal:1988:MGA

[GT88] A. Goyal and A. N. Tantawi.

A measure of guaranteed availability and its numerical evaluation. IEEE Transactions on Computers, 37(1):25–32, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75148.

Guerra:1986:SAL

[Gue86] C. Guerra. Systolic algorithms for local operations on images. IEEE Transactions on Computers, C-35(1):73–77, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676663.

Guibaly:1989:DAA

[Gui89] F. El Guibaly. Design and analysis of arbitration protocols. IEEE Transactions on Computers, 38(2):161–171, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16493.

Gannon:1984:ICC

[GV84] D. B. Gannon and J. Van Rosendale. On the impact of communication complexity on the design of parallel numerical algorithms. IEEE Transactions on Computers, C-33(12):1180–1194, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676393.

Gopal:1981:DAB

[GW81] G. Gopal and J. W. Wong. Delay analysis of broadcast routing in packet-switching networks. IEEE Transactions on Computers, C-30(12):915–922, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675730.

Guting:1984:FRI

[GW84] Ralf Hartmut Guting and Derick Wood. Finding rectangle intersections by divide-and-conquer. IEEE Transactions on Computers, C-33(7):671–675, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009341.

Gaudiot:1989:TRT

[GW89] J.-L. Gaudiot and Y.-H. Wei.

Token relabeling in a tagged token data-flow architecture. IEEE Transactions on Computers, 38(9):1225–1239, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29463.

Goodman:1986:CMM

[GY86] J. R. Goodman and H. C.

Young. Comments on “A Massive Memory Machine”. IEEE Transactions on Computers, C-35(10):907–910, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676682.

See [GMLV84].

Ghosh:1989:PSM

[GY89] S. Ghosh and M.-L. Yu. A preemptive scheduling mechanism for accurate behavioral simulation of digital designs. IEEE Transactions on Computers, 38(11):1595–1600, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42133.

Glasser:1988:CMC

[GZ88] L. A. Glasser and C. A.

Zukowski. Continuous models for communication density constraints on multiprocessor performance. IEEE Transactions on Computers, 37(6):652–656, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2204.

Gupta:1989:RMV

[GZR89] R. Gupta, A. Zorat, and I. V. Ramakrishnan. Reconfigurable multipipelines for vector supercomputers. IEEE Transactions on Computers, 38(9): 1297–1307, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29468.

Huang:1984:ABF

[HA84] Kuang-Hua Huang and J. A. Abraham. Algorithm-based fault tolerance for matrix operations. IEEE Transactions on Computers, C-33(6):518–528, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676475.

Hassan:1986:FTM

[HA86] A. S. Mahmudul Hassan and V. K. Agarwal. A fault-tolerant modular architecture for binary trees. IEEE Transactions on Computers, C-35(4):356–361, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676770.

Howells:1988:RSY

[HA88] M. C. Howells and V. K.

Agarwal. A reconfiguration scheme for yield enhancement of large area binary tree architectures. IEEE Transactions on Computers, 37(4):463–468, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2192.

Hagmann:1986:CRS

[Hag86] R. B. Hagmann. A crash recovery scheme for a memoryresident database system. IEEE Transactions on Computers, C-35(9):839–843, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676845.

Harrison:1986:EAP

[Har86] P. G. Harrison. An enhanced approximation by pair-wise analysis of servers for time delay distributions in queueing networks. IEEE Transactions on Computers, C-35(1):54–61, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676657.

Hassan:1984:STS

[Has84] Syed Zahoor Hassan. Signature testing of sequential machines. IEEE Transactions on Computers, C-33(8):762–764, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009368.

Hawkes:1985:RFT

[Haw85]  L. W. Hawkes.  A regu-

lar fault-tolerant architecture for interconnection networks. IEEE Transactions on Computers, C-34(7):677–680, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676608.

Hayes:1980:TMS

[Hay80] J. P. Hayes. Testing memories for single-cell pattern-sensitive faults. IEEE Transactions on Computers, C-29(3):249–254, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675556.

Hayes:1981:SSA

[Hay81] A. B. Hayes. Stored state asynchronous sequential circuits. IEEE Transactions on Computers, C-30(8):596–600, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675846.

Hayes:1984:FLF

[Hay84] J. P. Hayes. Foreword—liable and fault-tolerant computing. IEEE Transactions on Computers, C-33(6):465–466, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676469.

Hayes:1986:PBL

[Hay86a] J. P. Hayes. Pseudo-Boolean logic circuits. IEEE Transactions on Computers, C-35(7):602–612, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676801.

Hayes:1986:UEM

[Hay86b] J. P. Hayes. Uncertainty, energy, and multiple-valued logics. IEEE Transactions on Computers, C-35(2):107–114, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676728.

Hwang:1982:PMA

[HC82] Kai Hwang and Yeng-Heng Cheng. Partitioned matrix algorithms for VLSI arithmetic systems. IEEE Transactions on Computers, C-31 (12):1215–1224, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675945.

Huang:1986:DSF

[HC86] Kaiyuan Huang and Tinghuai Chen. On the diagnosis of system faults with propagation. IEEE Transactions on Computers, C-35(12):1082–1086, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676717.

Huang:1987:FTS

[HC87] Chun-Fu Huang and Wen-Tsuen Chen. Fault-tolerant singlestage interconnection networks. IEEE Transactions on Computers, C-36(5):637–640, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676951.

Hortensius:1989:ISI

[HCMP89] P. D. Hortensius, H. C. Card, R. D. McLeod, and W. Pries. Importance sampling for Ising computers using one-dimensional cellular automata. IEEE Transactions on Computers, 38(6):769– 774, June 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
24285.

Hennessy:1984:VPA

[Hen84] J. L. Hennessy. VLSI processor architecture. IEEE Transactions on Computers, C-33(12):1221–1246, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676395.

Herlihy:1987:EMT

[Her87] M. Herlihy. Extending multiversion time-stamping protocols to exploit type information. IEEE Transactions on Computers, C-36(4):443–448, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676926.

Herron:1982:GPH

[HFPS82] J. M. Herron, J. Farley, K. Preston, Jr., and H. Sellner. A general-purpose high-speed logical transform image processor. IEEE Transactions on Computers, C-31(8):795–800, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676086.

Hwang:1987:HCE

[HG87] Kai Hwang and Joydeep Ghosh.

Hypernet: a communicationefficient architecture for constructing massively parallel computers. IEEE Transactions on Computers, C-36(12):

1450–1466, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009497.

Halatsis:1983:ECC

[HGS83a] C. Halatsis, N. Gaitanis, and M. Sigala. Error-correcting codes in binary-coded radix-r arithmetic. IEEE Transactions on Computers, C-32(3):326–328, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676227.

Halatsis:1983:FET

[HGS83b] C. Halatsis, N. Gaitanis, and M. Sigala. Fast and efficient totally self-checking checkers for m-out-of-(2m ± 1) codes. IEEE Transactions on Computers, C-32(5):507–511, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676265.

Holgate:1980:AIF

[HI80] R. W. Holgate and R. N. Ibbett. An analysis of instructionfetching strategies in pipelined computers. IEEE Transactions on Computers, C-29(4):325–329, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675572.

Hseuh:1988:PMB

[HIT88] M. C. Hseuh, R. K. Iyer, and K. S. Trivedi. Performance modeling based on real data: a case study. IEEE Transactions on Computers, 37(4):478–484, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2195.

Harper:1987:VAP

[HJ87] David T. Harper and J. Robert Jump. Vector access performance in parallel memories using a skewed storage scheme. IEEE Transactions on Computers, C-36(12):1440–1449, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009496.

Hosseini:1984:DAD

[HKR84] S. H. Hosseini, J. G. Kuhl, and S. M. Reddy. A diagnosis algorithm for distributed computing systems with dynamic failure and repair. IEEE Transactions on Computers, C-33(3):223–233, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676419.

Hosseini:1987:DFT

[HKR87] S. H. Hosseini, J. G. Kuhl, and S. M. Reddy. Distributed fault-tolerance of tree structures. IEEE Transactions on Computers, C-36(11): 1378–1382, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009481.

Hosseini:1988:SFD

[HKR88] S. H. Hosseini, J. G. Kuhl, and S. M. Reddy. On self-fault diagnosis of the distributed systems. IEEE Transactions on Computers, 37(2):248–251, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2158.

Hartimo:1986:DDF

[HKSS86] I. Hartimo, K. Kronlof, O. Simula, and J. Skytta. DFSP: a data flow signal processor. IEEE Transactions on Computers, C-35(1):23–33, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676654.

Heuft:1980:CC

[HL80] R. W. Heuft and W. D. Little. Convolution computer. IEEE Transactions on Computers, C-29(8):738–740, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675658.

Heuft:1982:ITP

[HL82] R. W. Heuft and W. D. Little. Improved time and parallel processor bounds for Fortranlike loops. IEEE Transactions on Computers, C-31(1):78–81, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675890.

Heidelberger:1984:CPE

[HL84] P. Heidelberger and S. S. Lavenberg. Computer performance evaluation methodology. IEEE Transactions on Computers, C-33(12):1195–1220, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676394.

Hongyuan:1986:CSL

[HL86] Wang Hongyuan and S. C. Lee.

Comments on “Sign/Logarithm Arithmetic for FFT Implementation”. IEEE Transactions on Computers, C-35(5):482–484, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676792.

See [SCNS83].

Hlawiczka:1986:CTS

[Hla86]  A. Hlawiczka.  Compres-

sion of three-state data serial streams by means of a parallel LFSR signature analyzer.

IEEE Transactions on Computers, C-35(8):732–741, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676824.

Helbig:1989:DDM

[HM89] W. Helbig and V. Milutinovic. A DCFL E/D-MESFET GaAs experimental RISC machine. IEEE Transactions on Computers, 38(2):263–274, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16503.

Hortensius:1989:PRN

[HMC89] P. D. Hortensius, R. D. McLeod, and H. C. Card. Parallel random number generation for VLSI systems using cellular automata. IEEE Transactions on Computers, 38(10):1466–1473, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35843.

Hughes:1984:DTS

[HML84] J. L. A. Hughes, E. J. Mc-Cluskey, and D. J. Lu. Design of totally self-checking comparators with an arbitrary number of inputs. IEEE Transactions on Computers, C-33(6):546–550, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676478.

Hwang:1980:ROP

[HN80] K. Hwang and L. M. Ni. Resource optimization of a parallel computer for multiple vector processing. IEEE Transactions on Computers, C-29 (9):831–836, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675683.

Hakimi:1984:ASD

[HN84] S. L. Hakimi and K. Nakajima. On adaptive system diagnosis. IEEE Transactions on Computers, C-33(3):234–240, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676420.

Hopper:1988:CFR

[HN88] A. Hopper and R. M. Needham. The Cambridge fast ring networking system. IEEE Transactions on Computers, 37 (10):1214–1223, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5983.

Hiraki:1984:EAM

[HNS84] Kei Hiraki, Kenji Nishida, and Toshio Shimada. Evaluation of associative memory using parallel chained hashing. IEEE Transactions on Computers, C-33(9):851–855, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676502.

Hong:1981:SPG

[HO81] Se June Hong and D. L. Ostapko. A simple procedure to generate optimum test patterns for parity logic networks. IEEE Transactions on Computers, C-30(5):356–358, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675793.

Hockney:1983:CCO

[Hoc83]  R. W. Hockney.  Charac-

terizing computers and optimizing the FACR(ℓ) Poisson solver on parallel unicomputers. IEEE Transactions on Computers, C-32(10):933–941, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676137.

Hockney:1987:APD

[Hoc87]  R. Hockney.  Algorithmic

phase diagrams. IEEE Transactions on Computers, C-36 (2):231–233, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676886.

Hoffman:1985:SCP

[Hof85] D. Hoffman. The specification of communication protocols. IEEE Transactions on Computers, C-34(12):1102–1113, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hollaar:1982:DIA

[Hol82a]  L. A. Hollaar.  Direct im-

plementation of asynchronous control units. IEEE Transactions on Computers, C-31 (12):1133–1141, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675937.

Holzmann:1982:TPV

[Hol82b]  G. J. Holzmann.  A the-

ory for protocol validation. IEEE Transactions on Computers, C-31(8):730–738, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676079.

Hong:1981:EAS

[Hon81]  Sung Je Hong.  Existence

algorithms for synchronizing/ distinguishing sequences. IEEE Transactions on Computers, C-30(3):234–237, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675761.

Hong:1982:PRF

[Hon82] Se June Hong. Preface: Reliable and fault-tolerant computing. IEEE Transactions on Computers, C-31(7):575–577, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676053.

Hong:1985:ECR

[Hon85] Yang-Chang Hong. Efficient computing of relational algebraic primitives in a database machine architecture. IEEE Transactions on Computers, C-34(7):588–595, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676597.

Hosseini:1989:FTS

[Hos89]  S. H. Hosseini.  On fault-

tolerant structure, distributed fault-diagnosis, reconfiguration, and recovery of the array processors. IEEE Transactions on Computers, 38(7):932–942, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30846.

Hou:1987:FHT

[Hou87a] H. S. Hou. The Fast Hartley Transform algorithm. IEEE Transactions on Computers, C-36(2):147–156, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676877.

See correction [Hou87b].

Hou:1987:CFH

[Hou87b] Hsieh S. Hou. Correction to “The Fast Hartley Transform Algorithm”. IEEE Transactions on Computers, C-36(9): 1135–1136, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009549.

See [Hou87a].

Hafer:1982:ASD

[HP82] L. J. Hafer and A. C. Parker.

Automated synthesis of digital hardware. IEEE Transactions on Computers, C-31(2):93–109, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675955.

Hwu:1987:CRH

[HP87] Wen-Mei W. Hwu and Yale N.

Patt. Checkpoint repair for high-performance out-of-order execution machines. IEEE Transactions on Computers, C-36(12):1496–1514, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009500.

Homobono:1988:CII

[HP88] N. Homobono and C. Peyrat. Connectivity of Imase and Itoh digraphs. IEEE Transactions on Computers, 37(11): 1459–1461, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8718.

Hong:1989:PSR

[HP89] Y.-C. Hong and T. H. Payne. Parallel sorting in a ring network of processors. IEEE Transactions on Computers, 38(3):458–464, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21133.

Hochet:1989:SGE

[HQR89] B. Hochet, P. Quinton, and Y. Robert. Systolic Gaussian elimination over GF(p) with partial pivoting. IEEE Transactions on Computers, 38 (9):1321–1324, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29471.

Hariri:1987:SSR

[HR87] S. Hariri and C. S. Raghavendra. SYREL: a symbolic reliability algorithm based on path and cutset methods. IEEE Transactions on Computers, C-36(10):1224–1232, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676862.

Ha:1988:DPT

[HR88] Dong Sam Ha and S. M.

Reddy. On the design of pseudoexhaustive testable PLAs. IEEE Transactions on Computers, 37(4):468–472, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2193.

Huntsberger:1986:RUC

[HRJ86] T. L. Huntsberger, C. Rangarajan, and S. N. Jayaramamurthy. Representation of uncertainty in computer vision using fuzzy sets. IEEE Transactions on Computers, C-35(2):145–156, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676732.

Hsu:1984:VIR

[HRT^{+}84] In-Shek Hsu, I. S. Reed, T. K.

Truong, Ke Wang, Chiunn-Shyong Yeh, and L. J. Deutsch.

The VLSI implementation of a Reed–Solomon encoder using Berlekamp’s bit-serial multiplier algorithm. IEEE Transactions on Computers, C-33 (10):906–911, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676351.

Holt:1981:DSA

[HS81] C. S. Holt and J. E. Smith. Diagnosis of systems with asymmetric invalidation. IEEE Transactions on Computers, C-30 (9):679–690, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675868.

Hsiao:1984:AUR

[HS84] Ten-Chuan Hsiao and S. C. Seth. An analysis of the use of Rademacher–Walsh spectrum in compact testing. IEEE Transactions on Computers, C-33(10):934–937, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676357.

Holt:1985:SDD

[HS85] C. S. Holt and J. E. Smith. Self-diagnosis in distributed systems. IEEE Transactions on Computers, C-34(1):19–32, January 1985. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676512.

Hu:1986:TSD

[HS86] Mou Hu and K. C. Smith.

Ternary scan design for VLSI testability. IEEE Transactions on Computers, C-35 (2):167–170, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676735.

See comments [MA89].

Harden:1988:AYO

[HS88] J. C. Harden and N. R. Strader II. Architectural yield optimization for WSI. IEEE Transactions on Computers, 37(1):88–110, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75138.

Hill:1989:EAC

[HS89] M. D. Hill and A. J. Smith.

Evaluating associativity in CPU caches. IEEE Transactions on Computers, 38(12):1612–1630, December 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
40842.

Hemmati:1984:ABC

[HSE84] Farhad Hemmati, Donald L. Schilling, and George Eichmann. Adjacencies between the cycles of a shift register with characteristic polynomial (1 + x) ^{n}. IEEE Transactions on Computers, C-33(7):675–677, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009342.

Haviland:1980:CAP

[HT80] G. L. Haviland and A. A. Tuszynski. A CORDIC arithmetic processor chip. IEEE Transactions on Computers, C-29(2):68–79, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675529.

Heidelberger:1982:QNM

[HT82] P. Heidelberger and K. S. Trivedi. Queueing network models for parallel processing with asynchronous tasks. IEEE Transactions on Computers, C-31(11):1099–1109, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675926.

Heidelberger:1983:AQM

[HT83] P. Heidelberger and K. S. Trivedi. Analytic queueing models for programs with internal concurrency. IEEE Transactions on Computers, C-32(1):73–82, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676125.

Huang:1986:FSM

[HT86] Shing-Tsaan Huang and S. K.

Tripathi. Finite state model and compatibility theory: New analysis tools for permutation networks. IEEE Transactions on Computers, C-35(7):591–601, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676800.

Huang:1988:SRT

[HT88] S.-T. Huang and S. K. Tripathi. Self-routing technique in perfect-shuffle networks using control tags. IEEE Transactions on Computers, 37 (2):251–256, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2159.

Hsu:1988:CVA

[HTDR88] I. S. Hsu, T. K. Truong, L. J.

Deutsch, and I. S. Reed. A comparison of VLSI architecture of finite field multipliers using dual, normal, or standard bases. IEEE Transactions on Computers, 37(6):735–739, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2212.

Hwang:1989:OMP

[HTK89] K. Hwang, P.-S. Tseng, and D. Kim. An orthogonal multiprocessor for parallel scientific computations. IEEE Transactions on Computers, 38(1):47–61, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8729.

Hagiwara:1980:DMC

[HTOS80] H. Hagiwara, S. Tomita, S. Oyanagi, and K. Shibayama. A dynamically microprogrammable computer with low-level parallelism. IEEE Transactions on Computers, C-29(7):577–595, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675629.

Hong:1989:PAM

[HTT89] J. Hong, X. Tan, and D. Towsley. A performance analysis of minimum laxity and earliest deadline scheduling in a real-time system. IEEE Transactions on Computers, 38(12):1736–1744, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40851.

Huang:1983:FPM

[Hua83] C. H. Huang. A fully parallel mixed-radix conversion algorithm for residue number applications. IEEE Transactions on Computers, C-32(4):398–402, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676242;
 http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35226.

Hurst:1981:CDD

[Hur81]  S. L. Hurst.  Comments

on “Design of a Dynamically Programmable Logic Gate”. IEEE Transactions on Computers, C-30(12):986–987, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675740.

Hurst:1984:MVL

[Hur84] S. L. Hurst. Multiple-valued logic—its status and its future. IEEE Transactions on Computers, C-33(12):1160–1179, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676392.

Holliday:1987:EPE

[HV87] Mark A. Holliday and Mary K. Vernon. Exact performance estimates for multiprocessor memory and bus interference. IEEE Transactions on Computers, C-36(1):76–85, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009450.

Hwang:1987:CRL

[Hwa87] F. K. Hwang. Comments on “Reliable Loop Topologies for Large Local Computer Networks”. IEEE Transactions on Computers, C-36(3):383–384, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676914.

See [RGA85].

Hwang:1988:MNC

[HX88] K. Hwang and Z. Xu. Multipipeline networking for compound vector processing. IEEE Transactions on Computers, 37(1):33–47, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=75149.

Horowitz:1981:BTI

[HZ81] E. Horowitz and A. Zorat. The binary tree as an interconnection network: Applications to multiprocessor systems and VLSI. IEEE Transactions on Computers, C-30(4):247–253, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675772.

Horowitz:1983:DCP

[HZ83] E. Horowitz and A. Zorat.

Divide-and-conquer for parallel processing. IEEE Transactions on Computers, C-32(6):582–585, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676280.

Iyer:1982:SFL

[IBM82] R. K. Iyer, S. E. Butner, and E. J. Mccluskey. A statistical failure/load relationship: Results of a multicomputer study. IEEE Transactions on Computers, C-31(7):697–706, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676070.

Inkol:1980:DRT

[IC80] R. J. Inkol and S. G. Chamberlain. Design and realization of a two-level 64K byte CCD memory system for microcomputer applications. IEEE Transactions on Computers, C-29(2):195–199, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675543.

Irani:1982:MIC

[IC82] K. B. Irani and Kuo-Wei Chen. Minimization of interprocessor communication for parallel computation. IEEE Transactions on Computers, C-31(11):1067–1075, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675923.

Iyer:1986:APS

[IDH86] B. R. Iyer, L. Donatiello, and P. Heidelberger. Analysis of performability for stochastic models of fault-tolerant systems. IEEE Transactions on Computers, C-35(10):902–907, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676681.

Imase:1981:DMD

[II81] M. Imase and M. Itoh. Design to minimize diameter on building-block network. IEEE Transactions on Computers, C-30(6):439–442, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675809.

Imase:1983:DDG

[II83] M. Imase and M. Itoh. A design for directed graphs with minimum diameter. IEEE Transactions on Computers, C-32(8):782–784, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676323.

Ibaraki:1982:DFS

[IK82a] T. Ibaraki and T. Kameda.

Deadlock-free systems for a bounded number of processes. IEEE Transactions on Computers, C-31(3):188–193, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675974.

Irani:1982:MDC

[IK82b] K. B. Irani and N. G. Khabbaz. A methodology for the design of communication networks and the distribution of data in distributed supercomputer systems. IEEE Transactions on Computers, C-31(5):419–434, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676019.

Iyengar:1985:CFD

[IK85] V. S. Iyengar and L. L. Kinney.

Concurrent fault detection in microprogrammed control units.

IEEE Transactions on Computers, C-34(9):810–821, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676637.

Isoda:1983:GCH

[IKI83] S. Isoda, Y. Kobayashi, and T. Ishida. Global compaction of horizontal microprograms based on the generalized data dependency graph. IEEE Transactions on Computers, C-32(10):922–933, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676136.

Ibarra:1986:DSA

[IKP86] Oscar H. Ibarra, Sam M. Kim, and Michael A. Palis. Designing systolic algorithms using sequential machines. IEEE Transactions on Computers, C-35(6):531–542, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009430.

Ibaraki:1981:MTS

[IKT81] T. Ibaraki, T. Kameda, and S. Toida. On minimal test sets for locating single link failures in networks. IEEE Transactions on Computers, C-30(3):182–190, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675754.

Ilyas:1987:EEF

[IM87] M. Ilyas and H. T. Mouftah. End-to-end flow control in computer networks with noisy channels and quasi-cut-through switching. IEEE Transactions on Computers, C-36(11):1382–1386, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009482.

Irwin:1983:FDL

[IO83] M. J. Irwin and R. M. Owens.

Fully digit on-line networks. IEEE Transactions on Computers, C-32(4):402–406, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676243.

Irani:1984:CFS

[IO84] K. B. Irani and I. H. Onyuksel. A closed-form solution for the performance analysis of multiple-bus multiprocessor systems. IEEE Transactions on Computers, C-33 (11):1004–1012, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676374.

Ignizio:1982:MAS

[IPM82] J. P. Ignizio, D. F. Palmer, and C. M. Murphy. A multicriteria approach to supersystem architecture definition. IEEE Transactions on Computers, C-31(5):410–418, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676018.

Iyer:1986:MBM

[IR86] Ravishankar K. Iyer and David J. Rossetti. A measurementbased model for workload dependence of CPU errors. IEEE Transactions on Computers, C-35(6):511–519, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009428.

Irshid:1988:SMD

[Irs88] M. I. Irshid. A simple method for determining Hadamard sequency vectors. IEEE Transactions on Computers, 37(6):743–745, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2214.

Iyer:1984:DMI

[IS84] B. R. Iyer and J. B. Sinclair. Dynamic memory interconnections for rapid access. IEEE Transactions on Computers, C-33(10):923–927, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676354.

Imase:1985:CRD

[ISO85] M. Imase, T. Soneoka, and K. Okada. Connectivity of regular directed graphs with small diameters. IEEE Transactions on Computers, C-34(3):267–273, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676569.

Iyer:1984:REF

[Iye84] R. K. Iyer. Reliability evaluation of fault-tolerant systems—effect of variability in failure rates. IEEE Transactions on Computers, C-33(2):197–200, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676412.

Jain:1985:MTG

[JA85] S. K. Jain and V. D. Agrawal.

Modeling and test generation algorithms for MOS circuits. IEEE Transactions on Computers, C-34(5):426–433, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676582.

See corrections [Ano85d].

Jou:1988:FTF

[JA88] J.-Y. Jou and J. A. Abraham. Fault-tolerant FFT networks. IEEE Transactions on Computers, 37(5):548–561, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4606.

Janakiram:1988:RPB

[JAM88] V. K. Janakiram, D. P. Agrawal, and R. Mehrotra. A randomized parallel backtracking algorithm. IEEE Transactions on Computers, 37(12):1665–1676, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=9745.

Jansch:1988:DDS

[JC88] I. Jansch and B. Courtois. Definition and design of strongly language disjoint checkers. IEEE Transactions on Computers, 37 (6):745–748, June 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        2215.

Jenq:1981:DCA

[Jen81] Yih-Chyun Jenq. Digital convolution algorithm for pipelining multiprocessor systems. IEEE Transactions on Computers, C-30(12):966–973, December 1981. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675736.

Jenkins:1983:DEC

[Jen83]  W. K. Jenkins.  The de-

sign of error checkers for selfchecking residue number arithmetic. IEEE Transactions on Computers, C-32(4):388–396, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676240;
 http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35226.

Jesshope:1980:IFR

[Jes80a]  C. R. Jesshope.  The im-

plementation of fast radix 2 transforms on array processors. IEEE Transactions on Computers, C-29(1):20–27, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675452.

Jesshope:1980:SRC

[Jes80b] C. R. Jesshope. Some results concerning data routing in array processors. IEEE Transactions on Computers, C-29(7):659–662, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675638.

Jenkins:1980:COP

[JH80] J. H. Jenkins and J. A. Howard. Control overhead — A performance metric for evaluating control-unit designs. IEEE Transactions on Computers, C-29(4):300–308, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675568.

Johnsson:1989:OBP

[JH89] S. L. Johnsson and C.-T. Ho. Optimum broadcasting and personalized communication in hypercubes. IEEE Transactions on Computers, 38 (9):1249–1268, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29465.

Jha:1988:MSO

[Jha88] N. K. Jha. Multiple stuckopen fault detection in CMOS logic circuits. IEEE Transactions on Computers, 37(4):426–432, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2186.

Jian:1986:ASM

[Jia86]  Tang Jian.  An O(2^{0.304n})

algorithm for solving maximum independent set problem. IEEE Transactions on Computers, C-35(9):847–851, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676847.

Jansen:1980:DCM

[JK80] P. G. Jansen and J. L. W.

Kessels. The DIMOND: a component for the modular construction of switching networks. IEEE Transactions on Computers, C-29(10):884–889, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675471.

Jess:1982:DSP

[JK82] J. A. G. Jess and H. G. M.

Kees. A data structure for parallel L/U decomposition. IEEE Transactions on Computers, C-31(3):231–239, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675979.

Jagadish:1989:FNE

[JK89] H. V. Jagadish and T. Kailath.

A family of new efficient arrays for matrix multiplication. IEEE Transactions on Computers, 38(1):149–155, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8739.

Janus:1985:AMU

[JL85] Philip J. Janus and Edmund A. Lamagna. An adaptive method for unknown distributions in distributive partitioned sorting. IEEE Transactions on Computers, C-34(4):367–372, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009388.

Jafari:1980:SCR

[JLS80] H. Jafari, T. G. Lewis, and J. D. Spragins. Simulation of a class of ring-structured networks. IEEE Transactions on Computers, C-29(5):385–392, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675588.

Jahanian:1987:GTA

[JM87] Farnam Jahanian and Aloysius K.-L. Mok. A graphtheoretic approach for timing analysis and its implementation. IEEE Transactions on Computers, C-36(8):961–975, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009519.

Jagadish:1986:SPC

[JMKN86] H. V. Jagadish, R. G. Mathews, T. Kailath, and J. A. Newkirk. A study of pipelining in computing arrays. IEEE Transactions on Computers, C-35(5):431–440, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676785.

JaJa:1984:VSR

[JO84] Joseph Ja’Ja’ and Robert Michael Owens. VLSI sorting with reduced hardware. IEEE Transactions on Computers, C-33(7):668–671, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009340.

Johnson:1980:DQS

[Joh80]  E. L. Johnson.  A digi-

tal quarter square multiplier. IEEE Transactions on Computers, C-29(3):258–261, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675558.

Journeau:1986:NRR

[Jou86] M. Journeau. A note on the restricted range cutting algorithm. IEEE Transactions on Computers, C-35(1):73, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676662.

Jouppi:1989:NDI

[Jou89] N. P. Jouppi. The nonuniform distribution of instructionlevel and machine parallelism and its effect on performance. IEEE Transactions on Computers, 38(12):1645–1658, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=40844.

Jarwala:1988:TDM

[JP88] N. T. Jarwala and D. K. Pradhan. TRAM: a design methodology for high-performance, easily testable, multimegabit RAMs. IEEE Transactions on Computers, 37(10):1235–1250, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5985.

Jerrum:1984:FFD

[JS84] M. R. Jerrum and S. Skyum. Families of fixed degree graphs for processor interconnection. IEEE Transactions on Computers, C-33(2):190–194, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676410.

Jeng:1988:DAD

[JS88] M. Jeng and H. J. Siegel. Design and analysis of dynamic redundancy networks.

IEEE Transactions on Computers, 37(9):1019–1029, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2253.

Jaswa:1985:CCP

[JTP85] V. C. Jaswa, C. E. Thomas, and J. T. Pedicone. CPAC — concurrent processor architecture for control. IEEE Transactions on Computers, C-34(2):163–169, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676553.

Jullien:1980:IMM

[Jul80] G. A. Jullien. Implementation of multiplication, modulo a prime number, with applications to number theoretic transforms. IEEE Transactions on Computers, C-29(10):899–905, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675473.

Jaja:1989:NAR

[JW89a] J. Jaja and S.-M. Wu. A new approach to realizing partially symmetric functions. IEEE Transactions on Computers, 38(6):896–898, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24302.

Juang:1989:RSI

[JW89b] J.-Y. Juang and B. W. Wah. Resource sharing interconnection networks in multiprocessors. IEEE Transactions on Computers, 38(1):115–129, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8735.

Krishnamurthy:1984:CES

[KA84] Balakrishnan Krishnamurthy and Sheldon B. Akers. On the complexity of estimating the size of a test set. IEEE Transactions on Computers, C-33(8):750–753, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009364.

Kim:1987:MAT

[KA87] J. H. Kim and W. E. Alexander. A multiprocessor architecture for two-dimensional digital filters. IEEE Transactions on Computers, C-36(7):876–884, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676982.

Kim:1989:AFJ

[KA89] C. Kim and A. K. Agrawala. Analysis of the fork-join queue.

IEEE Transactions on Computers, 38(2):250–255, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16501.

Kung:1982:WAP

[KAGER82] Sun-Yuan Kung, K. S. Arun, R. J. Gal-Ezer, and D. V. Bhaskar Rao. Wavefront array processor: Language, architecture, and applications. IEEE Transactions on Computers, C-31(11):1054–1066, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675922.

Kak:1983:SRS

[Kak83] S. C. Kak. A structural redundancy in d-sequences. IEEE Transactions on Computers, C-32(11):1069–1070, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676160.

Kak:1985:EEC

[Kak85] Subhash C. Kak. Encryption and error-correction coding using D sequences. IEEE Transactions on Computers, C-34 (9):803–809, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676636.

Kallman:1983:FBC

[Kal83] R. Kallman. A faster 8-bit carry circuit. IEEE Transactions on Computers, C-32 (12):1209–1211, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676188.

Kamat:1980:CAF

[Kam80] P. S. Kamat. Comments on “An Algorithm for Fast Evaluation of Time Functions”. IEEE Transactions on Computers, C-29(4):333, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675575.

Kamal:1987:SLA

[Kam87] A. E. Kamal. Star local area networks: a performance study. IEEE Transactions on Computers, C-36(4):483–499, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676930.

Kanai:1981:IRM

[Kan81] T. Kanai. An improvement of reliability of memory system with skewing reconfiguration. IEEE Transactions on Computers, C-30(10):811–812, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675699.

Kaneda:1984:COW

[Kan84] Shigeo Kaneda. A class of odd-weight-column SEC–DED– SbED codes for memory system applications. IEEE Transactions on Computers, C-33(8):737–739, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009359.

Kant:1985:FIB

[Kan85]  K. Kant.  Finding inter-

ferences between rectangular paths. IEEE Transactions on Computers, C-34(11): 1045–1049, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676537.

Kapralski:1989:MMS

[Kap89] A. Kapralski. The maximum and minimum selector SEL-RAM and its application for developing fast sorting machines. IEEE Transactions on Computers, 38(11):1572–1577, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42127.

Karpovsky:1981:AED

[Kar81] M. Karpovsky. An approach for error detection and error correction in distributed systems computing numerical functions. IEEE Transactions on Computers, C-30(12):947–953, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675733.

Kartashev:1982:SCS

[Kar82] S. P. Kartashev. Supersystems: Current state-of-the-art Guest Editor’s introduction. IEEE Transactions on Computers, C-31(5):345–348, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676012.

Karpovksy:1983:UTD

[Kar83] M. Karpovksy. Universal tests for detection of input/output stuck-at and bridging faults. IEEE Transactions on Computers, C-32(12):1194–1198, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676182.

Karnin:1984:PAK

[Kar84a] E. D. Karnin. A parallel algorithm for the knapsack problem. IEEE Transactions on Computers, C-33(5):404–408, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676456.

Karp:1984:ELS

[Kar84b] A. H. Karp. Exponential and logarithm by sequential squaring. IEEE Transactions on Computers, C-33(5):462–464, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676464.

Karpovsky:1987:MLN

[Kar87] M. Karpovsky. Multilevel logical networks. IEEE Transactions on Computers, C-36 (2):215–226, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676884.

Katseff:1988:IH

[Kat88] H. P. Katseff. Incomplete hypercubes. IEEE Transactions on Computers, 37(5):604–608, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4611.

Koren:1984:AYC

[KB84] Israel Koren and Melvin A.

Breuer. On area and yield considerations for fault-tolerant VLSI processor arrays. IEEE Transactions on Computers, C-33(1):21–27, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009312.

Kwan:1985:PMM

[KB85] Sai Choi Kwan and Jean-Loup Baer. The I/O performance of multiway mergesort and tag sort. IEEE Transactions on Computers, C-34(4):383–387, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009392.

Kavi:1986:FDD

[KBB86] K. M. Kavi, B. P. Buckles, and U. N. Bhat. A formal definition of data flow graph models. IEEE Transactions on Computers, C-35(11):940–948, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676696.

Kleeman:1986:CRM

[KC86] L. Kleeman and A. Cantoni. Can redundancy and masking improve the performance of synchronizers? IEEE Transactions on Computers, C-35(7):643–646, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=1676806.

See comments and reply [EA89a].

Kleeman:1987:UMB

[KC87a] Lindsay Kleeman and Antonio Cantoni. On the unavoidability of metastable behavior in digital systems. IEEE Transactions on Computers, C-36(1):109–112, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009455.

Kurose:1987:LSS

[KC87b] James F. Kurose and Renu Chipalkatti. Load sharing in soft real-time distributed computer systems. IEEE Transactions on Computers, C-36(8):993–1000, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009522.

Kumar:1985:SSS

[KDJ85] M. Kumar, D. M. Dias, and J. R. Jump. Switching strategies in shuffle-exchange packetswitched networks. IEEE Transactions on Computers, C-34 (2):180–186, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676556.

Keedy:1983:ISE

[Kee83] J. L. Keedy. An instruction set for evaluating expressions. IEEE Transactions on Computers, C-32(5):476–478, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676258.

Kessels:1984:TDF

[Kes84] J. L. W. Kessels. Two designs of a fault-tolerant clocking system. IEEE Transactions on Computers, C-33(10):912–919, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676352.

Kandel:1980:PAF

[KF80] A. Kandel and J. M. Francioni. On the properties and applications of fuzzy-valued switching functions. IEEE Transactions on Computers, C-29 (11):986–994, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675493.

Kaneda:1982:SBE

[KF82] S. Kaneda and E. Fujiwara. Single byte error correcting– double byte error detecting codes for memory systems. IEEE Transactions on Computers, C-31(7):596–602, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676056.

Kameyama:1980:DDF

[KH80] M. Kameyama and T. Higuchi.

Design of dependent-failure-tolerant microcomputer system using triple-modular redundancy. IEEE Transactions on Computers, C-29 (2):202–206, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675545.

Kumar:1983:EIB

[KH83] M. Kumar and D. S. Hirschberg.

An efficient implementation of Batcher’s odd-even merge algorithm and its application in parallel sorting schemes. IEEE Transactions on Computers, C-32(3):254–264, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676217.

Krishnan:1986:ALM

[KH86] M. S. Krishnan and J. P.

Hayes. An array layout methodology for VLSI circuits. IEEE Transactions on Computers, C-35(12):1055–1067, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676713.

Kreutzer:1988:DDS

[KH88] S. E. Kreutzer and S. L. Hakimi. Distributed diagnosis and the system user. IEEE Transactions on Computers, 37(1):71–78, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=75141.

Khakbaz:1982:TSC

[Kha82]  J. Khakbaz.  Totally self-

checking checker for 1-out-of-n code using two-rail codes. IEEE Transactions on Computers, C-31(7):677–681, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676065.

Khakbaz:1984:TPD

[Kha84] Javad Khakbaz. A testable PLA design with low overhead and high fault coverage. IEEE Transactions on Computers, C-33(8):743–745, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009362.

Kang:1988:FDA

[KHS88] Y. J. Kang, J. H. Herzog, and J. Spragins. FISHNET: a distributed architecture for highperformance local computer networks. IEEE Transactions on Computers, 37(1):119–123, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75144.

Kim:1986:SDI

[Kim86]  M. Y. Kim.  Synchronized

disk interleaving. IEEE Transactions on Computers, C-35 (11):978–988, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676699.

Kitajima:1980:SCT

[Kit80] H. Kitajima. A symmetric cosine transform. IEEE Transactions on Computers, C-29(4):317–323, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675570.

Kumar:1986:PUS

[KJ86] Manoj Kumar and J. R.

Jump. Performance of unbuffered shuffle-exchange networks. IEEE Transactions on Computers, C-35(6):573–578, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009435.

Kung:1989:FTA

[KJC89] S.-Y. Kung, S.-N. Jean, and C.- W. Chang. Fault-tolerant array processors using single-track switches. IEEE Transactions on Computers, 38(4):501–514, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21143.

Kartashev:1980:PDS

[KK80a] S. I. Kartashev and S. P. Kartashev. Problems of designing supersystems with dynamic architectures. IEEE Transactions on Computers, C-29(12):1114–1132, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675517.

Kermani:1980:TSS

[KK80b] P. Kermani and L. Kleinrock. A tradeoff study of switching systems in computer communication networks. IEEE Transactions on Computers, C-29(12):1052–1060, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675510.

Kartashev:1982:DPS

[KK82] S. P. Kartashev and S. I. Kartashev. Distribution of programs for a system with dynamic architecture. IEEE Transactions on Computers, C-31(6):488–514, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676032.

Kartashev:1984:EIC

[KK84a] S. P. Kartashev and S. I.

Kartashev. Efficient internode communications in reconfigurable binary trees. IEEE Transactions on Computers, C-33(11):977–990, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676372.

Kartashev:1984:MAM

[KK84b] Svetlana P. Kartashev and Steven I. Kartashev. Memory allocations for multiprocessor systems that incorporate content-addressable memories. IEEE Transactions on Computers, C-33(1):28–44, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009313.

Kirrmann:1984:PPP

[KK84c] H. D. Kirrmann and F. Kaufmann. Poolpo — A pool of processors for process control applications. IEEE Transactions on Computers, C-33 (10):869–878, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676348.

Krawczyk:1985:AAD

[KK85] H. Krawczyk and M. Kubale. An approximation algorithm for diagnostic test scheduling in multicomputer systems. IEEE Transactions on Computers, C-34(9):869–872, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676647.

Krol:1986:CFT

[KK86] T. Krol and T. Krol. (N, K) concept fault tolerance. IEEE Transactions on Computers, C-35(4):339–349, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676767.

Kartashev:1987:ASD

[KK87] S. P. Kartashev and S. I. Kartashev. Analysis and synthesis of dynamic multicomputer networks that reconfigure into rings, trees, and stars. IEEE Transactions on Computers, C-36(7):823–844, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676979.

Krawczyk:1988:DMS

[KK88] H. Krawczyk and W. E. Kozlowski. On the diagnosability of multicomputer systems with homogeneous and incomplete tests. IEEE Transactions on Computers, 37(11):1419–1421, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8708.

Karmarkar:1989:IAD

[KK89] V. V. Karmarkar and J. G.

Kuhl. An integrated approach to distributed demand assignment in multiple-bus local networks. IEEE Transactions on Computers, 38(5):679–695, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24270.

Koren:1986:ACR

[KKS86] I. Koren, Z. Koren, and S. Y. H.

Su. Analysis of a class of recovery procedures. IEEE Transactions on Computers, C-35(8):703–712, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676821.

Kluge:1982:ORM

[KL82] W. E. Kluge and K. Lautenbach. The orderly resolution of memory access conflicts among competing channel processes. IEEE Transactions on Computers, C-31(3):194–207, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675975.

Kung:1987:OSD

[KLL87] Sun-Yuan Kung, Sheng-Chun Lo, and P. S. Lewis. Optimal systolic design for the transitive closure and the shortest path problems. IEEE Transactions on Computers, C-36(5):603–614, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676945.

Kluge:1983:CRM

[Klu83]  W. E. Kluge.  Cooperating

reduction machines. IEEE Transactions on Computers, C-32(11):1002–1012, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676151.

Koren:1981:CPN

[KM81] I. Koren and Y. Maliniak. On classes of positive, negative, and imaginary radix number systems. IEEE Transactions on Computers, C-30(5):312–317, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675788.

Kornerup:1983:FPR

[KM83a] Peter Kornerup and David W.

Matula. Finite precision rational arithmetic: An arithmetic unit. IEEE Transactions on Computers, C-32(4):378–388, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676239.

Krishnamurthy:1983:FID

[KM83b] E. V. Krishnamurthy and V. K. Murthy. Fast iterative division of p-adic numbers. IEEE Transactions on Computers, C-32(4):396–398, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676241.

Khakbaz:1984:STE

[KM84a] Javad Khakbaz and Edward J.

McCluskey. Self-testing embedded parity checkers. IEEE Transactions on Computers, C-33(8):753–756, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009365.

Kornerup:1984:CFP

[KM84b] Peter Kornerup and David W.

Matula. Correction to ‘‘Finite Precision Rational Arithmetic: An Arithmetic Unit”. IEEE Transactions on Computers, C-33(7):682, July 1984.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009346.

Kambayashi:1986:PWL

[KM86] Yahiko Kambayashi and Saburo Muroga. Properties of wired logic. IEEE Transactions on Computers, C-35(6):550–563, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009432.

King:1987:MSR

[KM87] P. J. B. King and I. Mitrani. Modeling a slotted ring local area network. IEEE Transactions on Computers, C-36(5):554–561, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676940.

Kennedy:1989:PSL

[KM89a] M. A. Kennedy and G. G. L. Meyer. The PMC system level fault model: cardinality properties of the implied faulty sets. IEEE Transactions on Computers, 38(3):478–480, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21136.

Kobayashi:1989:SGF

[KM89b] M. Kobayashi and M. H. Mac-Dougall. The stack growth function: cache line reference models. IEEE Transactions on Computers, 38(6):798–805, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24288.

Kasahara:1984:PMS

[KN84] H. Kasahara and S. Narita. Practical multiprocessor scheduling algorithms for efficient parallel processing. IEEE Transactions on Computers, C-33 (11):1023–1029, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676376.

Kopetz:1987:CSD

[KO87] Hermann Kopetz and Wilhelm Ochsenreiter. Clock synchronization in distributed real-time systems. IEEE Transactions on Computers, C-36(8):933–940, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009516.

Kobayashi:1983:DPI

[Kob83] M. Kobayashi. Dynamic profile of instruction sequences for the IBM System/370. IEEE Transactions on Computers, C-32(9):859–861, September 1983.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676334.

Kobayashi:1984:DCL

[Kob84] M. Kobayashi. Dynamic characteristics of loops. IEEE Transactions on Computers, C-33(2):125–132, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676404.

Kobayashi:1986:EST

[Kob86] M. Kobayashi. An empirical study of task switching locality in MVS. IEEE Transactions on Computers, C-35(8):720–731, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676823.

Kodek:1981:CEF

[Kod81] D. M. Kodek. Conditions for the existence of fast number theoretic transforms. IEEE Transactions on Computers, C-30(5):359–360, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675794.

Konard:1986:ECM

[Kon86] V. Konard. Efficient computation of the maximum of the sum of two sequences and applications. IEEE Transactions on Computers, C-35(7):651–653, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676809.

Koren:1986:CDA

[Kor86] I. Koren. Comments on “The Diogenes Approach to Testable Fault-Tolerant Arrays of Processors”. IEEE Transactions on Computers, C-35(1):93, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676668.

See [Ros83].

Klein:1980:AOC

[KP80a] P. Klein and M. S. Paterson.

Asymptotically optimal circuit for a storage access function. IEEE Transactions on Computers, C-29(8):737–738, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675657.

Kodandapani:1980:UBF

[KP80b] K. L. Kodandapani and D. K.

Pradhan. Undetectability of bridging faults and validity of stuck-at fault test sets. IEEE Transactions on Computers, C-29(1):55–59, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675457.

Koren:1987:MER

[KP87] I. Koren and D. K. Pradhan. Modeling the effect of redundancy on yield and performance of VLSI systems. IEEE Transactions on Computers, C-36(3):344–355, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676906.

Koubias:1988:FRP

[KP88] S. A. Koubias and G. D. Papadopoulos. Further results on the performance evaluation of the split channel reservation multiple access protocol ATP-2 for local area networks. IEEE Transactions on Computers, 37(3):376–383, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2178.

Kothari:1988:KNF

[KPR88] S. C. Kothari, G. M. Prabhu, and R. Roberts. The Kappa network with fault-tolerant destination tag algorithm. IEEE Transactions on Computers, 37(5): 612–617, May 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
4613.

Kamangar:1982:FAD

[KR82] F. A. Kamangar and K. R.

Rao. Fast algorithms for the 2-D discrete cosine transform. IEEE Transactions on Computers, C-31(9):899–906, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676108.

Kim:1989:DFT

[KR89a] J. H. Kim and S. M. Reddy.

On the design of fault-tolerant two-dimensional systolic arrays for yield enhancement. IEEE Transactions on Computers, 38(4):515–525, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21144.

Kumar:1989:FDP

[KR89b] V. P. Kumar and A. L. Reibman.

Failure dependent performance analysis of a fault-tolerant multistage interconnection network. IEEE Transactions on Computers, 38(12):1703–1713, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40848.

Kremla:1987:GCE

[Kre87] František Kremla. General criterion for essential nonfault locatability of logical functions. IEEE Transactions on Computers, C-36(5):623–629, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676948.

Krishnamurthy:1983:CHC

[Kri83]  E. V. Krishnamurthy.  On

the conversion of Hensel codes to Farey rationals. IEEE Transactions on Computers, C-32(4):331–337, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676233.

Krishnamurthy:1984:IMC

[Kri84] B. Krishnamurthy. An improved min-cut algorithm for partitioning VLSI networks. IEEE Transactions on Computers, C-33(5):438–446, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676460.

Krishnamurthy:1987:CTC

[Kri87] Balakrishnan Krishnamurthy. Constructing test cases for partitioning heuristics. IEEE Transactions on Computers, C-36(9):1112–1114, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009543.

Kruskal:1983:SMS

[Kru83]  C. P. Kruskal.  Searching,

merging, and sorting in parallel computation. IEEE Transactions on Computers, C-32 (10):942–946, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676138.

Karpovsky:1980:DLI

[KS80a] M. Karpovsky and S. Y. H.

Su. Detection and location of input and feedback bridging faults among input and output lines. IEEE Transactions on Computers, C-29(6):523–527, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675613.

Koren:1980:NAE

[KS80b] I. Koren and E. Sadeh. A new approach to the evaluation of the reliability of digital systems. IEEE Transactions on Computers, C-29(3):261–267, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675559.

Kini:1982:AGS

[KS82a] V. Kini and D. P. Siewiorek.

Automatic generation of symbolic reliability functions for processor-memory-switch structures. IEEE Transactions on Computers, C-31(8):752–771, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676082.

Kuck:1982:BSP

[KS82b] D. J. Kuck and R. A. Stokes. The Burroughs Scientific Processor (BSP). IEEE Transactions on Computers, C-31(5):363–376, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676014.

Kramer:1983:SIF

[KS83a] S. M. Kramer and D. P. Sidhu. Security information flow in multidimensional arrays. IEEE Transactions on Computers, C-32(12):1188–1191, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676180.

Kruskal:1983:PMI

[KS83b] C. P. Kruskal and M. Snir. The performance of multistage interconnection networks for multiprocessors. IEEE Transactions on Computers, C-32(12):1091–1098, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676169.

Kinoshita:1986:BTM

[KS86a] K. Kinoshita and K. K. Saluja.

Built-in testing of memory using an on-chip compact testing scheme. IEEE Transactions on Computers, C-35 (10):862–870, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676677.

Krishna:1986:STQ

[KS86b] C. M. Krishna and K. G.

Shin. On scheduling tasks with a quick recovery from failure. IEEE Transactions on Computers, C-35(5):448–455, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676787.

Kurose:1989:MAO

[KS89] J. F. Kurose and R. Simha.

A microeconomic approach to optimal resource allocation in distributed computer systems. IEEE Transactions on Computers, 38(5):705–717, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24272.

Krishna:1985:EFT

[KSB85] C. M. Krishna, K. G. Shin, and R. W. Butler. Ensuring fault tolerance of phase-locked clocks. IEEE Transactions on Computers, C-34(8):752–756, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676622.

Krishna:1987:PTD

[KSB87] C. M. Krishna, Kang G. Shin, and Inderpal S. Bhandari. Processor tradeoffs in distributed real-time systems. IEEE Transactions on Computers, C-36(9):1030–1040, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009534.

Kruskal:1988:DWT

[KSW88] Clyde P. Kruskal, Marc Snir, and Alan Weiss. The distribution of waiting times in clocked multistage interconnection networks. IEEE Transactions on Computers, 37(11):1337–1352, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8700.

Kljaich:1989:FVF

[KSW89] J. Kljaich, Jr., B. T. Smith, and A. S. Wojcik. Formal verification of fault tolerance using theorem-proving techniques.

IEEE Transactions on Computers, 38(3):366–376, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21123.

Kerkhoff:1981:MVL

[KT81] H. G. Kerkhoff and M. L. Tervoert. Multiple-valued logic charge-coupled devices. IEEE Transactions on Computers, C-30(9):644–652, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675862.

Kawaoka:1985:TPO

[KT85] T. Kawaoka and Y. Takahashi.

Test procedure optimization for layered protocol implementations. IEEE Transactions on Computers, C-34(1):94–97, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676521.

Kanakia:1987:DCL

[KT87] H. R. Kanakia and F. A. Tobagi. On distributed computations with limited resources. IEEE Transactions on Computers, C-36(5):517–528, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676936.

Krishnamurthy:1989:ITE

[KT89a] B. Krishnamurthy and I. G. Tollis. Improved techniques for estimating signal probabilities. IEEE Transactions on Computers, 38(7):1041–1045, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30854.

Kumar:1989:MAL

[KT89b] V. K. P. Kumar and Y.-C. Tsai. On mapping algorithms to linear and fault-tolerant systolic arrays. IEEE Transactions on Computers, 38(3):470–478, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21135.

Kubale:1982:CDP

[Kub82] M. Kubale. Comments on “Decomposition of Permutation Networks”. IEEE Transactions on Computers, C-31(3):265, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675983.

Kumar:1988:MPC

[Kum88] M. Kumar. Measuring parallelism in computation-intensive scientific/engineering applications. IEEE Transactions on Computers, 37(9):1088–1098, September 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
2259.

Kumar:1989:ALC

[Kum89]  A. Kumar.  Adaptive load

control of the central processor in a distributed system with a star topology. IEEE Transactions on Computers, 38(11):1502–1512, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42120.

Kuo:1987:GEP

[Kuo87] Y. S. Kuo. Generating essential primes for a Boolean function with multiple-valued inputs. IEEE Transactions on Computers, C-36(3):356–359, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676907.

Klatte:1985:ASS

[KUV85] R. Klatte, C. P. Ullrich, and J. W. Von Gudenberg. Arithmetic specification for scientific computation in ADA. IEEE Transactions on Computers, C-34(11):996–1005, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676532.

Kabat:1981:DVD

[KW81] W. C. Kabat and A. S. Wojcik. On the design of 4valued digital systems. IEEE Transactions on Computers, C-30(9):666–671, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675865.

Kabat:1985:ASC

[KW85] W. C. Kabat and A. S. Wojcik. Automated synthesis of combinational logic using theorem-proving techniques. IEEE Transactions on Computers, C-34(7):610–632, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676600.

Kim:1989:DER

[KW89] K. H. Kim and H. O. Welch. Distributed execution of recovery blocks: an approach for uniform treatment of hardware and software faults in real-time applications. IEEE Transactions on Computers, 38(5):626–636, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24266.

Keichafer:1988:MAD

[KWFT88] R. M. Keichafer, C. J. Walter, A. M. Finn, and P. M.

Thambidurai. The MAFT architecture for distributed fault tolerance. IEEE Transactions on Computers, 37(4):398–404, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2183.

Kushner:1982:IPZ

[KWR82] T. Kushner, A. Y. Wu, and A. Rosenfeld. Image processing on ZMOB. IEEE Transactions on Computers, C-31 (10):943–951, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675903.

Kulkarni:1982:SPI

[KY82] A. V. Kulkarni and D. W. L.

Yen. Systolic processing and an implementation for signal and image processing. IEEE Transactions on Computers, C-31(10):1000–1009, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675909.

Kim:1989:PIL

[KY89] K. H. Kim and S. M. Yang.

Performance impacts of lookahead execution in the conversation scheme. IEEE Transactions on Computers, 38(8):1188–1202, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30872.

Kita:1980:DBM

[KYSY80] Y. Kita, N. Yamaguchi, M. Sugie, and S. Yoshizama. The development of a bubble memory controller for low-cost file use. IEEE Transactions on Computers, C-29(2):89–96, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675531.

Lo:1985:GPB

[LA85] Hao-Yung Lo and Y. Aoki. Generation of a precise binary logarithm with difference grouping programmable logic array. IEEE Transactions on Computers, C-34(8):681–691, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676614.

Lee:1987:MSP

[LA87] Soo-Young Lee and J. K. Aggarwal. A mapping strategy for parallel processing. IEEE Transactions on Computers, C-36(4):433–442, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676925.

Lakani:1984:IDA

[Lak84] G. D. Lakani. An improved distribution algorithm for shortest paths problem. IEEE Transactions on Computers, C-33 (9):855–857, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676503.

Lam:1980:PBN

[Lam80] S. S. Lam. Packet broadcast networks — A performance analysis of the R-ALOHA protocol. IEEE Transactions on Computers, C-29(7):596–603, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675630.

Lam:1983:SDM

[Lam83] S. S. Lam. A simple derivation of the MVA and LBANC algorithms from the convolution algorithm. IEEE Transactions on Computers, C-32 (11):1062–1064, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676157.

Lange:1987:ITA

[Lan87] Eberhard Lange. Implementation and test of the ACRITH facility in a System/370. IEEE Transactions on Computers, C-36(9):1088–1096, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009539.

Li:1987:RDS

[LAS87] Zhiyuan Li and Walid Abu-Sufah. On reducing data synchronization in multiprocessed loops. IEEE Transactions on Computers, C-36 (1):105–109, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009454.

Lauther:1981:ACO

[Lau81] U. Lauther. Additional comments on “An Optimal Solution for the Channel-Assignment Problem”. IEEE Transactions on Computers, C-30(6):455, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675818.

See [Rub81, PTT81].

Lin:1988:TDE

[LB88] D. J. Lin and B. Bose. Theory and design of t-error correcting and d (d > t)-unidirectional error detecting (t-EC d-UED) codes. IEEE Transactions on Computers, 37(4):433–439, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2187.

Liu:1987:PAM

[LC87a] Steven S. Liu and Q. C.

Chow. Performance analysis of a multiprocessor-based packet switch in networks with linklevel sliding-window flow control. IEEE Transactions on Computers, C-36(9):1128–1132, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009547.

Lo:1987:HGA

[LC87b] Hao-Yung Lo and Jau-Ling Chen. A hardwired generalized algorithm for generating the logarithm base-k by iteration. IEEE Transactions on Computers, C-36(11):1363–1367, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009477.

Li:1989:LCE

[LCF89] C.-C. J. Li, P. P. Chen, and W. K. Fuchs. Local concurrent error detection and correction in data structures using virtual backpointers. IEEE Transactions on Computers, 38(11):1481–1492, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42119.

Long:1989:ERR

[LCS89] D. D. E. Long, J. L. Carroll, and K. Stewart. Estimating the reliability of regenerationbased replica control protocols. IEEE Transactions on Computers, 38(12):1691–1702, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=40847.

Lee:1981:CCS

[LCW81] D. T. Lee, Hsu Chang, and C. K. Wong. An on-chip compare/steer bubble sorter. IEEE Transactions on Computers, C-30(6):396–405, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675805.

Lee:1987:ATS

[LD87] Insup Lee and Susan B. Davidson. Adding time to synchronous process communications. IEEE Transactions on Computers, C-36(8):941–948, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009517.

Lee:1988:AOR

[LD88] D.-L. Lee and W. A. Davis. An O(n + k) algorithm for ordered retrieval from an associative memory. IEEE Transactions on Computers, 37(3):368–371, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2176.

Liu:1989:NOH

[LD89] L. C. Liu and H. C. Du.

A near-optimal heuristic algorithm for single-flow routing. IEEE Transactions on Computers, 38(4):603–608, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21154.

Lesser:1980:DIM

[LE80] V. R. Lesser and L. D. Erman. Distributed interpretation: a model and experiment. IEEE Transactions on Computers, C-29(12):1144–1163, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675519.

Lea:1986:LSB

[Lea86]  C.-T. A. Lea.  The load-

sharing banyan network. IEEE Transactions on Computers, C-35(12):1025–1034, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676710.

Lee:1980:QAG

[Lee80] Chin-Hwa Lee. Queueing analysis of global locking synchronization schemes for multicopy databases. IEEE Transactions on Computers, C-29(5):371–384, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675587.

Lee:1982:NNV

[Lee82] Der-Tsai Lee. On k-nearest neighbor Voronoi diagrams in the plane. IEEE Transactions on Computers, C-31(6):478–487, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676031.

Lee:1985:CCS

[Lee85a] D. Lee. Comparator with completion signal. IEEE Transactions on Computers, C-34 (9):855–857, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676642.

Lee:1985:RSP

[Lee85b] Kyungsook Yoon Lee. On the rearrangeability of 2(log_{2} N) − 1 stage permutation networks. IEEE Transactions on Computers, C-34(5):412–425, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676581.

Lee:1987:NBN

[Lee87] Kyungsook Yoon Lee. A new Benes network control algorithm. IEEE Transactions on Computers, C-36(6):768–772, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676970.

Lee:1989:MGR

[Lee89a] Corinna Lee. Multistep gradual rounding. IEEE Transactions on Computers, 38(4):595– 600, April 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://dlib.computer.org/
tc/books/tc1989/pdf/t0593.
pdf;   http://ieeexplore.
ieee.org/stamp/stamp.jsp?
tp=&arnumber=21152;   http:
//www.computer.org/tc/tc1989/
t0595abs.htm.

Lee:1989:PBM

[Lee89b] G. Lee. A performance bound of multistage combining networks. IEEE Transactions on Computers, 38(10):1387–1395, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35834.

Leiss:1984:DID

[Lei84] E. L. Leiss. Data integrity in digital optical disks. IEEE Transactions on Computers, C-33(9):818–827, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676498.

Leighton:1985:TBC

[Lei85a] Tom Leighton. Tight bounds on the complexity of parallel sorting. IEEE Transactions on Computers, C-34(4):344–354, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009385.

Leiserson:1985:FTU

[Lei85b] C. E. Leiserson. Fat-trees: Universal networks for hardwareefficient supercomputing. IEEE Transactions on Computers, C-34(10):892–901, October 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lenfant:1985:VMM

[Len85]  Jacques Lenfant.  A versa-

tile mechanism to move data in an array processor. IEEE Transactions on Computers, C-34(6):506–522, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009403.

Lenders:1988:GMP

[Len88] P. M. Lenders. A generalized message-passing mechanism for communicating sequential processes. IEEE Transactions on Computers, 37(6):646–651, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2203.

Lenahan:1980:PCL

[LF80] J. J. Lenahan and F. K.

Fung. Performance of cooperative loosely coupled microprocessor architectures in an interactive data base task. IEEE Transactions on Computers, C-29(2):161–180, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675540.

LeBlanc:1985:HMS

[LF85] T. J. LeBlanc and S. A. Friedberg. HPC: A model of structure and change in distributed systems. IEEE Transactions on Computers, C-34(12):1114–1129, December 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://academic.research.
microsoft.com/Publication/
713842/hpc-a-model-of-structure-
and-change-in-distributed-
systems.

Lowrie:1987:RTA

[LF87] M. B. Lowrie and W. K. Fuchs. Reconfigurable tree architectures using subtree oriented fault tolerance. IEEE Transactions on Computers, C-36(10):1172–1182, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676858.

Lee:1980:RCP

[LGH80] P. A. Lee, N. Ghani, and K. Heron. A recovery cache for the PDP-11. IEEE Transactions on Computers, C-29(6):546–549, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675618.

Lee:1988:ESG

[LH88] Kyungsook Yoon Lee and W. Hegazy. The extra stage Gamma network. IEEE Transactions on Computers, 37(11): 1445–1450, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8715.

Liu:1989:HCS

[LHC89] W. Liu, T. H. Hildebrandt, and R. Cavin III. Hamiltonian cycles in the shuffle-exchange network. IEEE Transactions on Computers, 38(5):745–750, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24277.

Loucks:1985:SPT

[LHPW85] W. M. Loucks, V. C. Hamacher, B. R. Preiss, and L. Wong. Short-packet transfer performance in local area ring networks. IEEE Transactions on Computers, C-34(11): 1006–1014, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676533.

Li:1985:FCD

[Li85] S.-Y. R. Li. Fast constant division routines. IEEE Transactions on Computers, C-34 (9):866–869, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676646.

Li:1987:PMT

[Li87] Victor O. K. Li. Performance models of timestamp-ordering concurrency control algorithms in distributed databases. IEEE Transactions on Computers, C-36(9):1041–1051, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009535.

Lincoln:1982:TDT

[Lin82]  N. R. Lincoln.  Technology

and design tradeoffs in the creation of a modern supercomputer. IEEE Transactions on Computers, C-31(5):349–362, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676013.

Liu:1982:AVD

[Liu82]  K. Y. Liu.  Architecture

for VLSI design of Reed– Solomon encoders. IEEE Transactions on Computers, C-31 (2):170–175, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675964.

Liu:1984:AVD

[Liu84a] Kuang Yung Liu. Architecture for VLSI design of Reed– Solomon decoders. IEEE Transactions on Computers, C-33 (2):178–189, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676409.

Liu:1984:RMP

[Liu84b] T. S. Liu. The role of a maintenance processor for a general-purpose computer system. IEEE Transactions on Computers, C-33(6):507–517, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676474.

Liu:1987:EMB

[LJ87] Yu-Cheng Liu and Chi-Jiunn Jou. Effective memory bandwidth and processor blocking probability in multiple-bus systems. IEEE Transactions on Computers, C-36(6):761–764, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676968.

Li:1989:RFT

[LJL89] H. F. Li, R. Jayakumar, and C. Lam. Restructuring for fault-tolerant systolic arrays. IEEE Transactions on Computers, 38(2):307–311, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16511.

Leinward:1980:DBA

[LL80] S. Leinward and T. Lamdan. Dynamic Boolean algebras. IEEE Transactions on Computers, C-29(10):950–953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675483.

Lam:1981:CCP

[LL81] S. S. Lam and Y. C. L. Lien. Congestion control of packet communication networks by input buffer limits — A simulation study. IEEE Transactions on Computers, C-30 (10):733–742, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675692.

Laning:1983:FAD

[LL83] L. J. Laning and M. S. Leonard. File allocation in a distributed computer communication network. IEEE Transactions on Computers, C-32(3):232–244, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676215.

Lee:1984:DCA

[LL84] D. T. Lee and Joseph Y-T. Leung. On the 2-dimensional channel assignment problem. IEEE Transactions on Computers, C-33(1):2–6, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009310.

Leighton:1985:WSI

[LL85] T. Leighton and C. E. Leiserson. Wafer-scale integration of systolic arrays. IEEE Transactions on Computers, C-34(5):448–461, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676584.

Lee:1988:ADM

[LL88] K. Y. Lee and D. Lee. On the augmented data manipulator network in SIMD environments. IEEE Transactions on Computers, 37(5):574–584, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4608.

Lam:1989:STA

[LLJ89] C. W. H. Lam, H. F. Li, and R. Jayakumar. A study of two approaches for reconfiguring fault-tolerant systolic arrays. IEEE Transactions on Computers, 38(6):833–844, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24292.

Lai:1982:LNC

[LM82a] Hung Chi Lai and S. Muroga.

Logic networks of carry-save adders. IEEE Transactions on Computers, C-31(9):870–882, September 1982. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1676102.

Levendel:1982:TGA

[LM82b] Y. H. Levendel and P. R. Menon. Test generation algorithms for computer hardware description languages. IEEE Transactions on Computers, C-31(7):577–588, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676054.

Lin:1983:TPC

[LM83] G. K. Lin and P. R. Menon. Totally preset checking experiments for sequential machines. IEEE Transactions on Computers, C-32(2):101–108, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676195.

Livny:1985:DCA

[LM85] M. Livny and U. Manber. Distributed computation via active messages. IEEE Transactions on Computers, C-34(12):1185–1190, December 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lui:1986:SST

[LM86] P. K. Lui and J. C. Muzio. Spectral signature testing of multiple stuck-at faults in irredundant combinational networks. IEEE Transactions on Computers, C-35(12):1088–1092, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676719.

Lai:1987:LNM

[LM87a] Hung Chi Lai and S. Muroga.

Logic networks with a minimum number of NOR(NAND) gates for parity functions of n variables. IEEE Transactions on Computers, C-36 (2):157–166, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676878.

Lee:1987:SSS

[LM87b] Edward Ashford Lee and David G.

Messerschmitt. Static scheduling of synchronous data flow programs for digital signal processing. IEEE Transactions on Computers, C-36(1):24–35, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009446.

LEcuyer:1988:COC

[LM88a] P. L’Ecuyer and J. Malenfant. Computing optimal checkpointing strategies for rollback and recovery systems. IEEE Transactions on Computers, 37(4):491–496, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2197.

Lin:1988:SHL

[LM88b] C. Lin and D. C. Marinescu. Stochastic high-level Petri nets and applications. IEEE Transactions on Computers, 37(7):815–825, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2227.

Li:1989:PTN

[LM89] H. Li and M. Maresca. Polymorphictorus network. IEEE Transactions on Computers, 38(9): 1345–1351, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29479.

Leblanc:1987:DPP

[LMC87] T. J. Leblanc and J. M. Mellor-Crummey. Debugging parallel programs with instant replay. IEEE Transactions on Computers, C-36(4):471–482, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676929.

Lampson:1984:IFU

[LMO84] Butler W. Lampson, Gene Mc-Daniel, and Severo M. Ornstein. An instruction fetch unit for a high-performance personal computer. IEEE Transactions on Computers, C-33(8):712–730, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009357.

Lacroix:1982:CAB

[LMP82] G. Lacroix, P. Marchegay, and G. Piel. Comments on “The Anomalous Behavior of Flip-Flops in Synchronizer Circuits”. IEEE Transactions on Computers, C-31(1):77–78, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675889.

Lo:1988:HAT

[Lo88]  Virginia Mary Lo.  Heuris-

tic algorithms for task assignment in distributed systems. IEEE Transactions on Computers, 37(11):1384–1397, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8704.

Lopriore:1984:CBT

[Lop84] L. Lopriore. Capability based tagged architectures. IEEE Transactions on Computers, C-33(9):786–803, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676495.

Lozier:1983:UFP

[Loz83] Daniel W. Lozier. The use of floating-point and interval arithmetic in the computation of error bounds. IEEE Transactions on Computers, C-32(4):411–417, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676245.

Labetoulle:1981:HTR

[LP81] J. Labetoulle and G. Pujolle. HDLC throughput and response time for bidirectional data flow with nonuniform frame sizes. IEEE Transactions on Computers, C-30(6):405–413, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675806.

Lozano-Perez:1983:SPC

[LP83] T. Lozano-Perez. Spatial planning: a configuration space approach. IEEE Transactions on Computers, C-32 (2):108–120, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676196.

Lee:1984:CGS

[LP84] D. T. Lee and F. P. Preparata. Computational geometry — A survey. IEEE Transactions on Computers, C-33(12):

1072–1101, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676388.

See correction [LP85].

Lee:1985:CCG

[LP85] D. T. Lee and Franco P.

Preparata. Correction to “Computational Geometry — A Survey”. IEEE Transactions on Computers, C-34(6):584, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009417.

See [LP84].

Lodi:1986:VSV

[LP86]  E. Lodi and L. Pagli.  A

VLSI solution to the vertical segment visibility problem. IEEE Transactions on Computers, C-35(10):923–928, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676685.

Luk:1988:FTM

[LP88] F. T. Luk and H. Park.

Fault-tolerant matrix triangularizations on systolic arrays. IEEE Transactions on Computers, 37(11):1434–1438, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8712.

Luk:1989:PCT

[LP89] F. T. Luk and H. Park. A proof of convergence for two parallel Jacobi SVD algorithms. IEEE Transactions on Computers, 38(6):806–811, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24289.

Laha:1988:ALC

[LPI88] S. Laha, J. H. Patel, and R. K. Iyer. Accurate low-cost methods for performance evaluation of cache memory systems. IEEE Transactions on Computers, 37(11):1325–1336, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8699.

Lotti:1980:AAA

[LR80] G. Lotti and F. Romani. Application of approximating algorithms to Boolean matrix multiplication. IEEE Transactions on Computers, C-29 (10):927–928, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675476.

Lesser:1980:EDT

[LS80] J. D. Lesser and J. J. Shedletsky. An experimental delay test generator for LSI logic. IEEE Transactions on Computers, C-29(3):235–248, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675555.

Lewis:1981:I

[LS81] T. G. Lewis and B. D.

Shriver. Introduction. IEEE Transactions on Computers, C-30(7):457–459, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675825.

Lee:1982:OIR

[LS82a] D. T. Lee and C. B. Silio, Jr.

An optimal illumination region algorithm for convex polygons. IEEE Transactions on Computers, C-31(12):1225–1227, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675946.

Leland:1982:DTG

[LS82b] W. E. Leland and M. H.

Solomon. Dense trivalent graphs for processor interconnection. IEEE Transactions on Computers, C-31(3):219–222, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675977.

Lee:1984:DEF

[LS84] Yann-Hang Lee and K. G. Shin. Design and evaluation of a fault-tolerant multiprocessor using hardware recovery blocks. IEEE Transactions on Computers, C-33(2):113–124, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676403.

See comments [CC89].

Lee:1988:SAR

[LS88a] I. Lee and D. Smitley. A synthesis algorithm for reconfigurable interconnection networks. IEEE Transactions on Computers, 37(6):691–699, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2207.

Lin:1988:MAI

[LS88b] L.-S. Lin and S. Sahni. Maximum alignment of interchangeable terminals. IEEE Transactions on Computers, 37 (10):1166–1177, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5978.

Lin:1989:SSC

[LS89a] H. P. Lin and H. E. Stoval, III.

Self-synchronizing communication protocols. IEEE Transactions on Computers, 38(5):609– 625, May 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
24265.

Lin:1989:FED

[LS89b] L. Lin and S. Sahni. Fair edge deletion problems. IEEE Transactions on Computers, 38(5): 756–761, May 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
24280.

Lipsky:1989:SPT

[LS89c] L. Lipsky and S. C. Seth. Signal probabilities in AND-OR trees. IEEE Transactions on Computers, 38(11):1558–1563, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42125.

Liu:1989:PTP

[LS89d] J.-C. Liu and K. G. Shin.

Polynomial testing of packet switching networks. IEEE Transactions on Computers, 38 (2):202–217, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16497.

Lin:1989:CFR

[LSD^{+}89] W. Lin, T.-L. Sheu, C. R. Das, T.-Y. Feng, and C.-L. Wu. A conflict free routing scheme on multistage interconnection networks. IEEE Transactions on Computers, 38(8):1086–1097, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30864.

Liaw:1981:TED

[LSM81] Chi-Chang Liaw, S. Y. H. Su, and Y. K. Malaiya. Testexperiments for detection and location of intermittent faults in sequential circuits. IEEE Transactions on Computers, C-30(12):989–995, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675743.

Lang:1985:SSM

[LSSS85] H.-W. Lang, M. Schimmler, H. Schmeck, and H. Schroder. Systolic sorting on a meshconnected network. IEEE Transactions on Computers, C-34(7):652–658, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676603.

Luk:1987:MAW

[LSW87] W. K. Luk, P. Sipala, and C. K. Wong. Minimum-area wiring for slicing structures. IEEE Transactions on Computers, C-36(6):745–760, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676967.

Lu:1982:WPS

[Lu82] D. J. Lu. Watchdog processors and structural integrity checking. IEEE Transactions on Computers, C-31(7):681–685, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676066.

Lundstrom:1987:ACS

[Lun87] Stephen F. Lundstrom. Applications considerations in the system design of highly concurrent multiprocessors. IEEE Transactions on Computers, C-36(11):1292–1309, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009469.

Lawrie:1982:PMS

[LV82] D. H. Lawrie and C. R. Vora.

The prime memory system for array access. IEEE Transactions on Computers, C-31(5):435–442, May 1982. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676020.

See comments [Ten83].

Lindstrom:1985:DAB

[LV85] E. E. Lindstrom and J. S. Vitter. The design and analysis of bucketsort for bubble memory secondary storage. IEEE Transactions on Computers, C-34(3):218–233, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676565.

Lang:1982:BCM

[LVA82] T. Lang, M. Valero, and I. Alegre. Bandwidth of crossbar and multiple-bus connections for multiprocessors. IEEE Transactions on Computers, C-31(12):1227–1234, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675947.

Lang:1983:RCM

[LVF83] T. Lang, M. Valero, and M. A. Fiol. Reduction of connections for multibus organization. IEEE Transactions on Computers, C-32(8):707–716, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676308.

Lindstrom:1985:IS

[LVW85] Eugene E. Lindstrom, Jeffrey Scott Vitter, and C. K. Wong. Introduction — sorting. IEEE Transactions on Computers, C-34(4):293–295, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009378.

Li:1985:DOS

[LW85] Guo-Jie Li and B. W. Wah.

The design of optimal systolic arrays. IEEE Transactions on Computers, C-34(1):66–77, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676516.

Li:1986:CAP

[LW86a] Guo-Jie Li and Benjamin W.

Wah. Coping with anomalies in parallel branch-and-bound algorithms. IEEE Transactions on Computers, C-35(6):568–573, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009434.

Lin:1986:RPP

[LW86b] Woei Lin and Chuan-Lin Wu.

Reconfiguration procedures for a polymorphic and partitionable multiprocessor. IEEE Transactions on Computers, C-35(10):910–916, October 1986.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676683.

Lin:1988:DRM

[LW88] W. Lin and C.-L. Wu. A distributed resource management mechanism for a partitionable multiprocessor system. IEEE Transactions on Computers, 37(2):201–210, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2149.

Lin:1989:FTM

[LW89] Woei Lin and C.-L. Wu. A fault-tolerant mapping scheme for a configurable multiprocessor system. IEEE Transactions on Computers, 38 (2):227–237, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16499.

Liu:1983:VAD

[LY83] P. S. Liu and T. Y. Young. VLSI array design under constraint of limited I/O bandwidth. IEEE Transactions on Computers, C-32(12):1160–1170, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676177.

Lee:1989:PIN

[LY89] K. Y. Lee and H. Yoon. The PM221 interconnection network. IEEE Transactions on Computers, 38(2):302–307, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16510.

Lee:1987:PTR

[LYI87] Yann-Hang Lee, Philip S. Yu, and Balakrishna R. Iyer. Progressive transaction recovery in distributed DB/DC systems. IEEE Transactions on Computers, C-36(8):976–987, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009520.

Yang:1988:DAF

[lYM88] Che liang Yang and G. M.

Masson. A distributed algorithm for fault diagnosis in systems with soft failures. IEEE Transactions on Computers, 37(11):1476–1480, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8723.

Lang:1985:ICL

[LZLH85] J. H. Lang, C. A. Zukowski, R. O. Lamaire, and Chae Han. Integrated-circuit logarithmic arithmetic units. IEEE Transactions on Computers, C-34(5):475–483, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676588.

Lu:1989:DAS

[LZM89] M. Lu, D. Zhang, and T. Murata. A design approach for self-diagnosis of fault-tolerant clock synchronization. IEEE Transactions on Computers, 38 (9):1337–1341, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29477.

Mangir:1982:FTD

[MA82] T. E. Mangir and A. Avizienis. Fault-tolerant design for VLSI: Effect of interconnect requirements on yield improvement of VLSI designs. IEEE Transactions on Computers, C-31(7):609–616, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676058.

McDowell:1986:PSL

[MA86] C. E. McDowell and W. F. Appelbe. Processor scheduling for linearly connected parallel processors. IEEE Transactions on Computers, C-35(7):632–638, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676804.

Molyneaux:1989:CTS

[MA89] R. F. Molyneaux and A. Albicki. Comments on “Ternary scan design for VLSI testability” by M. Hu and K. C. Smith. IEEE Transactions on Computers, 38(2):256–263, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16502.

See [HS86].

Maher:1980:FTE

[Mah80]  D. P. Maher.  On Fourier

transforms over extensions of finite rings. IEEE Transactions on Computers, C-29(4):331–333, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675574.

Majerski:1985:SRA

[Maj85] S. Majerski. Square-rooting algorithms for high-speed digital circuits. IEEE Transactions on Computers, C-34(8):724–733, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676618.

Manber:1980:SDR

[Man80] U. Manber. System diagnosis with repair. IEEE Transactions on Computers, C-29 (10):934–937, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675479.

Manner:1984:HTP

[Man84] Reinhard Manner. Hardware task/processor scheduling in a polyprocessor environment. IEEE Transactions on Computers, C-33(7):626–636, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009336.

Mannila:1985:MPO

[Man85] Heikki Mannila. Measures of presortedness and optimal sorting algorithms. IEEE Transactions on Computers, C-34(4):318–325, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009382.

Mandelbaum:1988:SAS

[Man88] D. M. Mandelbaum. On subsequences of arithmetic sequences. IEEE Transactions on Computers, 37(10):1314–1315, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5997.

Mandelbaum:1989:IAE

[Man89] D. M. Mandelbaum. On iterative arrays for the Euclidean algorithm over finite fields. IEEE Transactions on Computers, 38(10):1473–1478, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35844.

Markowsky:1981:STC

[Mar81]  G. Markowsky.  Syndrome-

testability can be achieved by circuit modification. IEEE Transactions on Computers, C-30(8):604–606, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675848.

Markenscoff:1984:DME

[Mar84] P. Markenscoff. A deterministic model for evaluating the performance of a multiple processor system with a shared bus. IEEE Transactions on Computers, C-33(3):281–285, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676426.

Marsaglia:1985:NPT

[Mar85] George Marsaglia. Note on a proposed test for random number generators. IEEE Transactions on Computers, C-34(8):756–758, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676623.

Maruoka:1986:CBP

[Mar86] A. Maruoka. Complexity based on partitioning of Boolean circuits and their relation to multivalued circuits. IEEE Transactions on Computers, C-35 (2):115–123, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676729.

Markowsky:1987:BSP

[Mar87] G. Markowsky. Bounding signal probabilities in combinational circuits. IEEE Transactions on Computers, C-36 (10):1247–1251, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676866.

Mudge:1984:MIM

[MAS84] T. N. Mudge and H. B. Al-Sadoun. Memory interference models with variable connection time. IEEE Transactions on Computers, C-33 (11):1033–1038, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676378.

Mudge:1985:SMM

[MAS85] T. N. Mudge and H. B. Al-Sadoun. A semi-Markov model for the performance of multiplebus systems. IEEE Transactions on Computers, C-34(10): 934–942, October 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mueller:1984:CHW

[MAV84] R. A. Mueller, V. H. Allan, and J. Varghese. The complexity of horizontal word encoding in microprogrammed machines. IEEE Transactions on Computers, C-33(10):938–939, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676358.

Maxwell:1988:CAD

[Max88]  P. C. Maxwell.  Compara-

tive analysis of different implementations of multiple-input signature analyzers. IEEE Transactions on Computers, 37(11):1411–1414, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8706.

Mazumder:1987:ECS

[Maz87] P. Mazumder. Evaluation of on-chip static interconnection networks. IEEE Transactions on Computers, C-36(3):365–369, March 1987. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676910.

Mathialagan:1980:OID

[MB80] A. Mathialagan and N. N. Biswas. Optimal interconnections in the design of microprocessors and digital systems. IEEE Transactions on Computers, C-29(2):145–149, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675538.

Marsan:1982:CPA

[MBC82] M. A. Marsan, G. Balbo, and G. Conte. Comparative performance analysis of single bus multiprocessor architectures. IEEE Transactions on Computers, C-31(12):1179–1191, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675942.

Marsan:1983:MBC

[MBCG83] M. A. Marsan, G. Balbo, G. Conte, and F. Gregoretti. Modeling bus contention and memory interference in a multiprocessor system. IEEE Transactions on Computers, C-32(1):60–72, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676124.

Milutinovic:1989:MSD

[MBH89] V. Milutinovic, M. Bettinger, and W. Helbig. Multiplier/shifter design tradeoffs in a 32-bit microprocessor. IEEE Transactions on Computers, 38(6):874–880, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24298.

McCluskey:1981:DAT

[MBN81] E. J. McCluskey and S. Bozorgui-Nesbat. Design for autonomous test. IEEE Transactions on Computers, C-30 (11):866–875, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675717.

Mikhail:1982:RMS

[MBR82] W. F. Mikhail, R. W. Bartoldus, and R. A. Rutledge. The reliability of memory with single-error correction. IEEE Transactions on Computers, C-31(6):560–564, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676042.

Mangin:1986:CPC

[MC86] J. L. Mangin and K. W.

Current. Characteristics of prototype CMOS quaternary logic encoder-decoder circuits. IEEE Transactions on Computers, C-35(2):157–161, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676733.

McColl:1981:PC

[McC81] W. F. McColl. Planar crossovers. IEEE Transactions on Computers, C-30(3):223–225, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675758.

McCluskey:1984:VTP

[McC84]  E. J. McCluskey.  Verifica-

tion testing—A pseudoexhaustive test technique. IEEE Transactions on Computers, C-33(6):541–546, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676477.

McCulloch:1985:WME

[McC85] C. M. McCulloch. Why merge? — an examination of disk sorting strategy. IEEE Transactions on Computers, C-34(4):387–391, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009393.

McGraw:1980:DFC

[McG80] J. R. McGraw. Data flow computing — software development. IEEE Transactions on Computers, C-29(12):1095–1103, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675515.

Maxemchuk:1986:ODS

[MD86] N. F. Maxemchuk and A. T.

Dahbura. Optimal diagnosable system design using fulldifference triangles. IEEE Transactions on Computers, C-35 (9):837–839, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676844.

Muntz:1989:BAR

[MDG89] R. R. Muntz, E. De Souza e Silva, and A. Goyal. Bounding availability of repairable systems. IEEE Transactions on Computers, 38(12): 1714–1723, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40849.

Melhem:1987:SDI

[Mel87] Rami G. Melhem. A study of data interlock in computational networks for sparse matrix multiplication. IEEE Transactions on Computers, C-36(9):1101–1107, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009541.

Melham:1989:SAI

[Mel89] R. Melham. A systolic accelerator for the iterative solution of sparse linear systems. IEEE Transactions on Computers, 38(11):1591–1595, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42132.

Mendez:1984:BJA

[Men84] R. H. Mendez. Benchmarks on Japanese and American supercomputers — preliminary results. IEEE Transactions on Computers, C-33(4):374, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676446.

Metzner:1982:CEM

[Met82] J. J. Metzner. Convolutionally encoded memory protection. IEEE Transactions on Computers, C-31(6):547–551, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676038.

Metzner:1983:PSL

[Met83]  J. J. Metzner.  A parity

structure for large remotely located replicated data files. IEEE Transactions on Computers, C-32(8):727–730, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676310.

Meyer:1980:EPD

[Mey80]  J. F. Meyer.  On evalu-

ating the performability of degradable computing systems. IEEE Transactions on Computers, C-29(8):720–731, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675654.

Meyer:1982:CFS

[Mey82] J. F. Meyer. Closed-form solutions of performability. IEEE Transactions on Computers, C-31(7):648–657, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676062.

Meyer:1984:DAB

[Mey84]  Gerard G. L. Meyer.  A

diagnosis algorithm for the BGM system level fault model. IEEE Transactions on Computers, C-33(8):756–758, August 1984. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009366.

Meyer:1988:FTF

[Mey88] F. J. Meyer. Flip-trees: faulttolerant graphs with wide containers. IEEE Transactions on Computers, 37(4):472–478, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2194.

Moldovan:1986:PMA

[MF86] D. I. Moldovan and J. A. B. Fortes. Partitioning and mapping algorithms into fixed size systolic arrays. IEEE Transactions on Computers, C-35(1):1–12, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676652.

Meyer:1980:PES

[MFW80] J. F. Meyer, D. G. Furchtgott, and L. T. Wu. Performability evaluation of the SIFT computer. IEEE Transactions on Computers, C-29(6):501–509, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675609.

Marsan:1982:MMM

[MG82] M. A. Marsan and M. Gerla.

Markov models for multiple bus multiprocessor systems. IEEE Transactions on Computers, C-31(3):239–248, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675980.

Mirsalehi:1986:CDI

[MG86] M. M. Mirsalehi and T. K.

Gaylord. Comments on “Direct Implementation of Discrete and Residue-Based Functions Via Optimal Encoding: A Programmable Array Logic Approach”. IEEE Transactions on Computers, C-35 (9):829–830, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676842.

See [Pap83].

McKinney:1988:MAP

[MG88] B. C. McKinney and F. El Guibaly. A multiple-access pipeline architecture for digital signal processing. IEEE Transactions on Computers, 37(3):283–290, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2165.

Martin:1980:CSE

[MH80] N. M. Martin and S. P. Hufnagel. Conditional-sum early completion adder logic. IEEE Transactions on Computers, C-29(8):753–756, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675663.

Myers:1981:USS

[MH81] G. J. Myers and D. G. Hocker. The use of software simulators in the testing and debugging of microprogram logic. IEEE Transactions on Computers, C-30(7):519–523, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675831.

Moitra:1985:MPB

[MI85] Abha Moitra and S. Sitharama Iyengar. A maximally parallel balancing algorithm for obtaining complete balanced binary trees. IEEE Transactions on Computers, C-34(6):563–565, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009411.

Masuyama:1989:TDL

[MI89] H. Masuyama and T. Ichimori. Tolerance of double-loop computer networks to multinode failures. IEEE Transactions on Computers, 38(5):738–741, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24275.

Miczo:1983:STH

[Mic83] A. Miczo. A self-test hardwired control section. IEEE Transactions on Computers, C-32(7):695–696, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676302.

Mili:1982:SSP

[Mil82] A. Mili. Self-stabilizing programs: The fault-tolerant capability of self-checking programs. IEEE Transactions on Computers, C-31(7):685–689, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676067.

See comments [BC85].

Mili:1985:AR

[Mil85]  A. Mili.  Author’s reply.

IEEE Transactions on Computers, C-34(1):98, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676523.

Miller:1988:DMS

[Mil88]  B. P. Miller.  DPM: a

measurement system for distributed programs. IEEE Transactions on Computers, 37 (2):243–248, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2157.

Miyakawa:1989:CSV

[Miy89] M. Miyakawa. Criteria for selecting a variable in the construction of efficient decision trees. IEEE Transactions on Computers, 38(1):130–141, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8736.

Madhavan:1983:COD

[MK83a] C. E. Veni Madhavan and S. Krishna. Comments on “Optimal Design of Distributed Information Systems”. IEEE Transactions on Computers, C-32(12):1200–1201, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676184.

Mitrani:1983:MSS

[MK83b] I. Mitrani and P. J. B. King. Multiserver systems subject to breakdowns: An empirical study. IEEE Transactions on Computers, C-32(1):96–98, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676127.

McPherson:1984:DPK

[MK84] J. A. McPherson and C. R.

Kime. Diagnosis in the presence of known faults. IEEE Transactions on Computers, C-33(10):943–947, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676360.

Matula:1985:FPR

[MK85] David W. Matula and Peter Kornerup. Finite precision rational arithmetic: Slash number systems. IEEE Transactions on Computers, C-34(1):3–18, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676511.

Muroga:1989:TMD

[MKLC89] S. Muroga, Y. Kambayashi, H. C. Lai, and J. N. Culliney. The transduction method-design of logic networks based on permissible functions. IEEE Transactions on Computers, 38 (10):1404–1424, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35836.

Montoye:1982:PAS

[ML82] R. K. Montoye and D. H. Lawrie. A practical algorithm for the solution of triangular systems on a parallel processing system. IEEE Transactions on Computers, C-31(11):1076–1082, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675924.

Min:1988:SFS

[ML88] Y. Min and J. Li. Strongly fault secure PLAs and totally self-checking checkers. IEEE Transactions on Computers, 37(7):863–867, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2233.

Milutinovic:1987:GBM

[MLB87] V. Milutinovic and N. Lopez-Benitez. A gaas-based microprocessor architecture for real-time applications. IEEE Transactions on Computers, C-36(6):714–727, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676964.

Ma:1982:TAM

[MLT82] Perng-Yi Richard Ma, E. Y. S. Lee, and M. Tsuchiya. A task allocation model for distributed computing systems.

IEEE Transactions on Computers, C-31(1):41–47, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675884.

Mallela:1980:DRH

[MM80] S. Mallela and G. M. Masson. Diagnosis without repair for hybrid fault situations. IEEE Transactions on Computers, C-29(6):461–470, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675605.

Masson:1983:ECN

[MM83a] G. M. Masson and S. B. Morris. Expected capacity of (m2)- networks. IEEE Transactions on Computers, C-32(7):649–657, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676296.

Miller:1983:SFS

[MM83b] D. M. Miller and J. C.

Muzio. Spectral fault signatures for internally unate combinational networks. IEEE Transactions on Computers, C-32(11):1058–1062, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676156.

Miller:1984:SFS

[MM84] D. M. Miller and J. C. Muzio. Spectral fault signatures for single stuck-at faults in combinational networks. IEEE Transactions on Computers, C-33(8):765–769, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009369.

Mahmood:1988:CED

[MM88] A. Mahmood and E. J. Mc-Cluskey. Concurrent error detection using watchdog processors — a survey. IEEE Transactions on Computers, 37 (2):160–174, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2145.

Moharir:1985:ESG

[Moh85] P. S. Moharir. Extending the scope of Golub’s method beyond complex multiplication. IEEE Transactions on Computers, C-34(5):484–487, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676590.

Moldovan:1982:ASV

[Mol82a] D. I. Moldovan. On the analysis and synthesis of VLSI algorithms. IEEE Transactions on Computers, C-31(11):

1121–1126, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675929.

Molloy:1982:PAU

[Mol82b] M. K. Molloy. Performance analysis using stochastic Petri nets. IEEE Transactions on Computers, C-31(9):913–917, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676110.

Morgan:1980:AFS

[Mor80] D. R. Morgan. Autocorrelation function of sequential Mbit words taken from an Nbit shift register (PN) sequence. IEEE Transactions on Computers, C-29(5):408–410, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675594.

See comments [Wus81].

Moraga:1986:DMV

[Mor86] C. Moraga. Design of a multiplevalued systolic system for the computation of the Chrestenson spectrum. IEEE Transactions on Computers, C-35 (2):183–188, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676739.

Mossberg:1987:VCM

[Mos87]  B. Mossberg.  Vectorization

of the calculation of a moving sum. IEEE Transactions on Computers, C-36(3):362–365, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676909.

McFarland:1983:AMB

[MP83a] M. C. McFarland and A. C. Parker. An abstract model of behavior for hardware descriptions. IEEE Transactions on Computers, C-32(7):621–637, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676294.

Mezzalama:1983:HDM

[MP83b] M. Mezzalama and P. Prinetto. A hierarchical description model for microcode. IEEE Transactions on Computers, C-32(5): 478–487, May 1983. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1676259.

Miller:1983:RNS

[MP83c] D. D. Miller and J. N. Polky. A residue number system implementation of the LMS algorithm using optical waveguide circuits. IEEE Transactions on Computers, C-32 (11):1013–1028, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676152;
http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35222.

Midkiff:1987:CAS

[MP87] Samuel P. Midkiff and David A.

Padua. Compiler algorithms for synchronization. IEEE Transactions on Computers, C-36(12):1485–1495, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009499.

Mazumder:1989:PTP

[MP89a] P. Mazumder and J. H. Patel. Parallel testing for patternsensitive faults in semiconductor random-access memories. IEEE Transactions on Computers, 38(3):394–407, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21126.

Meyer:1989:DTS

[MP89b] F. J. Meyer and D. K. Pradhan. Dynamic testing strategy for distributed systems.

IEEE Transactions on Computers, 38(3):356–365, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21122.

Meyer:1989:MDS

[MP89c] F. J. Meyer and D. K. Pradhan. Modeling defect spatial distribution. IEEE Transactions on Computers, 38(4):538–546, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21146.

Magenheimer:1988:IMD

[MPPZ88] Daniel J. Magenheimer, Liz Peters, Karl W. Peters, and Dan Zuras. Integer multiplication and division on the HP Precision Architecture. IEEE Transactions on Computers, 37(8):980–990, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2248.

Memmi:1982:SNR

[MR82] G. Memmi and Y. Raillard. Some new results about the (d, k) graph problem. IEEE Transactions on Computers, C-31(8):784–791, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676084.

Muzio:1986:IMV

[MR86] J. C. Muzio and I. C. Rosenberg. Introduction — multiplevalued logic. IEEE Transactions on Computers, C-35(2):97–98, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676726.

Malaiya:1981:RMH

[MS81] Y. K. Malaiya and S. Y. H.

Su. Reliability measure of hardware redundancy fault-tolerant digital systems with intermittent faults. IEEE Transactions on Computers, C-30(8):600–604, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675847.

McMillen:1982:RSA

[MS82] R. J. McMillen and H. J. Siegel.

Routing schemes for the augmented data manipulator network in an MIMD system. IEEE Transactions on Computers, C-31(12):1202–1214, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675944.

McMullen:1986:PIM

[MS86] C. McMullen and J. Shearer.

Prime implicants, minimum covers, and the complexity of logic simplification. IEEE Transactions on Computers, C-35(8):761–762, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676828.

McAnney:1988:BCC

[MS88a] W. H. McAnney and J. Savir. Built-in checking of the correct self-test signature. IEEE Transactions on Computers, 37 (9):1142–1145, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2268.

Miller:1988:EPC

[MS88b] R. Miller and Q. F. Stout. Efficient parallel convex hull algorithms. IEEE Transactions on Computers, 37(12): 1605–1618, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9737.

Miller:1988:SEP

[MS88c] R. Miller and Q. F. Stout. Simulating essential pyramids. IEEE Transactions on Computers, 37(12):1642–1648, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9741.

Miller:1989:MCA

[MS89] R. Miller and Q. F. Stout.

Mesh computer algorithms for computational geometry. IEEE Transactions on Computers, 38(3):321–340, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21120.

Melliar-Smith:1982:FSM

[MSS82] P. M. Melliar-Smith and R. L.

Schwartz. Formal specification and mechanical verification of SIFT: a fault-tolerant flight control system. IEEE Transactions on Computers, C-31(7):616–630, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676059.

McGough:1985:CRE

[MST85] J. McGough, M. Smotherman, and K. S. Trivedi. The conservativeness of reliability estimates based on instantaneous coverage. IEEE Transactions on Computers, C-34(7):602–609, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676599.

Mouftah:1980:TRM

[MSV80] H. T. Mouftah, K. C. Smith, and Z. G. Vranesic. Ternary rate-multipliers. IEEE Transactions on Computers, C-29 (10):929–931, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675477.

Melkemi:1987:CMP

[MT87] L. Melkemi and M. Tchuente. Complexity of matrix product on a class of orthogonally connected systolic arrays. IEEE Transactions on Computers, C-36(5):615–619, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676946.

Moret:1983:STB

[MTG83] B. M. E. Moret, M. G. Thomason, and R. C. Gonzalez. Symmetric and threshold Boolean functions are exhaustive. IEEE Transactions on Computers, C-32(12):1211–1212, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676189.

Mccall:1985:PCA

[MTG^{+}85] J. T. Mccall, J. G. Tront, F. G. Gray, R. M. Haralick, and W. M. McCormack. Parallel computer architectures and problem solving strategies for the consistent labeling problem. IEEE Transactions on Computers, C-34(11):973–980, November 1985. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676530.

Mackinnon:1985:OAA

[MTMA85] S. J. Mackinnon, P. D. Taylor, H. Meijer, and S. G. Akl. An optimal algorithm for assigning cryptographic keys to control access in a hierarchy. IEEE Transactions on Computers, C-34(9):797–802, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676635.

Mirchandaney:1989:AED

[MTS89] R. Mirchandaney, D. Towsley, and J. A. Stankovic. Analysis of the effects of delays on load sharing. IEEE Transactions on Computers, 38(11): 1513–1525, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42124.

Mukaidono:1986:RTL

[Muk86] M. Mukaidono. Regular ternary logic functions — ternary logic functions suitable for treating ambiguity. IEEE Transactions on Computers, C-35 (2):179–183, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676738.

Mukhopadhyay:1987:SPH

[Muk87] A. Mukhopadhyay. A solution to the polynomial Hensel code conversion problem. IEEE Transactions on Computers, C-36(5):634–637, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676950.

Mukherjee:1989:HAD

[Muk89] A. Mukherjee. Hardware algorithms for determining similarity between two strings. IEEE Transactions on Computers, 38(4):600–603, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21153.

Muller:1985:DBC

[Mul85] Jean-Michel Muller. Discrete basis and computation of elementary functions. IEEE Transactions on Computers, C-34 (9):857–862, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676643.

Murray:1981:SOE

[Mur81] N. V. Murray. Some observations on equivalence handling methods. IEEE Transactions on Computers, C-30(5):361–362, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675795.

Muzio:1980:CSA

[Muz80] J. C. Muzio. Composite spectra and the analysis of switching circuits. IEEE Transactions on Computers, C-29(8):750–753, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675662.

Meyer:1988:IWE

[MW88] J. F. Meyer and L. Wei. Influence of workload on error recovery in random access memories. IEEE Transactions on Computers, 37(4):500–507, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2199.

Mehra:1980:CSS

[MWM80] S. K. Mehra, J. W. Wong, and J. C. Majithia. A comparative study of some twoprocessor organizations. IEEE Transactions on Computers, C-29(1):44–49, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675455.

Nakamura:1986:AIA

[Nak86] S. Nakamura. Algorithms for iterative array multiplication. IEEE Transactions on Computers, C-35(8):713–719, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676822.

Nakamura:1987:IMN

[Nak87] K. Nakamura. Inverter-minimum networks. IEEE Transactions on Computers, C-36 (2):226–230, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676885.

Nakamura:1988:SCP

[NC88a] S. Nakamura and K.-Y. Chu. A single chip parallel multiplier by MOS technology. IEEE Transactions on Computers, 37(3):274–282, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2164.

Nicolaidis:1988:SCD

[NC88b] M. Nicolaidis and B. Courtois. Strongly code disjoint checkers. IEEE Transactions on Computers, 37(6):751–756, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2217.

Nicolau:1984:MPA

[NF84] A. Nicolau and J. A. Fisher.

Measuring the parallelism available for very long instruction word architectures. IEEE Transactions on Computers, C-33 (11):968–976, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676371.

Nikolos:1986:SEC

[NGP86] D. Nikolos, N. Gaitanis, and G. Philokyprou. Systematic t-error correcting/all unidirectional error detecting codes. IEEE Transactions on Computers, C-35(5):394–402, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676782.

Ni:1985:VRT

[NH85] L. M. Ni and Kai Hwang.

Vector-reduction techniques for arithmetic pipelines. IEEE Transactions on Computers, C-34(5):404–411, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676580.

Nash:1988:MFA

[NH88] J. G. Nash and S. Hansen. Modified Faddeeva algorithm for concurrent execution of linear algebraic operations. IEEE Transactions on Computers, 37 (2):129–137, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2142.

Nakatani:1989:WBS

[NHAT89] T. Nakatani, S.-T. Huang, B. W. Arden, and S. K. Tripathi. K-way bitonic sort. IEEE Transactions on Computers, 38(2):283–288, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=16506.

Nicolau:1989:RTD

[Nic89] A. Nicolau. Run-time disambiguation: coping with statically unpredictable dependencies. IEEE Transactions on Computers, 38(5):663–678, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24269.

Nilsson:1984:CRP

[Nil84] Arne A. Nilsson. Comments on “The Reliability of Periodically Repaired n − 1/n Parallel Redundant Systems”.

IEEE Transactions on Computers, C-33(7):681, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009345.

See [Can83, Wen85].

Nilsson:1986:AR

[Nil86] A. A. Nilsson. Author’s reply. IEEE Transactions on Computers, C-35(10):932, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676689.

Niznik:1983:QAM

[Niz83] C. A. Niznik. A quantization approximation for modeling computer network nodal queueing delay. IEEE Transactions on Computers, C-32(3):245–253, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676216.

Niznik:1984:PEC

[Niz84a] C. A. Niznik. Performance evaluation of the computer network dynamic congestion table algorithm. IEEE Transactions on Computers, C-33 (2):150–159, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=1676407.

See correction [Niz84b].

Niznik:1984:CPE

[Niz84b] Carol A. Niznik. Correction to “Performance Evaluation of the Computer Network Dynamic Congestion Table Algorithm”. IEEE Transactions on Computers, C-33(8):772, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009371.

See [Niz84a].

Nagpal:1983:PAT

[NJM83] H. K. Nagpal, G. A. Jullien, and W. C. Miller. Processor architectures for two-dimensional convolvers using a single multiplexed computational element with finite field arithmetic. IEEE Transactions on Computers, C-32(11):989–1001, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676150.

Nanya:1987:NSF

[NK87a] Takashi Nanya and Toshiaki Kawamura. A note on strongly fault-secure sequential circuits. IEEE Transactions on Computers, C-36(9):1121–1123, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009545.

Nanya:1987:EIT

[NK87b] Takashi Nanya and Toshiaki Kawamura. On error indication for totally self-checking systems. IEEE Transactions on Computers, C-36(11): 1389–1392, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009484.

Nageshwara:1988:CAP

[NK88a] R. V. Nageshwara and V. Kumar. Concurrent access of priority queues. IEEE Transactions on Computers, 37(12): 1657–1665, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9744.

Nanya:1988:ESP

[NK88b] T. Nanya and T. Kawamura.

Error secure/propagating concept and its application to the design of strongly fault-secure processors. IEEE Transactions on Computers, 37(1):14–24, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75147.

Nishioka:1980:AGA

[NKY^{+}80] I. Nishioka, T. Kurimoto, S. Yamamoto, T. Chiba, I. Shirakawa, and H. Ozaki. An approach to gate assignment and module placement for printed wiring boards. IEEE Transactions on Computers, C-29(8):681–688, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675649.

Nakamura:1982:LBC

[NM82] S. Nakamura and G. M. Masson. Lower bounds on crosspoints in concentrators. IEEE Transactions on Computers, C-31(12):1173–1179, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675941.

Nichols:1987:TSI

[NM87] K. M. Nichols and D. G. Messerschmitt. Traffic-specific interconnection networks for multicomputers. IEEE Transactions on Computers, C-36 (10):1183–1196, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676859.

Nath:1983:EVN

[NMB83] D. Nath, S. N. Maheshwari, and P. C. P. Bhatt. Efficient VLSI networks for parallel processing based on orthogonal trees. IEEE Transactions on Computers, C-32(6):569–581, June 1983. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676279.

Naclerio:1989:MPN

[NMN89] N. J. Naclerio, S. Masuda, and K. Nakajima. The via minimization problem is NP-complete. IEEE Transactions on Computers, 38(11):1604–1608, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42135.

Narasimhan:1986:ADF

[NN86] J. Narasimhan and K. Nakajima. An algorithm for determining the fault diagnosability of a system. IEEE Transactions on Computers, C-35(11):1004–1008, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676703.

Nishihara:1987:BSR

[NN87] Seiichi Nishihara and Hiroji Nishino. Binary search revisited: Another advantage of Fibonacci search. IEEE Transactions on Computers, C-36(9):1132–1135, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009548.

Noetzel:1989:IMU

[Noe89] A. S. Noetzel. An interpolating memory unit for function evaluation: analysis and design. IEEE Transactions on Computers, 38(3):377–384, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21124.

Naganuma:1988:HSC

[NOYK88] J. Naganuma, T. Ogura, S.-I. Yamada, and T. Kimura. Highspeed CAM-based architecture for a Prolog machine (ASCA). IEEE Transactions on Computers, 37(11):1375–1383, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8703.

Nikolos:1988:EDT

[NPP88] D. Nikolos, A. M. Paschalis, and G. Philokyprou. Efficient design of totally self-checking checkers for all low-cost arithmetic codes. IEEE Transactions on Computers, 37(7):807–814, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2226.

Nassimi:1981:SRB

[NS81] D. Nassimi and S. Sahni. A self-routing Benes network and parallel permutation algorithms.

IEEE Transactions on Computers, C-30(5):332–340, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675791.

Nassimi:1982:OBP

[NS82a] D. Nassimi and S. Sahni. Optimal BPC permutations on a cube connected SIMD computer. IEEE Transactions on Computers, C-31(4):338–341, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676004.

Nassimi:1982:PAS

[NS82b] D. Nassimi and S. Sahni. Parallel algorithms to set up the Benes permutation network. IEEE Transactions on Computers, C-31(2):148–154, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675960.

Norton:1987:PPA

[NS87] A. Norton and A. J. Silberger. Parallelization and performance analysis of the Cooley–Tukey FFT algorithm for shared-memory architectures. IEEE Transactions on Computers, C-36(5):581–591, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676943.

Nicol:1988:DRP

[NS88] D. M. Nicol and J. H. Saltz. Dynamic remapping of parallel computations with varying resource demands. IEEE Transactions on Computers, 37 (9):1073–1087, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2258.

Nelson:1988:AAF

[NT88] R. Nelson and A. N. Tantawi. Approximate analysis of fork/ join synchronization in parallel queues. IEEE Transactions on Computers, 37(6):739–743, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2213.

Nwachukwu:1985:AGA

[Nwa85] E. O. Nwachukwu. Address generation in an array processor. IEEE Transactions on Computers, C-34(2):170–173, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676554.

Ni:1982:MBO

[NWLP82] L. M. Ni, K. Y. Wong, D. T. Lee, and R. K. Poon. A microprocessor-based office image processing system. IEEE Transactions on Computers, C-31(10):1017–1022, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675911.

Ong:1983:BQC

[OA83] Shauchi Ong and D. E. Atkins.

A basis for the quantitative comparison of computer number systems. IEEE Transactions on Computers, C-32(4):359–369, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676237.

Oberman:1980:CMR

[Obe80] R. M. M. Oberman. Comments on “Modular Replacement of Combinational Switching Networks”. IEEE Transactions on Computers, C-29(1):59–60, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675458.

Oklobdzija:1982:LSR

[OE82] V. G. Oklobdzija and M. D.

Ercegovac. An on-line square root algorithm. IEEE Transactions on Computers, C-31(1):70–75, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675887.

Okano:1987:CMH

[OI87a] H. Okano and H. Imai. A construction method of highspeed decoders using ROM’s for Bose–Chaudhuri–Hocquenghem and Reed–Solomon codes. IEEE Transactions on Computers, C-36(10):1165–1171, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676857.

Owens:1987:AC

[OI87b] Robert Michael Owens and Mary Jane Irwin. The arithmetic cube. IEEE Transactions on Computers, C-36 (11):1342–1348, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009473.

Oikonomou:1987:AFS

[Oik87] K. N. Oikonomou. Abstractions of finite-state machines optimal with respect to single undetectable output faults. IEEE Transactions on Computers, C-36(2):185–200, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676881.

Ossfeldt:1980:RDC

[OJ80] B. E. Ossfeldt and I. Jonsson. Recovery and diagnostics in the central control of the AXE switching system. IEEE Transactions on Computers, C-29(6):482–491, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675607.

Owens:1985:PSS

[OJ85] Robert Michael Owens and Joseph Ja’Ja’. Parallel sorting with serial memories. IEEE Transactions on Computers, C-34(4):379–383, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009391.

Oikonomou:1983:ANL

[OK83] K. N. Oikonomou and R. Y.

Kain. Abstractions for node level passive fault detection in distributed systems. IEEE Transactions on Computers, C-32(6):543–550, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676276.

OLeary:1987:SAM

[O’L87]  Dianne P. O’Leary.  Sys-

tolic arrays for matrix transpose and other reorderings.

IEEE Transactions on Computers, C-36(1):117–122, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009457.

Oommen:1988:DLA

[OM88] B. J. Oommen and D. C. Y. Ma. Deterministic learning automata solutions to the equipartitioning problem. IEEE Transactions on Computers, 37(1):2–13, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75146.

Oruc:1987:PCP

[OO87] A. Y. Oruc and M. Y. Oruc. Programming cellular permutation networks through decomposition of symmetric groups. IEEE Transactions on Computers, C-36(7):802–809, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676977.

Oruc:1985:RAI

[OOB85] A. Y. Oruc, M. Y. Oruc, and N. Balabanian. Reconfiguration algorithms for interconnection networks. IEEE Transactions on Computers, C-34(8):773–776, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676628.

Oruc:1984:RAC

[OP84] A. Y. Oruc and D. Prakash.

Routing algorithms for cellular interconnection arrays. IEEE Transactions on Computers, C-33(10):939–942, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676359.

Ottmann:1982:DMV

[ORS82] T. A. Ottmann, A. L. Rosenberg, and L. J. Stockmeyer. A dictionary machine (for VLSI). IEEE Transactions on Computers, C-31(9):892–897, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676104.

Oruc:1984:CCC

[Oru84] A. Yavuz Oruc. A classification of cube-connected networks with a simple control scheme. IEEE Transactions on Computers, C-33(8):769–772, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009370.

ODonnell:1982:CPC

[OS82] M. J. O’Donnell and C. H.

Smith. A combinatorial problem concerning processor interconnection networks. IEEE Transactions on Computers, C-31(2):163–164, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675962.

OLeary:1987:DSA

[OS87] D. P. O’Leary and G. W. Stewart. From determinacy to systaltic arrays. IEEE Transactions on Computers, C-36(11):1355–1359, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009475.

Oruc:1989:SDC

[OT89] A. Y. Oruc and A. Thirumalai. A systematic design of cellular permutation arrays. IEEE Transactions on Computers, 38(10):1447–1451, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35838.

Owens:1983:TRI

[Owe83]  R. M. Owens.  Techniques

to reduce the inherent limitations of fully digit on-line arithmetic. IEEE Transactions on Computers, C-32(4):406–411, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676244.

Ozguner:1986:DFS

[Ozg86] F. Ozguner. Deductive fault simulation of internal faults of inverter-free circuits and programmable logic arrays. IEEE Transactions on Computers, C-35(1):70–73, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676661.

Pomper:1981:RMF

[PA81] G. Pomper and J. R. Armstrong. Representation of multivalued functions using the direct cover method. IEEE Transactions on Computers, C-30 (9):674–679, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675867.

Page:1980:MTR

[Pag80] E. W. Page. Minimally testable Reed–Muller canonical forms. IEEE Transactions on Computers, C-29(8):746–750, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675661.

Pal:1986:AOL

[Pal86] A. Pal. An algorithm for optimal logic design using multiplexers. IEEE Transactions on Computers, C-35(8):755–757, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676826.

Papachristou:1983:DID

[Pap83] C. A. Papachristou. Direct implementation of discrete and residue-based functions via optimal encoding: a programmable array logic approach. IEEE Transactions on Computers, C-32(10):961–968, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676143.

See comments [MG86].

Parker:1980:NSE

[Par80] D. S. Parker, Jr. Notes on shuffle/exchange-type switching networks. IEEE Transactions on Computers, C-29(3):213–222, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675553.

Parker:1981:CCM

[Par81]  D. S. Parker, Jr.  Correc-

tion to “Combinatorial Merging and Huffman’s Algorithm”. IEEE Transactions on Computers, C-30(6):454, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675815.

Park:1986:EMS

[Par86] Jong Won Park. An efficient memory system for image processing. IEEE Transactions on Computers, C-35(7):669–674, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676813.

Parhami:1987:CTL

[Par87] B. Parhami. On the complexity of table lookup for iterative division. IEEE Transactions on Computers, C-36 (10):1233–1236, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676863.

Parhami:1988:CFA

[Par88] B. Parhami. Carry-free addition of recoded binary signeddigit numbers. IEEE Transactions on Computers, 37(11): 1470–1476, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8722.

Parberry:1989:NNS

[Par89] I. Parberry. A note on nondeterminism in small, fast parallel computers. IEEE Transactions on Computers, 38(5):766–767, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24282.

Patel:1980:ADP

[Pat80]  J. H. Patel.  An alterna-

tive to the distributed pipeline. IEEE Transactions on Computers, C-29(8):736–737, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675656.

Patel:1981:PPM

[Pat81]  J. H. Patel.  Performance

of processor-memory interconnections for multiprocessors. IEEE Transactions on Computers, C-30(10):771–780, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675695.

Patel:1982:AMP

[Pat82] J. H. Patel. Analysis of multiprocessors with private cache memories. IEEE Transactions on Computers, C-31(4):296–304, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675995.

Polychoronopoulos:1987:PAH

[PB87] C. D. Polychoronopoulos and U. Banerjee. Processor allocation for horizontal and vertical parallelism and related speedup bounds. IEEE Transactions on Computers, C-36(4):410–420, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676923.

Park:1989:AAB

[PBL89] A. Park, K. Balasubramanian, and R. J. Lipton. Array access bounds for block storage memory systems. IEEE Transactions on Computers, 38(6):909–913, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24305.

Priester:1981:NMT

[PC81] R. W. Priester and J. B.

Clary. New measures of testability and test complexity for linear analog failure analysis. IEEE Transactions on Computers, C-30(11):884–888, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675719.

Peir:1989:MDM

[PC89] J.-K. Peir and R. Cytron. Minimum distance: a method for partitioning recurrences for multiprocessors. IEEE Transactions on Computers, 38(8):1203–1211, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30873.

Patel:1982:CED

[PF82] J. H. Patel and L. Y. Fung. Concurrent error detection in ALU’s by recomputing with shifted operands. IEEE Transactions on Computers, C-31(7):589–595, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676055.

Patel:1983:CED

[PF83] J. H. Patel and L. Y. Fung. Concurrent error detection in multiply and divide arrays. IEEE Transactions on Computers, C-32(4):417–422, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676246.

Patnaik:1986:DPE

[PGR86] L. M. Patnaik, R. Govindarajan, and N. S. Ramadoss. Design and performance evaluation of EXMAN: An EXtended MANchester data flow computer. IEEE Transactions on Computers, C-35(3):229–244, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676747.

Page:1986:IPB

[PH86] I. P. Page and J. Hagins. Improving the performance of buddy systems. IEEE Transactions on Computers, C-35(5):441–447, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676786.

Piestrak:1987:DFS

[Pie87]  S. J. Piestrak.  Design of

fast self-testing checkers for a class of Berger codes. IEEE Transactions on Computers, C-36(5):629–634, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676949.

Pincin:1989:NAM

[Pin89] A. Pincin. A new algorithm for multiplication in finite fields. IEEE Transactions on Computers, 38(7):1045–1049, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30855.

Pradhan:1980:URS

[PK80] D. K. Pradhan and K. L. Kodandapani. A uniform representation of single-and multistage interconnection networks used in SIMD machines. IEEE Transactions on Computers, C-29(9):777–791, September 1980. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675677.

Polychronopoulos:1987:GSS

[PK87] Constantine D. Polychronopoulos and David J. Kuck. Guided self-scheduling: a practical scheduling scheme for parallel supercomputers. IEEE Transactions on Computers, C-36(12):1425–1439, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009495.

Peizong:1988:SLA

[PK88] Lee Peizong and Z. M. Kedem. Synthesizing linear array algorithms from nested FOR loop algorithms. IEEE Transactions on Computers, 37 (12):1578–1598, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9735.

Padua:1980:HSM

[PKL80] D. A. Padua, D. J. Kuck, and D. H. Lawrie. High-speed multiprocessors and compilation techniques. IEEE Transactions on Computers, C-29 (9):763–776, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675676.

Polychronopoulos:1989:UML

[PKP89] C. D. Polychronopoulos, D. J.

Kuck, and D. A. Padua. Utilizing multidimensional loop parallelism on large scale parallel processor systems. IEEE Transactions on Computers, 38 (9):1285–1296, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29467.

Padmanabhan:1983:CRP

[PL83a] K. Padmanabhan and D. H.

Lawrie. A class of redundant path multistage interconnection networks. IEEE Transactions on Computers, C-32(12):1099–1108, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676170.

Parkinson:1983:MPH

[PL83b] D. Parkinson and H. M. Liddell. The measurement of performance on a highly parallel system. IEEE Transactions on Computers, C-32(1):32–37, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676121.

Preparata:1984:OTL

[PL84] F. P. Preparata and W. Lipski, Jr. Optimal three-layer channel routing. IEEE Transactions on Computers, C-33(5):427–437, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676459.

Probst:1988:ASS

[PL88] D. K. Probst and H. F. Li. Abstract specification of synchronous data types for VLSI and proving the correctness of systolic network implementations. IEEE Transactions on Computers, 37(6):710–720, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2209.

Padegs:1988:ISV

[PMSB88] A. Padegs, B. B. Moore, R. M. Smith, and W. Buchholz. The IBM System/370 vector architecture: design considerations. IEEE Transactions on Computers, 37(5):509–520, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4602.

Pfister:1985:HSC

[PN85] G. F. Pfister and V. A. Norton. “hot spot” contention and combining in multistage interconnection networks. IEEE Transactions on Computers, C-34(10): 943–948, October 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paschalis:1988:EMD

[PNH88] A. M. Paschalis, D. Nikolos, and C. Halatsis. Efficient modular design of TSC checkers for m-out-of-2m codes. IEEE Transactions on Computers, 37(3):301–309, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2167.

Polychronopoulos:1988:COE

[Pol88] C. D. Polychronopoulos. Compiler optimizations for enhancing parallelism and their impact on architecture design. IEEE Transactions on Computers, 37(8):991–1004, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2249.

Porter:1980:PRS

[Por80] W. A. Porter. Polylogic realization of switching functions. IEEE Transactions on Computers, C-29(7):657–659, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675637.

Porter:1982:ETD

[Por82] W. A. Porter. Error tolerant design of multivalued logic functions. IEEE Transactions on Computers, C-31(6):551–554, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676039.

Purdy:1987:IDL

[PP87] C. N. Purdy and G. B. Purdy. Integer division in linear time with bounded fan-in. IEEE Transactions on Computers, C-36(5):640–644, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676952.

Park:1988:TCM

[PP88] N. Park and A. C. Parker. Theory of clocking for maximum execution overlap of high-speed digital systems. IEEE Transactions on Computers, 37(6):678–690, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2206.

Parthasarathy:1981:TDI

[PR81a] R. Parthasarathy and S. M. Reddy. A testable design of iterative logic arrays. IEEE Transactions on Computers, C-30(11):833–841, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675714.

Prakash:1981:CVF

[PR81b] S. Prakash and V. V. Rao.

Comments on “Very Fast Fourier Transform Algorithms Hardware for Implementation”. IEEE Transactions on Computers, C-30(6):452, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675812.

Pradhan:1982:FTC

[PR82] D. K. Pradhan and S. M.

Reddy. A fault-tolerant communication architecture for distributed systems. IEEE Transactions on Computers, C-31 (9):863–870, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676101.

Parker:1984:GN

[PR84] D. S. Parker and C. S. Raghavendra. The Gamma network. IEEE Transactions on Computers, C-33(4):367–373, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676444.

Pradhan:1980:NCE

[Pra80] D. K. Pradhan. A new class of error-correcting/detecting codes for fault-tolerant computer applications. IEEE Transactions on Computers, C-29(6):471–481, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675606.

Pradhan:1983:SND

[Pra83]  D. K. Pradhan.  Sequen-

tial network design using extra inputs for fault detection. IEEE Transactions on Computers, C-32(3):319–323, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676225.

Pradhan:1985:DRF

[Pra85a] D. K. Pradhan. Dynamically restructurable fault-tolerant processor network architectures. IEEE Transactions on Computers, C-34(5):434–447, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676583.

Pradhan:1985:FTM

[Pra85b] D. K. Pradhan. Fault-tolerant multiprocessor link and bus network architectures. IEEE Transactions on Computers, C-34(1):33–45, January 1985.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676513.

See correction [Pra86a].

Pradhan:1986:CFT

[Pra86a] D. K. Pradhan. Correction to “Fault-Tolerant Multiprocessor Link and Bus Architectures”. IEEE Transactions on Computers, C-35(1):94, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676670.

See [Pra85b].

Pramanik:1986:PAD

[Pra86b]  S. Pramanik.  Performance

analysis of a database filter search hardware. IEEE Transactions on Computers, C-35(12):1077–1082, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676716.

Preparata:1983:MCA

[Pre83] F. P. Preparata. A meshconnected area-time optimal VLSI multiplier of large integers. IEEE Transactions on Computers, C-32(2):194–198, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676203.

Proskurowski:1981:MBT

[Pro81] A. Proskurowski. Minimum broadcast trees. IEEE Transactions on Computers, C-30(5): 363–366, May 1981. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675796.

Prohazka:1988:BMS

[Pro88]  C. G. Prohazka.  Bound-

ing the maximum size of a packet radio network. IEEE Transactions on Computers, 37 (10):1184–1190, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5980.

Prohazka:1989:DLS

[Pro89] C. G. Prohazka. Decoupling link scheduling constraints in multi-hop packet radio networks. IEEE Transactions on Computers, 38(3):455–458, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21132.

Patterson:1980:DCS

[PS80] D. A. Patterson and C. H. Sequin. Design considerations for single-chip computers of the future. IEEE Transactions on Computers, C-29 (2):108–116, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675534.

Pedar:1983:AOA

[PS83a] A. Pedar and V. V. S. Sarma.

Architecture optimization of aerospace computing systems. IEEE Transactions on Computers, C-32(10):911–922, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676135.

Pitchumani:1983:IAM

[PS83b] V. Pitchumani and E. P. Stabler. An inductive assertion method for register transfer level design verification. IEEE Transactions on Computers, C-32(12):1073–1080, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676167.

Papachristou:1985:IMD

[PS85a] C. A. Papachristou and N. B.

Sahgal. An improved method for detecting functional faults in semiconductor random access memories. IEEE Transactions on Computers, C-34 (2):110–116, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676547.

Pitchumani:1985:VRT

[PS85b] V. Pitchumani and E. P. Stabler. Verification of register transfer level parallel control sequences. IEEE Transactions on Computers, C-34(8):761–765, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676625.

Peng:1987:MCT

[PS87] D. Peng and K. G. Shin. Modeling of concurrent task execution in a distributed system for realtime control. IEEE Transactions on Computers, C-36(4):500–516, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676931.

Pitchumani:1988:FTG

[PS88] V. Pitchumani and S. S. Soman. Functional test generation based on unate function theory. IEEE Transactions on Computers, 37(6):756–760, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2218.

Pries:1986:GPC

[PTC86] W. Pries, A. Thanailakis, and H. C. Card. Group properties of cellular automata and VLSI applications. IEEE Transactions on Computers, C-35(12):1013–1024, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676709.

Persky:1981:COS

[PTT81] G. Persky, B. N. Tien, and B. S. Ting. Comments on “An Optimal Solution for the Channel-Assignment Problem”. IEEE Transactions on Computers, C-30(6):454, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675816.

See [Lau81, Rub81].

Parker:1981:SHD

[PW81] A. C. Parker and J. J. Wallace. Slide: An I/O hardware descriptive language. IEEE Transactions on Computers, C-30(6):423–439, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675808.

Qadah:1985:DMV

[QI85] G. Z. Qadah and K. B. Irani.

A database machine for very large relational databases. IEEE Transactions on Computers, C-34(11):1015–1025, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676534.

Quammen:1985:ESM

[QKS85] D. Quammen, J. P. Kearns, and M. L. Soffa. Efficient storage management for temporary values in concurrent programming languages. IEEE Transactions on Computers, C-34 (9):832–840, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676639.

Rauscher:1980:MTS

[RA80] T. G. Rauscher and P. M. Adams. Microprogramming: a tutorial and survey of recent developments. IEEE Transactions on Computers, C-29(1):2–20, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675451.

Raghavendra:1984:FTB

[RAE84] C. S. Raghavendra, A. Avizienis, and M. D. Ercegovac. Fault tolerance in binary tree architectures. IEEE Transactions on Computers, C-33(6):568–572, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676483.

Ramachandran:1983:SRE

[Ram83]  V. Ramachandran.  Single

residue error correction in residue number systems. IEEE Transactions on Computers, C-32(5):504–507, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676264;
 http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35227.

Ramachandran:1986:AAM

[Ram86a] V. Ramachandran. Algorithmic aspects of MOS VLSI switchlevel simulation with race detection. IEEE Transactions on Computers, C-35(5):462–475, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676789.

See corrections [Ram86b].

Ramachandran:1986:CAA

[Ram86b] V. Ramachandran. Corrections to “Algorithmic Aspects of MOS VLSI Switch-Level Simulation with Race Detection”. IEEE Transactions on Computers, C-35(9):851, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676848.

See [Ram86a].

Ramarao:1988:DSL

[Ram88] K. V. S. Ramarao. Distributed sorting on local area networks. IEEE Transactions on Computers, 37(2):239–243, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2156.

Ramarao:1989:DAN

[Ram89] K. V. S. Ramarao. Distributed algorithms for network recognition problems. IEEE Transactions on Computers, 38(9): 1240–1248, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29464.

Reed:1987:SPP

[RAP87] D. A. Reed, L. M. Adams, and M. L. Patrick. Stencils and problem partitionings: Their influence on the performance of multiple processor systems. IEEE Transactions on Computers, C-36(7):845–858, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676980.

Reeves:1980:EFI

[RB80] A. P. Reeves and J. D. Bruner. Efficient function implementation for bit-serial parallel processors. IEEE Transactions on Computers, C-29 (9):841–844, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675685.

Riley:1982:DES

[RB82] D. D. Riley and R. J. Baron.

Design and evaluation of a synchronous triangular interconnection scheme for interprocessor communications. IEEE Transactions on Computers, C-31(2):110–118, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675956.

Ramanatha:1983:DTU

[RB83a] K. S. Ramanatha and N. N.

Biswas. A design for testability of undetectable crosspoint faults in programmable logic arrays. IEEE Transactions on Computers, C-32(6):551–557, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676277.

Ramanatha:1983:LAL

[RB83b] K. S. Ramanatha and N. N.

Biswas. An on-line algorithm for the location of cross point faults in programmable logic arrays. IEEE Transactions on Computers, C-32(5):438–444, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676254.

Rao:1983:MWC

[RB83c] C. D. V. P. Rao and N. N. Biswas. On the minimization of wordwidth in the control memory of a microprogrammed digital computer. IEEE Transactions on Computers, C-32(9):863–868, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676336.

Rao:1983:DTS

[RB83d] K. V. S. S. Prasad Rao and D. Basu. Design of totally selfchecking circuits with an unrestricted stuck-at fault-set using redundancy in space and time domains. IEEE Transactions on Computers, C-32(5):464–475, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676257.

Ramanatha:1988:DCI

[RB88] K. S. Ramanatha and N. N. Biswas. Design of crosspointirredundant PLAs using minimal number of control inputs. IEEE Transactions on Computers, 37(9):1130–1134, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2265.

Reddy:1989:EMD

[RB89a] A. L. N. Reddy and P. Banerjee. An evaluation of multipledisk I/O systems. IEEE Transactions on Computers, 38(12): 1680–1690, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40846.

Roesler:1989:RDO

[RB89b] M. Roesler and W. A. Burkhard.

Resolution of deadlocks in object-oriented distributed systems. IEEE Transactions on Computers, 38(8):1212–1224, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30874.

Ravindran:1989:FTR

[RC89] K. Ravindran and S. T. Chanson. Failure transparency in remote procedure calls. IEEE Transactions on Computers, 38 (8):1173–1187, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30871.

Razouk:1980:MVC

[RE80] R. R. Razouk and G. Estrin. Modeling and verification of communication protocols in Sara: The X.21 interface. IEEE Transactions on Computers, C-29(12):1038–1052, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675509.

Redinbo:1987:FFF

[Red87a] G. R. Redinbo. Finite field fault-tolerant digital filtering architectures. IEEE Transactions on Computers, C-36 (10):1236–1242, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676864.

Redinbo:1987:FTD

[Red87b] G. Robert Redinbo. Faulttolerant decoders for cyclic error-correcting codes. IEEE Transactions on Computers, C-36(1):47–63, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009448.

Reeves:1980:SDB

[Ree80] A. P. Reeves. A systematically designed binary array processor. IEEE Transactions on Computers, C-29(4):278–287, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675566.

Reed:1984:PMN

[Ree84]  D. A. Reed.  The perfor-

mance of multimicrocomputer networks supporting dynamic workloads. IEEE Transactions on Computers, C-33(11): 1045–1048, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676381.

Regener:1988:TSG

[Reg88] E. Regener. A transition sequence generator for RAM fault detection. IEEE Transactions on Computers, 37(3):362–368, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2175.

Rennels:1984:FTC

[Ren84] D. A. Rennels. Fault-tolerant computing—concepts and examples. IEEE Transactions on Computers, C-33(12):1116–1129, December 1984. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1676390.

Raschid:1986:SFU

[RFLS86] L. Raschid, T. Fei, H. Lam, and S. Y. W. Su. A specialfunction unit for sorting and sort-based database operations. IEEE Transactions on Computers, C-35(12):1071–1077, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676715.

Ramakrishnan:1986:MHG

[RFS86] I. V. Ramakrishnan, D. S. Fussell, and A. Silberschatz. Mapping homogeneous graphs on linear arrays. IEEE Transactions on Computers, C-35(3):189–209, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676744.

Rios:1985:HMA

[RG85] M. Rios and N. D. Georganas. A hybrid multipleaccess protocol for data and voice-packet over local area networks. IEEE Transactions on Computers, C-34(1):90–94, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676520.

Raghavendra:1985:RLT

[RGA85] C. S. Raghavendra, M. Gerla, and A. Avizienis. Reliable loop topologies for large local computer networks. IEEE Transactions on Computers, C-34(1):46–55, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=1676514.

See comments [Hwa87].

Reddy:1987:NAD

[RH87] S. M. Reddy and Dong Sam Ha. A new approach to the design of testable PLA’s. IEEE Transactions on Computers, C-36(2):201–211, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676882.

Rhyne:1984:LCL

[Rhy84]  T. Rhyne.  Limitations on

carry lookahead networks. IEEE Transactions on Computers, C-33(4):373–374, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676445.

Rao:1989:IOA

[RI89] S. S. S. P. Rao and J. R.

Isaac. Interface optimization: an algorithm for the detection of data path redundancy and reconfigurability towards obtaining minimal bus interfaces. IEEE Transactions on Computers, 38(11):1577–1580, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42128.

Richards:1984:CSL

[Ric84] D. Richards. Complexity of single-layer routing. IEEE Transactions on Computers, C-33(3):286–288, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676428.

Rich:1986:SMM

[Ric86]  D. A. Rich.  A survey of

multivalued memories. IEEE Transactions on Computers, C-35(2):99–106, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676727.

Rahier:1980:DLM

[RJ80] M. C. Rahier and P. G. A. Jespers. Dedicated LSI for a microprocessor-controlled handcarried OCR system. IEEE Transactions on Computers, C-29(2):79–88, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675530.

Roeser:1982:FHT

[RJ82] P. R. Roeser and M. E. Jernigan. Fast Haar Transform algorithms. IEEE Transactions on Computers, C-31(2):175–177, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675965.

Rice:1989:HPA

[RJ89] T. A. Rice and L. H. Jamieson. A highly parallel algorithm for root extraction. IEEE Transactions on Computers, 38(3):443–449, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21130.

Raghavendra:1986:PII

[RK86] C. S. Raghavendra and V. K. Prasanna Kumar. Permutations on Illiac IV-type networks. IEEE Transactions on Computers, C-35(7):662–669, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676812.

Roberts:1989:CCP

[RK89] R. S. Roberts and S. C. Kothari.

On computing the combinatorial power of SW-banyan networks. IEEE Transactions on Computers, 38(5):761–765, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24281.

Raghavendra:1988:RAD

[RKH88] C. S. Raghavendra, V. K. P.

Kumar, and S. Hariri. Reliability analysis in distributed systems. IEEE Transactions on Computers, 37(3):352–358, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2173.

Requa:1983:PDF

[RM83] J. E. Requa and J. R. Mc-Graw. The piecewise data flow architecture: Architectural concepts. IEEE Transactions on Computers, C-32(5):425–438, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676253.

Ramamoorthy:1986:ORS

[RM86] C. V. Ramamoorthy and Y. W. E. Ma. Optimal reconfiguration strategies for reconfigurable computer systems with no repair. IEEE Transactions on Computers, C-35(3):278–280, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676751.

Rosenberger:1988:QMI

[RMCF88] F. U. Rosenberger, C. E. Molnar, T. J. Chaney, and T.-P. Fang. Q-modules: internally clocked delay-insensitive modules. IEEE Transactions on Computers, 37 (9):1005–1018, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2252.

Rego:1988:AMC

[RN88] V. Rego and L. M. Ni. Analytic models of cyclic service systems and their application to token-passing local networks. IEEE Transactions on Computers, 37(10):1224–1234, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5984.

Robinson:1983:AOG

[Rob83] J. P. Robinson. Addendum to “Optimum Golomb Rulers”. IEEE Transactions on Computers, C-32(2):201, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676205.

Robinson:1985:ST

[Rob85] J. P. Robinson. Segmented testing. IEEE Transactions on Computers, C-34(5):467–471, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676586.

Rosenberg:1983:DAT

[Ros83] A. L. Rosenberg. The Diogenes approach to testable faulttolerant arrays of processors. IEEE Transactions on Computers, C-32(10):902–910, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676134.

See comments [Kor86].

Rosberg:1985:PSC

[Ros85a] Z. Rosberg. Process scheduling in a computer system. IEEE Transactions on Computers, C-34(7):633–645, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676601.

Rosenberg:1985:HMF

[Ros85b]  Arnold L. Rosenberg.  A

hypergraph model for faulttolerant VLSI processor arrays. IEEE Transactions on Computers, C-34(6):578–584, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009416.

Rosenberg:1986:AR

[Ros86] A. L. Rosenberg. Author’s reply. IEEE Transactions on Computers, C-35(1):93–94, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676669.

Roth:1986:MDA

[Rot86] J. P. Roth. Minimization by the D algorithm. IEEE Transactions on Computers, C-35(5):476–478, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676790.

Reddy:1986:TRF

[RR86] S. M. Reddy and M. K. Reddy.

Testable realizations for FET stuck-open faults in CMOS combinational logic circuits. IEEE Transactions on Computers, C-35(8):742–754, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676825.

Raghavan:1983:SRR

[RS83a] R. Raghavan and S. Sahni.

Single row routing. IEEE Transactions on Computers, C-32(3):209–220, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676213.

Reed:1983:CPB

[RS83b] D. A. Reed and H. D. Schwetman. Cost-performance bounds for multimicrocomputer networks. IEEE Transactions on Computers, C-32(1):83–95, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676126.

Reischer:1984:GFB

[RS84] Corina Reischer and Dan A. Simovici. Graph functions of Boolean functions. IEEE Transactions on Computers, C-33(1):97–99, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009317.

Ranganath:1985:CAP

[RS85] H. S. Ranganath and S. G. Shiva. Correlation of adjacent pixels for multiple image registration. IEEE Transactions on Computers, C-34(7):674–677, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676607.

Reischer:1986:IPM

[RS86a] C. Reischer and D. A. Simovici. Iteration properties of multivalued switching functions. IEEE Transactions on Computers, C-35(2):173–178, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676737.

Rhyne:1986:SBS

[RS86b] T. Rhyne and N. R. Strader II. A signed bit-sequential multiplier. IEEE Transactions on Computers, C-35(10):896–901, October 1986. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676680.

See comments [Smi89a].

Robinson:1987:UVT

[RS87] John P. Robinson and Nirmal R. Saxena. A unified view of test compression methods. IEEE Transactions on Computers, C-36(1):94–99, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009452.

Ramanathan:1988:RBH

[RS88a] P. Ramanathan and K. G.

Shin. Reliable broadcast in hypercube multicomputers. IEEE Transactions on Computers, 37(12):1654–1657, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9743.

Rom:1988:RAD

[RS88b] R. Rom and N. Shacham. A reconfiguration algorithm for a double-loop token ring local area network. IEEE Transactions on Computers, 37 (2):182–189, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2147.

Reddy:1988:DCT

[RSK88] S. M. Reddy, K. K. Saluja, and M. G. Karpovsky. A data compression technique for built-in self-test. IEEE Transactions on Computers, 37(9): 1151–1156, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2271.

Rotem:1985:DS

[RSS85] Doron Rotem, Nicola Santoro, and Jeffrey B. Sidney. Distributed sorting. IEEE Transactions on Computers, C-34(4):372–376, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009389.

Ramamritham:1989:DST

[RSZ89] K. Ramamritham, J. A. Stankovic, and W. Zhao. Distributed scheduling of tasks with deadlines and resource requirements. IEEE Transactions on Computers, 38(8):1110–1123, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30866.

Rajski:1985:CAM

[RT85] Janusz Rajski and Jerzy Tyszer. Combinatorial approach to multiple contact faults coverage in programmable logic arrays. IEEE Transactions on Computers, C-34(6):549–553, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009407.

Rajski:1986:IMP

[RT86] J. Rajski and J. Tyszer. The influence of masking phenomenon on coverage capability of single fault test sets in PLA’s. IEEE Transactions on Computers, C-35(1):81–85, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676665.

Reed:1981:ANH

[RTB81] I. S. Reed, T. K. Truong, and B. Benjauthrit. Addendum to “A New Hybrid Algorithm for Computing a Fast Discrete Fourier Transform”. IEEE Transactions on Computers, C-30(6):453–454, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675814.

Reed:1986:VDE

[RTJH86] I. S. Reed, T. K. Truong, J. M. Jensen, and In-Shek Hsu. The VLSI design of an error-trellis syndrome decoder for certain convolutional codes. IEEE Transactions on Computers, C-35(9):781–789, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676838.

Ramamoorthy:1980:AR

[RTW80] C. V. Ramamoorthy, J. L. Turner, and B. W. Wah. Authors’ reply. IEEE Transactions on Computers, C-29(8):757, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675665.

Rashid:1988:MIV

[RTY^{+}88] R. Rashid, A. Tevanian, Jr., M. Young, D. Golub, R. Baron, D. Black, W. J. Bolosky, and J. Chew. Machine-independent virtual memory management for paged uniprocessor and multiprocessor architectures. IEEE Transactions on Computers, 37(8):896–908, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2242.

Rubin:1981:FCO

[Rub81] F. Rubin. Further comments on “An Optimal Solution for the Channel-Assignment Problem”. IEEE Transactions on Computers, C-30(6):455, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675817.

See [Lau81, PTT81].

Rudolph:1985:RSN

[Rud85] Larry Rudolph. A robust sorting network. IEEE Transactions on Computers, C-34(4):326–335, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009383.

Ramakrishnan:1984:MMM

[RV84] I. V. Ramakrishnan and P. J.

Varman. Modular matrix multiplication on a linear array. IEEE Transactions on Computers, C-33(11):952–958, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676369.

Raghavendra:1986:FTM

[RV86] C. S. Raghavendra and A. Varma.

Fault-tolerant multiprocessors with redundant-path interconnection networks. IEEE Transactions on Computers, C-35(4):307–316, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676763.

Ramamoorthy:1981:DMU

[RW81a] C. V. Ramamoorthy and B. W.

Wah. The degradation in memory utilization due to dependencies. IEEE Transactions on Computers, C-30 (10):813–818, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675702.

Ramamoorthy:1981:OAS

[RW81b] C. V. Ramamoorthy and B. W. Wah. An optimal algorithm for scheduling requests on interleaved memories for a pipelined processor. IEEE Transactions on Computers, C-30(10):787–800, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675697.

Rudin:1982:VTT

[RW82] H. Rudin and C. H. West. A validation technique for tightly coupled protocols. IEEE Transactions on Computers, C-31(7):630–636, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676060.

Ramamoorthy:1983:ISF

[RW83] C. V. Ramamoorthy and B. W. Wah. The isomorphism of simple file allocation. IEEE Transactions on Computers, C-32(3):221–232, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676214.

Rosenberger:1984:CAP

[RW84] F. U. Rosenberger and D. F.

Wann. A computer aided procedure for performing static loading validation of digital logic systems. IEEE Transactions on Computers, C-33(4):301–313, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676435.

Robinson:1982:MMM

[RY82] J. P. Robinson and Chia-Lung Yeh. A method for modulo-2 minimization. IEEE Transactions on Computers, C-31(8):800–801, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676087.

Somani:1985:EUV

[SA85] A. K. Somani and V. K.

Agarwal. An efficient unsorted VLSI dictionary machine. IEEE Transactions on Computers, C-34(9):841–852, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676640.

Somani:1987:GTS

[SAA87] A. K. Somani, V. K. Agarwal, and D. Avis. A generalized theory for system level diagnosis. IEEE Transactions on Computers, C-36(5):538–546, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676938.

Somani:1989:CSF

[SAA89] A. K. Somani, V. K. Agarwal, and D. Avis. On the complexity of single fault set diagnosability and diagnosis problems. IEEE Transactions on Computers, 38(2):195–201, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=16496.

Sahni:1984:SMM

[Sah84] Sartaj Sahni. Scheduling multipipeline and multiprocessor computers. IEEE Transactions on Computers, C-33(7):637–645, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009337.

Saluja:1980:SSM

[Sal80] K. K. Saluja. Synchronous sequential machines: a modular and testable design. IEEE Transactions on Computers, C-29(11):1020–1025, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675499.

Samet:1980:ELP

[Sam80] H. Samet. Efficient on-line proofs of equalities and inequalities of formulas. IEEE Transactions on Computers, C-29(1):28–32, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675453.

Sanders:1988:ADF

[San88a]  B. A. Sanders.  An asyn-

chronous, distributed flow control algorithm for rate allocation in computer networks. IEEE Transactions on Computers, 37(7):779–787, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2223.

Sanders:1988:ICF

[San88b]  B. A. Sanders.  An in-

centive compatible flow control algorithm for rate allocation in computer networks. IEEE Transactions on Computers, 37(9):1067–1072, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2257.

Sanchis:1989:MWN

[San89] L. A. Sanchis. Multiple-way network partitioning. IEEE Transactions on Computers, 38(1):62–81, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8730.

Sasao:1981:MVD

[Sas81]  T. Sasao.  Multiple-valued

decomposition of generalized Boolean functions and the complexity of programmable logic arrays. IEEE Transactions on Computers, C-30 (9):635–643, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675861.

Sasao:1984:IVA

[Sas84] T. Sasao. Input variable assignment and output phase optimization of PLA’s. IEEE Transactions on Computers, C-33(10):879–894, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676349.

Sasao:1985:ADC

[Sas85] T. Sasao. An algorithm to derive the complement of a binary function with multiple-valued inputs. IEEE Transactions on Computers, C-34(2):131–140, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676549.

Sasao:1989:ODM

[Sas89] T. Sasao. On the optimal design of multiple-valued PLAs. IEEE Transactions on Computers, 38(4):582–592, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21150.

Savir:1980:DSI

[Sav80a] J. Savir. Detection of single intermittent faults in sequential circuits. IEEE Transactions on Computers, C-29(7):673–678, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675642.

Savir:1980:STD

[Sav80b] J. Savir. Syndrome-testable design of combinational circuits. IEEE Transactions on Computers, C-29(6):442–451, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675603.

See correction [Ano80z].

Savir:1980:TSI

[Sav80c] J. Savir. Testing for single intermittent failures in combinational circuits by maximizing the probability of fault detection. IEEE Transactions on Computers, C-29(5):410–416, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675595.

Savir:1981:STS

[Sav81]  J. Savir.  Syndrome-testing

of “syndrome-untestable” combinational circuits. IEEE Transactions on Computers, C-30(8):606–608, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675849.

Savir:1983:GCO

[Sav83a] J. Savir. Good controllability and observability do not guarantee good testability. IEEE Transactions on Computers, C-32(12):1198–1200, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676183.

Savir:1983:NET

[Sav83b] J. Savir. A new empirical test for the quality of random integer generators. IEEE Transactions on Computers, C-32(10):960–961, October 1983.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676142.

Savage:1984:SDC

[Sav84] Carla Savage. A systolic design for connectivity problems. IEEE Transactions on Computers, C-33(1):99–104, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009318.

Savir:1986:BDL

[Sav86]  J. Savir.  The bidirectional

double latch (BDDL). IEEE Transactions on Computers, C-35(1):65–66, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676659.

Sutton:1980:MRL

[SB80] J. C. Sutton III and J. G.

Bredeson. Minimal redundant logic for high reliability and irredundant testability. IEEE Transactions on Computers, C-29(7):648–656, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675636.

Siomalas:1983:PCB

[SB83] K. O. Siomalas and B. A. Bowen. Performance of crossbar multiprocessor systems. IEEE Transactions on Computers, C-32(7):689–695, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676301.

Savir:1984:RPTb

[SB84] J. Savir and P. H. Bardell. On random pattern test length. IEEE Transactions on Computers, C-33(6):467–474, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676470.

Sinha:1985:NCS

[SB85] B. P. Sinha and B. B. Bhattacharya. On the numerical complexity of short-circuit faults in logic networks. IEEE Transactions on Computers, C-34(2):186–190, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676557.

Scott:1988:MME

[SB88] D. S. Scott and J. Brandenburg. Minimal mesh embeddings in binary hypercubes. IEEE Transactions on Computers, 37(10):1284–1285, October 1988. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5990.

Sinha:1986:PAC

[SBGS86] B. P. Sinha, B. B. Bhattacharya, S. Ghose, and P. K. Srimani. A parallel algorithm to compute the shortest paths and diameter of a graph and its VLSI implementation. IEEE Transactions on Computers, C-35(11):1000–1004, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676702.

Sarin:1985:SAP

[SBK85] S. K. Sarin, B. Blaustein, and C. Kaufman. System architecture for partition-tolerant distributed databases. IEEE Transactions on Computers, C-34(12): 1158–1163, December 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Swern:1987:ELF

[SBMM87] Frederic L. Swern, Salvatore J.

Bavuso, Anna L. Martensen, and Paul S. Miner. The effects of latent faults on highly reliable computer systems. IEEE Transactions on Computers, C-36(8):1000–1005, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009523.

Saavedra-Barrera:1989:MCB

[SBSM89] R. H. Saavedra-Barrera, A. J. Smith, and E. Miya. Machine characterization based on an abstract high-level language machine. IEEE Transactions on Computers, 38(12): 1659–1679, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=40845.

Shin:1987:PAD

[SC87] K. G. Shin and Ming-Syan Chen. Performance analysis of distributed routing strategies free of ping-pong-type looping. IEEE Transactions on Computers, C-36(2):129–137, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676875.

Shin:1989:LSD

[SC89] K. G. Shin and Y.-C. Chang. Load sharing in distributed realtime systems with state-change broadcasts. IEEE Transactions on Computers, 38(8):1124–1143, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30867.

Suarez:1981:AR

[SCA81] R. E. Suarez, O. Chang, and V. Adam. Authors’ reply.

IEEE Transactions on Computers, C-30(12):987, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675741.

Schmidt:1989:AMP

[Sch89] D. C. Schmidt. An analytic model of printed circuit wiring distributions. IEEE Transactions on Computers, 38(6):898– 903, June 1989. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
24303.

Swartzlander:1986:AR

[SCN86] E. E. Swartzlander, Jr., D. V. Satish Chandra, and H. T. Nagle, Jr. Authors’ reply. IEEE Transactions on Computers, C-35(5):484, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676793.

Swartzlander:1983:SLA

[SCNS83] E. E. Swartzlander, Jr., D. V. Satish Chandra, H. T. Nagle, Jr., and S. A. Starks. Sign/logarithm arithmetic for FFT implementation. IEEE Transactions on Computers, C-32(6):526–534, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/

                                        stamp.jsp?tp=&arnumber=1676274.

See comments [HL86].

Scott:1989:FRM

[Sco89] M. Scott. Fast rounding in multiprecision floating-slash arithmetic. IEEE Transactions on Computers, 38(7):1049–1052, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30856.

Sengupta:1981:RFT

[SCP^{+}81] A. Sengupta, D. K. Chattopadhyay, A. Palit, A. K. Bandyopadhyay, and A. K. Choudhury. Realization of faulttolerant machines — linear code application. IEEE Transactions on Computers, C-30(3):237–240, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675762.

Su:1980:HRR

[SD80] S. Y. H. Su and E. DuCasse. A hardware redundancy reconfiguration scheme for tolerating multiple module failures. IEEE Transactions on Computers, C-29(3):254–258, March 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675557.

Saluja:1986:ASD

[SD86a] K. K. Saluja and R. Dandapani. An alternative to scan design methods for sequential machines. IEEE Transactions on Computers, C-35(4):384–388, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676776.

Saluja:1986:TDS

[SD86b] K. K. Saluja and R. Dandapani. Testable design of single-output sequential machines using checking experiments. IEEE Transactions on Computers, C-35(7):658–662, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676811.

Stiles:1987:QCP

[SD87] G. S. Stiles and Dong-Lih Denq.

A quantitative comparison of the performance of three discrete distributed associative memory models. IEEE Transactions on Computers, C-36(3):257–263, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676898.

Scheurich:1989:DPM

[SD89a] C. Scheurich and M. Dubois. Dynamic page migration in multiprocessors with distributed global memory. IEEE Transactions on Computers, 38 (8):1154–1163, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30869.

Sharrock:1989:ECC

[SD89b] S. M. Sharrock and D. H.- C. Du. Efficient CSMA/CDbased protocols for multiple priority classes. IEEE Transactions on Computers, 38(7):943–954, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30847.

Savir:1984:RPTa

[SDB84] Jacob Savir, Gary S. Ditlow, and Paul H. Bardell. Random pattern testability. IEEE Transactions on Computers, C-33(1):79–90, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009315.

Sadayappan:1987:NNM

[SE87] Ponnuswamy Sadayappan and Fikret Ercal. Nearest-neighbor mapping of finite element graphs onto processor meshes. IEEE Transactions on Computers, C-36(12):1408–1424, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009494.

Seitz:1984:CVA

[Sei84] C. L. Seitz. Concurrent VLSI architectures. IEEE Transactions on Computers, C-33 (12):1247–1265, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676396.

Seznec:1987:NIN

[Sez87] A. Seznec. A new interconnection network for SIMD computers: The sigma network. IEEE Transactions on Computers, C-36(7):794–801, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676976.

Shen:1984:DET

[SF84] J. P. Shen and F. J. Ferguson. The design of easily testable VLSI array multipliers. IEEE Transactions on Computers, C-33(6):554–560, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676480.

Swartzlander:1980:AUH

[SG80] E. E. Swartzlander, Jr. and B. K. Gilbert. Arithmetic for ultra-high-speed tomography.

IEEE Transactions on Computers, C-29(5):341–353, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675584.

Swartzlander:1982:STA

[SG82] E. E. Swartzlander, Jr. and B. K. Gilbert. Supersystems: Technology and architecture. IEEE Transactions on Computers, C-31(5):399–409, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676017.

Sheraga:1983:EAM

[SG83] R. J. Sheraga and J. L. Gieser. Experiments in automatic microcode generation. IEEE Transactions on Computers, C-32(6):557–569, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676278.

Smith:1985:ICR

[SG85a] J. E. Smith and J. R. Goodman. Instruction cache replacement policies and organizations. IEEE Transactions on Computers, C-34(3):234–241, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676566.

Stelzer:1985:ICR

[SG85b] K. J. Stelzer and M. A. Gordon.

Implementation of a constrained regularization program (CON-TIN) on a desktop computer. IEEE Transactions on Computers, C-34(9):862–863, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676644.

Singh:1981:TSS

[SGA81] A. D. Singh, F. G. Gray, and J. R. Armstrong. Tree structured sequential multiple-valued logic design from universal modules. IEEE Transactions on Computers, C-30(9):671–674, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675866.

Schwan:1987:CKS

[SGB87] Karsten Schwan, Prabha Gopinath, and Win Bo. CHAOS-kernel support for objects in the realtime domain. IEEE Transactions on Computers, C-36(8):904–916, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009514.

Shimizu:1989:CCP

[SGI89] K. Shimizu, E. Goto, and S. Ichikawa. CPC (cyclic pipeline computer)-an architecture suited for Josephson and pipelined-memory machines. IEEE Transactions on Computers, 38(6):825–832, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24291.

Somenzi:1985:TST

[SGMP85] F. Somenzi, S. Gai, M. Mezzalama, and P. Prinetto. Testing strategy and technique for macro-based circuits. IEEE Transactions on Computers, C-34(1):85–90, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676519.

Smotherman:1986:PCA

[SGT86] M. Smotherman, R. M. Geist, and K. S. Trivedi. Provably conservative approximations to complex reliability models. IEEE Transactions on Computers, C-35(4):333–338, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676766.

Sridhar:1981:DET

[SH81a] T. Sridhar and J. P. Hayes. Design of easily testable bitsliced systems. IEEE Transactions on Computers, C-30 (11):842–854, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675715.

Sridhar:1981:FAT

[SH81b] T. Sridhar and J. P. Hayes.

A functional approach to testing bit-sliced microprocessors. IEEE Transactions on Computers, C-30(8):563–571, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675841.

Shen:1984:FTD

[SH84] J. P. Shen and J. P. Hayes.

Fault-tolerance of dynamic-full-access interconnection networks. IEEE Transactions on Computers, C-33(3):241–248, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676421.

Szymanski:1987:PCM

[SH87] T. H. Szymanski and V. C.

Hamacher. On the permutation capability of multistage interconnection networks. IEEE Transactions on Computers, C-36(7):810–822, July 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676978.

Shanthikumar:1981:BBP

[Sha81] J. G. Shanthikumar. On the buffer behavior with Poisson arrivals, priority service, and random server interruptions. IEEE Transactions on Computers, C-30(10):781–786, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675696.

Shiozaki:1982:SAE

[Shi82a]  A. Shiozaki.  Single asym-

metric error-correcting cyclic AN codes. IEEE Transactions on Computers, C-31(6):554–555, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676040.

Shively:1982:APD

[Shi82b] R. R. Shively. Architecture of a programmable digital signal processor. IEEE Transactions on Computers, C-31(1):16–22, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675880.

Shin:1987:ISI

[Shi87] Kang G. Shin. Introduction to the special issue on real-time systems. IEEE Transactions on Computers, C-36(8):901–902, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009512.

Siegel:1980:TUP

[Sie80]  H. J. Siegel.  The the-

ory underlying the partitioning of permutation networks. IEEE Transactions on Computers, C-29(9):791–801, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675678.

Siegel:1981:INP

[Sie81]  H. J. Siegel.  Interconnec-

tion networks for parallel and distributed processing: An overview. IEEE Transactions on Computers, C-30(4):245–246, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675771.

Siegel:1985:MSS

[Sie85a] Alan R. Siegel. Minimum storage sorting networks. IEEE Transactions on Computers, C-34(4):355–361, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009386.

Siegenthaler:1985:DCS

[Sie85b]  T. Siegenthaler.  Decrypt-

ing a class of stream ciphers using ciphertext only.

IEEE Transactions on Computers, C-34(1):81–85, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676518.

This paper breaks the cipher of [Gef73].

Signaevskii:1982:CMS

[Sig82] V. A. Signaevskii. Comments on “Multiprocessor Scheduling with Memory Allocation — A Deterministic Approach”. IEEE Transactions on Computers, C-31(9):898, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676105.

Silberman:1982:DFR

[Sil82] G. M. Silberman. Determining fault ratios in multilevel delayed staging storage hierarchies. IEEE Transactions on Computers, C-31(4):305–310, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675996.

Silberman:1983:DSH

[Sil83] G. M. Silberman. Delayedstaging hierarchy optimization. IEEE Transactions on Computers, C-32(11):1029–1037, November 1983. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1676153.

Sinclair:1988:OAB

[Sin88a] James B. Sinclair. Optimal assignments in broadcast networks. IEEE Transactions on Computers, 37(5):521–531, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4603.

Reviews: Computing Reviews, Vol. 30, No. 7.

Singh:1988:IRA

[Sin88b] A. D. Singh. Interstitial redundancy: an area efficient fault tolerance scheme for large area VLSI processor arrays. IEEE Transactions on Computers, 37(11):1398–1410, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8705.

Singhal:1989:HAA

[Sin89] M. Singhal. A heuristicallyaided algorithm for mutual exclusion in distributed systems. IEEE Transactions on Computers, 38(5):651–662, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24268.

Sips:1982:CPM

[Sip82]  H. J. Sips.  Comments on

“An O(n) Parallel Multiplier with Bit-Sequential Input and Output”. IEEE Transactions on Computers, C-31(4):325–327, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676000.

Sips:1984:BSA

[Sip84] Henk J. Sips. Bit-sequential arithmetic for parallel processors. IEEE Transactions on Computers, C-33(1):7–20, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009311.

Segel:1984:GEI

[SJ84] H. J. Segel and L. H. Jamieson. Guest Editors’ introduction parallel processing. IEEE Transactions on Computers, C-33 (11):949–951, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676367.

Saluja:1985:TPG

[SK85] K. K. Saluja and K. Kinoshita. Test pattern generation for API faults in RAM. IEEE Transactions on Computers, C-34(3):284–287, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676572.

Shenoy:1989:FBE

[SK89] A. P. Shenoy and R. Kumaresan. Fast base extension using a redundant modulus in RNS. IEEE Transactions on Computers, 38(2):292–297, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16508.

Saluja:1983:ETD

[SKF83] K. K. Saluja, K. Kinoshita, and H. Fujiwara. An easily testable design of programmable logic arrays for multiple faults. IEEE Transactions on Computers, C-32(11):1038–1046, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676154.

Skillicorn:1988:NCF

[Ski88] D. B. Skillicorn. A new class of fault-tolerant static interconnection networks. IEEE Transactions on Computers, 37(11):1468–1470, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8721.

Steiglitz:1988:ECO

[SKW88] K. Steiglitz, I. Kamal, and A. Watson. Embedding computation in one-dimensional automata by phase coding solitons. IEEE Transactions on Computers, 37(2):138–145, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2143.

Sedmak:1980:FTG

[SL80] R. M. Sedmak and H. L. Liebergot. Fault tolerance of a general purpose computer implemented by very large scale integration. IEEE Transactions on Computers, C-29(6):492–500, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675608.

Seth:1983:SMC

[SL83a] S. C. Seth and L. Lipsky. A simplified method to calculate failure times in fault-tolerant systems. IEEE Transactions on Computers, C-32(8):754–756, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676315.

Smith:1983:TTS

[SL83b] J. E. Smith and P. Lam. A theory of totally self-checking system design. IEEE Transactions on Computers, C-32 (9):831–844, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676332.

Shin:1984:EDP

[SL84] K. G. Shin and Yann-Hang Lee. Error detection process— model, design, and its impact on computer performance. IEEE Transactions on Computers, C-33(6):529–540, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676476.

Shin:1986:MAF

[SL86] K. G. Shin and Yann-Hang Lee.

Measurement and application of fault latency. IEEE Transactions on Computers, C-35(4):370–375, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676773.

Schwartz:1987:DMC

[SL87] Alan M. Schwartz and Michael C.

Loui. Dictionary machines on cube-class networks. IEEE Transactions on Computers, C-36(1):100–105, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009453.

Sciuto:1988:FTA

[SL88a] D. Sciuto and F. Lombardi. On functional testing of array processors. IEEE Transactions on Computers, 37(11): 1480–1484, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8724.

Shin:1988:MME

[SL88b] K. G. Shin and T.-H. Lin. Modeling and measurement of error propagation in a multimodule computing system. IEEE Transactions on Computers, 37(9):1053–1066, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2256.

Suzuki:1989:TPN

[SL89] I. Suzuki and H. Lu. Temporal Petri nets and their application to modeling and analysis of a handshake daisy chain arbiter. IEEE Transactions on Computers, 38(5):696–704, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24271.

Sha:1988:MCC

[SLJ88] L. Sha, J. P. Lehoczky, and E. D. Jensen. Modular concurrency control and failure recovery. IEEE Transactions on Computers, 37(2):146–159, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2144.

Shin:1987:OCR

[SLL87] Kang G. Shin, Tein-Hsiang Lin, and Yann-Hang Lee. Optimal checkpointing of realtime tasks. IEEE Transactions on Computers, C-36 (11):1328–1341, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009472.

Sloan:1985:CCA

[Slo85] K. R. Sloan, Jr. Comments on “A Computer Algorithm for Calculating the Product AB Modulo M”. IEEE Transactions on Computers, C-34(3):290–292, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676574.

See [Bla83].

Shin:1982:DHH

[SLS82] K. G. Shin, Yann-Hang Lee, and J. Sasidhar. Design of HM^{2}p — A hierarchical multimicroprocessor for generalpurpose applications. IEEE Transactions on Computers, C-31(11):1045–1053, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675921.

Schmidt:1980:AR

[SM80] D. C. Schmidt and G. Metze. Authors’ reply. IEEE Transactions on Computers, C-29(1):60–61, January 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675459.

Sowa:1982:DFC

[SM82] M. Sowa and T. Murata. A data flow computer architecture with program and token memories. IEEE Transactions on Computers, C-31(9):820–824, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676097.

Sakura:1983:PBA

[SM83] A. Sakura and S. Muroga. Parallel binary adders with a minimum number of connections. IEEE Transactions on Computers, C-32(10):969–976, October 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676144.

Serra:1987:TPL

[SM87] M. Serra and J. C. Muzio. Testing programmable logic arrays by sum of syndromes. IEEE Transactions on Computers, C-36(9):1097–1101, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009540.

Savir:1988:RPT

[SM88] J. Savir and W. H. McAnney.

Random pattern testability of delay faults. IEEE Transactions on Computers, 37(3):291–300, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2166.

Smith:1980:MEF

[Smi80a] J. E. Smith. Measures of the effectiveness of fault signature analysis. IEEE Transactions on Computers, C-29(6):510–514, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675610.

Smith:1980:CNP

[Smi80b]  R. G. Smith.  The Con-

tract Net Protocol: High-level communication and control in a distributed problem solver. IEEE Transactions on Computers, C-29(12):1104–1113, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/

                                        stamp.jsp?tp=&arnumber=1675516.

See correction [Smi81b].

Smith:1981:PML

[Smi81a] K. C. Smith. The prospects for multivalued logic: a technology and applications view. IEEE Transactions on Computers, C-30(9):619–634, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675860.

Smith:1981:CCN

[Smi81b] R. G. Smith. Correction to “The Contract Net Protocol: High-Level Communication and Control in a Distributed Problem Solver”. IEEE Transactions on Computers, C-30(5):372, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675799.

See [Smi80b].

Smith:1984:SUC

[Smi84] James E. Smith. On separable unordered codes. IEEE Transactions on Computers, C-33(8):741–743, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009361.

Smilauer:1985:GMM

[Smi85] B. Smilauer. General model for memory interference in multiprocessors and mean value analysis. IEEE Transactions on Computers, C-34(8):744–751, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676621.

Smith:1986:AR

[Smi86]  J. E. Smith.  Author’s re-

ply. IEEE Transactions on Computers, C-35(10):931, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676688.

Smith:1987:LBS

[Smi87] Alan Jay Smith. Line (block) size choice for CPU cache memories. IEEE Transactions on Computers, C-36(9):1063–1075, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009537.

Smith:1989:CSB

[Smi89a] S. G. Smith. Comments on “A signed bit-sequential multiplier” by T. Rhyne and N. R. Strader II. IEEE Transactions on Computers, 38(9): 1328–1330, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29473.

See [RS86b].

Smith:1989:ICS

[Smi89b]  S. G. Smith.  Incremental

computation of squares and sums of squares. IEEE Transactions on Computers, 38(9): 1325–1328, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29472.

Sholl:1982:MSR

[SMN82] H. A. Sholl, K. Morris, and J. Norris. A multimicroprocessor system for real-time classification of railroad track flaws. IEEE Transactions on Computers, C-31(10):1009–1017, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675910.

Savir:1987:FPT

[SMV87] J. Savir, W. H. McAnney, and S. R. Vecchio. Fault propagation through embedded multiport memories. IEEE Transactions on Computers, C-36(5):592–602, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676944.

Seth:1981:GMP

[SN81] S. C. Seth and K. Narayanaswamy. A graph model for patternsensitive faults in random access memories. IEEE Transactions on Computers, C-30 (12):973–977, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675737.

Snyder:1981:FMC

[Sny81]  L. Snyder.  Formal models

of capability-based protection systems. IEEE Transactions on Computers, C-30(3):172–181, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675753.

Shen:1989:SCP

[SO89] W. Shen and A. Y. Oruc. On systolic contractions of program graphs. IEEE Transactions on Computers, 38 (10):1451–1457, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35839.

Shimada:1985:NAC

[SOA85] R. Shimada, Y. Ohkura, and J. Aoe. Nonbinary arithmetic AN codes using odd radix expressions. IEEE Transactions on Computers, C-34 (11):1050–1056, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676538.

Shirakawa:1981:LSR

[SOH^{+}81] I. Shirakawa, N. Okuda, T. Harada, S. Tani, and H. Ozaki. A layout system for the random logic portion of an MOS LSI chip. IEEE Transactions on Computers, C-30(8):572–581, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675842.

Sohi:1989:CMO

[Soh89] G. S. Sohi. Cache memory organization to enhance the yield of high performance VLSI processors. IEEE Transactions on Computers, 38(4):484–492, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=21141.

Sorensen:1985:APP

[Sor85] D. C. Sorensen. Analysis of pairwise pivoting in Gaussian elimination. IEEE Transactions on Computers, C-34(3):274–278, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676570.

Shrivastava:1982:DRR

[SP82] S. K. Shrivastava and F. Panzieri. The design of a reliable remote procedure call mechanism. IEEE Transactions on Computers, C-31(7):692–697, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676069.

See comments [Cas86].

Smith:1988:IPI

[SP88] J. E. Smith and A. R. Pleszkun.

Implementing precise interrupts in pipelined processors. IEEE Transactions on Computers, 37(5):562–573, May 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4607.

Samatham:1989:BMN

[SP89] M. R. Samatham and D. K.

Pradhan. The de Bruijn multiprocessor network: a versatile parallel processing and sorting network for VLSI. IEEE Transactions on Computers, 38(4):567–581, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21149.

Suk:1980:TPC

[SR80] D. S. Suk and S. M. Reddy.

Test procedures for a class of pattern-sensitive faults in semiconductor random-access memories. IEEE Transactions on Computers, C-29(6):419–429, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675601.

Steinberg:1981:LSE

[SR81a] D. Steinberg and M. Rodeh. A layout for the shuffle-exchange network with O(N^{2}/ log^{3/2} N) area. IEEE Transactions on Computers, C-30(12):977–982, December 1981. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675738.

Suk:1981:MTF

[SR81b] D. S. Suk and S. M. Reddy. A march test for functional faults in semiconductor random access memories. IEEE Transactions on Computers, C-30 (12):982–985, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675739.

Strader:1982:CBS

[SR82] N. R. Strader and V. T. Rhyne. A canonical bit-sequential multiplier. IEEE Transactions on Computers, C-31(8):791–795, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676085.

Saxena:1986:ACT

[SR86] N. R. Saxena and J. P. Robinson. Accumulator compression testing. IEEE Transactions on Computers, C-35(4):317–321, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676764.

Shin:1987:CSL

[SR87] Kang G. Shin and P. Ramanathan. Clock synchronization of a large multiprocessor system in the presence of malicious faults. IEEE Transactions on Computers, C-36(1):2–12, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009444.

Shao:1988:VDP

[SR88a] H. M. Shao and I. S. Reed.

On the VLSI design of a pipeline Reed–Solomon decoder using systolic arrays. IEEE Transactions on Computers, 37 (10):1273–1280, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5988.

Shin:1988:TDH

[SR88b] K. G. Shin and P. Ramanathan.

Transmission delays in hardware clock synchronization. IEEE Transactions on Computers, 37(11):1465–1467, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8720.

So:1988:COM

[SR88c] K. So and R. N. Rechtschaffen. Cache operations by MRU change. IEEE Transactions on Computers, 37(6):700–709, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2208.

Stankovic:1985:EFT

[SRC85] J. Stankovic, K. Ramamritham, and S. Chang. Evaluation of a flexible task scheduling algorithm for distributed hard realtime systems. IEEE Transactions on Computers, C-34(12): 1130–1143, December 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Srinivas:1983:OPS

[Sri83] M. A. Srinivas. Optimal parallel scheduling of Gaussian elimination DAG’s. IEEE Transactions on Computers, C-32(12):1109–1117, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676171.

Srinivasan:1987:SIB

[Sri87] M. M. Srinivasan. Successively improving bounds on performance measures for single class product form queueing networks.

IEEE Transactions on Computers, C-36(9):1107–1112, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009542.

Sridhar:1989:FAT

[Sri89] M. A. Sridhar. A fast algorithm for testing isomorphism of permutation networks. IEEE Transactions on Computers, 38(6):903–909, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24304.

Shively:1984:CTG

[SRO84] R. R. Shively, W. V. Robinson, and D. E. Orton. Cascading transmission gates to enhance multiplier performance. IEEE Transactions on Computers, C-33(7):677–679, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009343.

Samari:1980:QTB

[SS80] N. K. Samari and G. M.

Schneider. A queueing theorybased analytic model of a distributed computer network. IEEE Transactions on Computers, C-29(11):994–1001, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675494.

Snow:1981:IPE

[SS81] E. A. Snow and D. P. Siewiorek. Implementation and performance evaluation of computer families. IEEE Transactions on Computers, C-30(6):443–447, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675810.

Snyder:1982:CAR

[SS82] W. E. Snyder and C. D. Savage. Content-addressable read/ write memories for image analysis. IEEE Transactions on Computers, C-31(10):963–968, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675905.

Seban:1984:SIP

[SS84] Robert R. Seban and Howard Jay Siegel. Shuffling with the Illiac and PM2I SIMD networks. IEEE Transactions on Computers, C-33(7):619–625, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009335.

Schmeck:1985:DMD

[SS85a] H. Schmeck and H. Schroder. Dictionary machines for different models of VLSI. IEEE Transactions on Computers, C-34(5):472–475, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676587.

Spencer:1985:LIT

[SS85b] T. H. Spencer and J. Savir.

Layout influences testability. IEEE Transactions on Computers, C-34(3):287–290, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676573.

Sengupta:1986:DGM

[SS86] A. Sengupta and A. Sen. On the diagnosability of a general model of system with threevalued test outcomes. IEEE Transactions on Computers, C-35(2):170–173, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676736.

Schuette:1987:PCF

[SS87] M. A. Schuette and J. P. Shen.

Processor control flow monitoring using signatured instruction streams. IEEE Transactions on Computers, C-36(3):264–276, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676899.

Saad:1988:TPH

[SS88a] Y. Saad and M. H. Schultz. Topological properties of hypercubes. IEEE Transactions on Computers, 37(7):867–872, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2234.

Shen:1988:FTM

[SS88b] L. Shen and S. Y. H. Su. A functional testing method for microprocessors. IEEE Transactions on Computers, 37 (10):1288–1293, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5992.

Santoro:1989:RTS

[SS89a] N. Santoro and E. Suen. Reduction techniques for selection in distributed files. IEEE Transactions on Computers, 38(6):891–896, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=24301.

Scherson:1989:PST

[SS89b] I. D. Scherson and S. Sen. Parallel sorting in two-dimensional VLSI models of computation. IEEE Transactions on Computers, 38(2):238–249, February 1989. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16500.

Sinha:1989:FPA

[SS89c] B. P. Sinha and P. K. Srimani.

Fast parallel algorithms for binary multiplication and their implementation on systolic architectures. IEEE Transactions on Computers, 38(3):424–431, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21128.

Sengupta:1986:SDP

[SSB86] A. Sengupta, A. Sen, and S. Bandyopadhyay. On system diagnosability in the presence of hybrid faults. IEEE Transactions on Computers, C-35(1):90–93, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676667.

Sengupta:1987:OFT

[SSB87] A. Sengupta, A. Sen, and S. Bandyopadhyay. On an optimally fault-tolerant multiprocessor network architecture. IEEE Transactions on Computers, C-36(5):619–623, May 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676947.

Simoncini:1980:DSD

[SSF80] L. Simoncini, F. Saheban, and A. D. Friedman. Design of self-diagnosable multiprocessor systems with concurrent computation and diagnosis. IEEE Transactions on Computers, C-29(6):540–546, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675617.

Siegel:1982:PPA

[SSF82] L. J. Siegel, H. J. Siegel, and A. E. Feather. Parallel processing approaches to image correlation. IEEE Transactions on Computers, C-31(3):208–218, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675976.

Siegel:1981:PPS

[SSK^{+}81] H. J. Siegel, L. J. Siegel, F. C. Kemmerer, P. T. Mueller, Jr., H. E. Smalley, Jr., and S. D. Smith. PASM: a partitionable SIMD/MIMD system for image processing and pattern recognition. IEEE Transactions on Computers, C-30 (12):934–947, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675732.

Segall:1983:IIE

[SSS^{+}83] Z. Segall, A. Singh, R. T.

Snodgrass, A. K. Jones, and D. P. Siewiorek. An integrated instrumentation environment for multiprocessors. IEEE Transactions on Computers, C-32(1):4–14, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676119.

Schmeck:1989:SWM

[SSS89] H. Schmeck, H. Schroder, and C. Starke. Systolic s^{2}- way merge sort is optimal. IEEE Transactions on Computers, 38(7):1052–1056, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30857.

Saeks:1981:DNC

[SSVV81] R. Saeks, A. Sangiovanni-Vincentelli, and V. Visvanathan. Diagnosability of nonlinear circuits and systems — Part II: Dynamical systems. IEEE Transactions on Computers, C-30(11):899–904, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675721.

Shen:1985:GMA

[ST85] Chien-Chung Shen and Wen-Hsiang Tsai. A graph matching approach to optimal task assignment in distributed computing systems using a minimax criterion. IEEE Transactions on Computers, C-34(3):197–203, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676563.

Snyers:1986:ASM

[ST86a] D. Snyers and A. Thayse. Algorithmic state machine design and automatic theorem proving: Two dual approaches to the same activity. IEEE Transactions on Computers, C-35(10):853–861, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676676.

Sousa:1986:CIC

[ST86b] M. Sousa and F. Taylor. Complex integer to complex residue encoding. IEEE Transactions on Computers, C-35(7):648–650, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676808.

Su:1987:MOM

[ST87] Stanley Y. W. Su and Arun K. Thakore. Matrix operations on a multicomputer system with switchable main memory modules and dynamic control. IEEE Transactions on Computers, C-36(12):1467–1484, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009498.

Schwiegelshohn:1988:SAA

[ST88a] U. Schwiegelshohn and L. Thiele.

A systolic array for the assignment problem. IEEE Transactions on Computers, 37(11): 1422–1425, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8709.

Stouraitis:1988:FPL

[ST88b] Thanos Stouraitis and Fred J.

Taylor. Floating-point to logarithmic encoder error analysis. IEEE Transactions on Computers, 37(7):858–863, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2232.

Stankovic:1981:TIV

[Sta81] J. A. Stankovic. The types and interactions of vertical migrations of functions in a multilevel interpretive system. IEEE Transactions on Computers, C-30(7):505–513, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675829.

Stankovic:1984:PDC

[Sta84] J. A. Stankovic. A perspective on distributed computer systems. IEEE Transactions on Computers, C-33(12):1102–1115, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676389.

Stankovic:1985:ABD

[Sta85]  J. A. Stankovic.  An ap-

plication of Bayesian decision theory to decentralized control of job scheduling. IEEE Transactions on Computers, C-34(2):117–130, February 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676548.

Stankovic:1987:IPD

[Sta87]  J. A. Stankovic.  Introduc-

tion — parallel and distributed computing. IEEE Transactions on Computers, C-36(4):385–386, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676919.

Staskaukas:1988:FSD

[Sta88a] Mark G. Staskaukas. The formal specification and design of a distributed electronic fundstransfer system. IEEE Transactions on Computers, 37(12): 1515–1528, December 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Staskauskas:1988:FSD

[Sta88b]  M. G. Staskauskas.  The

formal specification and design of a distributed electronic funds-transfer system. IEEE Transactions on Computers, 37(12):1515–1528, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9730.

Stankovic:1989:DDM

[Sta89] J. A. Stankovic. Decentralized decision-making for task reallocation in a hard real-time system. IEEE Transactions on Computers, 38(3):341–355, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21121.

Shao:1985:VDP

[STD^{+}85] H. M. Shao, T. K. Truong, L. J. Deutsch, J. H. Yuen, and I. S. Reed. A VLSI design of a pipeline Reed–Solomon decoder. IEEE Transactions on Computers, C-34(5):393–403, May 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676579.

Steinberg:1983:IPS

[Ste83] D. Steinberg. Invariant properties of the shuffle-exchange and a simplified cost-effective version of the Omega network. IEEE Transactions on Computers, C-32(5):444–450, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676255.

Stiffler:1980:FTC

[Sti80] J. J. Stiffler. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-29(6):417–419, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675600.

Stoffers:1980:TSC

[Sto80a]  K. E. Stoffers.  Test sets

for combinational logic — the edge-tracing approach. IEEE Transactions on Computers, C-29(8):741–746, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675660.

Stoughton:1980:MLM

[Sto80b] J. W. Stoughton. Modification of “A Least Mean Squares Cubic Algorithm for On-Line Differential of Sampled Analog Signals”. IEEE Transactions on Computers, C-29(10):953, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675484.

Stout:1983:MCC

[Sto83] Q. F. Stout. Mesh-connected computers with broadcasting. IEEE Transactions on Computers, C-32(9):826–830, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676331.

Stone:1984:DAF

[Sto84] Harold S. Stone. Database applications of the FETCH-AND-ADD instruction. IEEE Transactions on Computers, C-33(7):604–612, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009333.

Stone:1989:OSP

[Sto89] H. S. Stone. Optimal search policies for searches with I/O bound tasks. IEEE Transactions on Computers, 38(9): 1314–1320, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29470.

Strader:1982:CMB

[Str82] N. R. Strader II. Comments on “Magnetic Bubble Memory Architectures for Supporting Associative Searching of Relational Databases”. IEEE Transactions on Computers, C-31(3):265–266, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675984.

Shyu:1987:CIM

[STR87] H. C. Shyu, T. K. Truong, and I. S. Reed. A complex integer multiplier using the quadraticpolynomial residue number system with numbers of form 22n +

1. IEEE Transactions on Computers, C-36(10):1255–1258, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676868;
 http://ieeexplore.ieee.org/
xpl/RecentIssue.jsp?punumber=
12;   http://ieeexplore.ieee.
org/xpl/tocresult.jsp?isnumber=
35260.

Smith:1988:PAM

[STR88] R. M. Smith, K. S. Trivedi, and A. V. Ramesh. Performability analysis: measures, an algorithm, and a case study. IEEE Transactions on Computers, 37(4):406–417, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2184.

Sullivan:1988:FIA

[Sul88] G. F. Sullivan. A O(t^{3} + |E|) fault identification algorithm for diagnosable systems. IEEE Transactions on Computers, 37(4):388–397, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2182.

Surjaatmadja:1981:ASC

[Sur81] J. B. Surjaatmadja. An algebra for switching circuits. IEEE Transactions on Computers, C-30(8):609–613, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675851.

Susskind:1983:TVW

[Sus83]  A. K. Susskind.  Testing

by verifying Walsh coefficients. IEEE Transactions on Computers, C-32(2):198–201, February 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676204.

Sadayappan:1988:CSS

[SV88] P. Sadayappan and V. Visvanathan. Circuit simulation on shared-memory multiprocessors. IEEE Transactions on Computers, 37(12):1634–1642, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9740.

Six:1982:CRI

[SW82] H. W. Six and D. Wood. Counting and reporting intersections of d-ranges. IEEE Transactions on Computers, C-31(3):181–187, March 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675973.

Scheuermann:1984:HAB

[SW84] P. Scheuermann and G. Wu. Heuristic algorithms for broadcasting in point-to-point computer networks. IEEE Transactions on Computers, C-33 (9):804–811, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676496.

Stewart:1988:SSC

[SW88] W. K. Stewart and S. A. Ward. A solution to a special case of the synchronization problem. IEEE Transactions on Computers, 37(1):123–125, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75145.

Shenker:1989:OCH

[SW89] S. Shenker and A. Weinrib. The optimal control of heterogeneous queueing systems: a paradigm for load-sharing and routing. IEEE Transactions on Computers, 38(12):1724–1735, December 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=40850.

Swartzlander:1980:MA

[Swa80] E. E. Swartzlander, Jr. Merged arithmetic. IEEE Transactions on Computers, C-29 (10):946–950, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675482.

Stepoway:1984:MAG

[SWK84] S. L. Stepoway, D. L. Wells, and G. R. Kane. A multiprocessor architecture for generating fractal surfaces. IEEE Transactions on Computers, C-33(11):1041–1045, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676380.

Smith:1986:SSD

[SWP86] J. E. Smith, S. Weiss, and N. Y.

Pang. A simulation study of decoupled architecture computers. IEEE Transactions on Computers, C-35(8):692–702, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676820.

Schwab:1983:AMF

[SY83] T. F. Schwab and S. S. Yau. An algebraic model of faultmasking logic circuits. IEEE Transactions on Computers, C-32(9):809–825, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676330.

Saito:1989:CLS

[SYJ89] S. Saito, H. Yoshida, and T. L. Junii. The CrossoverNet LAN system using an intelligent head-end. IEEE Transactions on Computers, 38(8):1076–1085, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saito:1989:CNL

[SYK89] S. Saito, H. Yoshida, and T. L. Kunii. The Crossover Net LAN system using an intelligent head-end. IEEE Transactions on Computers, 38(8):1076–1085, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30863.

Thatte:1980:TGM

[TA80] S. M. Thatte and J. A. Abraham. Test generation for microprocessors. IEEE Transactions on Computers, C-29(6):429–441, June 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675602.

Treuer:1987:NBS

[TAF87] R. Treuer, V. K. Agarwal, and H. Fujiwara. A new built-in self-test design for PLA’s with high fault coverage and low overhead. IEEE Transactions on Computers, C-36(3):369–373, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676911.

Tamesada:1980:SMH

[Tam80] T. Tamesada. Sequential machines having quasi-stable states. IEEE Transactions on Computers, C-29(5):405–408, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675593.

Tanner:1984:FTM

[Tan84] R. M. Tanner. Fault-tolerant 256K memory designs. IEEE Transactions on Computers, C-33(4):314–322, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676436.

Tasaka:1983:SPR

[Tas83]  S. Tasaka.  Stability and

performance of the R-Aloha packet broadcast system. IEEE Transactions on Computers, C-32(8):717–726, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676309.

Taylor:1982:VRA

[Tay82]  F. J. Taylor.  A VLSI

residue arithmetic multiplier. IEEE Transactions on Computers, C-31(6):540–546, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676036.

Taylor:1983:OFR

[Tay83] F. J. Taylor. An overflowfree residue multiplier. IEEE Transactions on Computers, C-32(5):501–504, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676263.

Taylor:1982:PDS

[TB82] D. J. Taylor and J. P. Black. Principles of data structure error correction. IEEE Transactions on Computers, C-31(7):602–608, July 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676057.

Thomasian:1986:AQN

[TB86] A. Thomasian and P. F. Bay.

Analytic queueing network models for parallel processing of task systems. IEEE Transactions on Computers, C-35 (12):1045–1054, December 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676712.

Tang:1984:LTP

[TC84] D. T. Tang and Chin-Long Chen. Logic test pattern generation using linear codes. IEEE Transactions on Computers, C-33(9):845–850, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676501.

Truong:1986:TCD

[TCH^{+}86] T. K. Truong, J. J. Chang, I. S.

Hsu, D. Y. Pei, and I. S. Reed. Techniques for computing the discrete Fourier transform using the quadratic residue Fermat number systems. IEEE Transactions on Computers, C-35(11):1008–1012, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676704.

Chen:1988:PAM

[tCpS88] Wen tsuen Chen and Jang ping Sheu. Performance analysis of multistage interconnection networks with hierarchical requesting model. IEEE Transactions on Computers, 37(11):1438–1442, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8713.

Teng:1983:CPM

[Ten83] Meng-Hee Teng. Comments on “The Prime Memory Systems for Array Access”. IEEE Transactions on Computers, C-32(11):1072, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676162.

See [LV82].

Tanaka:1982:CEC

[TF82] E. Tanaka and King Sun Fu. Correction to “Error-Correcting Parsers for Formal Languages”. IEEE Transactions on Computers, C-31(4):327–328, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676001.

Taylor:1988:BLN

[TGJR88] F. J. Taylor, R. Gill, J. Joseph, and J. Radke. A 20 bit logarithmic number system processor. IEEE Transactions on Computers, 37(2):190–200, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2148.

Taylor:1982:ARM

[TH82] F. J. Taylor and C. H. Huang.

An autoscale residue multiplier. IEEE Transactions on Computers, C-31(4):321–325, April 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675999.

Thayse:1982:SOP

[Tha82]  A. Thayse.  Synthesis and

optimization of programs by means of P -functions. IEEE Transactions on Computers, C-31(1):34–40, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675883.

Thayse:1984:MFA

[Tha84a] A. Thayse. A matrix formalism for asynchronous implementation of algorithms. IEEE Transactions on Computers, C-33(4):289–300, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676434.

Thayse:1984:SAI

[Tha84b] A. Thayse. Synthesis and asynchronous implementation of algorithms using a generalized P -function concept. IEEE Transactions on Computers, C-33(10):861–868, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676347.

Townsend:1980:NMA

[THH80] M. Townsend, M. E. Hoff, Jr., and R. E. Holm. An NMOS microprocessor for analog signal processing. IEEE Transactions on Computers, C-29(2):97–102, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675532.

Thiebaut:1989:FDC

[Thi89] D. Thiebaut. On the fractal dimension of computer programs and its application to the prediction of the cache miss ratio. IEEE Transactions on Computers, 38(7):1012–1026, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30852.

Tao:1988:ECU

[THL88] D. L. Tao, C. R. P. Hartmann, and P. K. Lala. An efficient class of unidirectional error detecting/correcting codes. IEEE Transactions on Computers, 37(7):879–882, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2237.

Thompson:1983:FTV

[Tho83a] C. D. Thompson. Fourier transforms in VLSI. IEEE Transactions on Computers, C-32(11):1047–1057, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676155.

Thompson:1983:VCS

[Tho83b] C. D. Thompson. The VLSI complexity of sorting. IEEE Transactions on Computers, C-32(12):1171–1184, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676178.

Tilove:1980:SMC

[Til80] R. B. Tilove. Set membership classification: a unified approach to geometric intersection problems. IEEE Transactions on Computers, C-29 (10):874–883, October 1980.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675470.

Trevillyan:1986:GFA

[TJB86] L. Trevillyan, W. Joyner, and L. Berman. Global flow analysis in automatic logic design. IEEE Transactions on Computers, C-35(1):77–81, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676664.

Trees:1986:DEO

[TKK86] B. Trees, S. P. Kartashev, and S. I. Kartashev. Data exchange optimization in reconfigurable. IEEE Transactions on Computers, C-35(3):257–273, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676749.

Thomas:1986:CMI

[TKL86] J. J. Thomas, J. M. Keller, and G. N. Larsen. The calculation of multiplicative inverses over GF(P) efficiently where P is a Mersenne prime. IEEE Transactions on Computers, C-35(5):478–482, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676791.

Tabak:1980:MAD

[TL80] D. Tabak and G. J. Lipovski.

Move architecture in digital controllers. IEEE Transactions on Computers, C-29 (2):180–190, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675541.

Truong:1983:PPA

[TLR83] T. K. Truong, K. Y. Liu, and I. S. Reed. A parallel-pipeline architecture of the Fast Polynomial Transform for computing a two-dimensional cyclic convolution. IEEE Transactions on Computers, C-32(3):301–306, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676222.

Thanawastien:1981:IAS

[TN81] S. Thanawastien and V. P.

Nelson. Interference analysis of shuffle/exchange networks. IEEE Transactions on Computers, C-30(8):545–556, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675839.

Tanaka:1981:RPL

[TOM81] M. Tanaka, S. Ozawa, and S. Mori. Rewritable programmable logic array of current mode logic. IEEE Transactions on Computers, C-30(3):229–234, March 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675760.

Topkis:1989:AAB

[Top89] D. M. Topkis. All-to-all broadcast by flooding in communications networks. IEEE Transactions on Computers, 38(9): 1330–1333, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=29474.

Torng:1987:ISI

[Tor87] H. C. Torng. Introduction to the special issue on supercomputing. IEEE Transactions on Computers, C-36(12):1393–1394, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009491.

Towsley:1986:AMM

[Tow86] D. Towsley. Approximate models of multiple bus multiprocessor systems. IEEE Transactions on Computers, C-35(3):220–228, March 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676746.

Thomas:1987:IEC

[TP87] J. J. Thomas and S. R. Parker.

Implementing exact calculations in hardware. IEEE Transactions on Computers, C-36(6):764–768, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676969.

Taylor:1985:RFU

[TPSS85] F. J. Taylor, G. Papadourakis, A. Skavantzos, and A. Stouraitis. A radix-4 FFT using complex RNS arithmetic. IEEE Transactions on Computers, C-34(6):573–576, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009414.

Tung:1988:FAO

[TR88] C.-H. Tung and J. P. Robinson.

A fast algorithm for optimum syndrome space compression. IEEE Transactions on Computers, 37(2):228–232, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2153.

Truong:1988:PDF

[TRH^{+}88] T. K. Truong, I. S. Reed, I.- S. Hsu, H.-C. Shyu, and H. M. Shao. A pipeline design of a fast prime factor DFT on a finite field. IEEE Transactions on Computers, 37(3):266–273, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2163.

Trickey:1982:GLP

[Tri82] H. W. Trickey. Good layouts for pattern recognizers. IEEE Transactions on Computers, C-31(6):514–520, June 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676033.

Tsai:1989:AHR

[TRTN89] W. T. Tsai, C. V. Ramamoorthy, W. K. Tsai, and O. Nishiguchi. An adaptive hierarchical routing protocol. IEEE Transactions on Computers, 38(8):1059–1075, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30862.

Truong:1983:PAD

[TRYS83] T. K. Truong, I. S. Reed, C.- S. Yeh, and H. M. Shao. A parallel architecture for digital filtering using Fermat number transforms. IEEE Transactions on Computers, C-32 (9):874–877, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676339.

Tamir:1983:SMR

[TS83] Y. Tamir and C. H. Sequin. Strategies for managing the register file in RISC. IEEE Transactions on Computers, C-32(11):977–989, November 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676149.

Tamir:1984:DAS

[TS84a] Y. Tamir and C. H. Sequin. Design and application of self-testing comparators implemented with MOS PLA’s. IEEE Transactions on Computers, C-33(6):493–506, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676473.

Tuomenoksa:1984:TPS

[TS84b] D. L. Tuomenoksa and H. J.

Siegel. Task preloading schemes for reconfigurable parallel processing systems. IEEE Transactions on Computers, C-33 (10):895–905, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676350.

Taylor:1986:RSS

[TS86a] D. J. Taylor and C.-J. H.

Seger. Robust storage structures for crash recovery. IEEE Transactions on Computers, C-35(4):288–295, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676761.

See correction [TS87].

Tuomenoksa:1986:DOS

[TS86b] D. L. Tuomenoksa and H. J. Siegel. Determining an optimal secondary storage service rate for the PASM control system. IEEE Transactions on Computers, C-35(1):43–53, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676656.

Taylor:1987:CRS

[TS87] D. J. Taylor and C.-J. H. Seger. Correction to “Robust Storage Structures for Crash Recovery”. IEEE Transactions on Computers, C-36(2):256, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676893.

See [TS86a].

Thanawastien:1988:UCM

[TS88a] S. Thanawastien and P. K. Srimani. The universality of a class of modified singlestage shuffle/exchange networks. IEEE Transactions on Computers, 37(3):348–352, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2172.

Thuau:1988:OLM

[TS88b] G. Thuau and G. Saucier. Optimized layout of MOS cells. IEEE Transactions on Computers, 37(1):79–87, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=75140.

Tsao:1981:ECA

[Tsa81] Nai-Kuan Tsao. Error complexity analysis of algorithms for matrix multiplication and matrix chain product. IEEE Transactions on Computers, C-30(10):758–771, October 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675694.

Tsao:1983:SAE

[Tsa83] Nai-Kuan Tsao. A simple approach to the error analysis of division-free numerical algorithms. IEEE Transactions on Computers, C-32(4):343–351, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676235.

Tsin:1986:FLC

[Tsi86] Y. H. Tsin. Finding lowest common ancestors in parallel. IEEE Transactions on Computers, C-35(8):764–769, August 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676830.

Taylor:1988:AVR

[TSM88] G. F. Taylor, R. H. Steinvorth, and J. F. McDonald. An architecture for a video rate two-dimensional fast Fourier transform processor. IEEE Transactions on Computers, 37(9):1145–1148, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2269.

Thacker:1988:FMW

[TSS88] C. P. Thacker, L. C. Stewart, and E. H. Satterthwaite, Jr. Firefly: a multiprocessor workstation. IEEE Transactions on Computers, 37(8):909–920, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2243.

Tapia:1980:CSB

[TT80] M. A. Tapia and J. H. Tucker. Complete solution of Boolean equations. IEEE Transactions on Computers, C-29(7):662–665, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675639.

See correction [ATT81].

Tucker:1985:BIC

[TTB85] J. H. Tucker, M. A. Tapia, and A. W. Bennett. Boolean integral calculus for digital systems. IEEE Transactions on Computers, C-34(1):78–81, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676517.

Tokoro:1981:OM

[TTT81] M. Tokoro, E. Tamura, and T. Takizuka. Optimization of microprograms. IEEE Transactions on Computers, C-30(7):491–504, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675828.

Towsley:1982:WRA

[TV82] D. Towsley and G. Venkatesh.

Window random access protocols for local computer networks. IEEE Transactions on Computers, C-31(8):715–722, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676077.

Tang:1983:ETP

[TW83] D. T. Tang and L. S. Woo. Exhaustive test pattern generation with constant weight vectors. IEEE Transactions on Computers, C-32(12):1145–1150, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676175.

Twigg:1983:TMS

[Twi83] D. W. Twigg. Transposition of matrix stored on sequential file. IEEE Transactions on Computers, C-32(12):1185–1188, December 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676179.

Takagi:1987:LED

[TY87] Naofumi Takagi and Shuzo Yajima. On-line error-detectable high-speed multiplier using redundant binary representation and three-rail logic. IEEE Transactions on Computers, C-36(11):1310–1317, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009470.

Tyszer:1988:MFT

[Tys88] J. Tyszer. A multiple faulttolerant processor network architecture for pipeline computing. IEEE Transactions on Computers, 37(11):1414–1418, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8707.

Takagi:1985:HSV

[TYY85] N. Takagi, H. Yasuura, and S. Yajima. High-speed VLSI multiplication algorithm with a redundant binary addition tree. IEEE Transactions on Computers, C-34(9):789–796, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676634.

Tzeng:1988:RFT

[TYZ88] N.-F. Tzeng, P.-C. Yew, and C.- Q. Zhu. Realizing fault-tolerant interconnection networks via chaining. IEEE Transactions on Computers, 37(4):458–462, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2191.

Uhr:1982:CSC

[Uhr82]  L. Uhr.  Comparing serial

computers, arrays, and networks using measures of “active resources”. IEEE Transactions on Computers, C-31 (10):1022–1025, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675912.

Ulman:1983:SDI

[Ulm83]  Z. D. Ulman.  Sign detec-

tion and implicit-explicit conversion of numbers in residue arithmetic. IEEE Transactions on Computers, C-32(6):590–594, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676282.

Uyar:1988:DFR

[UR88] M. U. Uyar and A. P. Reeves. Dynamic fault reconfiguration in a mesh-connected MIMD environment. IEEE Transactions on Computers, 37 (10):1191–1205, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5981.

Upadhyaya:1988:ESD

[US88] S. J. Upadhyaya and K. K. Saluja. An experimental study to determine task size for rollback recovery systems. IEEE Transactions on Computers, 37(7):872–877, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2235.

Utku:1989:FPC

[USM89] S. Utku, M. Salama, and R. J.

Melosh. A family of permutations for concurrent factorization of block tridiagonal matrices. IEEE Transactions on Computers, 38(6):812–824, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24290.

Unger:1986:CSH

[UT86] S. H. Unger and Chung-Jen Tan. Clocking schemes for high-speed digital systems. IEEE Transactions on Computers, C-35(10):880–895, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676679.

Umeno:1987:NMR

[UT87] Hidenori Umeno and Shunji Tanaka. New methods for realizing plural near-native performance virtual machines. IEEE Transactions on Computers, C-36(9):1076–1087, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009538.

Uuger:1981:DET

[Uug81]  S. H. Uuger.  Double-edge-

triggered flip-flops. IEEE Transactions on Computers, C-30(6):447–451, June 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675811.

Uehara:1981:OLC

[UV81] T. Uehara and W. M. Vancleemput. Optimal layout of CMOS functional arrays. IEEE Transactions on Computers, C-30(5):305–312, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675787.

Vaishnavi:1982:CPE

[Vai82]  V. K. Vaishnavi.  Comput-

ing point enclosures. IEEE Transactions on Computers, C-31(1):22–29, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675881.

Vaishnavi:1984:MHB

[Vai84]  V. K. Vaishnavi.  Multidi-

mensional height-balanced trees. IEEE Transactions on Computers, C-33(4):334–343, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676438.

Vaishnavi:1986:HMH

[Vai86] V. K. Vaishnavi. On the height of multidimensional heightbalanced trees. IEEE Transactions on Computers, C-35 (9):773–780, September 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676837.

Vaishnavi:1989:MBB

[Vai89] V. K. Vaishnavi. Multidimensional balanced binary trees. IEEE Transactions on Computers, 38(7):968–985, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30849.

Valiant:1983:OTP

[Val83] L. G. Valiant. Optimality of a two-phase strategy for routing in interconnection networks. IEEE Transactions on Computers, C-32(9):861–863, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676335.

VanScoy:1980:PRC

[Van80] F. L. Van Scoy. The parallel recognition of classes of graphs. IEEE Transactions on Computers, C-29(7):563–570, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675627.

VanGils:1986:TMR

[Van86] W. J. Van Gils. A triple modular redundancy technique providing multiple-bit error protection without using extra redundancy. IEEE Transactions on Computers, C-35(7):623–631, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676803.

VanDam:1981:SHB

[VBH^{+}81] A. Van Dam, M. Barbacci, C. Halatsis, J. Joosten, and M. Letheren. Simulation of a horizontal bit-sliced processor using the ISPS architecture simulation facility. IEEE Transactions on Computers, C-30(7):513–519, July 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675830.

Vegdahl:1984:SPA

[Veg84] S. R. Vegdahl. A survey of proposed architectures for the execution of functional languages. IEEE Transactions on Computers, C-33(12):1050–1071, December 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676387.

Varshney:1984:SFD

[VH84] P. K. Varshney and C. R. P.

Hartmann. Sequential fault diagnosis of modular systems. IEEE Transactions on Computers, C-33(2):194–197, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676411.

Varshney:1982:AIT

[VHD82] P. K. Varshney, C. R. P. Hartmann, and J. M. De Faria, Jr. Application of information theory to sequential fault diagnosis. IEEE Transactions on Computers, C-31(2):164–170, February 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675963.

Velaardi:1984:SSF

[VI84] P. Velaardi and R. K. Iyer. A study of software failures and recovery in the MVS operating system. IEEE Transactions on Computers, C-33(6):564–568, June 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676482.

Volz:1987:TID

[VM87a] R. A. Volz and T. N. Mudge.

Timing issues in the distributed execution of Ada programs. IEEE Transactions on Computers, C-36(4):449–459, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676927.

Volz:1987:ILT

[VM87b] Richard A. Volz and Trevor N. Mudge. Instruction level timing mechanisms for accurate real-time task scheduling. IEEE Transactions on Computers, C-36(8):988–993, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009521.

Vasanthavada:1988:SFT

[VM88] N. Vasanthavada and P. N. Marinos. Synchronization of fault-tolerant clocks in the presence of malicious failures. IEEE Transactions on Computers, 37(4):440–448, April 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2188.

VonConta:1983:TON

[Von83] C. Von Conta. Torus and other networks as communication networks with up to some hundred points. IEEE Transactions on Computers, C-32(7):657–666, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676297.

Virupakshia:1983:SRT

[VR83] A. R. Virupakshia and V. C.

V. Pratapa Reddy. A simple random test procedure for detection of single intermittent fault in combinational circuits. IEEE Transactions on Computers, C-32(6):594–597, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676283.

Varman:1986:SOF

[VR86] P. J. Varman and I. V. Ramakrishnan. Synthesis of an optimal family of matrix multiplication algorithms on linear arrays. IEEE Transactions on Computers, C-35(11):989–996, November 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676700.

Varma:1989:FTR

[VR89a] A. Varma and C. S. Raghavendra. Fault-tolerant routing in multistage interconnection networks. IEEE Transactions on Computers, 38(3):385–393, March 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21125.

Varman:1989:OMM

[VR89b] P. J. Varman and I. V. Ramakrishnan. Optimal matrix multiplication on faulttolerant VLSI arrays. IEEE Transactions on Computers, 38 (2):278–283, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16505.

Varman:1984:RMM

[VRF84] P. J. Varman, I. V. Ramakrishnan, and D. S. Fussell. A robust matrix-multiplication array. IEEE Transactions on Computers, C-33(10):919–922, October 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676353.

Varanasi:1983:MPE

[VRP83] M. R. Varanasi, T. R. N. Rao, and Son Pham. Memory package error detection and correction. IEEE Transactions on Computers, C-32(9):872–874, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676338.

Vickers:1980:TGS

[VS80] V. E. Vickers and J. Silverman. A technique for generating specialized Gray codes. IEEE Transactions on Computers, C-29(4):329–331, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675573.

Vergis:1986:TCB

[VS86a] A. Vergis and K. Steiglitz.

Testability conditions for bilateral arrays of combinational cells. IEEE Transactions on Computers, C-35(1):13–22, January 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676653.

Vitter:1986:NCP

[VS86b] J. S. Vitter and R. A. Simons.

New classes for parallel complexity: a study of unification and other complete problems for P. IEEE Transactions on Computers, C-35(5):403–418, May 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676783.

Vassiliadis:1989:GPO

[VSH89] S. Vassiliadis, E. M. Schwarz, and D. J. Hanrahan. A general proof for overlapped multiplebit scanning multiplications. IEEE Transactions on Computers, 38(2):172–183, February 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=16494.

Vaidya:1982:DAI

[VSHM82] P. D. Vaidya, L. G. Shapiro, R. M. Haralick, and G. J. Minden. Design and architectural implications of a spatial information system. IEEE Transactions on Computers, C-31(10):1025–1031, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675913.

Vrsalovic:1988:PPC

[VSSG88] D. F. Vrsalovic, D. P. Siewiorek, Z. Z. Segall, and E. F. Gehringer. Performance prediction and calibration for a class of multiprocessors. IEEE Transactions on Computers, 37(11): 1353–1365, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8701.

Visvanathan:1981:DNC

[VSV81] V. Visvanathan and A. Sangiovanni-Vincentelli. Diagnosability of nonlinear circuits and systems — Part I: The dc case. IEEE Transactions on Computers, C-30(11):889–898, November 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675720.

Vu:1985:EIC

[Vu85] Thu Van Vu. Efficient implementations of the Chinese Remainder Theorem for sign detection and residue decoding. IEEE Transactions on Computers, C-34(7):646–651, July 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676602.

Vuillemin:1983:CLC

[Vui83]  J. Vuillemin.  A combina-

torial limit to the computing power of VLSI circuits. IEEE Transactions on Computers, C-32(3):294–300, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676221.

VanTilborg:1984:WSD

[VW84] A. M. Van Tilborg and L. D.

Wittie. Wave scheduling— decentralized scheduling of task forces in multicomputers. IEEE Transactions on Computers, C-33(9):835–844, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676500.

VanAken:1981:EPP

[VZ81] J. R. VanAken and G. L. Zick.

The expression processor: a pipelined, multiple- processor architecture. IEEE Transactions on Computers, C-30(8):525–536, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675837.

Vroomen:1989:CAS

[VZMBH89] L. Vroomen, P. Zsombor-Murray, P. Baracos, and R. Hudson. Comments on ‘algorithmic state machine design and automatic theorem proving: dual approaches to the same activity’. IEEE Transactions on Computers, 38(6):880–881, June 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24299.

Wang:1980:URM

[WA80] Ying Wang and A. A. Avizienis. A unified reliability model for fault-tolerant computers. IEEE Transactions on Computers, C-29(11):1002–1011, November 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675495.

Wong:1985:ABD

[WA85a] J. W. Wong and M. H. Ammar. Analysis of broadcast delivery in a Videotex system. IEEE Transactions on Computers, C-34(9):863–866, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676645.

Woo:1985:STS

[WA85b] Nam Sung Woo and A. Agrawala.

A symmetric tree structure interconnection network and its message traffic. IEEE Transactions on Computers, C-34(8):765–769, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676626.

Wah:1984:CSD

[Wah84] Benjamin W. Wah. A comparative study of distributed resource sharing on multiprocessors. IEEE Transactions on Computers, C-33(8):700–711, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009356.

Wainer:1988:GFL

[Wai88] M. Wainer. Generating fractallike surfaces on general purpose mesh-connected computers. IEEE Transactions on Computers, 37(7):882–886, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2238.

Walczak:1988:DFS

[Wal88] K. Walczak. Deductive fault simulation for sequential module circuits. IEEE Transactions on Computers, 37 (2):237–239, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2155.

Wang:1981:FTR

[Wan81] P. S.-P. Wang. Finite-turn repetitive checking automata and sequential/parallel matrix languages. IEEE Transactions on Computers, C-30(5):366–370, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675797.

Wang:1982:DTA

[Wan82]  Jinpo Wang.  Delay and

throughput analysis for computer communications with balanced HDLC procedures. IEEE Transactions on Computers, C-31(8):739–746, August 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676080.

Wang:1989:ADF

[Wan89] C. C. Wang. An algorithm to design finite field multipliers using a self-dual normal basis. IEEE Transactions on Computers, 38(10):1457–1460, October 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=35840.

Williams:1981:DLF

[WB81] T. W. Williams and N. C.

Brown. Defect level as a function of fault coverage. IEEE Transactions on Computers, C-30(12):987–988, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675742.

West:1988:CCN

[WB88] D. B. West and P. Banerjee.

On the construction of communication networks satisfying bounded fan-in of service ports. IEEE Transactions on Computers, 37(9):1148–1151, September 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2270.

Weste:1983:DTW

[WBA83] N. Weste, D. J. Burr, and B. D.

Ackland. Dynamic time Warp pattern matching using an integrated multiprocessing array. IEEE Transactions on Computers, C-32(8):731–744, August 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676311.

Winslow:1983:ADS

[WC83] L. E. Winslow and Yuan-Chieh Chow. The analysis and design of some new sorting machines. IEEE Transactions on Computers, C-32(7):677–683, July 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676299.

Wah:1984:PAD

[WC84] Benjamin W. Wah and Kuo-Liang Chen. A partitioning approach to the design of selection networks. IEEE Transactions on Computers, C-33(3):261–268, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676423.

Wagner:1987:PT

[WCM87] K. D. Wagner, C. K. Chin, and E. J. McCluskey. Pseudorandom testing. IEEE Transactions on Computers, C-36(3):332–343, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676905.

Wang:1989:RVW

[WCS89] M. Wang, M. Cutler, and S. Y. H. Su. Reconfiguration of VLSI/WSI mesh array processors with two-level redundancy. IEEE Transactions on Computers, 38(4):547–554, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21147.

Wold:1984:PPP

[WD84] E. H. Wold and A. M. Despain.

Pipeline and parallel-pipeline FFT processors for VLSI implementations. IEEE Transactions on Computers, C-33(5):414–426, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676458.

Watanuki:1983:EAC

[WE83] Osaaki Watanuki and Miloš D.

Ercegovac. Error analysis of certain floating-point on-line algorithms. IEEE Transactions on Computers, C-32(4):352–358, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676236.

Weglarz:1980:MSM

[Weg80]  J. Weglarz.  Multiprocessor

scheduling with memory allocation — a deterministic approach. IEEE Transactions on Computers, C-29(8):703–709, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675652.

Weglarz:1982:AR

[Weg82] J. Weglarz. Author’s reply. IEEE Transactions on Computers, C-31(9):898, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676106.

Wegner:1985:QEK

[Weg85] Lutz M. Wegner. Quicksort for equal keys. IEEE Transactions on Computers, C-34(4):362–367, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009387.

Weissberger:1980:LII

[Wei80] A. J. Weissberger. An LSI implementation of an intelligent CRC computer and programmable character comparator. IEEE Transactions on Computers, C-29(2):116–124, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675535.

Weide:1982:MUB

[Wei82] B. W. Weide. Modeling unusual behavior of parallel algorithms. IEEE Transactions on Computers, C-31(11): 1126–1130, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675930.

Weihl:1988:CBC

[Wei88] W. E. Weihl. Commutativitybased concurrency control for abstract data types. IEEE Transactions on Computers, 37 (12):1488–1505, December 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=9728.

Wensley:1985:FCR

[Wen85] J. H. Wensley. Further comments on “The Reliability of Periodically Repaired n − 1/n Parallel Redundant Systems”. IEEE Transactions on Computers, C-34(11):1068, November 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676541.

See [Can83, Nil84].

Wojcik:1980:DTV

[WF80a] A. S. Wojcik and Kwang-Ya Fang. On the design of threevalued asynchronous modules. IEEE Transactions on Computers, C-29(10):889–898, October 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675472.

Wu:1980:CMI

[WF80b] Chuan-Lin Wu and Tse-Yun Feng. On a class of multistage interconnection networks. IEEE Transactions on Computers, C-29(8):694–702, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675651.

Wu:1980:REI

[WF80c] Chuan-Lin Wu and Tse-Yun Feng. The reverse-exchange interconnection network. IEEE Transactions on Computers, C-29(9):801–811, September 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675679.

Wu:1981:USE

[WF81] Chuan-Lin Wu and Tse-Yun Feng. The universality of the shuffle-exchange network. IEEE Transactions on Computers, C-30(5):324–332, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675790.

Wann:1983:ACC

[WF83] D. F. Wann and M. A. Franklin. Asynchronous and clocked control structures for VSLI based interconnection networks. IEEE Transactions on Computers, C-32(3):284–293, March 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676220.

Wu:1982:SLN

[WFL82] Chuan-Lin Wu, Tse-Yun Feng, and Min-Chang Lin. Star: a local network system for real-time management of imagery data. IEEE Transactions on Computers, C-31(10):923–933, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675901.

Wagh:1980:NAD

[WG80] M. D. Wagh and H. Ganesh. A new algorithm for Discrete Cosine Transform of arbitrary number of points. IEEE Transactions on Computers, C-29(4):269–277, April 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675565.

Wei:1988:DDI

[WG88] Y.-H. Wei and J.-L. Gaudiot.

Demand-driven interpretation of FP programs on a data-flow multiprocessor. IEEE Transactions on Computers, 37(8):946–966, August 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2246.

Walters:1981:SDC

[WGT81] S. M. Walters, F. G. Gray, and R. A. Thompson. Selfdiagnosing cellular implementations of finite-state machines. IEEE Transactions on Computers, C-30(12):953–959, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675734.

Warlick:1980:HSS

[WH80a] W. W. Warlick, Jr. and J. E. Hershey. High-speed Msequence generators. IEEE Transactions on Computers, C-29(5):398–400, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675590.

Wing:1980:CMP

[WH80b] O. Wing and J. W. Huang. A computation model of parallel solution of linear equations. IEEE Transactions on Computers, C-29(7):632–638, July 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675634.

Watterson:1988:MRC

[WH88] J. W. Watterson and J. J. Hallenbeck. Modulo 3 residue checker: new results on performance and cost. IEEE Transactions on Computers, 37(5):608– 612, May 1988. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
4612.

Wada:1984:ATO

[WHT84] K. Wada, K. Hagihara, and N. Tokura. Area-time optimal fast implementation of several functions in a VLSI model. IEEE Transactions on Computers, C-33(5):455–462, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676463.

Wong:1984:LSS

[WI84] F. S. Wong and M. R. Ito. A loop-structured switching network. IEEE Transactions on Computers, C-33(5):450–455, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676462.

Wittie:1981:CSL

[Wit81] L. D. Wittie. Communication structures for large networks of microcomputers. IEEE Transactions on Computers, C-30(4):264–273, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675774.

Wah:1985:RSL

[WJ85] B. Wah and J. Juang. Resource scheduling for local computer systems with a multiaccess network. IEEE Transactions on Computers, C-34(12):1144–1157, December 1985. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wallach:1980:BPM

[WK80] Y. Wallach and V. Konrad. On block-parallel methods for solving linear equations. IEEE Transactions on Computers, C-29(5):354–359, May 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675585.

Wirsching:1981:CCN

[WK81] J. E. Wirsching and T. Kishi. CONET: a connection network model. IEEE Transactions on Computers, C-30(4):298–301, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675779.

Wu:1981:CSI

[WL81] S. B. Wu and M. T. Liu. A cluster structure as an interconnection network for large multimicrocomputer systems. IEEE Transactions on Computers, C-30(4):254–264, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675773.

Wang:1989:NIB

[WLC89] Y.-L. Wang, R. C. T. Lee, and J. S. Chang. The number of intersections between two rectangular paths. IEEE Transactions on Computers, 38(11):1564–1571, November 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=42126.

Wah:1984:MMA

[WM84] B. W. Wah and Y. W. E.

Ma. MANIP — A multicomputer architecture for solving combinatorial extremum-search problems. IEEE Transactions on Computers, C-33(5):377–390, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676453.

Wang:1985:LSD

[WM85] Yung-Terng Wang and R. J. T.

Morris. Load sharing in distributed systems. IEEE Transactions on Computers, C-34(3):204–217, March 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676564.

Wang:1986:CLF

[WM86] Laung-Terng Wang and E. J. McCluskey. Condensed linear feedback shift register (LFSR) testing — A pseudoexhaustive test technique. IEEE Transactions on Computers, C-35(4):367–370, April 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676772.

Wang:1988:LFS

[WM88a] L.-T. Wang and E. J. Mc-Cluskey. Linear feedback shift register design using cyclic codes. IEEE Transactions on Computers, 37(10):1302–1306, October 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5994.

Wong:1988:BSU

[WM88b] W. S. Wong and R. J. T. Morris. Benchmark synthesis using the LRU cache hit function. IEEE Transactions on Computers, 37(6):637–645, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2202.

Whiteside:1984:PJD

[WOH84] R. A. Whiteside, N. S. Ostlund, and P. G. Hibbard. A parallel Jacobi diagonalization algorithm for a loop multiple processor system. IEEE Transactions on Computers, C-33(5):409–413, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676457.

Worlton:1981:CPR

[Wor81]  J. Worlton.  Comments on

“Parallelism and Representation Problems in Distributed Systems”. IEEE Transactions on Computers, C-30(8):608–609, August 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675850.

See [FH80].

Williams:1982:DTS

[WP82] T. W. Williams and K. P. Parker.

Design for testability — A survey. IEEE Transactions on Computers, C-31(1):2–15, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675879.

Wu:1981:SMC

[WR81] A. Y. Wu and A. Rosenfeld. SIMD machines and cellular d-graph automata. IEEE Transactions on Computers, C-30(5):370–372, May 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675798.

Wang:1984:PST

[WR84] Guang Xing Wang and G. R. Redinbo. Probability of state transition errors in a finite state machine containing soft failures. IEEE Transactions on Computers, C-33(3):269–277, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676424.

Warpenburg:1982:SIR

[WS82a] M. R. Warpenburg and L. J. Siegel. SIMD image resampling. IEEE Transactions on Computers, C-31(10):934–942, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675902.

Wei:1982:EME

[WS82b] M. C. Wei and H. A. Sholl. An expression model for extraction and evaluation of parallelism in control structures. IEEE Transactions on Computers, C-31(9):851–863, September 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676100.

Weiss:1984:IIL

[WS84] S. Weiss and J. E. Smith. Instruction issue logic in pipelined supercomputers. IEEE Transactions on Computers, C-33 (11):1013–1022, November 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676375.

Woodbury:1988:PMM

[WS88] M. H. Woodbury and K. G. Shin.

Performance modeling and measurements of real time multiprocessors with time-shared buses. IEEE Transactions on Computers, 37(2):214–224, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2151.

White:1984:VBF

[WSR84] S. W. White, N. R. Strader II, and V. T. Rhyne. A VLSIbased I/O formatting device. IEEE Transactions on Computers, C-33(2):140–149, February 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676406.

Wang:1985:VAC

[WTS^{+}85] C. C. Wang, T. K. Troung, H. M. Shao, L. J. Deutsch, J. K. Omura, and I. S. Reed. VLSI architectures for computing multiplications and inverses in GF(2^{m}). IEEE Transactions on Computers, C-34(8):709–717, August 1985. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676616.

Wu:1987:TRF

[Wu87a] Chwan-Chia Wu. Time redundant fault-location in bitsliced ALU’s. IEEE Transactions on Computers, C-36 (11):1387–1389, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009483.

wu:1987:FDS

[wu87b]  I-Chen wu.  A fast 1-D

serial-parallel systolic multiplier. IEEE Transactions on Computers, C-36(10):1243–1247, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676865.

Wustmann:1981:CAF

[Wus81] G. Wustmann. Comments on “Autocorrelation Function of Sequential M-Bit Words Taken from an N-Bit Shift Register (PN) Sequence”. IEEE Transactions on Computers, C-30(3): 241, March 1981. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675763. See [Mor80].

Wustmann:1982:AFF

[Wus82]  G. Wustmann.  Autocorre-

lation function of filtered plevel maximal-length sequences. IEEE Transactions on Computers, C-31(1):75–77, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675888.

Wittle:1980:MDO

[WV80] L. D. Wittle and A. M. Van Tilborg. MICROS, A distributed operating system for Micronet, A reconfigurable network computer. IEEE Transactions on Computers, C-29 (12):1133–1144, December 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675518.

Wijshoff:1987:LSS

[WV87] H. A. G. Wijshoff and J. Van Leeuwen. On linear skewing schemes and d-ordered vectors. IEEE Transactions on Computers, C-36(2):233–239, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676887.

Wijshoff:1985:SPS

[WvL85] Harry A. G. Wijshoff and Jan van Leeuwen. The structure of periodic storage schemes for parallel memories. IEEE Transactions on Computers, C-34(6):501–505, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009402.

Wojciechowski:1983:ADM

[WW83] W. S. Wojciechowski and A. S. Wojcik. Automated design of multiple-valued logic circuits by automatic theorem proving techniques. IEEE Transactions on Computers, C-32 (9):785–798, September 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676328.

Whang:1984:SAP

[WWS84] Kyu-Young Whang, G. Wiederhold, and D. Sagalowicz. Separability —an approach to physical database design. IEEE Transactions on Computers, C-33(3):209–222, March 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676418.

Wu:1987:RSP

[WWSW87] Ying-Fung Wu, P. Widmayer, M. D. F. Schlag, and C. K. Wong. Rectilinear shortest paths and minimum spanning trees in the presence of rectilinear obstacles. IEEE Transactions on Computers, C-36(3):321–331, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676904.

Wang:1989:FFT

[WZ89] P. S. P. Wang and Y. Y. Zhang.

A fast and flexible thinning algorithm. IEEE Transactions on Computers, 38(5):741–745, May 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=24276.

Xu:1985:DIF

[XS85] Shiyi Xu and Stephen Y. H.

Su. Detecting I/O and internal feedback bridging faults. IEEE Transactions on Computers, C-34(6):553–557, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009408.

Yalamanchili:1987:CAP

[YA87] S. Yalamanchili and J. K. Aggarwal. A characterization and analysis of parallel processor interconnection networks. IEEE Transactions on Computers, C-36(6):680–691, June 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676961.

Yamamoto:1980:MMI

[Yam80]  M. Yamamoto.  A method

for minimizing incompletely specified sequential machines. IEEE Transactions on Computers, C-29(8):732–736, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675655.

Yang:1989:ACC

[YBL89] Q. Yang, L. N. Bhuyan, and B.- C. Liu. Analysis and comparison of cache coherence protocols for a packet-switched multiprocessor. IEEE Transactions on Computers, 38(8):1143–1153, August 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=30868.

Yamamoto:1988:RBV

[YF88] Y. Yamamoto and S. Fujita. Relationship between P - valued majority functions and P -valued threshold functions. IEEE Transactions on Computers, 37(11):1442–1445, November 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=8714.

Yu:1987:APD

[YGZ^{+}87] C. T. Yu, Keh-Chang Guh, Weining Zhang, M. Templeton, D. Brill, and A. L. P. Chen. Algorithms to process distributed queries in fast local networks. IEEE Transactions on Computers, C-36(10):1153–1164, October 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676856.

Yau:1984:POC

[YH84] S. S. Yau and Wonmo Hong.

Performance optimization of a CSMA protocol for local computer networks. IEEE Transactions on Computers, C-33 (9):812–817, September 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676497.

Yanney:1986:DRF

[YH86] R. M. Yanney and J. P.

Hayes. Distributed recovery in fault-tolerant multiprocessor networks. IEEE Transactions on Computers, C-35 (10):871–879, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676678.

Yoder:1989:SWR

[YJ89] M. A. Yoder and L. H. Jamieson.

Simulation of a word recognition system on two parallel architectures. IEEE Transactions on Computers, 38(9): 1269–1284, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=29466.

Yamaguchi:1982:PLP

[YK82] K. Yamaguchi and T. L. Kunil. PICCOLO logic for a picture database computer and its implementation. IEEE Transactions on Computers, C-31 (10):983–996, October 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675907.

Yu:1988:ODS

[YKL88] P. S. Yu, C. M. Krishna, and Y.- H. Lee. Optimal design and sequential analysis of VLSI testing strategy. IEEE Transactions on Computers, 37(3):339–347, March 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=2171.

Yew:1981:ECN

[YL81] Pen-Chung Yew and D. H. Lawrie. An easily controlled network for frequently used permutations. IEEE Transactions on Computers, C-30(4):296–298, April 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675778.

Young:1985:SMC

[YM85] Ming Huei Young and Saburo Muroga. Symmetric minimal covering problem and minimal PLA’s with symmetric variables. IEEE Transactions on Computers, C-34(6):523–541, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009404.

Yamakawa:1986:CMF

[YM86a] T. Yamakawa and T. Miki. The current mode fuzzy logic integrated circuits fabricated by the standard CMOS process. IEEE Transactions on Computers, C-35(2):161–167, February 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676734.

Yang:1986:FIA

[YM86b] Che-Liang Yang and Gerald M.

Masson. A fault identification algorithm for ti-diagnosable systems. IEEE Transactions on Computers, C-35(6):503–510, June 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009427.

Yang:1987:NMH

[YM87a] Che-Liang Yang and G. M.

Masson. A new measure for hybrid fault diagnosability.

IEEE Transactions on Computers, C-36(3):378–383, March 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676913.

Yang:1987:GHF

[YM87b] Che-Liang Yang and Gerald M. Masson. A generalization of hybrid fault diagnosability. IEEE Transactions on Computers, C-36(11):1369–1374, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009479.

Yamamoto:1988:MSC

[YM88a] Y. Yamamoto and M. Mukaidono. Meaningful special classes of ternary logic functions-regular ternary logic functions and ternary majority functions. IEEE Transactions on Computers, 37(7):799–806, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2225.

Yang:1988:HFD

[YM88b] C.-L. Yang and G. M. Masson. Hybrid fault diagnosability with unreliable communication links. IEEE Transactions on Computers, 37(2):175–181, February 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2146.

Yang:1986:FII

[YML86] Che-Liang Yang, G. M. Masson, and R. A. Leonetti. On fault isolation and identification in t1/t1-diagnosable systems. IEEE Transactions on Computers, C-35(7):639–643, July 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676805.

Yamada:1983:CDL

[YN83] T. Yamada and T. Nanya.

Comments on “Detection and Location of Input and Feedback Bridging Faults Among Input and Output Lines”. IEEE Transactions on Computers, C-32(5):511–512, May 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676266.

Yamada:1984:SFT

[YN84] Teruhiko Yamada and Takashi Nanya. Stuck-at fault tests in the presence of undetectable bridging faults. IEEE Transactions on Computers, C-33(8):758–761, August 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009367.

Yen:1982:MIS

[YPD82] D. W. L. Yen, J. H. Patel, and E. S. Davidson. Memory interference in synchronous multiprocessor systems. IEEE Transactions on Computers, C-31(11):1116–1121, November 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675928.

Yeh:1983:SCM

[YPD83] P. C. C. Yeh, J. H. Patel, and E. S. Davidson. Shared cache for multiple-stream computer systems. IEEE Transactions on Computers, C-32(1):38–47, January 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676122.

Yeh:1984:SMF

[YRT84] C.-S. Yeh, I. S. Reed, and T. K. Truong. Systolic multipliers for finite fields GF(2^{m}). IEEE Transactions on Computers, C-33(4):357–360, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676441.

Youn:1989:ILB

[YS89] H. Y. Youn and A. D. Singh. On implementing large binary tree architectures in VLSI and WSI. IEEE Transactions on Computers, 38(4):526–537, April 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=21145.

Yew:1987:DHS

[YTL87] Pen-Chung Yew, Nian-Feng Tzeng, and D. H. Lawrie. Distributing hot-spot addressing in large-scale multiprocessors. IEEE Transactions on Computers, C-36(4):388–395, April 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676921.

Yasuura:1982:PES

[YTY82] H. Yasuura, N. Takagi, and S. Yajima. The parallel enumeration sorting scheme for VLSI. IEEE Transactions on Computers, C-31(12):1192–1201, December 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675943.

Yuen:1980:NDC

[Yue80] C. K. Yuen. Negabinary A/D conversion. IEEE Transactions on Computers, C-29(8):740–741, August 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675659.

Ye-Wei:1986:CDF

[YWW86] Liang Ye-Wei and Jin Wei. Comments on “Detection of Faults in Programmable Logic Arrays”. IEEE Transactions on Computers, C-35(10):930–931, October 1986. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676687.

Yen:1985:DCP

[YYF85] W. C. Yen, D. W. L. Yen, and King-Sun Fu. Data coherence problem in a multicache system. IEEE Transactions on Computers, C-34(1):56–65, January 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676515.

Yang:1988:CPM

[YZ88] Q. Yang and S. G. Zaky. Communication performance in multiple-bus systems. IEEE Transactions on Computers, 37 (7):848–853, July 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        2230.

Zakharov:1984:PAP

[Zak84] Vasilii Zakharov. Parallelism and array processing. IEEE Transactions on Computers, C-33(1):45–78, January 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009314.

Zaks:1985:ODA

[Zak85] Shmuel Zaks. Optimal distributed algorithms for sorting and ranking. IEEE Transactions on Computers, C-34(4):376–379, April 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009390.

Zaccone:1987:ENP

[ZB87] Richard J. Zaccone and Jesse L.

Barlow. Eliminating the normalization problem in digit on-line arithmetic. IEEE Transactions on Computers, C-36(1):36–46, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009447.

Zurawski:1981:DHS

[ZG81] J. H. P. Zurawski and J. B.

Gosling. Design of highspeed digital divider units. IEEE Transactions on Computers, C-30(9):691–699, September 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675869.

Zurawski:1987:DHS

[ZG87] J. H. P. Zurawski and J. B.

Gosling. Design of a high-speed square root multiply and divide unit. IEEE Transactions on Computers, C-36(1):13–23, January 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009445.

Zak:1985:PDS

[ZH85] Stanislaw H. Zak and Kai Hwang. Polynomial division on systolic arrays. IEEE Transactions on Computers, C-34(6):577–578, June 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009415.

Zhiwei:1984:MLF

[Zhi84] Xu Zhiwei. Multivalued logic and fuzzy logic–their relationship, minimization, and application to fault diagnosis. IEEE Transactions on Computers, C-33(7):679–681, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009344.

Zhou:1988:NBS

[Zho88]  B. B. Zhou.  A new bit-

serial systolic multiplier over GF(2^{m}). IEEE Transactions on Computers, 37(6):749–751, June 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=2216.

Zarowski:1989:PCA

[ZMC89] C. J. Zarowski, R. D. McLeod, and H. C. Card. Primitive cellular automata, threshold decomposition, and ranked order operations. IEEE Transactions on Computers, 38(1):148–149, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8738.

Zeman:1980:HSM

[ZN80] J. Zeman and H. T. Nagle, Jr. A high-speed microprogrammable digital signal processor employing distributed arithmetic. IEEE Transactions on Computers, C-29(2):134–144, February 1980. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675537.

Zhao:1987:PSU

[ZRS87] Wei Zhao, Krithi Ramamritham, and John A. Stankovic. Preemptive scheduling under time and resource constraints. IEEE Transactions on Computers, C-36(8):949–960, August 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009518.

Zang:1987:RML

[ZW87] Wenlong Zang and J. K. Wolf. Rate 1/2 and 2/3 majority logic decodable binary burst error-correcting codes. IEEE Transactions on Computers, C-36(2):250–252, February 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1676891.

Zwaenepoel:1985:IPP

[Zwa85] W. Zwaenepoel. Implementation and performance of pipes in the V-System. IEEE Transactions on Computers, C-34(12): 1174–1178, December 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).